1

A Bibliography of Publications about Multithreading

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

31 August 2024 Version 3.200

Title word cross-reference

#4 [Pet00].

+ [BMV03]. 2 [TKHG04]. 3 [KSB^{+}08, PYP^{+}10]. _{cyclical} [YLLS16]. D^{3} [Evr01]. F^{2} [BCS11]. k [ZTN19]. LU [VD08]. N [ZJFA09]. π [III01]. QR [But13, GKK09, Myl22, VD08]. QZ [Myl22].

-based [Rót19]. -Calculus [III01]. -Machine [Evr01]. -way [ZJFA09].

.NET [Rob03, Tim03, DHR^{+}01, Rei01].

/multi [Taf13]. /multi-threaded [Taf13].

’01 [USE01].

1 [BM91, McM98a]. 1.0 [Sen19]. 1003.4 [GL91]. 1019 [Myl22]. 11 [ND16]. 11th [IEE94a, IEE94d]. ’12 [Hol12]. 16-20 [IEE92]. 162 [Stu95]. 1991 [Ano91, Ano94e]. 1993 [ACM93b]. 1994 [ACM94a, ACM94d, Hon94, IEE94c].

2 [BÇG14, DN94, Kan94, Kel94a, Kel94b, Mil95, Rei95, Ric91, Rod94, Sri93, WCW^{+}04b, WCW^{+}04c, WCW^{+}04d]. 2.0 [BO01, LPD^{+}11]. 2.6 [McM97]. 2000 [Ano99]. 2001 [ACM01]. 2003 [RM03, ACM03, AS14]. 2010 [Egg10]. 2011 [LCK11]. 2012 [Hol12]. 20th [IEE95]. 21st [ACM94b]. 22nd [ACM95b]. 25th [ACM98b, ACM98c]. 2k [USE00b]. 2nd [Ano94d, USE98a].

3.0 [Ano97c, Bra97, BRM03, MRGB91]. 32-Way [KAO05]. 35th [Gol94]. 3D [Ano97b, Loe97].

4 [BAM93, SKS^{+}92]. 4th [Ass96].

5 [FLR98]. 5th [Cha05].

64-bit [Kus15, SBKK99]. 6th [DLM99].

7th [USE00b].

80 [Bri89]. 821 [HBG02].

9 [Pre90, Yam96]. ’90 [IEE90, GOT03]. ’91 [Wat91, Ano94e]. 915 [Dav11]. ’92 [IEE92]. 925 [YFF^{+}12]. ’93 [ACM93c, IEE93]. ’94 [ACM94a, ACM94b, ACM94c, Hon94, IEE94a, IEE94c, IEE94d]. ’95 [ACM95b, KR01b]. ’96 [ACM96, Lak96, Ass96]. ’98 [ACM98b]. ’99 [ACM99a, Hol99b]. 9X [GMB93].

Abstract [CSS^{+}91b, CGSV93, DV99, KPP12, LMA^{+}16, MJF^{+}10, Ném00, CSS^{+}91a, CSS^{+}91c, Dil00, VDBN98, ZJFA09]. Abstraction [KI16, YDLW20, Bak95b, GSK^{+}18, GPR11, ZSJ06]. AC [BGK94a, BGK94b]. Accelerate [JLA16]. Accelerating [BAZ^{+}19, FNS^{+}22, LS11, MSU^{+}16, SMQP09, VGK^{+}10a, VGK^{+}10b]. acceleration [JSMP13, NBMM12]. Accelerators [MCS^{+}22, NTR16, SGLGL^{+}14]. Access [BP19, CmL21, Kle00, Spe94, VB00, AKSD16, APX12, CDD^{+}10, Hig97, KFG15, MVY05, Sch89]. access/execute [APX12]. accesses [DTK^{+}15]. accessibility [SSkP^{+}07]. Accounting [LMA^{+}16, EE09b]. Accuracy [SHK15, TO10]. Accurate [CPT08, VTSM12]. Achieving [AHW02, SP05, KGGK09, WTKW08]. ACM [ACM93b, RM03, IEE02, ACM98b, ACM99a]. ACM/IEEE [ACM98d]. across [ZP04].

Activation [KG94]. Activations [ABLL92, BVM19, DNR00, SS95]. Active [BKI06, BDJ06, Pla02, Ten98, Wei98a, SD95, WHJ^{+}95]. Actor [BFSK20]. actors [Bri89]. actually [Pra95c]. Ada [ACM93c, Bar09, Dil93, GMB93, KPPÉR06, KR01b]. ADAM [Far96]. adaptable [LLLC15]. Adaptation [CMBAN08]. Adaptative [ABN00]. Adaptive [ALHH08, HBTG98, HTDL18, KI95, LYH16, PM14, RCC12, STY99, SLG04, SLG06, SGS14, TLGM17, WXAL22, WYT^{+}20, ZWL15, BS06, Chr95a, Chr95b, Chr96, SLGZ99, TKHG04, ZLW^{+}16]. Adding [Ply89, Ric99, McM97]. Address [CLFL94, PWL^{+}11, CKZ12, Lie94]. Addresses [ZWGX22]. Addressing [WA08, CKD94, ZSB^{+}12]. ADL [BVL09]. Advanced [BGG95, BP19, GBG95, Hei03, BZ07, GBB^{+}05]. Advances [IEE97, JHM04, KKDV03, DLM99]. Advantage [Wei97]. Adversarial [FF10]. Affinity [HLH16, NAAL01]. Age [Cro98]. agent [Way95]. Agents [CWHB03, CR02, Way95, BDF98]. Aggregate [TGO99, TGO00]. Aggressive [Myl22]. AGNI [RBPM00]. Agnostic [SLJ^{+}18]. agreement [GMW09]. Aid [Wei97]. aided [MCRS10]. aids [Mat97]. Air [MPD04]. Al [TLA^{+}02]. Albuquerque [Ano94e]. Algebra [KLDB09, NBS^{+}15, PHCR09, YSY^{+}09]. Algebraic [ACM94c, Lak96, MR09, Wat91]. Algorithm [AT16, ABC^{+}09, CNZS17, HH11, MP13, Myl22, OR12, Rót19, TT03, WYT^{+}20, ZBS15, AG21, BKK17, GKK09, KGPH12, KNPS16, LCH^{+}08, Mah11, Mah13, SCG95, TKHG04, Dav11, HBG02, Myl22, YFF^{+}12]. Algorithmic [Lei97, BBH^{+}17]. Algorithms [BP05, EJRB13, FS96, LA93, MNG16, NSP^{+}14, Pan99, QOIM^{+}12, TTKG02, TCF22, YMR93b, Bar09, ÇFG^{+}12, CLRS09, FR95, GK05, Lei97, Lep95, NFBB17, QOQOV^{+}09, RRMJ12, YM92, YMR93a, Li05]. algorithms-by-blocks [QOQOV^{+}09]. Algorithms-by-Tiles [QOIM^{+}12]. aliasing [NA07]. Aligned [YWJ03]. alignment [KGPH12]. Allaire [Hig97]. Alleria [BP19]. Alleviate [BD00]. Alloc [KSU94]. Allocating [SEP96]. Allocation [MVZ93, Nak01, ZWL15, EFJM07, LLL10, Mic04, ZP04]. Allocations [LK20]. Allocator [BMBW00b, BMBW00a, BMBW00c]. almost [SV21]. Alpha [Ano00b]. alphabet [KNPS16]. alphabet-independent [KNPS16]. Alternating [CYYL18]. alternative [SV96c, SV96a, SV96b]. Alternatives [MB99, OA19, MKR02]. Alto [ACM01]. ALU [KDM^{+}98]. always [DWS^{+}12]. always-on [DWS^{+}12]. Amdahl [CN14, NZ17]. Among [CB16, HMC95, SJ95]. analysing [NJK16, PV06]. Analysis [AKS06, BCZY16, BVM19, BE12, BE13, BTL^{+}19, BBC^{+}00, BLG01, BNH01, CGS^{+}20, CC04, CH95, CGL92a, CGL92b, DSR15, EJRB13, Hai97b, Hol12, HLH16, LCK11, LML00, LHG^{+}16, NBM93, PFP^{+}22, REL00b, Rin01, RR99, SBCV90, SKV21, TAM^{+}08, VP16, Yoo96a, Zub02, AC09, ACC^{+}03, BGZ97, BBH^{+}17, BPSH05, BBM09, CHH^{+}03, CS12, CVJL08, Cor00, GBCS07, HEJ09, JPSN09, KTK12, KC09, Lei97, LBH12, LBE^{+}98, Met95, NWT^{+}07, PFH06, PL03, REL00a, REL00c, RS07, SR01a, SMK10, SRA06, SB80, TMC09, TR14, Wan94, WS06, WP10, WOKH96, WTH^{+}12, dB09, vPG03]. Analytic [Squ94, MAF19]. Analytical [DKF94, SV19, VT96, SBC91]. Analytics [JGS^{+}19, MCS^{+}22, LTL^{+}16]. analyze [LMC14]. analyzer [Fer13, HLB90]. Analyzing [ARA20, BB20, HRH08, Kor89, RHH10, TMCP10]. anatomy [Rei95]. Android [MKM14]. animation [WQLJ18]. Annotations [BM94, Wei98b, AGN09]. Annual [ACM93a, ACM98c, Gol94, Ass96, USE00a, ACM93b, USE96, USE98b]. anomalies [Sch89]. Anomaly [KW17]. antipatterns [BPSH05]. Antonio [USE92a]. any [Hig97, Mar07]. API [Ano00b, BFSK20, BDN02, DM98, LPD^{+}11, Van97a]. APL [CJ91]. applets [McM96c]. Application [AMRR98, CA20, HTDL18, KZTK15, KSU94, PG92, PLT^{+}15, SV19, TKA^{+}01, TAM^{+}08, Yas95, DWYB10, EJK^{+}96, HDT^{+}13, LVN10, LZ07, MRGB91, MKR10, Pha91, Pra95c, SE12, SS95, TKA^{+}02, ZJS^{+}11]. Application-Level [HTDL18, KSU94, PLT^{+}15, HDT^{+}13, LZ07, ZJS^{+}11]. Applications [Ano00c, AZG17, AKP99, BKI06, BMBW00b, BNH01, Cha05, Chl15a, DVAE18, DSAD^{+}18, DS16, Don02, Dru95, EV01, FF20, FURM00c, HC17, HMCP16, HWZ00, JYE^{+}16, JLA16, KMjC02, KEMC22, KRH98, LWSB19, LPK16, Lar97, MGI14, MG15, PCPS15, PFP^{+}22, PWL^{+}11, Pul00, RD96, DFC^{+}19, SGM^{+}97, Sod02, Ten02, Tet94, TSV12, TLGM17, VCM19, VP16, Vol93, WJA^{+}19, YG10, ZJS12, Ano92a, Ano92b, Ano94b, ASSS19, AAKK08, BWDZ15, BBFW03, BGZ97, BMBW00a, BMBW00c, BW97, DSEE13, BPSH05, BMV03, CB89, CB90, CSB00, CS12, CDP^{+}21, FM92, FURM00a, FURM00b, GS02, GCRD04, HLB90, ISS98, JSMP12, JSMP13, KVN^{+}09, LSW^{+}18, MLCW11, MKM14, MKIO04, MLC04, MT02a, MT02b, MT02c, MKK99, MKR10, NR06, Omm04, PJZA07, RČV^{+}10, Rei95, San04, SSN10, SKP^{+}02, TMC09, TMCP10, TPZ21]. applications [TP18, VIA^{+}05, VGK^{+}10a, VGK^{+}10b, WCZ^{+}07, WXAL22, WT10, WOKH96, XMN99, YZ14, kSYHX^{+}11, ZKR^{+}11, Len95]. Applied [CmL19a]. apply [NZ17]. Applying [VTSL12, MT02a, MT02b, MT02c]. Apprendre [Swi09]. Approach [AZG17, BBSG11, CJW^{+}15, ES97, FKT96, GMR98, KKW14, KS16, ND16, RCM^{+}16, TY97, VSDK09, WS08, Wei98b, YLLS16, BWDZ15, DHM^{+}12, GB20, LGN92, LZW17, LZSS19, LZL^{+}14, MS03, RCM^{+}12, SCZM00, TP18]. Approaches [BLPV04, MB07]. Approximate [HFV^{+}12, GEG07, GE08, KGPH12]. Apps [PCM16]. April [Ano00a, Ano03, USE01]. arbitrary [BGC14]. Arbitration [CFC^{+}22]. Arc [CNZS17]. Arc-Weighted [CNZS17]. ARCH [Ada98]. Architectural [ACM94d, HEMK17, IAD^{+}94, KC99, ME15, BS06, CMF^{+}13, Fan93, WHG07]. Architecture [ACM98c, BBD^{+}91, BVL09, BTE98, Car89b, CL95, DS09, DO95, EBKG01, For97, Gao93, GK94, GHG^{+}98, GV95, GN92, HTZ^{+}97, HMNN91, HHOM91, HHOM92, KBH^{+}04a, KBH^{+}04b, KIAT99, Man91, MM01, MB99, PVS^{+}17, PTMB09, PKB^{+}91, PS01, REL00b, RS08, SLJ^{+}18, SCL05, SHK15, SSYG97, SKK^{+}01, SZ02, TKA^{+}01, VK99, ZL10, ACC^{+}03, AAHF09, Ano97b, BT01, Bon13, CMF^{+}13, CL94, CHH^{+}03, Cho92, Don92, DGSB20, Dub95, Evr01, Far96, Fuj97, Gal94, GDSA^{+}17, GL98a, Gol96, HF88, HKN^{+}92, HMN^{+}92, I^{+}94, KHP^{+}95, KT99, Loi95, Mah13, MK12, Ném00, NPA92, PYP^{+}10, PDP^{+}13, PWD^{+}12, RGT17, REL00a, REL00c, RCDG06, SWYC94, Sod02, TPZ21, TNB^{+}95, Tsa97b, UZU00, Wan94, WCC^{+}07, YZ07, Yan97, CH04]. Architecture-Agnostic [SLJ^{+}18]. Architectures [AT16, ABLM19, Day92a, Day92b, HD02, GGB93a, GN00, HPA^{+}15, HMLB16, Hol98d, IXS18, IBST01, JLS99, KTR^{+}04, LLKS12, LB92, LH94, LG06, LDT^{+}16, MS02, MN00, MSU^{+}16, NGGA94, QOIM^{+}12, RLJ^{+}09, SGM^{+}97, TG99, THA^{+}12, Tra91, TJY98, TSV12, VCM19, WG94, XWG^{+}14, ZAK01, ABD^{+}12, ABC^{+}15, ABC^{+}09, BIK^{+}11, BS10a, BH95, CML00, ÇFG^{+}12, Cat94, DTR18, FTAB14, GGB93b, GK05, Gil94, GL98b, HFV^{+}12, ICH^{+}10, JMS^{+}10, LMC14, Lu94, MLCW11, MLC04, Mus09, OCRS07, PT91, PPA^{+}13, PJZA07, PHCR09, RHH10, RKBH11, SBCV90, Sch98, Sha95b, SLG06, Squ94, SMQP09, SKA01, TE94a, The95, TKHG04, ZTN19]. Area [AMPH09, FGT96, Par91]. Area-efficiency [AMPH09]. Aren’t [Sut99]. Ariadne [MR98]. arising [ARvW03]. Array [BVP^{+}19, GS06, LHS16, PDMM16]. Arrays [BWXF05, SHW19, AR19]. arrow [GE08]. arrow-type [GE08]. Art [MP13, I^{+}94]. artificial [KU17]. ASAT [SEP96]. Ashes [Thr99]. ASN [CJW^{+}15]. Aspects [SB80]. ASPLOS [ACM94d]. ASPLOS-VI [ACM94d]. Assembly [TCF22]. Asserting [BS10b]. assertion [ÁdBdRS05]. assertion-based [ÁdBdRS05]. assessment [Mah13]. Assignment [BC98, RCM^{+}16, KKT^{+}18, MCRS10, ORH93, RCM^{+}12]. assisted [Dub95]. associated [San04]. Associative [SW08]. Assume [BGP06]. Assume-guarantee [BGP06]. Assumptions [ES97]. ASSURE [SLP^{+}09, Dye98]. Asymmetric [MNU^{+}15, GA09, JSMP13, MWK^{+}06, RBK^{+}09, SCCP13, SMQP09]. Asynchronous [HH11, KFG15, KG07, KSD04, TP18, Yoo96a, GMR09, Kho97, KASD07]. Asynchrony [SRU98]. Athena [Egg10, Hud96]. ATL [SW97]. Atlanta [ACM99a]. Atomic [KKS^{+}08, RD06]. Atomicity [DELD18, MWP^{+}21, BLM06, BNS11a, BNS11b, BNS12, FF04, FFQ04, FF08, FFLQ08, FFY08, WS06]. atomics [ND13]. Atomizer [FF04, FF08]. Attacks [SBE^{+}19]. Audience1 [SBB96]. Augmented [GFJT19, LH09]. August [RM03, IEE99, USE93a, USE98a]. Austin [USE00b]. Austria [Hon94]. authoring [MCS15]. Auto [Pol90, RKHT17]. Auto-vectorization [RKHT17]. AutoDock [TO10]. Automata [ES97]. Automata-Theoretic [ES97]. Automated [BSSS14, DRV02, FF20, KZC15, TR14]. Automatic [BVL09, HBTG98, JJY^{+}03, KW17, Mou00, SSB^{+}22, SEP96, YLLS16, GJ11, JSB^{+}11, SLP^{+}09]. Automatically [NWT^{+}07, TG99, CJ91]. autotuning [CSV10]. Availability [SP07]. Avenue [Ano94d]. avoid [Pra95c]. avoidance [LC13, WLK^{+}09]. AVP [Ano00b]. Aware [AGJ18, BHP^{+}03, CCWY17, FSPD16, FSPD17, GVT^{+}17, HC17, Kim14, LZS^{+}08, LYH16, MNU^{+}15, PR05, SLJ^{+}18, EQT07, EE09a, HEJ09, KD22, LAH^{+}12, MR09, NB12, PAB^{+}14, PGB14, TAS07, XSaJ08, ZLW^{+}16]. Away [GBK^{+}09]. AWTEventMulticaster [Hol99b]. axiomatic [TVD10]. AXP [Ano97a].

B [Ano00c, DLZ^{+}13]. back [ECX^{+}12]. Backup [Ano00b]. Balance [SEP96]. balanced [CKZ12]. Balancers [KMAG01]. Balancing [HBTG98, KC98, KEMC22, KRH98, PGB16, THA^{+}12, WYT^{+}20, ZP04, Chr95a, Chr95b, Chr96, LTL^{+}16, MKIO04]. Baltimore [IEE02]. Bandwidth [FSPD16, LTL^{+}16]. Bandwidth-Aware [FSPD16]. Barcelona [ACM95a, ACM98c, DLM99]. Barnes [ZBS15]. Barrier [CJW^{+}15]. Barrier-Based [CJW^{+}15]. barriers [LZBW14, ZJFA09]. Base [VE93]. Based [Alf94, ARA20, AT16, AKP99, BVL09, BNH01, CJW^{+}15, CKRW99, CMBAN08, DSR15, EGP14, GHG^{+}98, GFJT19, HHOM91, HHOM92, KS16, KG05, KEL^{+}03, KW17, KS97, KRH98, KNE^{+}14, Kwo03, LLKS12, LG06, LS11, MWP^{+}21, MGQS^{+}08, MKC97, OB13, ROPdlT22, RSBN01, SG18, TESK06, WLM15, ÁdBdRS05, Ada98, AAHF09, Ama98, AKSD16, BKK17, CNQ13, CKD94, CKRW97a, CKRW97b, CNV^{+}06, DG99, DWYB10, EG11, GDSA^{+}17, GB20, GE08, HLGD19, JD08, JSMP13, KR01b, KKJ^{+}13, KI16, KBF^{+}12, LK15, LZW17, LLL10, Mus09, Myl22, NBMM12, NFBB17, PSG06a, PSG06b, PSG06c, PAdS^{+}17, PAB^{+}14, Ram94, RRP06, RWT21, Rót19, RS08, SKS^{+}92, TE94a, WCW^{+}04b, WCW^{+}04c, WCW^{+}04d, WQLJ18, YL16, Day92a, Day92b, RSB^{+}09]. Bases [GK94, Swi09]. basic [JJ91, KTLK13, Esp96]. Basis [AGK96]. Batching [LML^{+}19, DKG18]. Be [Pet03, Ano95a, Ano95b, Boe05, MMTW10]. Beach [USE92b]. beat [Gep00]. becoming [Ano92a]. Behavior [CA20, KLS92, LB17, REL00b, ACD^{+}18, DESE13, GS00, REL00a, REL00c]. Behavioral [Sch17]. Benchmark [BTE98, EHSU07, Mül03]. Benchmarking [HHOM92]. Benchmarks [CRE99]. Benefits [MHG95, LB95, LB96b, SD95]. benign [NWT^{+}07]. Berkeley [USE01]. Better [BDM98, Pla99, SV21]. Between [WG94, Pan99, SS95, Yam96, ZCSM02a, ZCSM02b]. Beyond [EKB^{+}92, Sen19]. biased [RD06]. Bibliography [Bee98]. Big [JLA16, AC09, CDL13, LTL^{+}16, LHS16]. BIGSAM [Ply89]. binary [BCCO10, KBF^{+}12, TJY^{+}11]. binding [RČV^{+}10]. Birthmarking [TLZ^{+}17, TLZ^{+}18]. bisection [RRMJ12]. bit [Kus15, SBKK99]. Black [Pla99]. BLAS [ARvW03]. BLIS [VSM^{+}16]. Block [ABLM19, CCWY17, KS97, ZM07, KTK12, KTLK13]. BlockChop [MK12]. Blocking [Ann96, GN00, Nak03, SB80]. Blocks [Pet03, QOQOV^{+}09]. Blue [GBB^{+}05]. Boltzmann [SKG^{+}11]. Bonn [Wat91]. Book [Lar97, Van97a, Vre04]. Bookshelf [Ano99, Cro98, Wil97, Wil00]. Boost.Threads [Kem02]. Boosting [AKSD16, APX12, MLC^{+}09, YZ07]. boosts [McM97]. Bootstrapping [KH18b]. Borland [Kel94a, Kel94b]. Borrowed [DC99, DC00]. Borrowed-virtual-time [DC99, DC00]. Boston [Ano94f]. Both [KZC15, CZSB16]. Bothnia [CCW^{+}11]. Bottle [DSEE13]. Bottleneck [JSMP12].

Bottlenecks [SU96, Zub02, DSEE13, CS12, DSG17]. Boulevard [ACM99b]. Bounded [ITF^{+}22, BMTZ21, LZTZ15, PAdS^{+}17]. Bounding [Lun97, Lun99, MQ07]. BowMapCL [NTR16]. Box [Ano00b]. Braids [BS06]. Branch [AKS06, EPAG16, IBST01, CTYP02, CPT08, GL98b, MTS10]. branches [UZU00]. breadth [LAH^{+}12]. breadth-first [LAH^{+}12]. Break [BVM19]. breakpoint [Ram94]. Bridge [Ano00b]. Bridging [RKBH11, VDBN98]. brief [Hay93]. Briefs [Gar01]. bring [Pra95b]. Bringing [Jon91]. Broadcast [SW08]. Broadcast/Reduction [SW08]. brokers [Sch98]. Browsing [HF96]. BSD [SS95]. BSDCon [USE02]. BSP [SYHL14]. BTRIMER [TJY^{+}11]. Buddy [ZWGX22]. buffered [DLZ^{+}13]. buffers [Koo93]. Bug [FF20, NBMM12]. bugs [BFSK20, JWTG11, VTSL12]. Build [Tro18, KSB^{+}08]. Building [Fon97, KS97, Pet03, ZM07, Omm04]. Building-Block [KS97]. bulk [RD06]. Bulldozer [BBSG11]. Bunka [Ano03]. Burrows [BVP^{+}19, LHS16, NTR16]. Bursty [HMCP16]. Bus [MKC97, Cat94, HHPV15]. Bus-Based [MKC97]. BVT [DC99, DC00]. Bytecode [ABH^{+}01, Coo02, GH03, A^{+}01, CAR08].

C [Kel94a, Kel94b, Lev97, Pla98, Pla99, Rod95a, Vre04, Ait96, AGEB08, Ano99, BM94, Bau92, Bed91, BYLN09, BPL07, BA08, CFK^{+}91, CGR92, Dug95, Eng95, Fin95, For95a, For95b, Gib94, Han97, HSD^{+}12, HSS^{+}14, HTZ^{+}97, HLGD19, HH97, Jon91, KD97, Laf00, Lea96, Man91, Mil95, Mix94, Nag21, ND13, ND16, Pet00, Pla93, Pom98, PS03, PS07, Pul00, Ric91, Rót19, SG18, SC17, Sch90, TB97a, TB97b, Vol93, Wal00, Yam95, Yam96]. C# [KPPÉR06, Stä05]. C-based [RSB^{+}09]. C-Stream [SG18]. C/C [Pla98, Pla99, BYLN09, ND13, ND16, Pet00, Pul00]. C3I [BTE98]. CA [ACM94d, IEE89, USE92b, Ass96, USE00a, USE01, USE02]. Cache [BCZY16, CmL21, CMX10, CCWY17, FJ08, GBP^{+}07, GL98a, HL08, HKSL96, KLS92, KET06a, LLD17, PEA^{+}96, PPG11, SLJ^{+}19, WG94, ZJS12, ZWL15, Car89b, Cho92, KHP^{+}95, KLH^{+}99, MKR10, PPGS20, Raj93, Sha95a, SSkP^{+}07, WCZ^{+}07, ZJS10, ZKR^{+}11]. Cache-conscious [GBP^{+}07]. Cache-oblivious [HL08]. CacheFlow [KET06a]. Cacheline [PBL^{+}17]. Caches [FJ08, PHBC18, KGGK09, ROA14]. Caching [DNT16, KC99, Boo93]. calculations [BD06]. calculi [LVS01]. Calculus [III01, ORH93]. Caldera [Ano97a]. Calif [ACM01]. California [ACM93b, ACM95b, ACM98b, IEE99, USE89, USE91a, USE93b, USE96, USE98b, USE01]. Call [GSC96, Hub01, ORH93, Xue12]. callbacks [VS96]. calling [TTY99]. calls [KASD07, TLZ^{+}16]. Cambridge [USE93a]. Can [Ber96b, Dye98, Pet03, Ano92a, Ber96a, Hig97]. Canada [Ano00b, BT01]. cannot [Boe05]. Cap [HC17]. Capabilities [VD08, Ply89]. capability [CKD94]. capability-based [CKD94]. capacity [SSkP^{+}07]. Capping [RCC12, CDP^{+}21]. capturing [BKC^{+}13]. Carlo [LTMK21]. Carolina [ACM93a]. Cascadia [ZL10]. Case [AH00, AGK96, Chl15a, EE14, LSB15, TAK^{+}00, TESK06, VK99, BDLM07, CASA14, CL94, HJT^{+}93, KPPÉR06, KI16, MSM^{+}11, MN03, SP05, Sod02, YN09, LPD^{+}11]. Cathedral [USE02]. causality [HH16]. causes [BFSK20]. cavity [RM99]. CD [Ano00b]. CDSChecker [ND13]. CE [Tim03]. Center [ACM98d, ACM99b, ACM00, Ano03, Hol12, IEE90]. Centers [JGS^{+}19]. Centric [BDN02, Bre02, Ham96, DHM^{+}12]. Certified [GSK^{+}18]. CFD [DK02]. CG [TAK^{+}00]. CGRAs [PJS15]. chain [LTMK21, SBC91]. Chaining [JY15, KFG15]. Challenge [Ano99]. Challenges [Ano99, GJ97, AG06]. Changing [Gar01]. channel [MN03]. Channels [EPAG16]. chant [HCM94, Ano94c]. Chapter [SKK^{+}01]. Characterization [Ano05, BCG^{+}08, DS09, LPM17, MR94, MMM^{+}05, DWYB10]. characterizations [GS00]. Characterizing [CA20, Gle91, OdSSP12, SSN10, MTPT12]. Charleston [ACM93a]. Chassis [Ano00b]. Chebyshev [Rót19]. Checker [FQS02, MWP^{+}21, FF04, FF08, FFY08]. CheckFence [BAM07]. Checking [ES97, ND16, AHK08, AAJ^{+}19, AD08, AGEB08, BAM07, BS10b, BNS11a, BNS11b, BNS12, CNQ13, Dil00, FFLQ08, Mil95, MQ08, ND13, PAdS^{+}17, Sto02, TVD10, VGR06]. Checkpoint [ZSA13]. Checkpoint/Restart [ZSA13]. Checkpointing [CS02, ZSJ06]. Chemkin [Ano97c, Bra97]. Chicago [Ano94d]. China [IEE97]. Chip [AGJ18, CmL19a, HHOM91, KST04, KML04, KU00, KKS^{+}08, LS07, LZS^{+}08, LKBK11, LMJ14, MTN^{+}00, MR09, PPG11, SV19, TESK06, VIA^{+}05, Wea08, ZM07, CSM^{+}05, DTK^{+}15, GA09, KT99, MWK^{+}06, SMK10, SKKC09, TEL95, TEL98a, TEL98b]. Chip-Multiprocessor [PPG11, KT99]. Chip-Multiprocessors [KU00, LMJ14]. Chips [Ano00a, Ano03, IEE99]. Chiron [TNB^{+}95]. Chiron-1 [TNB^{+}95]. Choice [III01, TEE^{+}96]. Cholesky [CIM^{+}17, VD08]. Chores [EJ93]. Chunking [WLM15]. CIL [CAR08]. Cilk [BJK^{+}95, BJK^{+}96, FLR98, Joe96, Mil95]. Cilk-5 [FLR98]. CIO [Ano94g]. Circuit [AMRR98]. City [Hol12]. CLAM [GMR98]. CLAP [HZD13]. Class [BS99, Cha02, Gib94, Rót19, VE93, CS00, MSLM91, Yam96]. Classes [Cal00, Fek08, How98, Lam95, SC17]. Classical [JSB^{+}12, JSB^{+}11]. Classics [Wil00]. Classification [CA20, KZC15, LMJ14, LCH^{+}08]. classifying [NWT^{+}07]. Claus [WP10]. Client [Day92a, Day92b, Sri95, Gol96]. client-server [Gol96]. Client/Server [Day92a, Day92b]. clients [CDL13]. climbing [CY09]. Clique [MP13]. Clojure [Hic20]. Closure [YMR93b, YM92, YMR93a]. Cloud [PFP^{+}22, FKS^{+}12, GDSA^{+}17]. clouds [FGG14]. Cluster [BNH01, CRE99, HD02, KKH03, Kwo03, SCD^{+}15, MWK^{+}06]. Clustered [GSL10]. Clustering [JY15, LK15, RVR04, TAS07]. Clusters [BWXF05, FA19, WG99, ZBS15, BMV03, FWL03, TMAG03]. CMP [TAS07, AMPH09, CWS06, ICH^{+}10, LLL10, SLJ^{+}18, SSkP^{+}07, ZJS10, ZJS12]. CMP-based [LLL10]. CMPs [GW10, JSMP13, SQP08a, SQP08b, SQP08c, YL16]. CNNs [ZDTM19]. Co [Goo97, SG18, AMPH09, BBH^{+}17]. co-design [BBH^{+}17]. co-optimization [AMPH09]. Co-processor [Goo97]. Co-routine-Based [SG18]. Coarse [NS97, SSK^{+}18, ZM07]. Coarse-Grain [ZM07]. Coarse-Grid [SSK^{+}18]. coated [Lep95]. Code [BBdH^{+}11, CFC^{+}22, Coo95, HYY^{+}15, JSB^{+}12, Kim14, KEL^{+}03, MS02, Nag21, NS97, ND16, PR98, Roh95, RNSB96, TGBS05, Tra91, Ann96, AU21, BB00, JSB^{+}11, SJ95]. Codes [CMBAN08, PHCR09, PT03]. Codesign [HPA^{+}15]. Codesigned [MKM17]. cognitive [MCS15, PWD^{+}12]. cognizant [LK13]. Coir [SG96]. Cold [Hig97, Hig97]. Collaborative [VSDK09]. Collection [AKP99, LB92, PUF^{+}04, PF01, QSaS^{+}16, KTK12]. Collections [Kle00, McM98a, McM98b]. collective [HMC95, SCB15]. collector [BBYG^{+}05, DL93, HL93, WK08a, WK08c, WK08b]. coloring [ÇFG^{+}12, GP05, SS10]. Colt [WN10]. Combinator [KLS92]. Combined [WYT^{+}20, UZU00]. Combining [HLGD19, KR01a, LZ07, CZSB16, ZLW^{+}16]. come [Pol90]. COMeT [RCC14]. Coming [LS07]. Commands [KD97]. Commercial [SBKK99, BEKK00, EJK^{+}96]. Commodity [FNS^{+}22, ZLJ16, LVN10, RPNT08]. Common [BVP^{+}19, Hol98a, Kuc92, BDF98, BDLM07, CL00, Küc91]. Communication [ABN00, BDJ06, DSR15, EHG95, FKT96, FGKT97, GMR98, HYY^{+}15, OA08a, OA08b, OA08c, Pan99, PWL^{+}11, Rod94, SKK^{+}01, TKA^{+}01, TCG95, BR92, DBRD91, GRS06, KASD07, Lam95, QSHI16, RR96, RR03, TG09, TKA^{+}02, VS96, WHJ^{+}95, ZCSM02a, ZCSM02b]. Communications [Ano03, BMN99, FJ08, SCB15, Sho97b, TP18]. Commutativity [AC09]. Compact [HEMK17]. compaction [WK08a, WK08b, WK08c]. Comparative [SKP^{+}02, Yoo96a, PL03]. Comparing [KPPÉR06, SV96c, SV96a, SV96b]. Comparison [ILFO01, SAC^{+}98, GL98b, KIM^{+}03, MKIO04, MMTW10]. Compass [PWD^{+}12]. Compatible [MM14, LBH12]. competition [YL16]. Competitive [MAH18]. Compilation [ACMA97, HLB94, BRRS10, GC92, HCD^{+}94, Tsa97b]. Compile [CS95a, CS95b, TSY99]. Compile-time [CS95a, CS95b]. Compile/run [TSY99]. Compile/run-time [TSY99]. Compiler [ATLM^{+}06, BD00, BF04, CHH^{+}03, CSS^{+}91b, CGSV93, DZKS12, JSB^{+}12, LEL^{+}99, Loc18, MCRS10, SCv91a, SCv91b, SYHL14, Sin99, TPZ21, TY97, TGBS05, YBL16, ZCSM02a, ZCSM02b, ZP11, BCG^{+}95, BAD^{+}10a, BAD^{+}10b, BVG97, CAR08, CSS^{+}91a, CSS^{+}91c, DC07, Dub95, Fon97, Gol97, Hop98, JSB^{+}11, MSM^{+}11, McM97, Mül03, RKCW98, Sch91, SKKC09, UZU00, WLG^{+}14]. compiler-assisted [Dub95]. Compiler-Controlled [CSS^{+}91b, SCv91a, SCv91b, CSS^{+}91a, CSS^{+}91c, Sch91]. Compiler-directed [DZKS12, TPZ21, SKKC09]. Compiler-Driven [YBL16]. compiler-managed [WLG^{+}14]. Compiler-Supported [ZP11]. Compilers [SS96]. Compiling [ABNP00, ABH^{+}01, TLA^{+}02, HTZ^{+}97, Sch91, Sha98, A^{+}01]. Complement [YFF^{+}12]. Complete [BR15, MWP^{+}21, Sch14, BW97, DWS^{+}12, FFY08, KGGK09, NV15]. Completion [AGK96, BGK96, Lun97, Man98, BGK94c]. Complex [SZM^{+}13]. Complexity [EG11, CMX10, SKA01]. complexity-effective [SKA01]. Compliant [BGK96, SP05, Hig97]. component [NFBB17]. component-based [NFBB17]. Components [Gon90, Sho97b]. Composable [MLGW18, SS10, FKS^{+}12]. Compositions [KS97]. Comprehensive [TAM^{+}08, BFSK20]. Compressed [PBL^{+}17]. Computation [ACM94c, BFA^{+}15, CWS06, HLB94, Hon94, HWW93, Kuc92, Lak96, OTY00, Wat91, BHKR95, Fan93, Fuj97, KKT^{+}18, KG07, Küc91, NJ00, Sha98, ST98, WHJ^{+}95]. computation-to-core [KKT^{+}18]. Computational [LNI^{+}19, PCPS15, Bar09]. Computations [BL98, FS96, KC98, KC99, WJ12, YWJ03, Blu92, BL93, BL94, BL99, Chr95a, Chr95b, Chr96]. Compute [BBSG11]. Computer [ACM98c, Ano94a, BVL09, CBN^{+}00, Gol94, BD06, DNB^{+}12, GK05, I^{+}94, PBDO92, WQLJ18]. Computers [Ano94e, SS96, BCM^{+}07, Boo93, LP09, SJ95]. Computing [ACM93b, ACM98a, ACM98d, ACM00, ABC^{+}93, Ama89, CGS^{+}20, CSM^{+}21, CT00, Den94, EJ93, FTP11, FGKT97, Gar01, GRS97, Ham96, Hol12, HG91, IEE94b, KR12, Kon00, LCK11, LFA96, ME17, ROPdlT22, SRU98, SKV21, SZ02, TKZ^{+}18, USE93a, Wea08, WN10, BGG95, BD06, Dan09, FWL03, GBG95, GS02, HF88, HG92, IEE97, Joe96, JCP17, KD22, Kim94, KU17, Lan97, Leg01, Lu95, Mar07, PWD^{+}12, SBCV90, Sen19, Sta90, SKA01, Tem97]. Concept [ÁMdBdRS02, BBFW02, KA97]. Concepts [McC97a]. Concrete [NSP^{+}14]. Concurrency [AN22, BM94, GMGZP14, MLR15, MQLR16, ME17, NFBB17, Tro18, ZWL15, BFSK20, BA08, But14, CBM10, DKG18, GCC15, HZD13, LZ07, NBMM12, NJK16, RR96, RR03, SK12, VTSL12, Yan02, ZLW^{+}16, dB09, SB80]. Concurrency-preserving [NFBB17]. Concurrent [BVM19, ILFO01, KD97, KCCD99, MSM^{+}16, NPT98, PCM16, PF01, SV19, TJY98, AGN09, BBYG^{+}05, Bar09, BO96, BC02, BCCO10, BAM07, Car89a, CVJL08, Cor00, DL93, FK12, GB20, GSK^{+}18, HZ12, HL93, JPS^{+}08, JP92, KIM^{+}03, KGGK09, LPD^{+}11, MSM^{+}10, MKIO04, Men91, NHFP08, Nev99, ND13, ORS^{+}06, STR16, San04, Sen08, ST05, Tsa97a, Tsa97b, WK08a, WK08b, WK08c, ZSJ06, Hay93]. Condensed [BIK^{+}11]. Condition [Hol98c, Yan02]. Conditional [IBST01, NA07]. Conditions [HM96]. Conference [ACM92, ACM93a, ACM93c, ACM94a, ACM94b, ACM94d, ACM95a, ACM95b, ACM96, ACM98b, ACM98d, ACM99a, ACM01, Ano90, Ano94a, AOV^{+}99, BT01, Hol12, IEE94b, IEE95, IEE96, IEE02, LCK11, USE89, USE91b, USE92a, USE93b, USE98b, USE00b, USE00a, Ano94d, Ano94f, Est93, KKDV03]. confidentiality [NSH14]. Confidently [Tro18]. Configuration [GPB^{+}17]. Confirmation [CJW^{+}15]. conflict [NJK16, vPG03]. conformant [Stu95]. Congress [Ano94d]. Conjugate [SSK^{+}18]. conjunction [Ano94e]. Connect [Ano00b]. conquer [FN17, TP18]. conscious [GBP^{+}07]. Consistency [ABH^{+}00, AB01, AB02, CH95, LB17, Rob03, WC99, AAJ^{+}19, BAM07, Cho93, DNB^{+}12, GS00, HT14, QSQ14, RMV22, SNM^{+}12]. consistent [NHFP08]. Consolidated [HC17]. Constrained [TLGM17, GW10, YN09]. Constraint [YDLW20, SCG95]. constraints [HB15]. Construction [KW17, LHS16]. constructs [BS06]. consumption [SCM05]. Contact [Nak03]. Contemporary [ZJS12, ZJS10]. Content [WLM15]. Content-Based [WLM15]. Contention [ALB^{+}18, XSaJ08, ALW^{+}15, DSG17, PGB14, TMCP10, ZKR^{+}11]. Contention-aware [XSaJ08]. Context [BMTZ21, TLA^{+}02, GN92, JLS99, FD95, LG04, MQ07, PAdS^{+}17, PFH06, SCB15, Yan97, LG04]. Context-bounded [BMTZ21, PAdS^{+}17]. context-sensitive [PFH06, LG04]. contexts [BGC14, TE94b, WW93]. Contextual [BGZ97, KH18a, NHFP08]. continuation [AAHF09]. continuation-based [AAHF09]. continuations [DBRD91, GRR06]. Continuing [Ano99]. Continuous [RCC14]. Continuously [DTLM14]. Contour [GFJT19]. Control [BP05, CmL19a, KW17, Lev97, PBR^{+}15, SU01, SZM^{+}13, SG96, CDD^{+}10, DKG18, FK12, FSYA09, GCC15, MLCW11, NT14, PPA^{+}13, PWWD18, Pol90, RPB^{+}09, UZU00, WLK^{+}09, Yoo96b]. control-flow [NT14]. Controlled [ALSJ09, BCG^{+}08, CSS^{+}91b, CGSV93, SCv91a, CSS^{+}91a, CSS^{+}91c, Luk01, MWP07, Sch91, SCv91b]. Controller [RLJ^{+}09]. controllers [KASD07]. controlling [AGN09, BKC^{+}13]. controls [McM96c]. Controversial [Gar01]. Convention [ACM98d, ACM99b, ACM00, Hol12]. Conventional [KET06b, HB92]. Convergence [RM03]. conversational [LG04]. Converse [BK96]. Convert [Vol93]. Converting [LEL^{+}97a, LEL^{+}97b]. convolutions [RB18]. convolver [Kep03]. Cool [Ano00a, Ano03, Wei97]. cooperation [BM07, SKBY07]. Cooperative [AMRR98, DNT16, ILFO01, LC13, KIM^{+}03, MKIO04, TCG95]. coordinated [KKJ^{+}13]. coordination [BDF98]. Coping [San04]. Coprocessor [LRZ16]. Coprocessors [CSM^{+}21]. copying [HL93]. CORBA [DHR^{+}01, PSCS01, SV96a, SV96b, VS96]. Core [CC18, CvdBC18, FMY^{+}15, FJ08, GBK^{+}09, IXS18, KST04, KTR^{+}04, MP01, MNU^{+}15, MM01, MB05, PVS^{+}17, PHBC18, PM14, QOIM^{+}12, VCM19, ABC^{+}15, AMPH09, ÇFG^{+}12, CSM^{+}05, DTR18, DWYB10, GW10, KKT^{+}18, KBF^{+}12, MLCW11, MLC^{+}09, MTPT12, Mus09, RGT17, SMQP09, TPZ21, TCF22, VPQ12, WCC^{+}07, YZ07]. CoreDet [BAD^{+}10a, BAD^{+}10b]. Cores [CSM^{+}21, CCK^{+}16, RRK11, TCS20, CWS06, MAF^{+}09, SW16]. coreSNP [GAC14]. Corner [SW97]. Corona [VSM^{+}08]. Corporation [Ano00b, Ano00b]. correct [DJLP10, SP00b, Shi00]. Correction [TLA^{+}02, HTDL18]. corrective [LG04]. Correctness [Ram94]. Correlation [SLT03, PFH06, SLT02]. cosimulator [LT97]. Cost [TY97, Bet73, DC07, Tsa97b]. cost-effective [Tsa97b]. Costs [MHG95]. COTS [RGG^{+}12]. counterexamples [NV15]. Counters [Wei98b]. Counting [Hol98c, Rec98]. County [ACM98d]. Coupled [MTN^{+}00]. coupling [AG21]. Course [BLPV04, BZ07, GL07, She98]. coverage [RRP06, YNPP12]. coverage-driven [YNPP12]. covering [BCG13]. Covert [EPAG16]. CPU [ASSS19, BSSS14, PGB16, DFC^{+}19]. CPUs [KS21, SKG^{+}11, SMD^{+}10, TCF22]. Craftworks [Ano97a]. Cray [BÇG14, Smi01, VTSM12]. Create [Ber96b, Ber96a, Len95]. Creating [Han97, Ten98]. Creation [Eng00, Rin99, Sin97]. Crisis [Ano99]. Critical [BLG01, CS12, OTY00, PFP^{+}22, AU21, DTLM14, DESE13, NM10, RGG^{+}12, San04, SMQP09, YL16]. Criticality [DESE13, NB12]. Cross [DSR17, Lam95, SHK15, BKC^{+}13, CZSB16]. Cross-Layer [SHK15]. Cross-platform [Lam95]. Cross-Thread [DSR17, BKC^{+}13, CZSB16]. cruncher [Yee20]. CS1 [GL07]. CSMT [GSL10]. CSP [Nev99]. CTS [ASSS19]. CUDA [LBH12, MM14, PAdS^{+}17, WJ12, YZ14]. CUDA-compatible [LBH12]. CUDA-NP [YZ14]. CUG306 [Col90a]. current [LVA^{+}13]. Curve [Rót19]. customizable [JP92]. cut [JEV04]. Cycle [LS11, EE09b]. Cycle-Level [LS11]. Cycles [CNZS17]. Cyclic [YLLS16, HKT93]. cyclone [Gro03]. Cyclops [ACC^{+}03]. Cyrus [HDT^{+}13].

D [KSB^{+}08, NTKA99, PYP^{+}10, TKHG04]. Daemon [Spe94]. DAG [LQ15]. Dallas [ACM00, USE91b]. Dame [IEE96]. dans [Zig96]. DARPA [Mat97]. Data [Ama89, ABNP00, DTLW16, EW96, FHM95a, GAC14, HMC97, HRH08, Hig97, HLH16, JMS^{+}10, JGS^{+}19, KZC15, KEL^{+}03, KET06a, KET06b, LPK16, LMJ14, LLD17, ME15, ME17, ROPdlT22, RCRH95, SBN^{+}97, SAC^{+}98, SSYG97, SG96, TKZ^{+}18, TKK21, Ten98, TESK06, VT96, Wil98, ZLJ16, ZAK01, AGEB08, AGN09, BAM07, CS95a, CS95b, CDL13, DHM^{+}12, Evr01, FHM95b, FK12, HL93, LTL^{+}16, LHS16, Mao96, MMN09, NWT^{+}07, ND13, PDMM16, PRB07, PHCR09, Pol90, PS03, PS07, PT03, Sha95a, SP00b, Shi00, Sin99, SKKC09, TPZ21, WDC^{+}13, YKL13, ZJS^{+}11]. data-centric [DHM^{+}12]. Data-Driven [DTLW16, KET06b, ME15, ME17, TESK06, Evr01]. Data-Parallel [ABNP00, ROPdlT22, SAC^{+}98, HMC97]. data-race [MMN09]. Database [BAZ^{+}19, KD97, MM14, YM92, YMR93b, Hig97, LBE^{+}98, YMR93a]. Databases [AOV^{+}99, GDSA^{+}17, HL08, MIGA18]. Dataflow [CVJL08, FA19, GGB93a, Gao93, HPB11, HKSL96, LH94, NBM93, RSBN01, SRU98, DFC^{+}19, Tra91, YMR93b, BGG95, GGB93b, GBG95, HG92, JHM04, KHP^{+}95, PT91, SKS^{+}92, Sch91, YMR93a]. Dataflow-Based [RSBN01]. dataflow/von [HG92]. datarace [CLL^{+}02, CVJL08]. Datarol [KA97]. Datarol-II [KA97]. Dawning [Cro98]. DC [IEE94c, ACM92, Ano90]. DCE [RD96, Yam95, Yam96]. DDOS [HBCG13]. Deadlock [Hol98a, Mou00, Ver97, ABF^{+}10, GB20, SR14, WLK^{+}09]. Deadlocks [CC14, CJW^{+}15, CZWC13, JPSN09, PRB07]. dealiased [RB18]. Deallocation [LPE^{+}99]. dearth [Len95]. debate [Bak95b]. Debug [MWP^{+}21, PT03]. debugger [CB89, CB90]. Debugging [Ano98b, Caz02, HWZ00, MQLR16, PHK91, SJB92a, SJB92b, BGZ97, MLR15, WOKH96]. decentralized [RPB^{+}09]. Decision [LFA96, LQ15]. Decisions [JGS^{+}19]. decomposition [JEV04]. Decompression [PBL^{+}17]. Decoupled [DO95, IXS18, APX12, Evr01, RVOA08, RCDG06, SKA01, VS96]. decoupling [KGGK09, PG01]. Decoy [MIGA18]. Deductive [ÁdBdRS08, BK13]. Deep [SM19]. Deeply [GKCE17]. Defect [OB13]. Deflation [Myl22]. Defragmentation [PVS^{+}17]. Delaunay [ABC^{+}09]. Delivering [SCCP13]. DeLorean [MCT08]. Demand [KKJ^{+}13]. Demand-based [KKJ^{+}13]. Demus [Sri93]. Demus-2 [Sri93]. dense [ABD^{+}12, MM07]. Dependable [SUF^{+}12]. Dependence [CZS^{+}17]. Dependences [DSR17, BKC^{+}13, CZSB16]. dependencies [NPC06]. Deployment [GARH14]. DepSpawn [FA19]. Depth [McM96a, McM96b, McM96c, McM98a, McM98b]. Derivation [Kim14, SV19]. Derivative [TT03]. describes [Yam96]. Design [ACM94a, ACM99a, Ano94c, BRM03, BC94, CL95, GMB93, GRS97, GMR98, Hai97b, JGS^{+}19, KHP^{+}95, Laf00, LML^{+}19, MB99, NBM93, Raj93, RCDG06, Sch17, STW93, Sha95a, SWYC94, SBKK99, The95, TAM^{+}08, Ven98, ZBS15, AMPH09, BBH^{+}17, BO96, Car89b, FWL03, HCM94, Hud96, KU17, KGGK09, Mah11, Met95, Moo95, Moo96, MKR02, Ném00, OKID92, OCRS07, RSB^{+}09, SB80, Sri93, Ver97, WLG^{+}14, Wan94, WCV^{+}98, Xue12]. designed [San04]. Designing [CSM^{+}21, Dru95, GKZ12, RR93, Rei95, TSV12, Hai97a, TCG95]. Designs [SM19]. Desktop [Ano97a, FURM00c, FURM00a, FURM00b, Mar07, Pra95b, WSKS97]. desktops [Ano94b]. despite [Len95]. Destructing [Pet00]. destructive [FF10]. Desupport [DHR^{+}01]. Detailed [MKR02, ACC^{+}03]. Details [FMY^{+}15]. Detect [CNZS17, DS16, CZWC13]. Detecting [DSR15, RBK^{+}09, SK97, FF10, JPSN09]. Detection [ABF^{+}10, CC14, FF20, HTDL18, KUCT15, KW17, LS18, LLS06, Mou00, TLZ^{+}17, TLZ^{+}18, ZLJ16, AFF06, CLL^{+}02, CVJL08, FF09, HR16, LLLC15, LTHB14, MKM14, MMN09, NBMM12, NAW06, NA07, PS03, PS07, PFH06, RVS13, RM00, SR14, Sch89, TLZ^{+}16, TDW03, WDC^{+}13, ZKR^{+}11, DWS^{+}12]. Detection/Correction [HTDL18]. Detector [SBN^{+}97, SLG06]. determined [Kub15]. determinism [BS10b, LWV^{+}10, LZW^{+}13]. Deterministic [DK02, KRBJ12, LB17, LSS12, VSDL16, BAD^{+}10a, BAD^{+}10b, BAD^{+}09, Bon13, DLCO09, DNB^{+}12, LZBW14, MAAB14, OAA09, QSHI16]. Deterministically [MCT08]. DetLock [MAAB14]. develop [Fek08]. Developer [IEE96]. developers [Way95]. Developing [SP00b, Shi00, TKA^{+}01, OT95]. Development [Ano97a, Ano98b, Ano99, Gil88, Sri95, Tet94, ARvW03, Hig97, Pom98, TNB^{+}95]. Device [MSFC21]. devices [Xue12]. diagnosing [CS12]. diagnostics [GBB^{+}05]. diagrams [SK12]. Diego [ACM93b, ACM98b, USE89, USE93b, USE98b, USE00a]. differences [BFSK20, Yam96]. Different [BLPV04, GLC99]. Differential [BTL^{+}19, Loe97, MQLR16, MLR15]. Difficult [CTYP02]. Difficult-path [CTYP02]. Diffusions [LTM^{+}17]. Digital [SS91]. Digraph [CNZS17]. dimension [NJ00]. dimensional [AR19]. DIMM [ALSJ09]. Direct [PR98, KS21]. Direct-threaded [PR98]. Directed [LPE^{+}99, STR16, AR19, DZKS12, Fan93, Sen08, SKKC09, TPZ21]. directory [QSQ14, HR10]. DISC [Don92]. disciplines [Bar09]. Discrete [WYT^{+}20, Leg01, TKHG04, WLK^{+}09]. discussion [Sho97a, Sho97b]. Disintermediated [BDJ06]. Disjoint [SJA12]. Dispo [MGK^{+}00]. Disruptions [IMMP20]. Dissecting [ACC^{+}03]. Distance [BCZY16, KZTK15, SV19, KNPS16]. distinguish [HL93]. Distinguished [ABH^{+}01, TKA^{+}01]. Distributed [ABNP00, ABH^{+}01, BBD^{+}91, BWXF05, BHKR95, BC94, CV98, CJK95, DKA16, FSS06, GJ97, Jen95, MGK^{+}00, PG92, Pra95a, RLJ^{+}09, RBPM00, RW97, RCRH95, SUF^{+}12, TDW03, USE92b, VS96, Yas95, Ano96, A^{+}01, BCG^{+}95, CML00, Car89a, Gol96, GKK09, Gun97, HB92, HMC95, HWW93, HBCG13, IEE97, ISS98, Leg01, MS03, MLC04, MGL95, MKK99, Ong97, Pha91, Ply89, QSQ14, Sen19, Sto02, Tod95]. Distributed-Memory [RCRH95, BCG^{+}95, HWW93]. Distributed-sum [TDW03]. Distribution [SSYG97, ZAK01, CY09]. divergence [MTS10]. Divergent [WJA^{+}19]. divide [FN17, TP18]. Divisors [Kuc92, Küc91]. DMP [DLCO09]. DNA [LZL^{+}20]. Do [Cri98b, Cri98a, RPNT08, Ber96a, Ber96b, YLLS16]. Dock [BCS11]. Docking [BCS11, TO10]. document [JCP17]. documentation [HF96]. Does [Hag02, RKK15, ZJS10, San04]. doing [Yam96]. domains [LAK09]. DomLock [AN22]. données [Swi09]. Don’t [HHPV15]. DOSThread [VE93]. DOTMIX [RB22]. DOTMIX-Pro [RB22]. DoubleVision [Ano00b]. downdating [VV11]. Downturn [Gar01]. DRAM [LLKS12, kSYHX^{+}11]. DRAMs [ALSJ09]. drf [MSM^{+}16]. DRFX [MSM^{+}10]. Drinking [CZSB16]. Driven [DTLW16, For95a, For95b, HLB94, KET06a, KET06b, LWSB19, ME15, ME17, TESK06, YBL16, CSV10, Evr01, RVS13, RSB^{+}09, SLP08, SQP08a, SQP08b, SQP08c, YNPP12]. driver [CCW^{+}11]. DSLs [RKHT17]. DSM [ABH^{+}00, AB01, AB02, BDF98, KKH04]. DSM-PM [AB02]. DSM-PM2 [AB01]. DSMs [FBF01]. DTS [BHKR95]. Dual [BBC^{+}00, EHG95, KST04, DK02, MB05, WS08, CCW^{+}11, FRL18]. Dual-Core [KST04, MB05]. Dual-Level [BBC^{+}00, DK02]. dual-mode [FRL18]. dual-personality [CCW^{+}11]. Dual-Processor [EHG95]. Dual-Thread [MB05, WS08]. Duplex [KG05]. Duplication [Kwo03]. Dynamic [BPSH05, CJW^{+}15, CmL19a, FSYA09, GPB^{+}17, HSS^{+}14, Hig97, IMMP20, KMAG01, KPC96, KC98, KC99, KUCT15, KEMC22, LK20, MVZ93, MTS10, Nak01, PBL^{+}17, RCRH95, RS08, SBN^{+}97, SLG04, SKK^{+}01, Sta90, SMP19, SG96, WHG07, XMN99, ZKW15, ZKR^{+}11, ZL10, AR17, CAR08, Chr95a, Chr95b, Chr96, Don92, FF04, FF08, FFY08, FF09, HSD^{+}12, JPSN09, KBF^{+}12, LSS12, MK12, Mic04, NHFP08, RB22, SCB15, SLG06, TJY^{+}11, WXAL22, WW96, BK13]. dynamic-multithreading [LSS12, RB22, WXAL22]. Dynamically [PGB12, TLGM17, DMBM16, Kep03]. dynamically-typed [DMBM16]. Dynamics [LNI^{+}19]. DyPO [GPB^{+}17].

e6500 [BGH^{+}12]. Early

[GL91, Myl22, PBL^{+}17, SLP08]. EARTH [HTZ^{+}97, HMT^{+}96, Sod02, TAK^{+}00, TKA^{+}01, TKA^{+}02, TMAG03, Nak03]. EARTH-MANNA [HMT^{+}96, Sod02]. Easy [FA19, Har99]. Easysoft [Ano00b]. ECMA [Stu95]. ECMA-162 [Stu95]. economics [Bar09]. Edge [CSM^{+}21, ROPdlT22]. Edge-Computing [CSM^{+}21]. Edinburgh [AOV^{+}99]. edit [KNPS16]. Editors [GGB93a, GJ97]. Education [Gar01]. Effect [PFP^{+}22, BAD^{+}09, GL98b, YSY^{+}09]. Effective [ABLL92, DN94, GH03, GMGZP14, NAW06, NSH14, PGB16, RVS13, Sat02, TMC09, TY97, WLT19, CBM10, JSB^{+}11, MMN09, MTC^{+}07, SKA01, Tsa97b]. Effectiveness [PR05, TE94b]. Effects [Cho93, HRH08, KLH^{+}99, KRBJ12, NHFP08]. Efficent [TTKG02]. Efficiency [AJK^{+}12, Ano05, IXS19, THA^{+}12, AMPH09, FGG14, GA09, MMM^{+}05, MWK^{+}06, Pra95b, RCG^{+}10, SP05]. Efficient [AD08, ALSJ09, Alf94, ABN99, AU21, BCZY16, BGDmWH12, BJK^{+}96, BL98, BMN99, CZS^{+}17, CFC^{+}22, CYYL18, CLL^{+}02, DMBM16, Gao93, GJT^{+}12, GRS97, GS06, GN96, HMCP16, HSS^{+}14, HR10, HEMK17, KPC96, KASD07, LS18, Lem02, LHG^{+}16, LZBW14, MB07, MAAB14, NB99, PS03, RWT21, SP07, TY97, TGBS05, Tro18, ZLJ16, ZTN19, ZWGX22, ATLM^{+}06, AG21, BL93, BJK^{+}95, BHK^{+}04, EKKL90, FWL03, FF09, GB99, HSD^{+}12, KSB^{+}08, KNPS16, KSD04, LK13, LWV^{+}10, LML^{+}19, LHS16, LZW^{+}13, MSM^{+}10, NLK09, OAA09, Pan99, PSG06a, PSG06b, PSG06c, PRS14, PS07, PPGS20, RL14, RB22, Sch91, SRA06, SP00b, Shi00, SGS14, SQP08a, SQP08b, SQP08c, TO10, Wei98a, kSYHX^{+}11, ZLW^{+}16, FSYA09]. Efficiently [KBF^{+}12, MCT08, SW16, Blu95, BKC^{+}13]. eigenproblems [ABD^{+}12]. eigenvalue [BIK^{+}11]. Elastic [SG18]. Electronic [Ano00b, BB00]. Elegant [Hub01]. Element [HBTG98, MS02, TCF22]. elementary [HKN^{+}92]. elide [MLS15]. Eliminating [DSG17, OCT14, RD06, MTPT12]. elimination [MK12]. elision [NM10]. Elliptic [Loe97]. EM-4 [BAM93, SKS^{+}92]. Embedded [BVM19, BGH^{+}12, DS09, Dru95, GKCE17, KG05, KE15, MS15, WM03, ZDTM19, DCK07, KVN^{+}09, KASD07, KBF^{+}12, LLLC15, LBvH06a, LBvH06b, LBvH06c, RSB^{+}09, SKP^{+}02, Xue12]. Embedded-Systems [Dru95]. Embedding [Pul00]. emergencies [MTPT12]. Emerging [VSM^{+}08, WJA^{+}19, GBP^{+}07, HFV^{+}12]. empirical [LC13]. employing [CWS06]. Employment [Gar01]. Empowering [JSB^{+}12]. Enabling [CC18, Pan99, SMZ18, JMS^{+}10, VGK^{+}10a, VGK^{+}10b]. End [SNM^{+}12]. End-to-end [SNM^{+}12]. Energy [ALSJ09, AJK^{+}12, GJT^{+}12, GKCE17, KE15, LK13, LML^{+}19, LMA^{+}16, PR05, RL14, AAC^{+}15, CIM^{+}17, GA09, KSB^{+}08, MAF19, NB12, PJZA07]. Energy-Aware [PR05]. Energy-Effectiveness [PR05]. Energy-Efficient [GJT^{+}12, LK13, LML^{+}19, RL14]. energy-performance [PJZA07]. enforcement [GWM07, SCCP13]. Engine [SG18, SMP19, CNQ13]. Engineering [GJ97, LSB15, WCV^{+}98]. engines [HB15]. England [ACM94c]. Enhance [FSPD17, FJ08]. Enhanced [Ano00b, EJ93]. Enhancing [KKT^{+}18, OL02a, OL02b, OL02c, TKK21, HWW93, RHH10]. Environment [ABNP00, BC00, CdOS01, EC98, KKH03, PG92, BK96, DSH^{+}10, GCRD04, GCC15, GBB^{+}05, HMC97, Hud96, KG07, Lan97, Pha91, SWYC94, Sta90, Tem97, WCC^{+}07]. Environments [AKP99, BDN02, KG05, SP00a, EJK^{+}96, RGG^{+}12, Sam99, Ver96, Way95]. equality [AD08]. Equalization [TLGM17]. Equations [Loe97]. equivalence [AAJ^{+}19]. equivalent [Pra95c]. Eraser [SBN^{+}97]. Errata [Ano01, Ano05]. Error [EUVG06, OA19, SSN10]. Errors [SK97, VACG09]. escape [SR01a]. Esterel [LBvH06a, LBvH06b, LBvH06c, LvH12]. Estimating [PCPS15]. etc [Hol98a]. European [DLM99]. EuroPVMMPI [KKDV03]. Evaluate [EE14]. Evaluating [BL96, CML00, NPT98, PSCS01, RPNT05, Sch98, SD95, TG09]. Evaluation [Aru92, Boo93, BTE98, CL95, CBN^{+}00, EJK^{+}96, Eic97, GLC99, HN91, RNSB96, SCD^{+}15, TT03, ZL10, BGDmWH12, BLCD97, Car89b, Cho92, Don92, LZ07, Mah11, MKR02, NFBB17, RGG^{+}12, RCDG06, SWYC94, SKP^{+}02, SMS^{+}03, TGO00, TKA^{+}02, WLG^{+}14, WZSK19]. Evaluations [MM14, Roh95]. evaluator [SP00b, Shi00]. even [Ano94b]. événements [Swi09]. Event [Ber96b, CKRW99, For95a, For95b, WYT^{+}20, Ber96a, CKRW97a, CKRW97b, GWM07, KCCD99, KBP^{+}03, Leg01, LTMK21, RVS13]. Event-Based [CKRW99, CKRW97a, CKRW97b]. event-chain [LTMK21]. Event-Driven [For95a, For95b, RVS13]. event-handling [KBP^{+}03]. Events [BDN02, LZ07, Van97b]. Evolutionary [TAK^{+}00, KU17]. Evolving [MS87, MS89]. Exact [Sch17]. examines [Yam96]. Examining [Kan94, Ric91, Rod95a, Tim03]. Example [BLPV04]. Exception [DH98, Lea96]. Exceptions [ÁdBdRS08, KR01b]. exclusion [BRE92]. exclusiveness [Lie94]. execute [APX12]. Executing [Blu95, BS99]. Execution [ABH^{+}01, BTL^{+}19, CC18, CJ91, Coo02, EC98, Far96, GMGZP14, GS06, HMCP16, HEMK17, HZ12, IXS19, KS16, KLG08, KI95, KG94, ME15, MGK^{+}00, MCT08, NBM93, NS97, PR05, RG03, RKK15, RSBN01, STY99, VSDL16, Ann96, A^{+}01, BAD^{+}10a, BAD^{+}10b, BGC14, Dil93, JWTG11, LVN10, Luk01, PAB^{+}14, PG03, SBC91, SJA12, SGS14, SQP08a, SQP08b, SQP08c, SMQP09, SMS^{+}03, TSY99, TSY00, TDW03, UZU00, WCT98, XIC12, XSaJ08]. Executions [CdOS01, HZD13, Roh95, STR16]. Exemplar [BLCD97]. Existing [Ric99]. EXOCHI [WCC^{+}07]. expansion [YKL13]. Expectation [SC17]. Expectation-Maximisation [SC17]. expediting [YL16]. Experience [BMR94, HLB90, Jon86, Yas95, RM03, GL91, Yam96]. Experiences [BHK^{+}04, EHG95, PST^{+}92, SGM^{+}97, USE92b]. Experimental [BLCD97, EGC02, Kno14, YMR93b, GRS06, Pha91, WCW^{+}04b, WCW^{+}04c, WCW^{+}04d, YMR93a]. Experiments [DV99, GMR98, SZM^{+}13, VSM^{+}16, VV00]. Explicit [DV99, VDBN98, BM07, URŠ02b, URŠ03, VV00]. explicitly [MT02a, MT02b, MT02c]. exploit [Ano92a]. exploitation [KVN^{+}09, PSG06a, PSG06b, PSG06c]. Exploiting [AACK92, CmL21, EUVG06, FFQ04, KDM^{+}98, KOE^{+}06, Kwo03, LGN92, MG99, NAAL01, QSaS^{+}16, ROPdlT22, SP07, TLZ^{+}16, TEE^{+}96]. Exploration [PTMB09, Sch17]. Exploring [AAKK08, BS10a, LPM17, SE12, WWW^{+}02]. Expressions [Hei03]. Extended [BLG01, DV99, Rót19, VDBN98]. Extending [BF08, Mar03, PMCP22, RMV22, TCS20]. Extensible [CdOS01]. Extension [RCC14, CCW^{+}11, Lan97, PDP^{+}13, Tem97]. Extensions [Sch90, Bau92]. external [LWV^{+}10]. Extracting [GP95]. Extremal [MNG16].

FAB [YWJ03]. Facility [KSU94]. Facing [KML04]. Factorization [ABLM19, But13, CYYL18, CIM^{+}17, Dav11]. Factorizations [VD08]. failing [STR16]. failure [CZ02, LC13]. failure-inducing [CZ02]. failures [HZD13]. Fair [MQ08, FSPD17]. Fairness [ES97, FSPD17, GWM07, IMMP20, VS11a, SCCP13, WTKW08]. false [LTHB14]. farms [MR98]. Fast [BCS11, BRE92, GSC96, HN91, LDT^{+}16, SSB^{+}22, SMZ18, STY99, SLF14, ST05, VTSM12, ZSA13, ZCO10, BDLM07, CKD94, Kep03, Kus15, SV21, TT03, TTKG02]. Faster [PCM16, BDM98, RB22]. FastTrack [FF09]. Fault [BVM19, HTDL18, OA19, RRP06, RM00, VPC02]. Faulty [BVM19]. FCRC [ACM96]. Fe [Gol94]. Feature [LH09]. Features [GMB93, BDM98]. Featuring [RRK11]. February [USE89, USE00b, USE02]. Feedback [SQP08a, SQP08b, SQP08c, TGO99, ALHH08]. Feedback-driven [SQP08a, SQP08b, SQP08c]. Felix [Ano00c]. Fernandez [Ano00c]. fetch [EE09a, TEE^{+}96, AGJ18]. FFTs [MJF^{+}10]. Fiber [GDSA^{+}17]. Fiber-based [GDSA^{+}17]. fibers [BS06]. Fibonacci [GFJT19]. FIFO [HHOM91, HHOM92, QSaS^{+}16]. fifth [ACM93b, AOV^{+}99]. File [FG91, GJT^{+}12, KS97, Pea92, WLM15, BLCD97, DZKS12]. Files [RRK11, CCC12, kSYHX^{+}11]. filtering [Kep03]. final [HCM94]. Finding [MNG16]. Fine [AZG17, BBG^{+}10, BSSS14, But13, CSS^{+}91a, CSS^{+}91b, CSS^{+}91c, HG91, KG94, LKBK11, LVS01, LFA96, MKM17, NS97, PBR^{+}15, DFC^{+}19, TY97, TAK^{+}00, YSS^{+}17, BGK94c, Dub95, Gol97, KDM^{+}98, Kim94, Loi95, MLC^{+}09, Met95, PL03, RPB^{+}09, TKHG04, Wei98a, kSYHX^{+}11]. Fine-Grain [AZG17, CSS^{+}91b, HG91, KG94, LFA96, CSS^{+}91a, CSS^{+}91c, TY97, KDM^{+}98, Kim94, Loi95, MLC^{+}09, Met95, PL03, TKHG04]. Fine-Grained [BBG^{+}10, BSSS14, But13, LKBK11, PBR^{+}15, DFC^{+}19, TAK^{+}00, YSS^{+}17, LVS01, BGK94c, Dub95, Gol97, RPB^{+}09, Wei98a, kSYHX^{+}11]. Finite [HBTG98, MS02, TCF22, Cor00]. Finite-Element [MS02]. finite-state [Cor00]. firmware [ABB^{+}15]. First [MSLM91, Wei97, LAH^{+}12, MHW02, Hon94]. First-class [MSLM91]. FL [ACM94a]. flat [PPGS20]. FlexBFS [LAH^{+}12]. Flexible [ABG^{+}08, KS97, Lem02, MSM^{+}16, SP00a, Sam99, SCM05, WW93]. Floating [LWSB19]. Floating-Point [LWSB19]. Florida [ACM98d]. Flow [AT16, Ama89, HH11, PBR^{+}15, FSYA09, JD08, KBH^{+}03, NT14, Pol90, RM99, RPB^{+}09, SV98]. fluid [JD08]. FluidCheck [KS16]. fly [CWS06, PS03, PS07, Sch89]. Focus [EHP^{+}07]. Forces [FTP11]. Forecasting [Ano98b]. fork [ALS10]. fork-join [ALS10]. FORM [TV10]. Formal [LPD^{+}11, Stä05, BVM19, WP10]. formalising [RMV22]. formation [FSYA09]. forms [BIK^{+}11]. FORTH [Jon86]. FORTH-like [Jon86]. Fortran [Ano97a, Ano97c, Bra97, AS14, GOT03, HBG01, HBG02, Nag01]. forum [Sho97a, Sho97b]. Forwardflow [GW10]. foundation [MCS15, RBF^{+}89]. Foundations [BA08, Gol94]. Four [CH95, MTN^{+}00, KNPS16]. Four-Russians [KNPS16]. Four-Way [MTN^{+}00]. Fourier [TT03, TTKG02, BCS11, HN91]. fourth [USE96]. FPGA [DFC^{+}19]. FPGAs [ROPdlT22]. fragment [APX12, MAF19]. fragments [LG04]. Framework [BMF^{+}16, BVL09, BF04, BP19, CV98, DHR^{+}01, EFG^{+}03, KC98, KF97, LCS04, LMJ14, Loe97, NSP^{+}14, Rei01, DFC^{+}19, VSM^{+}16, Yam95, ZDTM19, AMC^{+}03, AL21, BDF98, EHSU07, GJ11, Hop98, PV06]. France [FR95]. Francisco [ACM95b, USE02]. Free [DELD18, Way95, AR19, DTLM14, GP08, MLS15, Mic04, PMCP22, ST05]. free-lunch [DTLM14]. FreeBSD [Ano00b, Bal02]. freeness [AHK08]. Freescale [BGH^{+}12].

French [Zig96]. frequent [GBP^{+}07]. Friendly [Nag21]. Fthreads [Nag01]. Fukuoka [Ano91]. Full [MHW02, GB99]. Full-system [MHW02]. fully [RD99, Stu95]. Function [Hub01, LLKS12, Rót19, TO10]. Functional [Coo95, DCK07, GS06, Kim94, KIAT99, LP94, SSP99, Gun97, RKBH11, TAN04, VGR06, WZWS08, ZSJ06]. Functions [Bed91, KI16]. Further [GV95]. Fusion [PWL^{+}11, Hig97]. futex [BF08]. Future [Jon91, TAM^{+}08]. futures [TTY99].

GALAHAD [GOT03]. GAMBIT [CBM10]. Game [DHR^{+}01]. GAMESS [BB00]. Garbage [AKP99, LB92, PUF^{+}04, PF01, QSaS^{+}16, BBYG^{+}05, DL93, HL93]. Garcia [Ano00c]. Gateway [Yas95]. Gating [RRK11]. Gaussian [SC17]. GbA [LZW17]. GC [HHPV15]. Geant4 [SCD^{+}15]. GEMM [SLJ^{+}19]. Gene [GBB^{+}05]. Gene/L [GBB^{+}05]. General [Ber96b, BF04, HSS^{+}14, Man98, YKL13, ZSA13, Ber96a, Car89a, DC99, DC00, HSD^{+}12, MQW95, SKA01]. General-Purpose [Ber96b, HSS^{+}14, Man98, Ber96a, DC99, DC00, HSD^{+}12]. generalized [ABD^{+}12, BCM^{+}07, FTAB14]. Generated [BD00, MJF^{+}10]. Generating [AZG17]. Generation [ARB^{+}02, BVL09, Coo95, EFN^{+}01, EEL^{+}97, HEMK17, HYY^{+}15, NBS^{+}15, RNSB96, SSK^{+}18, TGBS05, Tra91, TSV12, ABC^{+}09, EFN^{+}02, GJ11, Gep00, KI16, LK13, LSS12, LZSS19, Way95, CH04]. generational [DL93, WK08a, WK08b, WK08c]. generations [Roh95]. generators [SLF14, SV21]. Generic [ABH^{+}00, AB02, Fer13]. Genetic [NSP^{+}14]. genome [LHS16]. GeoFEM [Nak03]. Geometric [Caz02]. Georgia [ACM99a]. Germany [RM03, Wat91]. getting [MAH18]. ghosts [TVD14]. Gigabit [AHW02]. Gigabit/sec [AHW02]. Gilgamesh [SZ02]. glasses [CZSB16].

Global [HH11, PWL^{+}11, Ten02, FWL03, LZBW14, OCT14, OA08a, OA08b, OA08c, Ano98b]. globally [CZWC13]. gmm diag [SC17]. gmm full [SC17]. GNAT [dlPRGB99]. Go [Mia90]. Going [Bak95b]. Goldilocks [EQT07]. good [Mat03]. GPGPU [CCWY17, LLKS12, YZ14]. GPGPUs [LSB15, ZWL15]. GPS [TVD14]. GPU [APX12, Bon13, DTR18, FTP11, IXS19, KI17, LWSB19, LTL^{+}16, LML^{+}19, LHG^{+}16, LAH^{+}12, Sen19, WLG^{+}14, WJA^{+}19, XWG^{+}14, YSS^{+}17, YSS^{+}19, ZCO10]. GPU-Oriented [LHG^{+}16]. GPUDet [Bon13]. GPUMixer [LWSB19]. GPUs [CSV10, DNT16, LBH12, SKG^{+}11, VD08, WJ12]. Grace [BYLN09]. Gradient [SSK^{+}18, MAF19]. Grain [AZG17, CSS^{+}91b, HG91, KG94, LFA96, MKM17, NS97, ZM07, CSS^{+}91a, CSS^{+}91c, KDM^{+}98, Kim94, Loi95, MLC^{+}09, Met95, PL03, TY97, TKHG04]. Grained [BBG^{+}10, BSSS14, But13, LKBK11, PBR^{+}15, DFC^{+}19, TAK^{+}00, YSS^{+}17, BGK94c, Dub95, Gol97, LVS01, RPB^{+}09, Wei98a, kSYHX^{+}11]. Grande [ACM01]. Grande/ISCOPE [ACM01]. Granularity [KI95]. Graph [ARA20, ÇFG^{+}12, CL95, EJRB13, FNS^{+}22, HPA^{+}15, KS93, KLS92, MCS^{+}22, MM14, LK15, LZW17, RVR04]. graph-based [LZW17]. GraphCT [EJRB13]. Graphical [ACR01]. Graphics [SSB^{+}22, BGDmWH12, CCW^{+}11, FSYA09, PYP^{+}10]. Graphs [HPB11, Nik94, OB13, AD08, ABG^{+}08, DSEE13]. grass [MMTW10]. Greatest [Kuc92, Küc91]. Green [SKP^{+}02]. greener [MMTW10]. Grid [KEL^{+}03, SSK^{+}18]. Grid-Based [KEL^{+}03]. GRIDiron [MCS15]. grids [SKG^{+}11]. Griffin [Ano00c]. Gröbner [AGK96]. Group [BNH01, DLM99, QSHI16]. Group-Based [BNH01]. Grouping [OR12, WC99]. groups [WZSK19]. Grove [IEE89]. Growth06 v2 [Dan09]. Guarantee [Hag02, BGP06]. Guarantees [PSM01, YWJ03, GPS14, MTC^{+}07, PSM03, ZHCB15]. Guarded [Sim97]. Guest [GGB93a, GJ97]. GUI [Tet94]. Guide [Ano99, BBD^{+}91, LB96a, Wil97, BW97, ND96, RR96, Sun95]. guided [NB12]. Guidelines [RD96]. GUIs [Mia90]. Gyrokinetic [KEL^{+}03, PWL^{+}11].

Hagenberg [Hon94]. Hagenberg/Linz [Hon94]. Halide [DKA16]. Hamilton [Ric91]. Handles [Rec98]. Handling [DH98, LSB15, SK97, BM91, KCCD99, Koo93, KBP^{+}03, Lea96, Met95]. Hands [Tro18]. Hands-on [Tro18]. Harbor [BBC^{+}00]. Hardware [AGJ18, BVM19, BAZ^{+}19, CKD94, CSS^{+}91b, DVAE18, FNA^{+}18, KE15, KH18b, LLS06, MWP07, MKM17, Men91, ROPdlT22, SW08, SSK^{+}18, ZLJ16, ABC^{+}09, BMF^{+}19, CWS06, CSS^{+}91a, CSS^{+}91c, ECX^{+}12, FSYA09, GP05, LT97, MLS15, MQW95, OCT14, PAB^{+}14, PRS14, RPNT05, SE12, TE94b, DWS^{+}12]. hardware-aware [PAB^{+}14]. Hardware-Based [ROPdlT22]. Hardware/Software [MKM17, LT97]. harmful [NWT^{+}07]. Harmony [KTK12]. Harness [Ama98, EBKG01]. Hash [GK05, VB00, RWT21]. hash-based [RWT21]. Hash-join [GK05]. Hashing [SMZ18, MIGA18]. having [YFF^{+}12]. Head [Mia90]. healing [SLP^{+}09]. Heaps [DGK^{+}03, GFJT19, Man99, Ste01]. help [Len95]. Helper [ALS10, WCW^{+}04b, WCW^{+}04c, WCW^{+}04d, WCW^{+}04a]. Here [Ano92a, Pra95c]. Hessenberg [BKK17]. Hessenberg-triangular [BKK17]. Heterogeneity [CCK^{+}16, Kwo03, RKBH11]. Heterogeneous [AT16, AACK92, FBF01, GPB^{+}17, KTR^{+}04, LPM17, Lu95, MSFC21, NTR16, SM19, THA^{+}12, ZDTM19, FKS^{+}12, GKZ12, KD22, LK13, SJ95, TPZ21, WCC^{+}07]. Heuristic [HH11, Mah11, OCRS07]. Heuristics [MGI14]. Hewlett [BLCD97]. HFS [KS97]. hiding [BR92]. Hierarchical [GJT^{+}12, JY15, KC98, KG94, BMV03, DZKS12, LK13, LQ15, RCDG06]. Hierarchies [AN22, BCZY16, TAM^{+}08]. Hierarchy [ZM07, BGDmWH12, PMCP22]. High [ACM98a, ACM98d, ACM00, Ano00a, Ano03, BGH^{+}12, CGS^{+}20, CFC^{+}22, CT00, FF20, FGKT97, Gar01, Hol12, HG91, IEE94b, LCK11, LG06, LMJ14, LML^{+}19, LBH12, LHG^{+}16, LCH^{+}08, MR94, MSM^{+}16, MPD04, ME17, MCS^{+}22, NBS^{+}15, PH97, RG03, SRS98, Sch17, SLJ^{+}19, TCI98, VV11, WG99, WN10, WLT19, ZDTM19, AL21, CIM^{+}17, GS02, HG92, Kim94, Lan97, RRP06, Rei95, Sen19, SQP08a, SQP08b, SQP08c, Tem97]. high- [RRP06]. High-Level [CFC^{+}22, MCS^{+}22, Sch17, FF20]. High-Performance [ACM98a, BGH^{+}12, FGKT97, Gar01, IEE94b, NBS^{+}15, RG03, SLJ^{+}19, TCI98, WN10, LML^{+}19, LCH^{+}08, VV11, AL21, CIM^{+}17, Kim94, SQP08a, SQP08b, SQP08c]. high-powered [Rei95]. High-Speed [Ano00a, Ano03, HG91, SRS98, HG92]. High-Utility [WLT19]. Higher [CJK95, NV15]. Higher-Order [CJK95, NV15]. highly [BGDmWH12, Kub15, KGGK09, MAAB14]. Hill [CY09, USE02]. Hill-climbing [CY09]. Hilton [IEE90]. HippogriffDB [LTL^{+}16]. Hist [Gar01]. history [Ano97b, Hic20]. Hoard [BMBW00a, BMBW00b, BMBW00c]. Hoare [KI17]. HoME [OKID92]. Homogeneous [CC18, JGS^{+}19]. Hood [Ven97]. HoPE [PBL^{+}17]. Hopping [MSFC21]. Hot [IEE99, PBL^{+}17, Gle91]. Hot-Cacheline [PBL^{+}17]. Hotel [Ano94d, USE02]. Householder [BKK17, VV11]. Householder-based [BKK17]. Householder-like [VV11].

Houston [Cha05]. HP [Ano95a, Ano95b, Yam96]. HP-UX [Ano95a, Ano95b, Yam96]. HPC [GKK09, KC09, PLT^{+}15, SLJ^{+}18]. HPF [BMV03, CM98]. HTM [KGGK09]. HTMT [Gar01]. HTTP [Zha00]. Hut [ZBS15]. HW [ZDTM19]. HW/SW [ZDTM19]. Hybrid [BBG^{+}10, Gao93, JYE^{+}16, LH09, MS02, NBM93, YZ07, GKK09, HG92, LZSS19, MK12, MTC^{+}07, SKS^{+}92, Sha95b, kSYHX^{+}11]. Hybridizing [CZS^{+}17]. Hyper [PFP^{+}22]. Hyper-Threading [PFP^{+}22]. Hyperion [A^{+}01]. Hyperobjects [LS18]. hyperscalar [Raj93, Sha95a]. Hyperthreading [HRH08, KM03].

I-WAY [FGT96]. i.e [USE98b]. I/O [RM03, Ano95a, Ano95b, ABB^{+}15, BDN02, KSU94, LTL^{+}16, Man98, MG15, Yoo96a]. IBM [ABB^{+}15, CJB^{+}15, KST04, LSF^{+}07, WZWS08]. Id [Nik94]. IDA* [Mah11]. Ideas [JLA16]. idempotency [KOE^{+}06]. identical [LSW^{+}18]. Identification [SBE^{+}19, JSMP12]. Identify [BTL^{+}19]. Identifying [BCZY16, LZL^{+}20, SU96, DESE13]. IEEE [ACM98d]. IFIP [BT01]. Igniting [ACM03]. II [HCD^{+}94, IEE89, JJ91, KA97, KR01a, McM96b, Wal95]. III [Ano00a, USE92b]. Illinois [GHG^{+}98]. Illinois-Intel [GHG^{+}98]. Illuminating [BLPV04]. ILP [OCRS07, RLJ^{+}09]. im [HL93]. Image [WN10, BÇG14, Kep03, RKHT17]. Impact [KLG08, LK20, SCL05, TE94a, ZAK01, Div95, Met95, RGG^{+}12, RPNT05]. Impaired [Wei97]. imperative [SV98]. implement [DBRD91]. implementable [TEE^{+}96]. Implementation [ACM94a, ACM99a, Alf94, AB01, AKP99, BBD^{+}91, BHP^{+}03, BRM03, CWHB03, DSH^{+}10, FLR98, Hai97b, KA97, MS02, Nik94, STW93, TKA^{+}02, TMAG03, BK96, BB00, BMV03, CMX10, DL93, FGT96, GCC99, GB99, IAD^{+}94, KASD07, Lev97, Li05, LZ07, LAH^{+}12, NFBB17, OKID92, RGT17, Stu95, Tod95, YZYL07, Ano95a, Ano95b]. Implementations [Han97, SAC^{+}98, Ram94, SKG^{+}11, Sha95b]. implemented [Boe05, KEL^{+}03]. Implementing [ABH^{+}00, AB02, BP05, CB89, CB90, Day92a, Day92b, DPZ97, GMB93, GSC96, HPA^{+}15, KR01b, KBA08, KIAT99, Pra95a, TY97, TAN04, BHK^{+}04, Lie94]. Implications [RM03, BS96, VSM^{+}08, CSM^{+}05]. Implicit [BAM93, MS02]. Implicitly [ACMA97, PFV03, SAC^{+}98, RB18]. Implicitly-multithreaded [PFV03]. Improve [FSPE20, GV95, KH18b, QSaS^{+}16, RKK15, Sin99]. Improved [BR92, GMGZP14, LLS06, Smi06]. Improving [AJK^{+}12, AN22, BDN02, CCWY17, DKG18, FT96, FM92, FBF01, GA09, IXS19, IBST01, LYH16, Man99, MEG03, Nak01, PG01, PAB^{+}14, MCRS10, TO10]. In-Memory [BAZ^{+}19, RWT21]. In-Order [RRK11]. In-place [SGLGL^{+}14, SCM05]. in-situ [LSW^{+}18, RGK99]. IN-Tune [RGK99]. includes [SJ95]. incomplete [HR16]. incompressible [RM99]. Incorrectly [SCL05]. Increasing [DELD18, PHCR09]. Incremental [BFA^{+}15, Caz02, Lar95, LB92, BBYG^{+}05]. Independent [DS09, EW96, FSS06, USE93a, KNPS16, MEG94, PG03, WZSK19]. Independently [ALSJ09]. indexing [MIGA18, MLS15]. induced [MTPT12]. inducing [CZ02]. Industrial [KW17, Kon00]. Industry [DM98]. Industry-Standard [DM98]. inference [FFLQ08]. inflation [OdSSP12]. InfoDock [Ano97a]. Information [BS96, PBR^{+}15, CML00, KKT^{+}18, KBH^{+}03, RPB^{+}09, SV98]. Informix [Ger95]. Initial [BTE98]. Injection [SBE^{+}19]. Inline [GH03, DJLP10, EKKL90]. Inline-Threaded [GH03]. Inlining [PR98, LQ15]. innovating [JD08]. Innovation [ACM03]. innovations [ABB^{+}15]. Input [BCG13, MP89, Tan87]. Input-covering [BCG13]. input/output [MP89]. Insertion [CmL21]. Insight [IEE02]. Instruction [DV99, HMNN91, LEL^{+}97a, LEL^{+}97b, MCFT99, RYSN04, RS08, TCS20, AMC^{+}03, Aru92, Cho92, HKN^{+}92, HMN^{+}92, KBF^{+}12, LGN92, Mis96, OA08a, OA08b, OA08c, PYP^{+}10, Raj93, SD13, SMS^{+}03, TEE^{+}96, VS11b, VDBN98, VV00]. Instruction-Level [LEL^{+}97a, LEL^{+}97b, MCFT99, LGN92, SD13]. instruction-systolic [PYP^{+}10]. instructions [PPA^{+}13]. instrumentation [RS07, XMN99]. Integer [Ger18, GH98]. integral [Küc91]. integrated [CCW^{+}11, MTS10, RD99]. Integrating [Cal00, CM98, DNR00, DTLW16, FKT96, TTY99, Tsa97b]. Integration [BWXF05, KSD04, KASD07, SD13]. integrity [NT14]. Intel [ARB^{+}02, CCW^{+}11, GHG^{+}98, PDP^{+}13, RMV22, SCD^{+}15]. Intel-x86 [RMV22]. intensity [BD06]. Intensive [LK20, TKA^{+}01, AAKK08, TKA^{+}02, YSY^{+}09]. Interaction [Hei03, HF96, Pan99]. Interactions [WG94, WSKS97]. Interactive [FURM00c, PTMB09, SSB^{+}22, WOKH96, CSB00, FURM00a, FURM00b, HJT^{+}93, KG07, Lan97, MCS15, Tem97]. Interconnection [NGGA94, RR93, SMK10]. Interface [Chl15a, HBG01, HTDL18, KKDV03, MS89, Met95, PS01, SW97, Ada98, DLM99, HBG02, Li05, MQW95, MS87, MEG94, TNB^{+}95, FGT96]. Interfaces [Han97, HF96, LG04]. Interference [BTL^{+}19]. Interleaving [LGH94, YN09]. Intermediate [McC97a]. Internals [MM01, Wea08]. International [ACM92, ACM94c, ACM94d, ACM95a, ACM96, ACM98c, Ano91, Ano94a, Ano94d, Ano00a, Ano03, AOV^{+}99, Cha05, EV01, Hol12, Hon94, Lak96, LCK11, Wat91, FR95]. Internationalization [Ano98b]. Internet [Ano96, Hig97, SBB96, van95]. Interoperability [DHR^{+}01, Way95]. interplay [MLS15]. Interpretation [GH03, LG04]. interpreter [OCT14]. Interprocedural [NR06]. Interprocess [Rod94]. Interrupts [KE95]. Interval [AN22, Kub15]. Intra [MKR10]. Intra-application [MKR10]. Introducing [GL07]. Introduction [CLRS09, Dra96, GGB93a, GJ97, Mas99, Bir89, GC92, Hay93, She98]. Intrusive [Caz02]. INUX [DNR00]. invasive [RGK99]. Inverse [HMLB16, GEG07]. inverses [GE08]. inversion [JCP17]. Invocation [SKK^{+}01]. IPC [EE14, Koo93]. IPs [Sch17]. iReplayer [LSW^{+}18]. IRREGULAR [FR95, FF20, TSV12, ZAK01, TP18]. irregularly [FR95]. ISA [KTR^{+}04, MNU^{+}15]. ISCOPE [ACM01]. Isolating [CZ02, JWTG11]. Isolation [FSPE20, CMX10, MTC^{+}07, SKBY07]. Isomigration [ABNP00]. ISSAC [ACM94c, Lak96, Wat91]. Issue [KU00, RYSN04, Ano94e, GGB93b, TEE^{+}96]. Issues [GMB93, PS01, ARvW03, Ann96, GC92, HCD^{+}94, IAD^{+}94, TCG95]. Issuing [HMNN91, HKN^{+}92, HMN^{+}92]. Itanium [MB05, WCW^{+}04b, WCW^{+}04c, WCW^{+}04d]. Itanium-2 [WCW^{+}04b, WCW^{+}04c, WCW^{+}04d]. Itemset [WLT19]. iterations [UZU00]. Iterative [MQ07, Nak03, AAC^{+}15]. iThreads [BFA^{+}15]. IUknown [SW97]. Ivan [Ano00c]. IXP [ARB^{+}02, LCH^{+}08]. IXP2800 [AHW02].

J.UCS [KU00]. January

[ACM94b, ACM95b, ACM98b, Ano90,

USE89, USE91b, USE93b, ACM93a]. Japan [Ano91, Ano00a, Ano03]. JaRec [Chr01, GCRD04]. Jason [Ano00c]. Java [ACM98a, ACM01, Ano97a, USE01, AFF06, ÁMdBdRS02, AddS03, ÁdBdRS05, ÁdBdRS08, Ait96, Ano96, Ano97c, Ano98b, ABH^{+}00, ABH^{+}01, A^{+}01, AG96, ACR01, ABG^{+}08, BZ07, Ber96b, BVG97, BAD^{+}09, BR15, BPSH05, BHK^{+}04, BS00, Bra97, BP05, BLPV04, Cal02, CV98, CKRW97a, CKRW97b, CKRW99, CWHB03, CC04, CCH11, Chr01, CT00, Coo02, Cor00, Cri98b, Cri98a, DJLP10, DH98, DRV02, DLZ^{+}13, DS09, Dil00, DGK^{+}03, Dra96, DHR^{+}01, Dye98, EFN^{+}01, EFN^{+}02, EFG^{+}03, EQT07, FSS06, FWL03, Fek08, Fer13, FFLQ08, GH03, GCRD04, GS00, GEG07, GE08, GLC99, Hag02, Ham96, Hei03, Hol98d, Hol98a, Hol98b, Hol98c, Hol99a, Hol99b, Hol00, Hyd00, KPPÉR06, KBP^{+}03, LB00, LCS04, Loc18, Loe97, Man96, MP01, McM96a, McM96b, McM96c, McM98b, McM97, Mit96, MC06, NAW06]. Java [NM10, NR06, Nev99, OW97, OW99, PSM01, PSM03, PRB07, Pet03, PUF^{+}04, PV06, PG03, RKCW98, San04, SE12, Sat02, Sch14, Sho97a, Sho97b, SBE^{+}19, Sto02, SMP19, SKP^{+}02, Van97a, Ven97, Ver97, WN10, Whi03, XSaJ08, Xue12, Yan02, van95]. Java-like [DJLP10]. JavaBeans [Van97b]. javar [BVG97]. JavaScript [PCM16, VP16]. Javier [Ano00c]. Jersey [MT93]. JIT [McM97]. job [EE10, EE12, ST00a]. Jobscheduling [ST00c, ST00b, STV02]. John [Ano00c]. Joho [Ano03]. join [ALS10, GK05]. Joint [FTP11, KS21]. Jones [Ano00c]. Jorgenson [Ano00c]. Jose [ACM94d]. Journeyman [Bec00]. JPF [WKG17]. JPR [WKG17]. Jr [ACM99b]. Julia [BB20, Kno14, Sen19]. July [ACM92, ACM94c, ACM95a, ACM98c, EV01, IEE96, Lak96, Ass96, USE96, Wat91]. June [ACM94a, ACM98c, ACM01, Ano94f, USE92a, USE00a]. JUnit [Goe01]. just [KBF^{+}12]. just-in-time [KBF^{+}12]. JVM [Lan02, McM97, USE01, WKG17].

K-Java [BR15]. KAI [Ano98b]. Kaikan [Ano00a]. Karlsruhe [RM03]. Kaspersky [Ano00b]. Kendo [OAA09]. Kernel [Alf94, ABLL92, Bal02, DNR00, EBKG01, EKB^{+}92, Kor89, MM01, MSFC21, ZSA13, Ano95a, Ano95b, BF08, JJ91, MP89, SS95]. Kernel-Based [Alf94]. Kernels [KI17, dlPRGB99, GLC99]. Kiel [LvH12]. Kikai [Ano00a]. Kikai-Shinko-Kaikan [Ano00a]. kinds [San04]. kinematical [BD06]. Kinematics [HMLB16]. King [ACM99b]. Kingdom [ACM94c]. Kitsune [HSD^{+}12, HSS^{+}14]. Klessydra [CSM^{+}21]. Klessydra-T [CSM^{+}21]. Knoxville [IEE94b]. Kroll [Ano00c]. KUMP [NTKA99]. KUMP/ [NTKA99].

L [DNR00, GBB^{+}05]. L1 [PHBC18]. L2 [SLP08]. L2-miss-driven [SLP08]. L3 [FJ08]. Lab [Ano00b]. labeling [D’H92]. Lafayette [EV01]. Lake [Hol12]. lambda [ORH93]. Laminar [PBR^{+}15, RPB^{+}09]. LAN [Yas95]. LAN/WAN [Yas95]. Landing [TAK^{+}00]. Language [ACM94a, ACM99a, ACMA97, BS06, FLR98, GS06, KN17, Kno14, KIAT99, Sat02, BO96, CFK^{+}91, ECX^{+}12, GPS14, Jon86, KN19, LT97, Man96, Mil95, Ong97, PRB07, RL14, SV98, Smi06, TMAG03, VGR06]. Languages [ACM93a, ACM94b, ACM94d, ACM95b, ACM98b, Coo95, MSM^{+}16, NPT98, OTY00, SCv91a, SS96, TY97, DMBM16, HL93, JP92, JHM04, MSM^{+}10, Sch91, SCv91b, ST98, TAN04]. LAPACK [ARvW03]. Laptops [Ano00c]. Large [AOV^{+}99, CC14, CJW^{+}15, GN92, JLA16, JCP17, LA93, Rót19, SKV21, BCM^{+}07, Boo93, GOT03, Koo93, SMK10, WCV^{+}98]. Large-Scale [CC14, CJW^{+}15, LA93, SKV21, BCM^{+}07, GOT03, SMK10]. Latencies [Sch17, BS06]. Latency [BD00, BAZ^{+}19, Fan93, ÖCS01, PFP^{+}22, SW08, Smi01, SKK^{+}01, WWW^{+}02, YLLS16, ASSS19, BR92, DC99, DC00, Jef94, Luk01, MVY05, PG01, TK98]. Latency-Critical [PFP^{+}22]. Latency-directed [Fan93]. Latency-Resistant [YLLS16]. latency-sensitive [ASSS19, DC99, DC00]. Latency-Tolerant [ÖCS01]. lattice [SKG^{+}11]. Law [Gar01, NZ17, CN14]. Layer [SHK15, CDD^{+}10]. layers [GSK^{+}18]. layout [DZKS12, HB15]. Lazy [GSC96, Gol97, ITF^{+}22, LP94]. LB4OMP [KEMC22]. LCMT [LKBK11]. leadfoot [HHPV15]. Leakage [Mus09, SYHL14]. Leakage-saving [Mus09]. leaks [ZJS^{+}11]. Learned [HPA^{+}15]. Learning [CmL19a, CYYL18, DS16, LPM17, ROA14, GB20, PWWD18]. learning-based [GB20]. least [FTAB14]. least-squares [FTAB14]. lecture [Egg10]. Lenient [SCv91a, Sch91, SCv91b]. Lepp [RRMJ12]. Lepp-bisection [RRMJ12]. Lessons [RM03, HPA^{+}15]. Letters [DHR^{+}01, TLA^{+}02]. letting [AC09]. Level [ABLL92, BBC^{+}00, CFC^{+}22, FURM00c, GP95, HTDL18, JYE^{+}16, JLS99, DK02, KSU94, LS11, LEL^{+}97a, LEL^{+}97b, MG99, MR94, MGI14, MCS^{+}22, PLT^{+}15, RR93, Ric99, Sch17, SLT03, YBL16, BBH^{+}17, CCC12, DG99, EE09a, FF20, FURM00a, FURM00b, GMW09, GPS14, GRR06, HDT^{+}13, JEV04, KDM^{+}98, KVN^{+}09, KC09, Lan97, LGN92, LZ07, MSLM91, MT02a, MT02b, MT02c, MQW95, MCFT99, OT95, OCRS07, PO03, PT03, QOQOV^{+}09, RGT17, STY99, SD13, SLT02, SCZM00, Tem97, WS08, YZYL07, YZ14, ZJS^{+}11]. Level-2 [Ric99]. Leveraging [PRS14, AU21]. LFTHREADS [GP08]. Libraries [Ano00c, BCR01, CGS^{+}20, GF00, Jon91, MLGW18, MM14, ARvW03, CBM10]. Library [Ano98b, ABN00, BFA^{+}15, CGR92, EHG95, Gib94, GHG^{+}98, Kem02, KEMC22, Man91, Rót19, WN10, Yas95, Ada98, Boe05, CS00, GP08, GOT03, Mix94, Ong97, TB97a, TB97b, Yam96, Lev97]. life [KU17]. Light [IXS19, Way95, LZTZ15]. Light-Weight [IXS19, Way95]. Lightweight [AGN09, Col90b, Don02, Est93, Fin95, Hai97b, SLJ^{+}18, CASA14, Hai97a, LVN10, MMN09, MEG94, VACG09, WSKS97, LKBK11]. like [DJLP10, Jon86, VV11, Kor89]. limit [ROA14]. limitations [Gal94, HL08]. limited [Bri89]. Limits [LB95, LB96b, AAKK08]. Line [Ano00c, FSPD16, FdL02]. Linear [KLDB09, Loe97, MR09, AAC^{+}15, Bak95a, ÇM20, MM07, YSY^{+}09]. Link [Ano00b]. Linked [WJ12]. links [WW96]. LinkScan [Ano00b]. LINQits [CDL13]. Lint [Kor89]. Lint-like [Kor89]. Linux [Ano97a, Ano00b, Ano00c, Ano97a, RGK99, SKP^{+}02, WTKW08, ZSA13]. Linux/AXP [Ano97a]. Linux/FreeBSD [Ano00b]. Linz [Hon94]. liquid [KRBJ12]. Lisp [Nor90]. List [DV99, WJ12, VV00]. LiteRace [MMN09]. little [CDL13]. liveness [BMTZ21, GMR09]. LLCs [PBL^{+}17]. Load [HBTG98, HR10, KMAG01, KC98, KEMC22, KRH98, PGB16, VPQ12, WYT^{+}20, Chr95a, Chr95b, Chr96, MKIO04, TKHG04]. load-adaptive [TKHG04]. Load-Balancing [KC98, PGB16, Chr96]. Load-Load [HR10]. Loadable [ZSA13]. Loading [PCM16]. Local [DGK^{+}03, IEE95, Whi03, ZWGX22, HZD13, LTMK21, RWT21, ZLW^{+}16]. localities [CS95a, CS95b]. Locality [BS96, CCWY17, PEA^{+}96, TKK21, Wei98b, HWW93, LK13, PSG06a, PSG06b, PSG06c, Sin99, SD95]. locality-cognizant [LK13]. Localization [OB13]. Location [USE93a, KKT^{+}18]. Location-Independent [USE93a]. Lock [ALB^{+}18, EFJM07, MNU^{+}15, NM10, PGB14, AR19, CS12, GP08, MLS15, MCRS10, Mic04, ST05, TMCP10, ZLW^{+}16]. lock-free [AR19, GP08, MLS15, Mic04, ST05]. Lock manager [Hol98b]. Locking [Bal02, LDT^{+}16, AFF06, Lie94, MMTW10, RWT21, RD06, ZLW^{+}16]. Locks [ACR01, ALS10, MT93, OCT14]. LOCKSMITH [PFH06]. LOGFLOW [NTKA99]. Logic [Bre02, KI17, TAN04, BK13]. Logic-Centric [Bre02]. Logical [CR02]. LOIS [KT17]. Long [CmL21]. Long-Term [CmL21]. longer [XHB06]. Longest [BVP^{+}19]. LOOG [IXS19]. Looking [ECX^{+}12]. lookup [KNPS16]. Loop [BB20, RLJ^{+}09, SSP99, JMS^{+}10, KVN^{+}09, UZU00]. loop-level [KVN^{+}09]. loops [D’H92, FN17]. Low [ABLM19, Ano00a, Ano03, BGH^{+}12, PHBC18, RGT17, SM19, TCF22, ZHCB15, GPS14, PPGS20, RRP06]. Low-level [RGT17, GPS14]. Low-order [TCF22]. Low-overhead [ZHCB15, RRP06]. Low-Power [Ano00a, Ano03, BGH^{+}12, PHBC18, SM19]. Low-Rank [ABLM19]. LPVM [ZG98]. LRU [CmL21]. Ltd [Ano00b]. lunch [DTLM14]. Luther [ACM99b]. LXM [SV21]. Lyon [FR95].

M [Ano00c, USE01, FKD^{+}97, MSU^{+}16]. M-Machine [FKD^{+}97]. MA [Ano94f]. Mach [USE91a, CB89, CB90, Hol99b, Koo93, MRGB91, RBF^{+}89]. Machine [Ama89, CmL19a, CSS^{+}91b, DS16, FKD^{+}97, KA97, KKDV03, Laf00, LPM17, USE01, CSS^{+}91a, CSS^{+}91c, DLM99, Gle91, MEG94, Ném00, Pra95c, SKS^{+}92, Ven97, CGSV93, Evr01, PRB07]. Machines [BSSS14, CYYL18, Den94, GH98, GBK^{+}09, RCRH95, STY99, BBM09, DKF94, GKZ12, GC92, Kus15, MRG17, TSY99, TSY00, VPQ12]. macromolecular [ABC^{+}15]. Made [Har99]. Magiclock [CC14]. main [AKSD16, BBH^{+}17, ZTN19]. main-memory [ZTN19]. maintenance [TNB^{+}95]. Major [PFP^{+}22]. makes [Van97a]. Making [BDLM07, LFA96, Low00, Pla93, PLT^{+}15, YCW^{+}14]. malloc [Kus15]. Mambo [WZWS08]. MAMPO [GJ11]. managed [WLG^{+}14]. Management [ALB^{+}18, ABLL92, GMGZP14, HC17, HRH08, KG94, LG06, LLS06, RSBN01, STY99, VCM19, ZP11, Bak95a, BM91, DBRD91, HCD^{+}94, ICH^{+}10, Jef94, KKH04, RCG^{+}10, SS95]. Manager [Ano00b, PDMM16, Ply89]. Managing [ARA20, Blu92, FGKT97, MVY05, PJZA07, SEP96, VS11b, ROA14, WSKS97]. MANNA [HMT^{+}96, Sod02]. manual [MS87, PO03]. Many [FMY^{+}15, GBK^{+}09, PVS^{+}17, PHBC18, VCM19, DTR18, MLCW11, MTPT12, San04, TPZ21]. Many-Core [FMY^{+}15, GBK^{+}09, PVS^{+}17, PHBC18, VCM19, DTR18, MLCW11, MTPT12, TPZ21]. Many-Thread [GBK^{+}09]. Manycore [BMF^{+}16, KS16, BWDZ15, HFV^{+}12]. Maple [YNPP12]. Mapping [CCK^{+}16, HLH16, LBvH06a, LBvH06b, LBvH06c, NTR16, WK08a, WK08c, WK08b]. Mappings [Lun97]. MapReduce [IXS18]. Maps [BC94]. March [IEE97, USE92b]. Mark [Ano00c]. Markerless [LH09]. Markov [SBC91]. Martin [ACM99b]. MASA [HF88]. Masking [BAZ^{+}19]. Massachusetts [USE93a]. Massive [EJRB13, OR12, SMZ18, Mus09, RČV^{+}10]. Massively [BÇG14, KR12, MSU^{+}16, TSV12, BS10a, ÇFG^{+}12, CDD^{+}10, Lu94, NJ00, NPA92, ROA14, RWT21, WT10, WOKH96]. master [TJY^{+}11]. master-slave [TJY^{+}11]. Match [TKK21]. Matching [HPA^{+}15, OR12, DGSB20, HFV^{+}12, KGPH12]. Mathematica [Tam95]. mathematical [KI16]. Matlab [Ano97c, Bra97]. Matrices [But13, SGLGL^{+}14]. Matrix [NBS^{+}15, QOIM^{+}12, YFF^{+}12, CSV10, DTR18, KS21, QOQOV^{+}09]. matrix-vector [CSV10, KS21]. matter [ZJS10]. maxflow [BÇG14]. Maximal [HH16, HR16]. Maximisation [SC17]. maximize [RCG^{+}10]. Maximizing [LKBK11, TEL95, TEL98a, TEL98b]. Maximum [AT16, HH11, MP13, GJ11]. May [ACM93b, ACM96, ACM99a, Cha05, IEE94a, IEE94b, IEE94d, SS96, MMTW10, Pra95c]. MBTAC [FRL18]. MD [IEE02]. MDMA [Spe94]. measured [ECX^{+}12]. Measurement [LLD17, TMC09]. measurements [JFL98]. Measuring [FMY^{+}15, DTLM14]. Mechanising [Loc18]. mechanism [FD95, GCC15, PWWD18, WHJ^{+}95]. Mechanisms [BB20, KPC96, KC99, SK97, TVB^{+}13, Loe05, Men91, PT03]. Media [Ano03, Van97a]. medium [CDD^{+}10]. Meeting [DLM99]. meets [Tam95]. Member [BS99]. Memories [HKSL96, KHP^{+}95]. Memory [ALSJ09, AJK^{+}12, BS96, BMBW00b, BD00, BP19, BAZ^{+}19, CFC^{+}22, CH95, DM98, EJ93, EE09a, FMY^{+}15, GMR98, GMGZP14, GH98, HG91, HL07, IXS18, JLA16, KZTK15, KZC15, KKH04, KUCT15, LK20, LSB15, LB92, LB17, LML^{+}19, MSM^{+}16, MVZ93, MCT08, Nak01, RCC14, Rob03, RCRH95, SCL05, STY99, SLT03, SZ02, TAM^{+}08, Thr99, TCF22, Tro18, VCM19, Ver96, WJA^{+}19, WC99, XWG^{+}14, YMR93b, ZM07, ZLJ16, ATLM^{+}06, AKSD16, AAKK08, BS06, BGDmWH12, BCG^{+}95, BMTZ21, BBH^{+}17, BMBW00a, BMBW00c, BLM06, BDLM07, BA08, BB00, Boo93, BAM07, CMF^{+}13, Cha05, Cho93, CNV^{+}06, DLZ^{+}13, DLCO09, DPZ97, EKKL90, EV01, FF10, GCC15, Gle91, GL98a, GS00, GKK09, HB92, HWW93, HG92, HHPV15, ISS98, KFG15, Luk01, MLS15, MCRS10, MSM^{+}10, MLC04, MMTW10, MTS10, Mic04, MTC^{+}07]. memory [MVY05, NPC06, NAAL01, OCT14, RMV22, RWT21, SLT02, TSY99, TSY00, TPZ21, TVD10, TVD14, VTSL12, WK08a, WK08b, WK08c, XHB06, YMR93a, YSY^{+}09, YN09, kSYHX^{+}11, ZKW15, ZHCB15, ZTN19]. Memory-Divergent [WJA^{+}19]. memory-intensive [YSY^{+}09]. Memory-level [EE09a]. Memory-safe [Tro18]. Memristor [KNE^{+}14]. Memristor-Based [KNE^{+}14]. MemSAT [TVD10]. mergesort [AL21]. Merlot [MTN^{+}00]. mesh [ABC^{+}09, Mus09]. mesh-based [Mus09]. Meshes [HBTG98, Lep95]. Message [BWXF05, HLB94, KKDV03, PH97, Ada98, BCM^{+}07, DLM99, FM92, Met95, PRS14, SCM05, FGT96, PS01]. message-handling [Met95]. message-passing [BCM^{+}07, FM92]. messages [Koo93, SD95, WHJ^{+}95]. meta [FKS^{+}12]. meta-scheduler [FKS^{+}12]. Metering [LMA^{+}16]. Method [CYYL18, LPK16, LHG^{+}16, MAF19, SKG^{+}11]. Methodology [Sri95]. Methods [CMK00, FGKT97]. Methylation [LZL^{+}20]. Metrics [EE14, VS11a]. Metro [Ano00b]. Metro-X [Ano00b]. Mexico [Ano94e, Gol94]. MFC [Oni97]. MICE [BK96]. Michael [Yam96]. Michigan [Ano94d]. Micro [Mat97]. Microarchitectural [FMY^{+}15, LS11, WHG07]. Microarchitecture [KM03, AMPH09, LSF^{+}07, Wil98]. Microarray [GAC14]. microbenchmark [BO01]. Microbenchmarking [FMY^{+}15]. Microcontroller [BP05, PUF^{+}04, KBP^{+}03]. microkernel [BO96]. Microprocessor [KE15, SU96, Aru92, CJB^{+}15, Gul95]. Microprocessors [Gep00, KET06b, CGL92a, CGL92b, HL07, RCG^{+}10]. microthreading [CSK^{+}99]. microthreads [CTYP02]. Mid [MSFC21]. Mid-Kernel [MSFC21]. Middleware [RBPM00, KBH^{+}03]. Migrant [MR98]. Migrating [PG92, BDF98]. Migration [ABN99, Sat02, WG99, CWS06, CSM^{+}05, HWW93, ISS98, Pha91]. migrations [PGB14]. MIMD [FSYA09, Gle91]. MiMPI [GCC99]. Min [JEV04]. Min-cut [JEV04]. Minimal [BMR94, CSS^{+}91b, IMMP20, Lun97, TY97, CSS^{+}91a, CSS^{+}91c]. Minimizing [SPDLK^{+}17]. Mining [OB13, WLT19, GBP^{+}07]. Mining-Based [OB13]. Minneapolis [IEE92, IEE95]. Minnesota [IEE92, IEE95]. MIPS [Aru92, Swe07]. miss [SLP08]. mitigate [ASSS19]. Mitigating [EPAG16, OdSSP12]. Mitigation [PHBC18, AU21]. Mitosis [MGQS^{+}08]. Mix [TKK21]. Mixed [XIC12]. mixture [SC17]. ML [BCL^{+}98, DL93, MT93]. Mobile [BDF98, USE93a, APX12]. Mobility [CWHB03, BHK^{+}04, SJ95]. mode [AR19, FRL18]. mode-directed [AR19]. Model [AHK08, ACMA97, CC18, Chl15b, CSV10, CBN^{+}00, DTLW16, Dil00, ES97, FG91, Gao93, Loc18, MSM^{+}16, ND16, SAC^{+}98, Sto02, TESK06, VK99, WC99, AAJ^{+}19, ABG^{+}08, BA08, BMV03, CNQ13, Car89a, CYZ98, Chr95a, Chr95b, Chr96, DLZ^{+}13, Dil93, DSH^{+}10, DC07, GKZ12, JPS^{+}08, JD08, LZW^{+}13, MSM^{+}10, MQ08, PAdS^{+}17, PG03, RSB^{+}09, Stä05, TMAG03]. Model-Checking [ES97, Sto02]. Model-driven [CSV10, RSB^{+}09]. Modeling [KMjC02, KE15, PPG11, Rót19, TAM^{+}08, WJA^{+}19, AMC^{+}03, AG21, CIM^{+}17, DKF94, EE10, EE12, Mao96, SBC91, Squ94, TR14]. Models [CMK00, CH95, Den94, HYY^{+}15, KZC15, Kim14, KW17, LB17, ROPdlT22, ST98, VT96, BAM07, But14, Cho93, Cor00, Gil94, SC17, TVD10, VDBN98, XIC12, ZKW15]. Modern [DFC^{+}19, GK05, GBP^{+}07, HL07, NJK16, ZJS10]. modes [WZWS08]. Modular [Chl15a, FQS02, FFQS05, JBK18, Kuc92, NT14, SZM^{+}13, FK12, GBCS07, MJF^{+}10, ZSJ06]. modularity [LK15]. Module [ALSJ09, ZSA13]. modulo [LQ15]. Molecular [LNI^{+}19, MAF19]. monad [FKS^{+}12]. monadic [LZ07]. Monitoring [BBFW02, BBFW03, DJLP10, MC06, NFBB17, VGK^{+}10a, VGK^{+}10b]. Monitors [Bec01, SS91, KPPÉR06]. Monsoon [NCA93]. Monte [LTMK21]. Montecito [MB05]. Monterey [USE91a, Ass96, USE96, USE01]. Mosaic [Ano94d]. Most [PLT^{+}15]. mostly [BBYG^{+}05]. Motifs [LZL^{+}20]. Moving [Ait96, Sim97]. MP [Pea92, TTY99]. MPD [PHK91]. MPEG [BC00]. MPI [PS01, Vre04, Ada98, ALW^{+}15, ALB^{+}18, BBG^{+}10, BK96, BBC^{+}00, BRM03, CRE99, DSG17, HD02, DLM99, FGT96, GCC99, IEE96, LPD^{+}11, MS02, Pla02, SCB15, STY99, SPH96, TSY99, TSY00, TG09]. MPI-based [Ada98]. MPI-OpenMP [MS02]. MpSoCs [GPB^{+}17]. MrBayes [LHG^{+}16]. MS [Wil94a, Wil94b]. MS-DOS [Wil94a, Wil94b]. MSFV [HHOM91, HHOM92]. MSparc [MN00, MD96]. MT [EC98, TJY^{+}11]. MT-BTRIMER [TJY^{+}11]. MTA [Mat97, Smi01]. MTAC [For97]. MTB [AGJ18]. MTB-Fetch [AGJ18]. Mth [MKM17]. MTraceCheck [LB17]. MTS [Gal94]. MUCH [WLM15]. MulTEP [WM03]. Multi [Ada98, AMRR98, AACK92, AGK96, AN22, AR19, ABN00, BC98, Bed91, BBH^{+}17, BC00, BGK94a, BGK94b, BGK96, CV98, CL95, CKRW99, CWHB03, CdOS01, CCC12, CCK^{+}16, CC18, CvdBC18, CFC^{+}22, cC91, Chr01, CR02, Coo95, CNZS17, DV99, DS16, DTLW16, EBKG01, FMY^{+}15, FNS^{+}22, FD96, FdL02, FF20, FJ08, GVT^{+}17, GK94, Gil93, GS06, GH98, HC17, HG91, IXS18, ITF^{+}22, III01, JY15, Jon91, JLS99, KI95, KW17, Kno14, KRH98, Kuc92, KTR^{+}04, LK15, LK20, LB92, Leg01, LKBK11, LZL^{+}20, MLGW18, MNU^{+}15, Mas99, MTN^{+}00, McC97a, McC97b, MS15, MP13, MG15, MCS^{+}22, MCFT99, MGK^{+}00, Myl22, Nag21, NJ00, OR12, PCPS15, PTMB09, PWWD18, PKB^{+}91, PM14, Pul00, PGB16, RR93, RCC14, RBPM00, RKCW98, RVR04, RS08, SV19, SP00a, STW93, Sch90]. Multi [SKG^{+}11, SMZ18, Sei98, Sei99, Smi92, Ste01, SBKK99, TGO99, Tan87, TCS20, Tra91, TCF22, TLGM17, VSDK09, VS11a, VB00, VCM19, VK99, Wal00, YDLW20, YLLS16, AG21, ABD^{+}12, ASSS19, BWDZ15, Bak95a, BK13, BM07, BIK^{+}11, DSEE13, CNQ13, CIM^{+}17, ÇFG^{+}12, CASA14, CKRW97a, CKRW97b, CSB00, CYZ98, CL00, CSM^{+}05, DWYB10, Don92, DGSB20, EFG^{+}03, EHSU07, FTAB14, FWL03, FGG14, GCRD04, GCC15, GPR11, HLGD19, JCP17, KHP^{+}95, KDM^{+}98, KKH04, KD22, Kep03, Küc91, KBF^{+}12, Lan97, LBvH06a, LBvH06b, LBvH06c, LVA^{+}13, LZW^{+}13, MLCW11, MLC^{+}09, MS03, MKK99, Mus09, NFBB17, NH09, NSH14, OA08a, OA08b, OA08c, PYP^{+}10, PMCP22, RČV^{+}10, RGT17, RKM^{+}10a, RKM^{+}10b, RGK99, SCB15, Sam99, SC17, SE12, SV98, Smi06, Sto02, SQP08a, SQP08b, SQP08c, SMQP09, ST05, Tem97]. multi [TCG95, TMAG03, TJY^{+}11, VIA^{+}05, VDBN98, VV00, VPQ12, WCC^{+}07, WCV^{+}98, YZ07, Yan97, Yee20, YSY^{+}09, YN09, kSYHX^{+}11, YKL13, ZKR^{+}11, dB09, vPG03, Ano97b, CH04, Mix94]. Multi- [FMY^{+}15]. multi-ALU [KDM^{+}98]. Multi-C [Mix94]. multi-context [Yan97]. Multi-Core [CC18, CvdBC18, FJ08, IXS18, KTR^{+}04, MNU^{+}15, PM14, TCF22, ÇFG^{+}12, CSM^{+}05, DWYB10, KBF^{+}12, MLC^{+}09, Mus09, RGT17, SMQP09, WCC^{+}07, YZ07]. Multi-Cores [CCK^{+}16]. Multi-CPU [PGB16]. Multi-dimensional [AR19]. multi-engine [CNQ13]. Multi-Interval [AN22]. Multi-Level [RR93, CCC12]. Multi-Level-Context [JLS99]. multi-physics [AG21]. multi-process [WCV^{+}98]. Multi-Processing [MLGW18]. Multi-Processor [SV19, VIA^{+}05, YN09]. Multi-protocol [ABN00]. Multi-shift [Myl22]. Multi-Tasking [CvdBC18]. Multi-Thread [HG91, LZL^{+}20, MTN^{+}00, AMRR98, PKB^{+}91, SKG^{+}11, Tan87, Tra91, DWYB10, Don92, ST05, TCG95]. Multi-Threaded [AGK96, BC98, Bed91, BGK94a, BGK94b, BGK96, CL95, CKRW99, CFC^{+}22, Coo95, DV99, FdL02, GVT^{+}17, GK94, Gil93, III01, JY15, Jon91, KW17, Kuc92, LK20, LB92, Mas99, MG15, MCS^{+}22, MGK^{+}00, PCPS15, Pul00, RKCW98, SV19, STW93, Sei99, Smi92, Ste01, SBKK99, TLGM17, VSDK09, VS11a, VB00, VCM19, YDLW20, Ada98, AACK92, BBH^{+}17, BC00, CV98, CWHB03, CdOS01, cC91, Chr01, CR02, DS16, EBKG01, FD96, FF20, GS06, GH98, HC17, ITF^{+}22, KI95, KRH98, LK15, Leg01, PWWD18, RBPM00, RS08, SP00a, Sei98, VK99, Wal00, AG21, ABD^{+}12, ASSS19, BWDZ15, BK13, BIK^{+}11, DSEE13, CIM^{+}17, CASA14, CKRW97a, CKRW97b, CSB00, CYZ98, CL00, DGSB20, EFG^{+}03, EHSU07, FTAB14, FGG14, GCRD04, GCC15, GPR11, HLGD19, JCP17]. multi-threaded [KHP^{+}95, KKH04, Kep03, Küc91, Lan97, LBvH06a, LBvH06b, LBvH06c, LVA^{+}13, MLCW11, MS03, MKK99, NFBB17, NH09, NSH14, OA08a, OA08b, OA08c, PYP^{+}10, PMCP22, RČV^{+}10, RKM^{+}10a, RKM^{+}10b, RGK99, SCB15, Sam99, SC17, SE12, SV98, Smi06, Sto02, SQP08a, SQP08b, SQP08c, Taf13, Tem97, TMAG03, TJY^{+}11, VV00, Yee20, YSY^{+}09, ZKR^{+}11, dB09, vPG03, Ano97b]. Multi-Threading [CvdBC18, CNZS17, FNS^{+}22, LKBK11, MLGW18, McC97a, McC97b, MS15, MP13, Nag21, OR12, PTMB09, RCC14, Sch90, SMZ18, TGO99, TCS20, YLLS16, DTLW16, Kno14, MCFT99, NJ00, RVR04, Bak95a, BM07, FWL03, KD22, LZW^{+}13, MLC^{+}09, VDBN98, kSYHX^{+}11, YKL13, CH04]. multiagent [Bar09]. Multicomputer [FKD^{+}97]. multicomputers [BCG^{+}95].

Multicore [ALSJ09, ABLM19, BCZY16, CCH11, CB16, DVAE18, FSPE20, GJ11, HEMK17, KLDB09, LS11, LMA^{+}16, LYH16, LDT^{+}16, MR09, NBMM12, PGB16, RCM^{+}16, RRK11, SLJ^{+}18, SHK15, SMD^{+}10, THA^{+}12, ZBS15, CNQ13, CN14, CMX10, LK13, LLLC15, NZ17, RCG^{+}10, RKBH11, SCCP13, SE12, ZSB^{+}12, ZTN19]. Multicore/Multithreaded [RCM^{+}16]. Multicores [FSPD16, FSPD17, Ger18, RKK15, DTK^{+}15, GARH14, SSN10]. Multifrontal [ABLM19, But13, Dav11]. Multigrain [AZG17]. Multigrid [SSK^{+}18, RM99]. Multilevel [PPG11, Cat94, JJY^{+}03, LK15]. Multimedia [Spe94, Est93, Gol96]. multimethod [FGT96]. MultiMotifMaker [LZL^{+}20]. Multiple [CB16, FGKT97, HW92, HKT93, NTR16, OR12, CS95a, CS95b, FD95, HKN^{+}92, LT97, TE94b, TFG10, TAN04, WCT98]. multiple-context [FD95]. multiplication [DTR18, KS21]. multiply [CSV10]. Multiprocessing [EKB^{+}92, Len95, NV94, Wal95, DLCO09, MT93, Pra95b, RGK99]. Multiprocessor [AACK92, AKP99, BC00, Cat94, EHG95, GHG^{+}98, HN91, KMAG01, MCT08, Pre90, PPG11, SZ92, SEP96, USE92b, WC99, Zub02, Cho93, DCK07, EKKL90, HB92, KT99, LVN10, LWV^{+}10, PJZA07, Ano94b]. multiprocessor/multithreaded [Cat94]. MultiProcessors [BMV03, AGJ18, BS96, BL96, BLG01, CH95, GMR98, KU00, KKS^{+}08, LS07, LMJ14, LPM17, LA93, MVZ93, MKC97, NS97, TESK06, YMR93b, BR92, GA09, HT14, LGH94, Mao96, Men91, MWK^{+}06, QSQ14, SMK10, Sha98, SKKC09, TAS07, Yoo96b, YMR93a]. Multiprogram [EE14]. Multiprogrammed [MVZ93, TSY99, WXAL22]. Multiprogramming [BHP^{+}03, JJ91, CGL92a, CGL92b]. MultiRace [PS07]. Multistring [BVP^{+}19].

Multitasking [Col90b, Gib94, Gon90, JJ91]. Multithread [BVP^{+}19, LCS04, RRMJ12, SSB^{+}22, SYHL14, CS95a, CS95b, DSH^{+}10, GCC99, JD08, SWYC94, ZG98, Zig96]. multithread-safe [GCC99]. Multithreaded [AddS03, ÁdBdRS08, ABC^{+}93, AT16, Ama98, ALB^{+}18, Ano92a, Ano92b, Ano94e, Ano94g, Ano98a, Ano98b, Ano01, ABH^{+}00, ABH^{+}01, AB01, AB02, AG96, AZG17, ACMA97, ABN00, AKP99, Bal02, BBFW02, BCR01, BBdH^{+}11, BVL09, BKI06, BMBW00b, BF04, BJK^{+}96, BL98, BB00, BMN99, BDN02, BP05, BLG01, BTE98, BNH01, BD06, BGH^{+}12, BBSG11, BH95, CC14, CJW^{+}15, CS02, CmL19a, CmL19b, CGK06, CC04, CSM^{+}21, Chl15a, CH95, Chr95a, Chr95b, Chr96, CT00, CW98, CBN^{+}00, CMBAN08, Dan09, DNR00, DVAE18, DH98, DRV02, DTR18, DO95, EFN^{+}01, EFN^{+}02, EJRB13, EHP^{+}07, EC98, EGP14, FSS06, FT96, FS96, FTP11, FNA^{+}18, FQS02, For97, FLR98, GGB93a, GRS97, GMR98, Goo97, GN00, GN92, HPA^{+}15, HMLB16, HTZ^{+}97, HMNN91, HHOM91, HHOM92, HLB94, HH11, HWZ00]. Multithreaded [HPB11, HYY^{+}15, Hud96, HMT^{+}96, I^{+}94, JBK18, JYE^{+}16, JSB^{+}12, KA97, KKW14, KMAG01, KST04, KML04, KC98, KC99, KMjC02, KR12, KU00, KE15, KG94, Kim14, KU17, KAO05, KEMC22, Kor89, KTR^{+}04, LS07, LG06, LH09, LG04, LB96a, LB98, LB00, LLS06, LvH12, LTMK21, LTM^{+}17, LYH16, LPE^{+}99, Loc18, Loe97, Lun97, Lun99, MGQS^{+}08, MP01, MAG21, MS89, MB99, MD96, MAF19, Moo95, Moo96, MR09, Nak01, NPT98, NGGA94, NTKA99, Nik94, OB13, OTY00, PBDO92, PUF^{+}04, PG92, PG96, PG99, PF01, PHK91, PFP^{+}22, PWL^{+}11, PS01, QOIM^{+}12, RCM^{+}16, RW97, RCC12, REL00b, Rin01, RB18, RNSB96, RSBN01, RRK11, RBAA05, RR99, SPDLK^{+}17, SRS98, SR14, SBN^{+}97, SCD^{+}15, SCL05, SAC^{+}98, She98, SU96, SU01, SZM^{+}13, SGM^{+}97, SMD^{+}10, SR01b, SSYG97, SKK^{+}01, SKV21, Spe94]. Multithreaded [Sri95, SZ02, SUF^{+}12, Sut99, TG99, Ten02, TKA^{+}01, TCI98, TT03, TTKG02, TGBS05, TLZ^{+}17, TLZ^{+}18, TJY98, TSV12, URŠ02a, VTSM12, Vol93, VE93, Wan94, WS08, Wea08, WJ12, Wil97, WLM15, WG94, WC99, Yas95, YWJ03, Yoo96a, YMR93b, ZSA13, Zha00, ZJS12, ZBS15, ZP11, ZAK01, Zub02, ÁdBdRS05, ACD^{+}18, Aga89, Aga91, Aga92, ABF^{+}10, ABC^{+}15, AAC^{+}15, ACC^{+}03, AGEB08, Ann96, Ano94b, Ano95a, Ano95b, A^{+}01, ABC^{+}09, AR17, AR19, Aru92, BGDmWH12, BBFW03, BRRS10, BMTZ21, BGZ97, BCHS00, BAD^{+}10a, BAD^{+}10b, BCG13, BGC14, BMBW00a, BMBW00c, BYLN09, Blu92, BL93, BL94, BJK^{+}95, Blu95, BL99, BS10a, BÇG14, BEKK00, BPSH05, BS10b, BNS11a, BNS11b, BNS12, CZWC13, CS00, CMS03, Car89b, CB89, CB90, ÇFG^{+}12, Cat94, CL94, CN14, CS12, CDD^{+}10, CLL^{+}02, Cho93]. multithreaded [Cho92, CDP^{+}21, ÇM20, CGL92a, CGL92b, CJB^{+}15, DJLP10, DSG17, Dav11, DL93, DKF94, EJK^{+}96, Eic97, EG11, Est93, Evr01, Fan93, Far96, Fer13, FF04, FFQS05, FF08, FFY08, FRL18, Fuj97, GMW09, Gal94, GJ11, GGB93b, GK05, GPS14, GB20, GL98b, GL98a, Gol96, GRS06, GRR06, GA09, GLC99, HMC97, HFV^{+}12, HF88, HLB90, Hig97, HMN^{+}92, Hop98, JMS^{+}10, JWTG11, JFL98, JSMP12, JSMP13, Joe96, JSB^{+}11, KGPH12, KR01a, KR01b, KNPS16, KS21, KBP^{+}03, Kub15, Kus15, LLLC15, Lea96, Lei97, Len95, LGN92, Lev97, LLL10, LCH^{+}08, LMC14, LSW^{+}18, LBE^{+}98, LT97, Lu94, Lu95, LC13, Mah11, Mah13, MEG03, MS87, Mil95, Mis96, Mix94, MC06, MKR10, MQ07, NB12, NR06, Ném00, NPA92, ND96, NZ17, Omm04, Par91, PFV03, PJZA07, Pha91, Ply89, PDP^{+}13, PS03, PS07]. multithreaded [Pra95c, PT03, PPGS20, RGG^{+}12, RCM^{+}12, Raj93, RCG^{+}10, RHH10, REL00a, REL00c, Rei95, ROA14, Roh95, RWT21, RS07, SBCV90, SBC91, SR01a, SV96c, SV96a, SV96b, Sch98, SRA06, Sha95a, Sha95b, Sha98, She02, SLG06, SP00b, Shi00, SP05, Sim97, ST00a, ST00b, Sod02, SSN10, Squ94, Sri93, Sta90, Sun95, SMS^{+}03, TMC09, TMCP10, TPZ21, TR14, TV10, TG09, TP18, TE94a, The95, TKA^{+}02, TB97a, TB97b, TKHG04, TLZ^{+}16, Tod95, Tsa97a, TDW03, UZU00, VGR06, Ver97, Ver96, VGK^{+}10a, VGK^{+}10b, WS06, WCC^{+}07, Way95, WT10, XIC12, XSaJ08, Yan02, Yan97, YZYL07, Yoo96b, YM92, YMR93a, YNPP12, ZJS10, ZP04, WM03, LP09]. Multithreading [ÁMdBdRS02, AH00, AGJ18, Ano99, Ano05, BBG^{+}10, BB20, BWXF05, Bec00, Bee98, BW97, BD00, BL96, BPL07, Bre02, BLPV04, But13, CCH11, CCK^{+}16, Cro98, Dug95, EEL^{+}97, Eng00, Eng95, Esp96, EKB^{+}92, FBF01, FKT96, GHG^{+}98, GV95, Gul95, Gun97, GSL10, Har99, HBTG98, HTDL18, ILFO01, IBST01, KPC96, Kel94a, Kel94b, Kho97, KF97, KNE^{+}14, KLH97, Kwo03, KET06a, KET06b, LPS07, LH94, LEL^{+}97a, LEL^{+}97b, LEL^{+}99, LRZ16, MB07, Man91, MHG95, MN00, MKC97, Nag01, Oni97, OA19, ÖCS01, PJS15, PT91, PST^{+}92, Pea92, Pra97, RLJ^{+}09, RG03, ROPdlT22, RD96, SSP99, SPY^{+}93, SW08, SCv91a, SP07, SLG04, SHW19, SRU98, DFC^{+}19, Sin97, Smi01, ST00c, SKA01, TY97, Ten98, TAK^{+}00, TESK06, VT96, WWW^{+}02, WCW^{+}04a, Wei97, YG10, ZL10, Zig96, AAHF09, AU21, AAKK08]. multithreading [ABB^{+}15, BCM^{+}07, BGG95, BR92, Boo93, CHH^{+}03, CCC12, Div95, DN94, Dub95, Dye98, EE09a, FM92, Fis97, Fon97, GWM07, GBG95, Gea98, GEG07, GE08, Gro03, HB92, HCD^{+}94, Hol98a, HH97, IAD^{+}94, KIM^{+}03, KCCD99, Kim94, KG07, KT99, KLH^{+}99, LK13, LGH94, LSS12, LZW17, LZSS19, LB95, LB96b, LZL^{+}14, Loi95, LVS01, LZBW14, Luk01, MIGA18, MWP07, Mao96, MKIO04, MGL95, MMM^{+}05, McM97, Met95, MKR02, MAAB14, OAA09, Ong97, PSG06a, PSG06b, PSG06c, PG01, PHCR09, PV06, Pra95b, RM00, RB22, RR96, RPNT05, San04, Sch91, SCv91b, Sen19, Sin99, SW16, STV02, Swi09, TK98, TSCH99, TO10, Tsa97b, TEL95, TEE^{+}96, Tul96, TEL98a, TEL98b, URŠ02b, URŠ03, VPC02, WLG^{+}14, WW93, WCW^{+}04b, WCW^{+}04c, WCW^{+}04d, WXAL22, YCW^{+}14, Lar97]. multithreading-based [GE08]. must [NA07]. mutable [HL93]. Mutex [Hol98b]. mutual [BRE92]. My [BVM19]. Mysteries [Hol99b].

name [ORH93]. Nanophotonic [VSM^{+}08]. Narrow [EUVG06, YSS^{+}17, YSS^{+}19]. NAS [CRE99, GH98]. native [SJ95]. navigating [TVD14]. NDP [Ano97a]. Near [TKZ^{+}18]. Nearest [JY15]. Nearest-Neighbor [JY15]. Need [SLG04, RPNT08]. Neighbor [JY15]. Nelson [Ano00c]. Nested [EW96, NB99, TGO99, TGO00, YZ14]. Net [Ham96]. Net-Centric [Ham96]. Netburst [KM03]. Nets [KMjC02, MKC97]. Network [ACM98a, RM03, ARB^{+}02, Chl15a, Don02, GRS97, HH11, KML04, KRH98, LZS^{+}08, NGGA94, YG10, ZP11, BDM98, GL07, KGPH12, LZ07, LLL10, LCH^{+}08, OCRS07, RČV^{+}10, RPNT05, Sta90, ZP04, PH97]. Network-Facing [KML04]. Network-I [RM03]. Network-I/O [RM03]. Network-on-Chip [LZS^{+}08]. Networked [CT00, FGKT97]. Networking [ACM98d, ACM00, Hol12, LCK11, DWYB10]. Networks [IEE95, KLH97, Lu98, RR93, SM19, PWWD18, SMK10]. Neumann [HG92]. Neural [SM19, PWWD18]. Neurons [LTM^{+}17]. newly [Ano95a, Ano95b]. NewOS [TLA^{+}02, Gei01]. Newport [USE92b].

News [Ano97c, Bra97, Gar01, Mat97, McM97]. Newton [AG21, CYYL18]. Next [ARB^{+}02, EEL^{+}97, TSV12, CH04]. Next-Generation [EEL^{+}97, TSV12, CH04]. Nexus [FKT96]. NFS [Ano95a, Ano95b]. NFV [GDSA^{+}17]. Niagara [KAO05]. NLM [Day92a, Day92b]. NLM-Based [Day92a, Day92b]. NoC [YL16]. node [TK98]. Nodes [EHG95]. noise [GA09]. Non [Caz02, Coo95, JLS99, KIAT99, LB17, MWP^{+}21, SGM^{+}97, Tra91, Ann96, RMV22, RGK99, SCG95, SKG^{+}11]. non-blocking [Ann96]. Non-Deterministic [LB17]. Non-Intrusive [Caz02]. non-invasive [RGK99]. Non-numeric [SGM^{+}97]. Non-preemptive [JLS99]. Non-Serializable [MWP^{+}21]. Non-Strict [Coo95, Tra91, KIAT99, SCG95]. non-temporal [RMV22]. non-uniform [SKG^{+}11]. Nonblocking [HH11]. nondestructive [AD08]. nondeterminism [HBCG13]. Nondeterministic [DSAD^{+}18, LPS07]. Noninterference [BC02, Smi06]. noninterruptible [AAHF09]. Nonlinear [Nak03, GOT03, Kub15]. nonoperational [GS00]. nonuniformity [WA08]. norm [Ano92a]. Note [AKP99, Lie94]. NOTES [Gil88]. notification [BF08]. Notifiers [Pom98]. Notre [IEE96]. Novel [HG91, GKK09]. November [ACM98d, ACM99b, ACM00, ACM03, Ano91, Ano94e, Gol94, Hol12, IEE90, IEE92, IEE93, IEE94c, IEE02, LCK11, USE91a]. NOWs [SLGZ99]. NP [YZ14]. NPB [EGC02]. NT [Ano98b, Hig97, PG96, Pra95c, Pra95b, TCI98, USE98a, Wil94a, Wil94b, Yam96]. NT-Style [Wil94a, Wil94b]. NUMA [LMC14, ZLW^{+}16]. NUMA-aware [ZLW^{+}16]. number [LSS12, SLF14, SV21]. Numeric [MLGW18, SGM^{+}97]. Numerical [MR09]. Numerics [Ano97a]. NY [SS96].

O [RM03, Ano95a, Ano95b, ABB^{+}15, BDN02, KSU94, LTL^{+}16, Man98, MG15, Yoo96a]. Object [Ano99, BBD^{+}91, BC94, GK94, HH97, KC99, Kim14, NPT98, SJ95, SG96, Ada98, Car89a, CYZ98, CLL^{+}02, FWL03, FL90, JPS^{+}08, LLLC15, Sch98, Wei98a, Yan02, dB09, vPG03]. Object-Oriented [Ano99, BBD^{+}91, BC94, Kim14, NPT98, SG96, HH97, Ada98, Car89a, CYZ98, CLL^{+}02, FL90, JPS^{+}08, Wei98a, Yan02, dB09, vPG03]. Objects [ACR01, CJK95, CR02, Low00, Pra95a, Ric99, Ten02, Yas95, Bak95a, Bri89, DMBM16]. objet [Swi09]. Oblivious [UALK17, UALK19, HL08, HZ12]. Observer [Hol99b]. occupancy [PAB^{+}14]. Ocean [SAC^{+}98]. OCTET [BKC^{+}13]. October [ACM94d, Ano94d, BT01, IEE95]. ODBC [Ano00b, Hig97]. ODBC-compliant [Hig97]. ODBC-ODBC [Ano00b]. ODE [Ano97c, Bra97]. Off [MHG95, AAC^{+}15, DTK^{+}15, Gep00]. off-beat [Gep00]. off-chip [DTK^{+}15]. Off-the-Shelf [MHG95]. offs [Par91]. Old [Wil00]. On-Chip [CmL19a, LKBK11, ZM07, SMK10, TEL95, TEL98a, TEL98b]. On-Line [Ano00c, FSPD16, FdL02]. On-the-fly [Sch89, CWS06, PS03, PS07]. once [Bak95a]. one [QSHI16]. one-sided [QSHI16]. Online [Ger95, OTY00, RCC14, Sei98, Sei99, SRA06, TGO99, HF96, LWV^{+}10, RS07, VGK^{+}10a, VGK^{+}10b]. only [Dil00, MJF^{+}10, NM10, ZJFA09]. onto [LBvH06a, LBvH06b, LBvH06c]. Open [Ano00c, BMF^{+}16, Hai97b, BMF^{+}19, KR01a, KR01b, RBF^{+}89]. Open-Source [Ano00c]. OpenGL [Rót19]. OpenGL- [Rót19]. OpenMP [Cha05, ARvW03, Ano97c, BHP^{+}03, BBC^{+}00, Bra97, BMV03, BO01, CRE99, CDK^{+}01, CM98, DM98, HD02, EV01, JJY^{+}03, KKH03, Lu98, MS02, Mar03, MLC04, MPD04, Mat03, MG15, MM14, Mül03, NAAL01, RBAA05, SLGZ99, Thr99, TGBS05, Vre04, RM99]. OpenMP-oriented [MLC04]. OpenOpt [NSP^{+}14]. OpenPiton [BMF^{+}16, BMF^{+}19]. OpenSPARC [Wea08]. Operand [SP07]. Operating [ACM94d, CLFL94, TLA^{+}02, Gei01, IEE89, IEE94a, MS87, REL00b, SEP96, Ano92a, Ano92b, ASSS19, BDM98, DBRD91, IEE94d, Jef94, Jen95, LVN10, LAK09, Ply89, RBF^{+}89, REL00a, REL00c, She98, Way95]. operation [DKG18, RHH10]. Operational [CKRW99, CKRW97a, CKRW97b]. Operations [KKS^{+}08, KLDB09, SCL05, HMC95, RD06]. operators [RWT21]. Opportunistic [YL16]. Opportunities [GJ97, HL08, Mus09]. OPR [QSHI16]. Optimal [AAJ^{+}19, AT16, GPB^{+}17, Lar95, RCM^{+}12, Lep95, LML00]. Optimisations [TCF22]. Optimistic [WHJ^{+}95, CZSB16, DKG18, VPQ12]. Optimization [BLG01, CmL19b, CvdBC18, GN96, RNSB96, SYHL14, TJY98, TLGM17, WJ12, AMC^{+}03, AMPH09, CDP^{+}21, DZKS12, GOT03, Koo93, RKCW98, Sin99, TPZ21, TO10, ZCSM02a, ZCSM02b]. Optimizations [HYY^{+}15, JSB^{+}12, KET06a, LEL^{+}99, Sut99, ZM07, ABC^{+}09, JSB^{+}11, OA08a, OA08b, OA08c, Roh95, Sen19]. Optimized [Sin97]. Optimizing [DTK^{+}15, KZTK15, PR98, PSCS01, WCZ^{+}07, GS02]. Orange [ACM98d]. orbital [MAF19]. Orchestration [GVT^{+}17]. Order [CJK95, HLGD19, IXS19, RRK11, SMP19, TCS20, NV15, SJA12, SW16, TCF22, ZKW15]. Ordering [DELD18, HR10]. Ordering-Free [DELD18]. Oregon [ACM94b, ACM99b, IEE93]. Organization [HG91, HG92, PPGS20]. organizing [LAK09]. Oriented [Ano99, BBD^{+}91, BC94, Kim14, KS97, LHG^{+}16, NPT98, SG96, Ada98, Car89a, CYZ98, CLL^{+}02, DWYB10, FL90, HH97, JPS^{+}08, MLC04, Wei98a, WP10, Yan02, dB09, vPG03]. Origami [AL21]. Orlando [ACM94a, ACM98d]. OS/2 [DN94, Kan94, Kel94a, Kel94b, Rei95, Ric91, Rod94]. oscillations [BD06]. OSF [BM91]. OSF/1 [BM91]. Other [SPY^{+}93, MMTW10]. Ottawa [BT01]. Out-of-Core [QOIM^{+}12, ABC^{+}15]. Out-Of-Order [IXS19, TCS20, SJA12, SW16]. output [MP89]. Outstanding [LSB15]. Overall [SEP96]. Overcome [SW08]. overflow [KOE^{+}06]. Overhead [DSR15, RRP06, YL16, ZHCB15]. Overheads [SHK15]. Overlay [DFC^{+}19]. overview [Li05]. Own [BS99, Sho97a, Sho97b]. Oxford [ACM94c].

P [Ano00b, Nik94, PR05]. P-RISC [Nik94]. P-STAT [Ano00b]. P-Thread [PR05]. Pacbio [LZL^{+}20]. Pacific [IEE89]. Pacifier [QSQ14]. Package [Ano94c, FL90, HCM94]. packages [GOT03, OT95, PL03]. Packaging [RR93]. Packard [BLCD97]. Packet [AHW02, LCH^{+}08, MVY05, WCZ^{+}07]. page [CNV^{+}06]. page-based [CNV^{+}06]. PageRank [KG07]. Pages [JLA16]. Paging [FD96, FdL02, Sei98, Sei99]. Pagoda [YSS^{+}17, YSS^{+}19]. PaiLisp [KI95]. pain [Gus05]. Pajé [CdOS01, CSB00]. Palo [ACM01]. panel [Ano94e, Bak95b, HCD^{+}94, IAD^{+}94]. Paper [ABH^{+}01, TKA^{+}01]. papers [ACM93a, ACM94b, ACM95b, ACM98b, KKDV03, Cha05]. par-monad [FKS^{+}12]. ParADE [KKH03]. Paradigm [EW96, JD08, LK15, PPA^{+}13, BCG^{+}95]. Paradigms [CM98, HD02, YMR93b, YMR93a]. Parallel [ABC^{+}93, AMRR98, ARA20, Ama89, ABNP00, ACMA97, Bau92, BC00, BFA^{+}15, BE13, BBC^{+}00, BTE98, CZS^{+}17, CL95, CDK^{+}01, CBN^{+}00, DS16, Den94, DSR17, EJ93, FHM95a, Gil94, GSC96, GJ97, GAC14, HMLB16, Hon94, HN91, JY15, KTLK13, KI95, KEL^{+}03, Kon00, KKDV03, Kwo03, LK20, Len95, LHS16, LFA96, Mah11, MS02, Mar07, MAG21, MG15, MRG17, Nak03, NS97, Pan99, QSaS^{+}16, ROPdlT22, Sch17, SCv91a, SAC^{+}98, SRU98, Tro18, WYT^{+}20, WC99, YFF^{+}12, ARvW03, ALS10, BBYG^{+}05, BCM^{+}07, BAD^{+}09, BB00, Boo93, BE12, BGK94c, CAR08, CFK^{+}91, Cha05, CSB00, Chr95a, Chr95b, Chr96, ÇM20, DLM99, DESE13, EV01, FHM95b, FD95, Fuj97, GC92, Gol97, GKK09, GEG07, GE08, GB99, HMC97, HF88, Hop98, HWW93, IEE97, JMS^{+}10, Joe96, KTK12, Kep03]. parallel [Kim94, LSS12, Lu94, MT02a, MT02b, MT02c, MR98, Mis96, NJ00, NPA92, OdSSP12, RČV^{+}10, RHH10, SBCV90, Sch91, SCv91b, Sha98, SWYC94, ST98, SGS14, Taf13, TCG95, VPQ12, VGK^{+}10a, VGK^{+}10b, WZSK19, WK08a, WK08b, WK08c, WOKH96, WTH^{+}12, YCW^{+}14, FR95, Vre04, WN10]. Parallel-Based [ARA20]. Parallel-Multithreaded [WC99]. Parallelism [AACK92, ABLL92, BAM93, CSS^{+}91b, DV99, EW96, FKP15, FURM00c, GVT^{+}17, GP95, DK02, LKBK11, LEL^{+}97a, LEL^{+}97b, MG99, MR94, Mar03, MCFT99, NB99, RBAA05, SSP99, SMD^{+}10, SG96, Thr99, WS08, YBL16, Yoo96b, ALHH08, AKSD16, CSS^{+}91a, CSS^{+}91c, EE09a, FN17, FURM00a, FURM00b, HDT^{+}13, KRBJ12, KDM^{+}98, KVN^{+}09, KC09, LGN92, LAH^{+}12, MAH18, QOQOV^{+}09, SLGZ99, SD13, TEL95, TEL98a, TEL98b, VDBN98, VV00, Wei98a, XSaJ08, YZ14, Zig96]. parallelism-aware [LAH^{+}12]. parallèlisme [Zig96]. Parallelization [CRE99, DSAD^{+}18, KC09, LVA^{+}13, RM99, WZWS08, YLLS16, AC09, DC07, JJY^{+}03, MAF19, PO03, RKM^{+}10a, RKM^{+}10b, RRMJ12, TFG10]. parallelized [CJ91]. Parallelizing [BM91, WDC^{+}13, KBF^{+}12]. ParaLog [VGK^{+}10a, VGK^{+}10b].

Parameterized [BCR01, FK12]. Parametric [Ano98b, FRT95]. Paravirtualization [YSY^{+}09]. PARC [Ong97]. Pareto [GPB^{+}17]. Pareto-Optimal [GPB^{+}17]. Parsing [BC00, Lar95, PCM16]. Part [Ano92a, Ano92b, KR01a, McM98b, Hol98a, Hol98b, Hol98c, Hol99a, Hol99b, McM96b, McM98a]. Partial [Loe97, SMP19, RRP06, SP00b, Shi00, ZKW15, HLGD19]. partition [LZW17]. Partitioning [AMRR98, Coo95, D’H92, EW96, SLJ^{+}19, TG99, DCK07, LZL^{+}14, MKR10, SCG95, WW96]. Partitioning-Independent [EW96]. Pascal [Hay93]. PASCO [Hon94]. Passing [BWXF05, TLA^{+}02, FGT96, KKDV03, PH97, PS01, Ada98, BCM^{+}07, DLM99, FM92, PRS14]. Path [BLG01, TAK^{+}00, CTYP02, WCT98]. pathfinder [KPP12, SMP19]. Paths [OTY00, Ano95a, Ano95b]. Pattern [Ano97b, EGP14, OR12, EG11, GBP^{+}07, SCM05]. Pattern-Based [EGP14, EG11]. Pattern-recognition [Ano97b]. Patterns [CmL21, DS16, LPK16]. PC [Mia90, CFK^{+}91]. PCM [AKSD16]. PCM-based [AKSD16]. PCs [CRE99, NV94]. PDE [Chr95a, Chr95b, Chr96, JD08]. PDES [LTM^{+}17]. pedagogical [CMS03]. PegaSoft [Ano00b]. Pennsylvania [ACM96]. Pentium [RGK99]. peptides [MIGA18]. Per-node [TK98]. Per-Thread [Cha02, EE09b]. Perf [FSPD17]. Performance [ACM98a, ACM98d, ACM00, Aga89, Aga91, Aga92, ARA20, ABLM19, BS96, BL96, BRM03, BLG01, BNH01, BGH^{+}12, BBSG11, Cal97, CRE99, CGS^{+}20, CCH11, CCK^{+}16, CCWY17, CH95, Cho92, CT00, CSM^{+}05, CBN^{+}00, CMBAN08, DVAE18, DWYB10, EGC02, EE14, FT96, FSPD17, FBF01, FURM00c, FGKT97, Gal94, Gar01, GN00, HRH08, Hol12, HN91, IEE94b, JFL98, KZTK15, KH18b, KS97, KTR^{+}04, LWSB19, LCK11, LG06, Lep95, LMJ14, LHG^{+}16, LYH16, Mah13, Man99, Mao96, MSM^{+}16, MPD04, ME17, MWK^{+}06, MKC97, MM14, NCA93, NBS^{+}15, NGGA94, Par91, PH97, PS01, QSaS^{+}16, RG03, RVOA08, RKK15, SCD^{+}15, SKV21, SLJ^{+}19, TCI98, TT03, Tsa97a, TLGM17, VP16, Wei98b, WG99, WN10, YWJ03, ZL10, ZAK01, Zub02, AAC^{+}15, AL21, APX12, AAKK08, BGDmWH12, BS10a, BBM09, BMV03]. performance [CML00, Car89b, CIM^{+}17, Cho93, CDP^{+}21, Div95, Don92, DKF94, ECX^{+}12, FL90, FM92, Fis97, FURM00a, FURM00b, GS02, GEG07, GLC99, HLB90, ICH^{+}10, Kim94, KLH^{+}99, LML^{+}19, LB95, LB96b, LBH12, LCH^{+}08, LMC14, LBE^{+}98, MLC^{+}09, Mah11, MCRS10, McM97, PJZA07, PGB12, RGK99, SE12, Sen19, SSkP^{+}07, SQP08a, SQP08b, SQP08c, SKP^{+}02, TMC09, TR14, TG09, The95, VV11, Wan94, WCZ^{+}07, WOKH96, YZ07, YM92, ZJS10]. Performance-area [Par91]. Performance-Driven [LWSB19]. performance-energy [AAC^{+}15]. Performance-Oriented [KS97]. performance-prediction [BMV03]. Performance/Power [RKK15]. performs [Ven97]. perils [Dye98]. Perl [TLA^{+}02]. Perl/Tk [TLA^{+}02]. Persistence [WYT^{+}20, BHK^{+}04]. persistency [RMV22]. personality [CCW^{+}11]. perspective [AG06]. Perspectives [PLT^{+}15]. pessimistic [CZSB16]. Petaflops [SZ02]. Peter [Ano00c]. Petri [KMjC02, MKC97]. PGI [Ano00b]. pH [ACMA97]. Phase [CA20]. Phi [SCD^{+}15]. Philadelphia [ACM96]. Phoenix [ACM03]. Photomosaics [TLA^{+}02]. Phylogenetic [LHG^{+}16, LBH12]. physical [AMPH09, KKT^{+}18]. physics [AG21]. pi [Yee20]. pi-program [Yee20]. PIC [BMV03]. PicoServer [KSB^{+}08]. picture [AC09]. Piecing [Ano97b]. Pipelining [GV95, RVOA08]. PIRATE [ICH^{+}10]. Pitfalls [Hol98a, SPY^{+}93, CL00, San04]. place [SCM05, SGLGL^{+}14]. placement [NLK09, TE94a]. Plagiarism [TLZ^{+}17, TLZ^{+}18, TLZ^{+}16]. Plan [DLZ^{+}13, Pre90]. PlanICS [NSP^{+}14]. Planning [NSP^{+}14]. plans [GARH14]. plastic [MCS15]. Platform [AB01, AB02, CT00, DTLW16, EEL^{+}97, FSS06, BMF^{+}19, Lam95, MT93, PG03, WCW^{+}04b, WCW^{+}04c, WCW^{+}04d]. Platform-Independent [FSS06]. Platforms [LS11, PWL^{+}11, CNQ13, LSS12, RB22]. PLDI [ACM94a, ACM99a]. Plug [DHR^{+}01]. Plug-in [DHR^{+}01]. plus [Ano95a, Ano95b]. PM [AB02]. PM2 [ABN99, AB01]. Point [LWSB19]. Pointer [RR99, SR01a]. pointers [Sim97, WW96]. Points [CC04, CHH^{+}03, SLP^{+}09]. Points-to [CC04, CHH^{+}03]. policies [Eic97, EE09a, KPPÉR06]. Policy [LLKS12, MVZ93]. Polling [Pla02]. Pollution [MPD04]. Polynomial [Kuc92, Küc91]. Pool [PSCS01, LML00]. Pools [Cal97]. POPL [ACM94b, ACM95b, ACM98b]. Port [Koo93]. Portability [VSM^{+}16, SP05]. Portable [AB01, ABN00, BBFW02, Eng00, KF97, LDT^{+}16, Yas95, CS00, GCRD04, Mix94, MT93, MAAB14, TB97a, TB97b]. Portals [BRM03]. Porting [JJ91, Yam96]. Portland [ACM94b, ACM99b, IEE93]. Ports [Man98, Yam96]. posium [USE01]. POSIX [Ano00c, Alf94, BMR94, But97, GL91, GF00, GMB93, HBG01, HBG02, SP05, dlPRGB99]. POSIX-compliant [SP05]. Post [LB17]. Post-Silicon [LB17]. Pot [VSDL16]. Potential [CC14, EGC02, LLKS12, Mou00, DG99]. potentials [ABF^{+}10]. POwer [GJ11, AKS06, Ano00a, Ano03, BCZY16, BGH^{+}12, CMBAN08, MB07, MR09, PHBC18, RCC12, RKK15, RRK11, SYHL14, TVB^{+}13, TLGM17, CDP^{+}21, ECX^{+}12, GW10, MLCW11, MWK^{+}06, Pra95b, PPGS20, Ric91, SM19, SQP08a, SQP08b, SQP08c, CMF^{+}13]. Power-aware [MR09]. Power-Constrained [TLGM17, GW10]. Power-Efficient [BCZY16, SQP08a, SQP08b, SQP08c]. Power-Performance [CMBAN08]. POWER5 [BCG^{+}08, MMM^{+}05, KST04, Ano05]. POWER6 [LSF^{+}07]. powered [Rei95]. PowerPC [BEKK00, SBKK99]. PowerRAC [Ano00b]. Practical [HW92, LMJ14, MNG16, ND16, PBR^{+}15, RR96, TGBS05, BCCO10, LPD^{+}11, RD99, RPB^{+}09]. PRAM [For97, Lep95]. Pre [PR05, Luk01]. Pre-Execution [PR05, Luk01]. Precise [HR16, KUCT15, CLL^{+}02, FF09, WTH^{+}12]. Precomputation [MGQS^{+}08, WWW^{+}02]. Preconditioning [Nak03, GEG07]. PREDATOR [LTHB14]. Predicate [GPR11, How00]. Predictable [BBdH^{+}11]. Predicting [Lun99]. Prediction [AKS06, CMBAN08, DVAE18, IBST01, PBL^{+}17, BWDZ15, BMV03, CTYP02, CPT08, GL98b, RRP06, TFG10, WHG07]. Prediction-Based [CMBAN08, RRP06]. predictive [LTHB14, SRA06]. Predictors [EPAG16]. preemptive [JLS99]. prefetch [AMC^{+}03]. Prefetcher [LYH16]. Prefetching [AGJ18, BL96, GK94, MKC97, SLT03, VT96, LB95, LB96b, Mao96, SLT02, SKKC09]. Prefix [BVP^{+}19, WJ12]. Preliminaries [NBM93]. Preliminary [EHG95]. Preparation [GH03]. preprocessor [Fon97, Mil95]. prescient [AMC^{+}03]. Presentation [Kub15]. presented [ACM93a, ACM94b, ACM95b, ACM98b]. preserving [MSM^{+}11, NFBB17]. pressure [DTLM14, SLP08]. preventing [PRB07]. prevention [GB20]. Price [Ano98b]. Pricing [TT03]. Primer [LB96a, Wil97]. Primitive [Low00]. primitives [BBH^{+}17, LZ07, NLK09]. principle [LAK09]. Principles [ACM93a, ACM94b, ACM95b, ACM98b, TLA^{+}02]. print [Van97a]. priorities [MAH18, STV02]. Prioritization [CmL21, FD95]. Priority [BCG^{+}08, NBMM12, SCCP13, ST05]. priority-based [NBMM12]. Private [Man99]. privatization [HZ12]. Pro [Ano97a, RB22]. Proactive [FJ08]. Probabilistic [EE10, EE12, CHH^{+}03, Smi06]. Problem [HH11, Lee06, YFF^{+}12, BIK^{+}11, Mit96]. Problems [DK02, Nak03, AG21, AR17, Bar09, FTAB14, FR95]. procedure [BGK94c, KASD07, LQ15]. procedures [MCS15]. Proceedings [ACM94c, ACM98d, ACM99a, ACM01, Ano90, Ano94a, Ano94d, AOV^{+}99, Gol94, Hol12, IEE89, IEE90, IEE92, IEE93, IEE94a, IEE95, IEE96, IEE02, Lak96, LCK11, USE89, USE91a, USE91b, USE92a, USE93a, USE93b, USE96, USE98b, USE98a, USE00b, USE01, USE02, ACM92, ACM95a, ACM96, EV01, IEE97, Wat91, ACM93b, ACM98c, RM03, Ano91, DLM99, IEE94b, IEE94c, FR95]. Process [FT96, FG91, BM91, HF96, LVS01, MR98, Ply89, WP10, WCV^{+}98]. process-oriented [WP10]. Processes [CB16, III01, SPY^{+}93, ZSA13, YZYL07, Zig96]. Processing [AHW02, GAC14, MLGW18, RW97, SG18, SS91, WN10, How98, MVY05, Par91, PYP^{+}10, RKHT17, WCZ^{+}07]. Processor [ABC^{+}93, Ano00b, BVM19, BCG^{+}08, BGH^{+}12, EHG95, GV95, HMNN91, HHOM91, HHOM92, KST04, KML04, KAO05, LvH12, MGQS^{+}08, MG99, MTN^{+}00, MVZ93, MB05, SV19, SW08, Sin97, ST00c, SZ02, SBKK99, SUF^{+}12, UALK17, UALK19, WS08, AAHF09, APX12, BEKK00, BH95, CL94, CY09, Cho92, EE10, Fis97, FRL18, Fuj97, Goo97, HF88, HKN^{+}92, HMN^{+}92, KDM^{+}98, Kho97, KBA08, LBvH06a, LBvH06b, LBvH06c, LCH^{+}08, Lu94, MK12, Met95, Moo95, Moo96, OCRS07, Raj93, Sha95a, SJA12, Sin99, ST00a, ST00b, STV02, Squ94, Sri93, Tsa97a, Tsa97b, TEE^{+}96, VIA^{+}05, WCW^{+}04b, WCW^{+}04c, WCW^{+}04d, YN09, ZP04]. processor-based [WCW^{+}04b, WCW^{+}04c, WCW^{+}04d]. Processor-In-Memory [SZ02]. Processor-Oblivious [UALK17, UALK19]. Processors [ARB^{+}02, AH00, Ano01, BF04, CmL19a, CmL19b, EEL^{+}97, FT96, FSPE20, FNS^{+}22, FJ08, GJT^{+}12, GSL10, JGS^{+}19, KS16, KLG08, KU00, KLDB09, LPE^{+}99, MHG95, MCFT99, MR09, ÖCS01, PF01, RCM^{+}16, RYSN04, RRK11, SU01, SR01b, URŠ02a, VS11a, YG10, ZP11, Aga89, Aga91, Aga92, AAC^{+}15, BGDmWH12, BWDZ15, CS95a, CS95b, CN14, CDD^{+}10, DWYB10, Div95, Eic97, EE09a, EE09b, EE12, FD95, GMW09, GBP^{+}07, KBF^{+}12, LLL10, LBE^{+}98, Luk01, MN03, MEG03, MTPT12, Mis96, NB12, NZ17, PFV03, PAB^{+}14, PPGS20, RGG^{+}12, RCM^{+}12, RPNT08, SLP08, SMS^{+}03, URŠ02b, URŠ03, ZSB^{+}12, WM03]. processus [Zig96]. Procs [MT93]. Products [Ano97a, Ano00b, Ano97c, Bra97]. Professional [Ano00b]. Profile [BMR94, SV19]. profiler [DTLM14]. Profiling [BP19, DG99]. Program [BVM19, Chl15a, DSR15, EFN^{+}01, GN96, KKW14, NBM93, PF01, PS01, SHK15, TSY00, TLZ^{+}17, TLZ^{+}18, TJY98, YDLW20, YLLS16, AC09, BGC14, BD06, Cal02, Dan09, Dub95, EFN^{+}02, FRT95, JEV04, JPSN09, Yee20]. Programmability [THA^{+}12]. programmable [PYP^{+}10]. programmation [Swi09]. programmed [PPA^{+}13]. Programmer [Cro98, Wil00, MS87, San04, Swi09]. Programming [ACM93a, ACM94a, ACM94b, ACM94d, ACM95b, ACM98b, ACM99a, BBG^{+}10, BO17, BTE98, But97, CMK00, CV98, CDK^{+}01, Chl15b, CT00, CW98, DM98, FHM95a, FTP11, FA19, HCD^{+}94, Hol98d, Hol98a, Hol98b, Hol98c, Hol99a, Hol99b, ILFO01, KKH03, KN17, KSS95, KSS96, Kno14, KIAT99, LB96a, LB00, LvH12, Mas99, MAG21, NBF96, Nor96, PG99, QOQOV^{+}09, QOIM^{+}12, Rod95b, SBB96, TCI98, Vre04, Wil97, YFF^{+}12, dlPRGB99, van95, ALS10, AR17, AG96, ABG^{+}08, BCHS00, BO96, BYLN09, Bir89, CFK^{+}91, Car89a, CS00, CMS03, Cha05, CYZ98, DSH^{+}10, DMSM18, EV01, FHM95b, GKZ12, Gil94, Gol97, GL07, HMC97, Hyd00, JPS^{+}08, JHM04, KIM^{+}03, Kim94, KN19, LB98, LPD^{+}11, LP09, Man96, MSM^{+}10, MKIO04, MR98, Mix94, NHFP08, Nev99, NBF98, ND96, PG96, Pra97, RR96, RR03, SKS^{+}92]. programming [SV96c, SV96a, SV96b, Sen19, She98, She02, Sun95, TB97a, TB97b, TMAG03, Wal00, WCC^{+}07, Yan02]. Programs [ABNP00, BBFW02, BE13, BLG01, CC14, CJW^{+}15, CRE99, CS02, CC04, CdOS01, Chr01, DRV02, DSR17, EGP14, FQS02, GKCE17, HLB94, ITF^{+}22, JBK18, KH18a, Kri98, LK20, LCS04, Lun97, Lun99, MS89, MGK^{+}00, OB13, PHK91, Rin01, RD96, RR99, SPDLK^{+}17, SBN^{+}97, SYHL14, Ste01, TGBS05, Tra91, Vol93, VE93, ABF^{+}10, BRRS10, BMTZ21, BK13, BCG13, BGC14, Blu95, BE12, BC02, BS10b, BNS11a, BNS11b, BNS12, CZWC13, CJ91, CL00, CLL^{+}02, CVJL08, Cor00, DJLP10, Dil00, DESE13, EFG^{+}03, EG11, EHSU07, FK12, Fer13, FF04, FFQS05, FF08, FFY08, GMR09, GB20, GRS06, GPR11, HLGD19, HZ12, JPS^{+}08, JWTG11, JFL98, KC09, LQ15, Lea96, LMC14, LC13, MS03, MS87, MC06, MQ07, NR06, NH09, NSH14, NV15, OdSSP12, ORS^{+}06, PAdS^{+}17, PDP^{+}13, PS03, PS07]. programs [RVS13, Rei95, RS07, SR01a, SCG95, SRA06, Sen08, SP00b, Shi00, SP05, SGS14, Sto02, Taf13, TR14, TLZ^{+}16, WS06, WTH^{+}12, XSaJ08, YCW^{+}14, YNPP12, ZJS10, ZSJ06, dB09, vPG03]. Progress [FSPD17, TLGM17, ZHCB15]. Progress-Aware [FSPD17]. Progressive [BBdH^{+}11, TGO00]. Project [Ano99]. projection [SSkP^{+}07]. Projections [MQLR16, MLR15]. proliferating [Ano94b]. Prolog [AR17, AR19, EC98, KA97, MGK^{+}00]. Promises [Gar01]. Proof [AddS03, ÁdBdRS08, FKP15, ÁdBdRS05, GLPR12]. properties [BMTZ21, KTLK13, Van97b]. proposal [GP05]. Proposed [GV95]. protect [San04]. Protecting [ZWGX22, ZJS^{+}11]. Protection [CLFL94]. Protein [BCS11]. Protein-Protein [BCS11]. Protocol [GRS97, III01, ABN00, KASD07, QSQ14, RGT17]. Protocols [AB01, AB02, GRR06, TVD14]. Prototype [BMR94, HHOM91, HHOM92, BK96, BVG97, Far96]. prototyping [PDP^{+}13]. Provably [DJLP10, GB99]. provide [Way95]. provides [Hig97]. Providing [PSM01, PSM03]. proving [Taf13]. Provisioning [BSSS14, FGG14]. Pruning [WLT19]. pseudorandom [SLF14, SV21]. PSO [HH16]. PTF [Yam96]. Pthreads [NBF98, Yam96, LB98, AS14, NBF96]. Publications [Bee98]. Publishing [Ano00b, Hig97]. purity [FFQ04]. Purpose [Ber96b, HSS^{+}14, Man98, ZSA13, Ber96a, DC99, DC00, HSD^{+}12, SKA01]. Put [Wal95]. PVM [DLM99, DPZ97, Pla02, ZG98]. PVM/MPI [DLM99]. PyOMP [MAG21]. Python [Swi09, How98, MAG21, Pul00].

Q [Ber96b, Cri98a]. Q&A [Cri98b, Hag02]. QoS [ICH^{+}10, PSM01]. QR [Dav11]. quality [PSM03]. Quantitative [LPK16, NBM93]. Quasi [Pla02]. Quasi- [Pla02]. Queries [TGO99, TGO00]. query [GARH14]. QUERYFLEX [Ano97a]. querying [HF96]. Queue [Cri98b, Cri98a]. queues [DGSB20, SCM05, ST05]. Queuing [VK99, KPPÉR06]. Quick [Ano00b]. QuickRec [PDP^{+}13]. quicksort [Mah13].

R [Nag21]. R-Friendly [Nag21]. R3000 [Aru92]. Race [HM96, KUCT15, LS18, MKM14, SBN^{+}97, Sen08, Yan02, ZLJ16, AFF06, AHK08, EQT07, FF09, HR16, HHPV15, MMN09, NAW06, NA07, PS03, PS07, PFH06, RVS13, WDC^{+}13, XHB06, DWS^{+}12]. race-freeness [AHK08]. RaceFree [LZW^{+}13]. Races [KZC15, FF10, NWT^{+}07, PRB07, PT03, RBK^{+}09]. racy [SRJ15]. RADISH [DWS^{+}12]. Ramada [Ano94d]. Ramada-Congress [Ano94d]. random [LSS12, Sen08]. random-number [LSS12]. Randomization [ZWGX22]. Randomized [Sei98, Sei99, JPSN09]. Rank [AJK^{+}12, ABLM19, Dav11]. rank-revealing [Dav11]. Ranking [DV99, VV00]. Raphson [AG21]. Rapid [DVAE18]. ray [Tod95]. RCDC [DNB^{+}12]. RCU [CKZ12]. Re [ZWGX22]. Re-Randomization [ZWGX22]. Reachability [LCS04, LQ15]. reachability-modulo-theories [LQ15]. Reactions [LTM^{+}17]. Reactive [LvH12, LNI^{+}19, DMSM18]. Reactivity [BDN02]. read [NM10]. read-only [NM10]. Reads [LZL^{+}20, AAJ^{+}19]. reads-from [AAJ^{+}19]. ready [Ano92b, DFC^{+}19]. Real [BC94, IEE94a, IEE94d, JLS99, Kim14, KBP^{+}03, MN00, PSM01, PUF^{+}04, PSCS01, SZG91, SM19, SUF^{+}12, Tet94, WLG^{+}14, dlPRGB99, CZWC13, CMX10, Hol98d, Hol98a, Hol98b, Hol98c, Hol99a, Hol99b, Jen95, JPSN09, MKK99, OT95, PSM03, RPNT05, San04, SZ92, SJB92a, SJB92b]. Real-Time [IEE94a, JLS99, Kim14, MN00, PUF^{+}04, PSCS01, SUF^{+}12, Tet94, dlPRGB99, IEE94d, KBP^{+}03, PSM01, SZG91, SM19, Jen95, MKK99, OT95, PSM03, San04, SZ92, SJB92a, SJB92b]. Real-Time-and-Distributed [BC94]. Real-world [WLG^{+}14]. Reality [LH09]. realizing [YZ14]. Realtime [BMR94]. reasoning [FK12]. rebiasing [RD06]. recognition [Ano97b, LG04]. Reconfigurable [ROPdlT22, SSK^{+}18, DSH^{+}10, LP09]. ReconOS [LP09]. reconstructive [MCS15]. Record [Chr01, UALK17, UALK19, ACM93a, ACM94b, ACM95b, ACM98b, GCRD04, HDT^{+}13, HT14, LSW^{+}18, PDP^{+}13, QSQ14, RD99]. record-and-replay [LSW^{+}18]. record-replay [HDT^{+}13]. Record/Replay [Chr01, GCRD04, RD99]. Recording [MCT08, NPC06, HZD13, LZTZ15, XHB06]. recoverable [LAK09]. Recovery [LAK09, VPC02, WCV^{+}98, YZYL07]. RecPlay [RD99]. rectangular [SGLGL^{+}14]. Recursively [BE13, BE12]. Red [Pla99]. Red-Black [Pla99]. Redesign [KSU94]. RedThreads [HTDL18]. Reduce [DSR15, CCC12, Cor00, KOE^{+}06]. reduced [GA09]. Reducer [LS18]. Reducing [LK20, SLP08, SYHL14, PGB12]. Reduction [Ama89, CL95, HLH16, KLS92, SW08, SMP19, BKK17, HH16, XHB06, YL16, ZKW15, HLGD19]. Reductions [ZAK01]. Redundant [CCK^{+}16, CvdBC18, HTDL18, KS16, MB07, AU21, MKR02, PSG06a, PSG06b, PSG06c, RRP06, WLG^{+}14]. ReEnact [PT03]. Reentrant [ÁMdBdRS02]. Refactoring [Ten02]. Reference [Rec98, Sch14, KOE^{+}06]. Reference-Counting [Rec98]. refinement [GPR11, KPPÉR06, KI16]. Reflection [OT95, Bak95a]. region [KBF^{+}12, WZSK19]. region-based [KBF^{+}12]. Regions [DELD18, AU21, GPS14]. RegionTrack [MWP^{+}21]. Register [GJT^{+}12, LPE^{+}99, RRK11, WW93, CCC12, HKT93, SLP08, kSYHX^{+}11, ZP04]. regulated [XHB06]. Relabeling [HH11]. related [Bar09, RD06, TLZ^{+}16]. relational [HB15]. relative [Bet73]. Relatively [NV15]. relaxed [BAM07, DNB^{+}12, HT14, QSQ14, ZKW15]. relaxed-consistency [HT14, QSQ14]. Relaxing [CZS^{+}17]. RelaxReplay [HT14]. Relay [Zha00]. Release [AB02, PST^{+}92, SLP08, EKB^{+}92, Pea92]. Reliability [CCK^{+}16, CvdBC18, OL02a, OL02b, OL02c]. Reliable [KS16, NBS^{+}15, RG03, YZYL07, YCW^{+}14]. relocation [WW93]. remains [Ano94b]. remedies [ALW^{+}15]. remote [TK98, ZLW^{+}16]. Remove [CNZS17]. Rendering [SSB^{+}22]. reordering [DKG18]. Reorganizing [TKK21]. Replay [Chr01, UALK17, UALK19, GCRD04, HDT^{+}13, HT14, LVN10, LWV^{+}10, LZTZ15, LSW^{+}18, NWT^{+}07, PDP^{+}13, QSQ14, QSHI16, RD99]. Replaying [MCT08, WKG17]. Replica [AT16, FRL18]. Replication [AKP99, BKI06, AU21, VACG09]. Replication-Based [AKP99]. Report [Ano97a, HCM94]. reproduce [HZD13]. request [Sch98]. Requirements [PCPS15, GL98a]. rescue [SLP^{+}09]. Research [BMF^{+}16, USE01, AG06, BMF^{+}19, RPNT08]. réseau [Swi09]. Reservation [LZS^{+}08]. Resilience [SHK15]. Resistance [Gar01]. Resistant [YLLS16]. resize [Mit96]. resolution [Evr01]. Resource [CmL19a, HC17, LG06, LZS^{+}08, LHG^{+}16, RSBN01, YSS^{+}17, CY09, HCD^{+}94, VS11b]. Resource-Efficient [LHG^{+}16]. Resources [LSB15, PFP^{+}22, RGG^{+}12, ZSB^{+}12]. Respec [LWV^{+}10]. Response [BBC^{+}00, Smi01]. responses [BS06]. Responsive [SUF^{+}12]. Restart [ZSA13]. Restating [EE14]. Restore [Ano00b]. restricted [ABG^{+}08]. restructuring [BVG97]. Results [GV95, GRS06]. Retentive [RRK11]. Rethinking [Xue12, Len95]. retrieval [CML00]. Retrospective [TEL98a]. Return [ZWGX22]. Reusable [Han97]. Reuse [BCZY16, KZTK15, LPK16, SV19, JSB^{+}11, NAAL01, PHCR09]. revealing [Dav11]. Reverse [Coo02, LSB15, WCV^{+}98]. Review [Lar97, Van97a, Vre04]. Reviews [Ano97c, Bra97]. Revised [Cha05]. revisionist [PT91]. Reviving [TLZ^{+}17, TLZ^{+}18]. revolutions [ECX^{+}12]. Rewriting [BGK94a, BGK94b]. RHEED [BD06]. right [MAH18]. RISC [Cho92, GV95, MHG95, Men91, Nik94, SBKK99]. rise [Len95]. Robin [CmL19b]. Robot [Lev97]. Robust [CMF^{+}13, LG04]. Rockefeller [IEE90]. Rogue [Ano00b]. Role [BC94, KZTK15]. rollback [YZYL07]. root [BFSK20, CMX10]. Ropes [HMC95]. Round [CmL19b]. routine [SG18]. Row [KZTK15]. RP3 [CJ91]. RPC [Tod95]. RPPM [DVAE18]. RPython [MRG17]. RTOSS [IEE94a, IEE94d]. RTR [XHB06]. Ruby [OCT14]. rules [GLPR12]. Run [EJ93, LFA96, Swe07, SS96, Pra95c, TNB^{+}95]. Run-Time [EJ93, LFA96, SS96, TSY99, TNB^{+}95]. Running [SV19, Cal02, MLCW11, SSN10]. runs [Hig97]. Runtime [ABN99, ABNP00, ABH^{+}00, ABN00, BJK^{+}96, BMN99, CZS^{+}17, DNR00, FSS06, GB20, KPC96, MSFC21, NPT98, NS97, QOIM^{+}12, SSP99, WS06, YSS^{+}19, ZWGX22, ATLM^{+}06, ALW^{+}15, BAD^{+}10a, BAD^{+}10b, BJK^{+}95, EQT07, Gol97, Ong97, TSY00, TMAG03]. runtimes [RL14]. Russians [KNPS16]. Rust [BO17, KN17, KN19, Tro18].

SA [SHW19]. SableSpMT [PV06]. SAC [GS06]. Safe [BCL^{+}98, Kle00, Loc18, Low00, NH09, Pla02, AFF06, BYLN09, DMBM16, DMSM18, Fek08, GCC99, GOT03, Gro03, NHFP08, Nev99, Rin99, Tro18]. Safe-for-Space [BCL^{+}98]. Safety [Hag02, Pla98, Ric99, SP00a, GPS14, Sam99, San04, SRA06, Taf13, Van97b, Ven98, Yan02]. safety-critical [San04]. Salt [Hol12]. sample [LZSS19]. Sampled [JYE^{+}16]. sampling [MMN09]. San [ACM93b, ACM94d, ACM95b, ACM98b, USE89, USE92a, USE93b, USE98b, USE00a, USE02]. Santa [Gol94, WP10]. SAT [VSDK09]. Save [Pla93, Dye98]. saving [Mus09]. SC-preserving [MSM^{+}11]. SC’11 [LCK11]. SC2000 [ACM00]. SC2002 [IEE02]. SC2003 [ACM03]. SC98 [ACM98d, ACM98d]. SC’99 [ACM99b]. Scalability [ABLM19, CCH11, GVT^{+}17, Nak01, VP16, BWDZ15, DSEE13, MWK^{+}06, RVOA08, VIA^{+}05]. Scalability-Aware [GVT^{+}17]. Scalable [BMBW00b, CC14, CH04, CKZ12, IEE94b, KUCT15, LMJ14, LNI^{+}19, MLCW11, Mic04, SSB^{+}22, SS96, ZLW^{+}16, BMBW00a, BMBW00c, GW10, LZ07, Mao96, PWD^{+}12, SCZM00, WZSK19]. Scalably [DELD18]. scalar [GL98b, ZCSM02a, ZCSM02b]. Scale [CC14, CJW^{+}15, HC17, LA93, PWL^{+}11, AG06, BCM^{+}07, GOT03, JCP17, SMK10, SKV21, KBA08]. scale-out [AG06]. Scaling [HC17, AR17, ECX^{+}12, KTLK13, SW16]. Scaling-Aware [HC17]. SCALO [GVT^{+}17]. scene [RVR04]. Schedulability [Kim14]. Schedulability-Aware [Kim14]. Schedule [MQLR16, MLR15, NAAL01, WTH^{+}12]. Scheduler [ABLL92, BDN02, FSPD17, GJT^{+}12, QSaS^{+}16, SRS98, SS95, ASSS19, DC99, DC00, FKS^{+}12, GP05, HZ12, WTKW08, XSaJ08]. Scheduler-Centric [BDN02]. scheduler-oblivious [HZ12]. schedulers [NBMM12]. schedules [BCG13, CZ02]. Scheduling [BB20, BL94, BL98, BL99, CCWY17, FS96, FSPD16, FSPE20, GRS06, JLS99, KLDB09, LLKS12, MNU^{+}15, NB99, PEA^{+}96, PM14, RS08, SM19, SLG04, YDLW20, YWJ03, BL93, CS95a, CS95b, CCC12, DC99, DC00, EE10, EE12, FD95, FKS^{+}12, GA09, HL07, JSMP12, KD22, KKJ^{+}13, KBP^{+}03, Mis96, OA08a, OA08b, OA08c, PAB^{+}14, Pol90, ROA14, SCCP13, SLG06, ST00a, TAS07, WHJ^{+}95, WXAL22, ZSB^{+}12]. Scheme [ABN99, PJS15, SKKC09]. Schur [YFF^{+}12]. Science [Gol94]. Scientific [CMBAN08, HLB94, LWSB19, WN10, BT01, BD06, Dan09, NJ00, Ano97c, Bra97]. scoring [TO10]. Scotland [AOV^{+}99]. SCP [SLJ^{+}19]. Scratchpad [VCM19, TPZ21]. Scratchpad-Memory [VCM19]. Scriptics [Ano00b]. Scripting [RBPM00]. Scripts [TLA^{+}02]. SDC [AU21]. Seamless [CV98]. Search [AMRR98, BCCO10, LAH^{+}12, Mah11]. searches [TCG95]. Seattle [ACM93c, IEE94a, IEE94d, LCK11, USE98a]. sec [AHW02]. Second [IEE89, IEE96, FR95]. Section [DSR15, MNU^{+}15, CS12, DTLM14, SMQP09, YL16]. Section-Aware [MNU^{+}15]. Section-Based [DSR15]. sections [NM10]. Secure [SV98]. Security [BRRS10, MS03, Way95]. sedition [Bak95b]. SEDMS [USE92b]. See [Swe07, AC09]. segmentation [BÇG14]. Select [KKDV03]. selected [Cha05]. Selection [AT16, CmL19b, GPB^{+}17, PR05, Sta90, ZTN19]. Selections [BAZ^{+}19]. Selective [Nak03, PR98, VACG09, AU21, MCRS10]. Self [LLLC15, Pet00, SEP96, BDF98, SLP^{+}09]. Self-Allocating [SEP96]. self-healing [SLP^{+}09]. self-migrating [BDF98]. Sema [Kor89]. Semantic [ARA20, BNS11a, BNS11b, BNS12]. Semantics [BR15, CKRW99, HEJ09, MP01, BLM06, CKRW97a, CKRW97b, KT17, RMV22, ZHCB15]. Semantics-aware [HEJ09]. Semaphore [Hol98b, Kor89]. Semaphores [Hol98c]. semiconductor [Ano97b]. Semidefinite [YFF^{+}12]. Seminar [Nev99]. sense [Bak95b]. Sensible [LMA^{+}16]. Sensitive [CC04, RYSN04, ASSS19, DC99, DC00, PFH06, ZJS^{+}11, LG04]. Separation [SCG95, TFG10, TVD14]. September [ACM93c, AOV^{+}99, DLM99, FR95, Hon94, IEE89, USE98b]. Sequences [GH03, FTAB14]. Sequential [CV98, TLZ^{+}17, TLZ^{+}18, AAJ^{+}19, CKRW97a, CKRW97b, ORS^{+}06, SCG95, SNM^{+}12]. Sequentialization [ITF^{+}22, HLGD19]. sequentialization-based [HLGD19]. Serializable [DELD18, MWP^{+}21]. serialization [BHK^{+}04]. Server [Ano00b, Cal97, Day92a, Day92b, Smi92, VB00, Zha00, CASA14, Est93, Gol96, Hig97, MEG03, SBB96, Sho97b, Sta90]. server-side [SBB96]. Servers [PHBC18, RCC12, BDM98, BBYG^{+}05, BEKK00, KSB^{+}08, RPNT05, SV96c, SV96a, SV96b]. Service [CGK06, GMW09, Hig97, PSM03]. services [LZ07]. session [Bak95b, HCD^{+}94, IAD^{+}94, VGR06]. sessions [Ano94e]. set [Aru92, KBF^{+}12]. Sets [MNG16]. Seven [But14]. several [FGG14]. shader [PYP^{+}10]. shallow [LVA^{+}13]. Shanghai [IEE97]. shape [Cor00, GBCS07]. SharC [AGEB08]. Shared [BWXF05, BS96, DM98, EJ93, FJ08, GMR98, GH98, IXS18, LB92, MVZ93, MCT08, STY99, SLJ^{+}19, Thr99, VB00, WC99, YMR93b, BMTZ21, BB00, Boo93, DLCO09, DPZ97, EKKL90, EV01, Gle91, ISS98, Jef94, MLC04, MKR10, NPC06, RGG^{+}12, TSY99, TSY00, YMR93a, YN09, ZSB^{+}12, dB09, Cha05]. Shared-Memory [BS96, DM98, EJ93, IXS18, MVZ93, MCT08, Thr99, WC99, BMTZ21, EKKL90, TSY00, YN09]. shared-variable [dB09]. Sharing [CLFL94, CB16, LLD17, RKK15, SP00a, Wei98b, ZJS12, AGEB08, AGN09, LTHB14, Sam99, SS95, TAS07, TE94a, Ver96, VPQ12, ZJS10]. sharing-aware [TAS07]. sharing-based [TE94a]. Shelf [MHG95]. shell [Ric91]. Shift [Ham96, Myl22]. Shifting [TVB^{+}13]. Shinko [Ano00a]. Shootdown [PHBC18]. Shop [Bec00]. short [CPT08, Lie94]. shortage [Ano94b]. Should [EHP^{+}07]. SICStus [EC98]. side [MMTW10, SBB96]. sided [QSHI16]. SIGACT [ACM93a, ACM94b, ACM95b, ACM98b]. SIGCOMM [RM03]. Signal [Eng00, BM91]. Signals [GRR06]. Significance [ZJS12]. SIGPLAN [ACM94a, ACM93a, ACM94b, ACM95b, ACM98b, ACM99a]. SIGPLAN-SIGACT [ACM93a, ACM94b, ACM95b, ACM98b]. Silicon [LB17, THA^{+}12]. SIMD [FSYA09, SW08]. Simple [AKS06, Chl15b, WS08, BDLM07, CL00, MSM^{+}10]. SimpleGraphics [MKK99]. simplify [PO03]. Simplifying [Pom98]. SIMT [CC18, LPK16, TCS20]. SIMT-X [TCS20]. simulate [MAF^{+}09]. Simulation [For97, GV95, HPB11, JYE^{+}16, MPD04, SLJ^{+}18, VTSM12, WYT^{+}20, WG94, Ano97b, BBH^{+}17, KD22, KBF^{+}12, Leg01, Lep95, MHW02, SWYC94, Sri93]. Simulations [HEMK17, LNI^{+}19, LS11, SCD^{+}15, ABC^{+}15, KU17, LVA^{+}13, VPQ12]. Simulator [SRS98, PWD^{+}12, TSCH99, WZWS08, Nak03]. Simulators [BVL09]. Simulink [HYY^{+}15]. Simultaneous [Ano05, CSK^{+}99, EEL^{+}97, GSL10, HMNN91, LEL^{+}97a, LEL^{+}97b, LPE^{+}99, LEL^{+}99, LRZ16, MCFT99, REL00b, SP07, SLG04, SHW19, SU01, ST00c, TEL95, Tul96, TEL98b, WS08, YG10, ABC^{+}09, AAKK08, ABB^{+}15, CCC12, EE09a, Fis97, HKN^{+}92, HMN^{+}92, LBE^{+}98, Luk01, Mah13, MMM^{+}05, MEG03, PHCR09, RCG^{+}10, REL00a, REL00c, RM00, RPNT05, SLG06, SW16, ST00a, ST00b, STV02, SMS^{+}03, TSCH99, TEE^{+}96, VPC02, TEL98a]. Single [CLFL94, Dub95, EHP^{+}07, FT96, HHOM91, JBK18, KH18b, KTR^{+}04, MNU^{+}15, MTN^{+}00, TCS20, CSM^{+}05, MLC^{+}09, Pra95c, VIA^{+}05, YZ07, YSY^{+}09]. Single-Address-Space [CLFL94]. single-and [YSY^{+}09]. Single-Chip [HHOM91, MTN^{+}00]. Single-Instruction [TCS20]. Single-ISA [KTR^{+}04, MNU^{+}15]. Single-Process [FT96]. Single-program [Dub95]. Single-Thread [KH18b, MLC^{+}09]. Single-Threaded [EHP^{+}07, JBK18, Pra95c, VIA^{+}05, YZ07]. Singleton [Cha02, Rin99]. SISC [RGT17]. situ [LSW^{+}18, RGK99]. sixth [USE98b, ACM94d]. size [LML00]. skyline [WZSK19]. slave [TJY^{+}11]. slice [PSG06a, PSG06b, PSG06c]. slice-based [PSG06a, PSG06b, PSG06c]. Slices [MGQS^{+}08, PF01]. Slicing [Kri98, FRT95, NR06]. SlicK [PSG06a, PSG06b, PSG06c]. slower [Pra95c]. Small [JLA16, Koo93, MM07]. Smalltalk [Bri89]. Smalltalk-80 [Bri89]. Smart [ROPdlT22, Sim97]. SMP [BWXF05, BNH01, CRE99, HD02, KKH03, KKJ^{+}13, Pra95c, TAS07, TMAG03]. SMPs [WG99]. SMT [Ano05, AH00, CY09, EE09b, EE10, EE12, FSPD16, FSPD17, FSPE20, HR10, KLG08, KH18b, KI16, MG99, MMM^{+}05, NSP^{+}14, PAdS^{+}17, PAB^{+}14, PLT^{+}15, RYSN04, RPNT08, SLP08, SHW19, TAS07, TVB^{+}13, VS11b, WA08]. SMT-based [KI16, PAdS^{+}17, PAB^{+}14]. SMT-Directory [HR10]. SMT-SA [SHW19]. SMTp [CH04]. Snippet [Nag21]. SoC [ZDTM19]. SOFRITAS [DELD18]. Soft [EUVG06, OA19, PSM01, PSM03, SSN10, VACG09]. Software [Ano97a, Ano98b, Ano99, Ano00b, BVM19, BCR01, BCG^{+}08, Gar01, Gon90, GJ97, HB92, Han97, HSS^{+}14, IEE94a, KE15, LPE^{+}99, MKM17, PJS15, SZM^{+}13, SD13, TVB^{+}13, TLZ^{+}17, TLZ^{+}18, Tro18, XWG^{+}14, YBL16, ATLM^{+}06, AC09, Ano97c, ABC^{+}09, BT01, Bra97, CDD^{+}10, DPZ97, GLPR12, Hai97a, HSD^{+}12, IEE94d, KKH04, KSD04, KASD07, LT97, Luk01, MWP07, MCRS10, MGL95, MEG03, NHFP08, OAA09, OL02a, OL02b, OL02c, PV06, RKM^{+}10a, RKM^{+}10b, RVOA08, San04, SP05, SLP^{+}09, SB80, TNB^{+}95, WCZ^{+}07, WCV^{+}98, YSY^{+}09, ZHCB15, DWS^{+}12]. Software-Controlled [BCG^{+}08, Luk01]. Software-Directed [LPE^{+}99]. Solaris [Cat94, Lun97, Lun99, MM01, McM97, Pra95b, Sun95]. Solution [Ano98b, SSK^{+}18, SBC91, WP10]. Solutions [Ano00b]. solve [Bar09, MM07]. Solver [YFF^{+}12, ÇM20, Kub15, RM99]. Solvers [MR09, Nak03, SSK^{+}18, AAC^{+}15, ZCO10]. Solving [ABD^{+}12, FTAB14, Loe97, VSDK09]. SONET [AHW02]. Sort [GH98, RHH10]. Sorting [Ger18]. Sound [MWP^{+}21, SMP19, WTH^{+}12, DWS^{+}12, FFY08, NFBB17, WQLJ18]. Source [Ano00c, BMF^{+}16, BMF^{+}19]. sources [SJ95]. South [ACM93a, Ano94d]. Space [BCL^{+}98, BL93, BL98, CLFL94, CB16, Eng00, GRS97, GN96, NB99, PWL^{+}11, Sch17, FWL03, KNPS16, KASD07, Lie94, LHS16]. Space-Efficient [BL98, NB99, BL93, KNPS16, KASD07, LHS16]. Spacecraft [SRS98]. Spaces [FKP15, Rót19, CKZ12, KGGK09]. Spain [ACM95a, DLM99, ACM98c]. SPARC [Cat94, KAO05, MD96]. Sparcle [ABC^{+}93]. Sparse [But13, YFF^{+}12, CSV10, ÇM20, Dav11, DTR18, KS21, MM07, PHCR09]. spatial [WZSK19]. spatially [PPA^{+}13]. spatially-programmed [PPA^{+}13]. Special [Ano94e, GGB93b, KU00]. specialization [WTH^{+}12]. specialize [CWS06]. Specialized [dlPRGB99]. Specific [Ste01, SP00b, Shi00]. specification [LPD^{+}11, Stä05]. specifications [TVD10]. Specifying [BNS11a, BNS11b, BNS12, LPD^{+}11]. spectroscopy [KC09]. spectrum [DKF94, Sha95b]. Speculated [SCL05]. Speculation [MGI14, SU01, WS08, YBL16, DG99, GB99, JEV04, LWV^{+}10, MT02a, MT02b, MT02c, NB12, PO03, PT03, RGT17, SCZM00]. Speculative [AH00, Ano01, Ano02, BF04, IBST01, KLG08, MGQS^{+}08, MG99, MT02a, MT02b, MT02c, RKM^{+}10a, RKM^{+}10b, SR01b, TFG10, WWW^{+}02, ZJFA09, ZL10, CHH^{+}03, DC07, Dub95, KOE^{+}06, KT99, LZW17, LZSS19, LZL^{+}14, NB12, OL02a, OL02b, OL02c, PV06, SMS^{+}03, VS11b, XIC12, ZCSM02a, ZCSM02b]. speech [LG04]. Speed [Ano00a, Ano03, GV95, HG91, MR09, HG92, Pra95b, SRS98, TO10]. Speed-up [MR09]. Speedup [Lun99]. Spin [LLS06]. SPIRAL [MJF^{+}10]. SPIRAL-generated [MJF^{+}10]. splittable [SLF14, SV21]. spots [Gle91]. spreading [CWS06]. SPSM [Dub95]. SQL [CGK06]. squares [FTAB14]. squash [MK12]. SR [BO96]. SRAM [kSYHX^{+}11]. SSMT [CSK^{+}99]. Stabilizers [ZSJ06]. Stabilizing [BCM^{+}07]. stable [YCW^{+}14]. Stacey [Ano00c]. Stack [Eng00, Xue12]. Stackable [Loe05]. stacking [KSB^{+}08]. Stackless [MS15]. Stacks [ZWGX22, DESE13]. StackThreads [TTY99]. StackThreads/MP [TTY99]. Standard [DM98, FSS06, WKG17, Ano97c, BCL^{+}98, Bra97, MT93, Pla98, Pla99]. standardization [Bet73]. Standards [Thr99, TTY99]. Standing [TLA^{+}02]. Stanford [IEE99]. STAT [Ano00b]. State [Laf00, LP94, MP13, RRK11, Wei98b, Cor00, I^{+}94, TFG10, WHG07]. State-of-the-Art [MP13]. State-Retentive [RRK11]. Statechart [KW17]. Statechart-Based [KW17]. stateless [AAJ^{+}19, MQ08]. Static [GPS14, Kri98, Lun97, SCB15, SBE^{+}19, WW96, vPG03, Fer13, NAW06, NA07, AFF06, FFLQ08]. Static/dynamic [SCB15]. Statistical [Ano00b, RCM^{+}16, Lan97, RCM^{+}12, Tem97].

Stay [GBK^{+}09]. Stealing [SKV21, ALHH08, BL94, BL99, RL14, WYT^{+}20]. Step [Sho97a, Sho97b, ZG98]. Steroids [JLA16]. Stethoscope [Caz02]. Stochastic [DK02, LTM^{+}17]. Storage [AT16, Hol12, LCK11, ZWGX22, Bak95a, Blu92, DZKS12, KOE^{+}06, MM07, PDMM16, PPGS20]. stores [RMV22, TAN04]. strand [RČV^{+}10]. strata [NPC06]. Strategies [PSCS01, WLT19, AGEB08, FGG14]. Strategy [BGK96]. Stream [KSU94, SG18, SG18]. Streaming [HHOM91, HHOM92, KEL^{+}03]. Streaming/FIFO [HHOM91, HHOM92]. Streams [Pre90, SPY^{+}93]. Strength [Kon00]. Strict [Coo95, FS96, Tra91, KIAT99, SCG95]. Strictly [Ano00c]. Strong [CWHB03, KZC15, AG21, MTC^{+}07, ZHCB15]. Structural [CKRW99]. structure [BB00, YKL13]. Structured [TCI98, FR95]. Structures [RCRH95, AGN09, Gol97, ND13]. students [Fek08]. Study [AGK96, Chl15a, EGC02, Ger18, HMT^{+}96, LSB15, Sat02, TAK^{+}00, VK99, WG94, YMR93b, BFSK20, Bri89, CASA14, CL00, Fis97, HJT^{+}93, HF96, KPPÉR06, LPD^{+}11, MGL95, SP05, Sod02, Tsa97a, YM92, YMR93a]. Style [Wil94a, Wil94b]. subdivision [MTS10]. subordinate [CSK^{+}99, CTYP02]. Subsetting [AJK^{+}12]. Substrate [ACMA97, Hai97a, JP92]. Subsumption [Man91]. Subtleties [BLM06]. Suffix [OR12, LHS16]. SugarCubes [BS00]. Suite [BTE98, BO01, TG09]. Suites [SPDLK^{+}17]. SuiteSparseQR [Dav11]. sum [TDW03]. summary [I^{+}94]. Summer [Ano94f, USE92a]. Sun [McM97]. SunOS [Cat94, PKB^{+}91]. super [Kus15]. Supercomputer [VTSM12, Gil94]. Supercomputing [ACM92, ACM95a, ACM96, Ano91, Ano94e, IEE90, IEE92, IEE93, IEE94c]. SuperLU [Li05]. SuperMalloc [Kus15]. Superscalar [SU96, Div95, Fis97, Gul95, Loi95, Men91]. Superthreading [Tsa97b]. Support [ACM94d, ABLL92, BBG^{+}10, CZS^{+}17, CSS^{+}91b, EJ93, GHG^{+}98, KC99, Kno14, MKM17, ME15, MS89, NS97, PTMB09, SSP99, TY97, ZSA13, ATLM^{+}06, BS06, BO96, CMF^{+}13, CKD94, CHH^{+}03, CSS^{+}91a, CSS^{+}91c, Evr01, Fan93, HMC95, MWP07, MEG03, MS87, Men91, TSY99, TSY00, TNB^{+}95, WK08a, WK08b, WK08c]. Supported [AddS03, ZP11]. Supporting [RCRH95, Sam99, SP00a, DC99, DC00, TDW03]. suppression [JWTG11]. Surface [Rót19]. surgery [MCS15]. Surprises [BC98]. Survey [CA20, Man96, OA19, ZSB^{+}12, Cat94, URŠ02b, URŠ03]. Survival [Ano99]. Surviving [Ano99]. Svelto [MCS^{+}22]. SVGs [SSB^{+}22]. SVR4 [SPY^{+}93]. SW [ZDTM19]. swap [MLS15]. Swing [Gea98]. Switch [GN00, Eic97, GWM07, TVB^{+}13]. Switching [MSFC21]. Switzerland [Lak96]. Sy [USE01]. Symantec [Rod95a]. symbiosis [Bri89, EE10, EE12]. Symbiotic [FSPE20, ST00a, ST00c, ST00b, STV02]. Symbolic [ACM94c, BGC14, Hon94, Lak96, Wat91, BHKR95, Fuj97, HF88, HLGD19]. Symmetric [BMV03, NV94, BIK^{+}11, Pra95b, RGK99, Sha98]. Symmetry [ES97]. Symposium [ACM93a, ACM93b, ACM94b, ACM94c, ACM95b, ACM98b, ACM98c, Ano91, Ano94a, Ano00a, Ano03, Gol94, Hon94, Lak96, USE91a, USE92b, USE93a, USE98a, Wat91]. symptoms [BFSK20]. Synapsys [Col90a]. Synchronization [Bec01, DSR17, Hei03, LA93, Rec98, SLJ^{+}18, DHM^{+}12, DESE13, MT02a, MT02b, MT02c, MTPT12, NLK09, PRS14, RD06, Ven97]. Synchronization-Aware [SLJ^{+}18]. synchronization-induced [MTPT12]. synchronization-related [RD06]. Synchronizing [McM96a, McM96b, CZWC13].

Synchronous [BM07, HPB11]. SynchroTrace [SLJ^{+}18]. Synergy [ZDTM19]. syntax [KT17]. Synthesis [CFC^{+}22, FN17, FF20, HB15, LNI^{+}19, MCS^{+}22, Sch17, MP89, SR14, STR16, WQLJ18]. Synthesizing [GLPR12, Kim14, SRJ15]. synthetic [GJ11]. System [AddS03, ÁdBdRS08, AJK^{+}12, Ano98a, Ano00b, ABN99, ABH^{+}00, BMR94, BBD^{+}91, BJK^{+}96, BTE98, CLFL94, CC18, DNR00, FG91, Gei01, HMT^{+}96, KMAG01, KS97, MS89, NPT98, PH97, PST^{+}92, Pea92, PFP^{+}22, PLT^{+}15, QOIM^{+}12, REL00b, SEP96, Sri93, SG96, TCI98, VSM^{+}08, Yam96, YSS^{+}19, ÁdBdRS05, AAC^{+}15, Ano96, Ano97b, A^{+}01, AR17, ASSS19, BBFW03, BDM98, BCHS00, BAD^{+}10a, BAD^{+}10b, BJK^{+}95, BAD^{+}09, BLCD97, Cat94, ÇM20, Gil88, Hig97, Joe96, JCP17, Lan02, MHW02, MS87, Met95, MTC^{+}07, MC06, OCRS07, PRB07, Ply89, Pom98, REL00a, REL00c, RD99, She02, TKA^{+}02, TLZ^{+}16, TMAG03, WCC^{+}07, WZWS08, WZSK19, TLA^{+}02, EKB^{+}92, MS87, Pea92]. System- [PLT^{+}15]. system-level [OCRS07]. systematic [MQ07]. SystemC [RSB^{+}09]. SystemC/C [RSB^{+}09]. SystemC/C-based [RSB^{+}09]. Systems [ACM94d, AG06, Ano00b, ABN00, BMN99, Bre02, BC94, CCH11, CvdBC18, Dru95, FMY^{+}15, FGKT97, GHG^{+}98, GJ97, HRH08, HKSL96, IEE89, IEE94a, KR12, KKH03, KG05, KUCT15, KW17, LLS06, LMA^{+}16, LYH16, MS15, MSFC21, PPG11, PGB16, RW97, RR03, SUF^{+}12, SS96, USE92b, Wal95, WC99, Zub02, Ano92a, Ano92b, BCM^{+}07, BC02, Cat94, DCK07, DWYB10, DZKS12, DSH^{+}10, DBRD91, GJ11, Gol96, GKK09, HJT^{+}93, Hop98, HWW93, HBCG13, IEE94d, ISS98, JD08, Jef94, Jen95, KKH04, Kub15, LVN10, LLLC15, Leg01, LAK09, LVA^{+}13, MLC^{+}09, MGL95, MM07, NFBB17, PBDO92, PMCP22, RČV^{+}10, RBF^{+}89, RSB^{+}09, RVR04, SCCP13, She98, SP05, Sim97, SJB92a, SJB92b, ST05, Wei98a, WCV^{+}98, Ano98b]. Systolic [SHW19, PYP^{+}10].

T [Ano00c, CSM^{+}21, NPA92]. T/TCP [Ano00c]. T1 [Wea08]. T1/T2 [Wea08]. T2 [Wea08]. Table [VB00, KNPS16]. tabling [AR17, AR19]. Tabu [AMRR98]. Tail [DGSB20, ASSS19]. taint [ZJS^{+}11]. TaintEraser [ZJS^{+}11]. Take [Wei97]. taking [Ano92b]. Talking [Ano94c, HCM94]. TAM [CGSV93]. Taming [Hol00, HBCG13, HHPV15]. TapeWare [Ano00b]. Target [MIGA18]. targeting [LGH94]. Task [CCK^{+}16, GP95, GFJT19, Kwo03, Mar03, Mis96, Myl22, PM14, ABG^{+}08, CASA14, DCK07, KD22, OdSSP12, RCM^{+}12]. Task-Based [GFJT19, Myl22]. Task-Level [GP95]. Tasking [CvdBC18, Dil93, KR01a]. Tasks [Fin95, PVS^{+}17, TKK21, YSS^{+}17, YSS^{+}19, FGG14]. Taxonomy [HM96, SPH96]. TC2 [BT01]. TC2/WG2.5 [BT01]. Tcl [Ass96, USE96, USE98b, USE00b, Ama98, MKK99, SBB96]. Tcl-based [Ama98]. Tcl/2k [USE00b]. Tcl/Tk [Ass96, USE96, USE98b, USE00b, MKK99]. TCP [Ano00c, Ano00c]. Teaching [Fek08, CS00, She02]. TeamWork [CZWC13]. Tech [Ano97b, Gar01]. Technical [USE00a, Cat94]. Technique [JSB^{+}12, KG94, Lem02, ÖCS01, PGB16, JSB^{+}11, JPSN09, LGH94, MIGA18, RS07, UZU00, VACG09, WCV^{+}98]. Techniques [ARA20, CA20, DS16, EKKL90, GS02, Han97, NLK09, PWL^{+}11, TGBS05, Zig96, BR92, GEG07, OCRS07, Pra97, RCG^{+}10, SV96c, SV96a, SV96b, ZSB^{+}12]. Technologies [Ano00b, Ano98b]. Technology [Ano97c, Bra97, KM03, LB00, USE01, VSM^{+}08, KSB^{+}08, Tsa97b]. TeleNotes [WSKS97]. temperature [CCC12]. Template [Cal00, How98].

Temporal [CmL21, RMV22]. Ten [Ano99]. Tennessee [IEE94b]. Tera [BTE98, Mat97]. Terabytes [IEE02]. Term [BGK94a, BGK94b, BGK96, CmL21]. Termination [JBK18, TDW03]. Test [Ama98, EFN^{+}01, GRS97, SPDLK^{+}17, TG09, EFN^{+}02, KI16, SR14]. test-case [KI16]. Testing [BBdH^{+}11, Goe01, KH18a, LCS04, RCC14, SK12, BGP06, CBM10, EFG^{+}03, EHSU07, MQ07, Sen08, YNPP12]. tests [SRJ15]. Texas [USE92a, USE00b]. TFlux [DTLW16]. tgMC [LHG^{+}16]. Their [YWJ03, Gil94]. them [Ano92a, Ano94b]. Theoretic [ES97]. theories [LQ15]. Theory [ACM93b, LLD17, NFBB17, WLK^{+}09]. there [Ano94b]. thermal [WA08]. though [Ano94b]. Thread [Ano00c, ABN99, ABNP00, Bet73, BTL^{+}19, BS99, CNQ13, Cal97, CmL19b, CC04, Cha02, CCWY17, Col90a, DSR15, DELD18, DGK^{+}03, Don02, DMSM18, DSR17, Eng00, FSPE20, FD95, FURM00a, FURM00c, FURM00b, GF00, GJT^{+}12, GP05, GBCS07, GBK^{+}09, Hag02, Hei03, HG91, ISS98, KG05, Kim14, Kle00, KH18b, KBH^{+}03, KBH^{+}04a, KBH^{+}04b, LPM17, LML^{+}19, LZL^{+}20, LLL10, LYH16, LEL^{+}97a, LEL^{+}97b, Low00, LLD17, Man99, MG99, MNU^{+}15, MGI14, MTN^{+}00, MB05, MCFT99, ND96, Pan99, PR05, PEA^{+}96, Pla02, Pla98, Pra95b, PGB12, PSCS01, RČV^{+}10, RCM^{+}16, RCG^{+}10, Rec98, Ric99, Rin99, RYSN04, Rod95b, SKS^{+}92, Sat02, STY99, SLG04, Sin97, SKK^{+}01, SLT03, Ste01, TAS07, TLGM17, Wei98b, WG99, Wei97, Whi03, YBL16, ZP11, ZWGX22, AMRR98, ABG^{+}08, BKC^{+}13, BHK^{+}04, BC02, CZSB16, CZ02, CSM^{+}05]. thread [DMBM16, DG99, DWYB10, Don92, DBRD91, Eic97, EE09b, Fek08, GP08, GOT03, GLC99, Hyd00, JEV04, KDM^{+}98, KC09, KBA08, KSD04, KASD07, LK13, LZW17, Lie94, LML00, LZL^{+}14, Loe05, MLC^{+}09, MT02a, MT02b, MT02c, MC06, OT95, PAB^{+}14, PRS14, PKB^{+}91, PO03, PT03, PGB14, QOQOV^{+}09, RGT17, SKG^{+}11, Sha95b, SLG06, SP00b, Shi00, SPH96, SS95, SD13, SLT02, Stä05, SJ95, SCZM00, ST05, SS10, Tan87, TE94a, TLZ^{+}16, TCG95, Tra91, Van97b, Ven97, Ven98, WS08, YZ14, SKP^{+}02]. Thread-Aware [LYH16]. Thread-Based [KG05, CNQ13, SKS^{+}92]. Thread-Level [LEL^{+}97a, LEL^{+}97b, MG99, MGI14, YBL16, FURM00a, FURM00b, MCFT99, WS08, DG99, JEV04, KC09, MT02a, MT02b, MT02c, PO03, PT03, QOQOV^{+}09, RGT17, SCZM00, YZ14]. Thread-Local [DGK^{+}03, Whi03, ZWGX22]. Thread-management [RCG^{+}10]. Thread-modular [GBCS07]. Thread-Private [Man99]. thread-related [TLZ^{+}16]. Thread-Safe [Kle00, Pla02, DMSM18, Rin99, DMBM16, Fek08, GOT03]. Thread-Sensitive [CC04, RYSN04]. Thread-Specific [Ste01, SP00b, Shi00]. thread-switch [Eic97]. threadbare [Bak95b]. Threaded [AGK96, BBG^{+}10, BC98, Bed91, BGK94a, BGK94b, BGK96, CL95, CKRW99, CFC^{+}22, Coo95, CSS^{+}91b, DV99, EHG95, EHP^{+}07, FdL02, GH03, GVT^{+}17, GK94, Gil93, III01, JBK18, JY15, Jon91, KW17, Kri98, Kuc92, KIAT99, LK20, LB92, Mas99, MG15, MCS^{+}22, MGK^{+}00, MSU^{+}16, NS97, PCPS15, Pul00, RKCW98, SV19, STW93, Sei99, Smi92, Ste01, SBKK99, TLGM17, VSDK09, VS11a, VB00, VCM19, WCT98, YDLW20, Ada98, AG21, ABD^{+}12, AACK92, Ano97b, ASSS19, BWDZ15, BK13, BBH^{+}17, BC00, BIK^{+}11, DSEE13, CV98, CIM^{+}17, CASA14, CKRW97a, CKRW97b, CWHB03, CSB00, CdOS01, CYZ98, cC91, CL00, Chr01, CR02, CSS^{+}91a, CSS^{+}91c, DS16, DGSB20, EFG^{+}03, EBKG01, EHSU07, FTAB14, FD96, FF20, FGG14, GCRD04, GCC15, GS06, GH98, GPR11, HC17]. threaded [HLGD19, ITF^{+}22, JCP17, KHP^{+}95, KI95, KKH04, Kep03, KRH98, Küc91, LK15, Lan97, Leg01, LBvH06a, LBvH06b, LBvH06c, LVA^{+}13, MLCW11, MS03, MKK99, NFBB17, NH09, NSH14, OA08a, OA08b, OA08c, PYP^{+}10, PMCP22, PR98, PWWD18, Pra95c, RČV^{+}10, RKM^{+}10a, RKM^{+}10b, RBPM00, RGK99, RS08, SCB15, Sam99, SP00a, SC17, SE12, Sei98, Sho97a, Sho97b, SV98, Smi06, Sto02, SQP08a, SQP08b, SQP08c, Taf13, TSY99, TSY00, Tem97, TMAG03, TJY^{+}11, VIA^{+}05, VV00, VK99, Wal00, Wil98, XMN99, YZ07, Yee20, YSY^{+}09, ZKR^{+}11, dB09, vPG03, CGSV93]. Threading [BFA^{+}15, CGS^{+}20, CvdBC18, CNZS17, DHR^{+}01, FNS^{+}22, Hol98d, KS16, LKBK11, MLGW18, McC97a, McC97b, MS15, MP13, Nag21, Nor90, OR12, PTMB09, PFP^{+}22, RCC14, Rei01, Sch90, SMZ18, TGO99, TCS20, YLLS16, Bak95a, BM07, DTLW16, FWL03, KD22, Kno14, LZW^{+}13, MLC^{+}09, MCFT99, NJ00, RRP06, RVR04, SQP08a, SQP08b, SQP08c, VDBN98, kSYHX^{+}11, YKL13, CH04]. Threading-Based [KS16]. ThreadMentor [CMS03, She02]. Threads [Alf94, Ano94c, ACR01, Ber96b, BCL^{+}98, Boe05, BLPV04, BAZ^{+}19, Cal00, CGR92, Col90b, Cri98b, Cri98a, TLA^{+}02, FHM95a, For95a, For95b, GMB93, GSC96, GN96, Gus05, Hai97b, HW92, HBG01, Hol00, How00, HLH16, JLS99, KSS95, LP94, Lee93, Lee06, LB96a, LFA96, Man98, MKM17, MP89, McM96c, Nor96, PSM01, Pet00, Pet03, Pla93, Pra95c, San04, SEP96, TG99, WCW^{+}04a, Wil94a, Wil94b, Wil97, Yam95, Yam96, dlPRGB99, Ano02, Bak95b, BZ07, Ber96a, BW97, BDF98, Bir89, BS00, But14, But97, CZWC13, Cal02, CPT08, Dra96, DESE13, DC99, DC00, FHM95b, FL90, GP05, Gol97, HCM94, HMC95, Hai97a, HBG02, HJT^{+}93, HKT93, HKN^{+}92, Hol98d, Hol98a, Hol98b, Hol98c, Hol99a, Hol99b, Kan94, KE95, KSS96, Lan02, LZ07, MSLM91, MR98, MQW95, McM96a, McM96b, McM98a]. threads [McM98b, Men91, Mit96, MEG94, OW97, OW99, OL02a, OL02b, OL02c, ORS^{+}06, PSM03, Pan99, PG03, PL03, RR03, Sch91, SCG95, SZG91, SZ92, SCM05, SKP^{+}02, TAN04, WCW^{+}04b, WCW^{+}04c, WCW^{+}04d, Wei98a, WCV^{+}98, WW96, ZCSM02a, ZCSM02b, ZP04, ALW^{+}15, Van97a]. Threads.h [Ano00b, TB97a, TB97b]. ThreadScope [WT10]. Three [YMR93b, YMR93a]. Throttling [LG06]. Throttling-Based [LG06]. Throughput [GJT^{+}12, Wea08, ZDTM19]. Tightly [MTN^{+}00, LZTZ15]. TileDB [PDMM16]. Tiles [QOIM^{+}12]. Time [BC94, CIM^{+}17, EJ93, GN96, IEE94a, JLS99, Kim14, LFA96, Lun97, MN00, PUF^{+}04, PSCS01, SUF^{+}12, SS96, Tet94, dlPRGB99, CS95a, CS95b, DC99, DC00, GB99, IEE94d, Jef94, Jen95, KBP^{+}03, KASD07, KBF^{+}12, MKK99, ND96, OT95, OdSSP12, PSM01, PSM03, RGG^{+}12, San04, SZG91, SZ92, SM19, SJB92a, SJB92b, TSY99, TNB^{+}95]. time- [KASD07]. time-critical [RGG^{+}12]. time-efficient [GB99]. time-shared [Jef94]. timely [NH09]. Timers [Hol99a, GRR06]. times [LTMK21]. Timethread [BC94]. Timethread-Role [BC94]. Timing [SK97, MHW02]. timing-first [MHW02]. tiny [Xue12]. Tip [Pet00]. Tips [Mit96, Pet00]. Tk [Ass96, USE96, USE98b, USE00b, TLA^{+}02, MKK99]. TLB [PHBC18]. together [Ano97b, Pol90]. Tokyo [Ano00a]. Tolerance [EUVG06, OA19, MTS10, PG01, RRP06]. Tolerant [ÖCS01]. Tolerating [Luk01, RBK^{+}09, SKK^{+}01]. Tool [AddS03, Ano98b, Goe01, Kor89, LZL^{+}20, TAM^{+}08, ACD^{+}18, CMS03, CSB00, Hig97, LMC14, RGK99, YNPP12]. Tool-Supported [AddS03]. Toolbox [Ano97c, Bra97]. Toolkit [MSU^{+}16, SZM^{+}13]. Tools [Ano98b, Cha05, EV01, WWW^{+}02, EHSU07, Len95]. Tools.h [Ano00b]. Toolset [Ano97a]. Top [Ano99, AB02, DNR00, ZTN19]. top- [ZTN19]. Topaz [MS87]. topics [BGG95, GBG95]. Toroidal [KEL^{+}03]. Totally [DHR^{+}01]. Trace [MWP^{+}21, RS08, HEJ09]. Trace-Based [MWP^{+}21, RS08]. Traces [HEMK17, MWP^{+}21, SLJ^{+}18, WKG17, HR16]. Tracing [Lem02, EKKL90, Tod95]. Tracking [CZS^{+}17, LH09, CZSB16, GB20, ZJS^{+}11]. trade [AAC^{+}15, Par91, KUCT15]. trade-off [AAC^{+}15]. trade-offs [Par91]. Tradeoff [SHK15]. tradeoffs [Aga89, Aga91, Aga92, Ann96, PJZA07]. Traffic [HLH16, TCF22]. training [MCS15]. Tranquilizer [PGB12]. Transaction [LZS^{+}08, RW97, SS91, DKG18, EQT07, Ver96]. Transaction-Aware [LZS^{+}08, EQT07]. Transactional [GMGZP14, KUCT15, MWP^{+}21, RG03, VSDL16, XWG^{+}14, ZLJ16, ATLM^{+}06, BLM06, BDLM07, CMF^{+}13, CNV^{+}06, GCC15, MLS15, MCRS10, MMTW10, MTC^{+}07, OCT14, VTSL12, ZHCB15]. Transactions [Ano00c, DTLW16, FNA^{+}18, SKBY07, BD06, Dan09, KR01a, KR01b, KGGK09, RKM^{+}10a, RKM^{+}10b]. transfer [TPZ21]. Transform [BVP^{+}19, HN91, LHS16, TKHG04, TT03, TTKG02]. transformation [TSY00]. transformations [AC09, D’H92, JMS^{+}10, VV11]. Transient [RM00, VPC02]. Transient-fault [VPC02]. Transitive [YMR93b, XHB06, YM92, YMR93a]. Translation [SSB^{+}22, KBF^{+}12]. translator [TJY^{+}11]. Transparency [GKCE17, KBH^{+}03]. Transparent [ABN99, LVN10, MSFC21, SLGZ99, ZSA13]. Transparently [CB16, JSB^{+}12]. Transport [GRS97]. transposed [KS21]. transposition [SGLGL^{+}14]. trap [Ram94, GRS97]. trap-based [Ram94]. Tree [Pla99, BCCO10]. Trees [GFJT19, AD08, CKZ12]. Trends [Gar01]. TRI [ACM93c]. TRI-Ada [ACM93c]. triangular [BKK17, ÇM20]. Trick [Eng00]. Tridia [Ano00b]. tridiagonal [ZCO10]. trigger [Kho97]. Triggered [PPA^{+}13]. Troy [SS96]. TSGL [ACD^{+}18]. TSO [HH16]. Tumbler [PGB16]. Tune [RGK99]. tuned [Ano95a, Ano95b, Kub15]. Tuning [LWSB19, LEL^{+}99, CSB00, RGK99]. Tunnelling [Don02]. Tutorial [Taf13]. Twentieth [ACM93a]. Twenty [AOV^{+}99, ACM93b]. Twenty-fifth [AOV^{+}99, ACM93b]. Two [BBH^{+}17, CM98, JYE^{+}16, STY99, GLC99]. Two-Level [JYE^{+}16, BBH^{+}17, STY99]. TX [Cha05, ACM00, USE91b]. TxRace [ZLJ16]. Type [Gro03, Loc18, VGR06, BAD^{+}09, GE08, Lan02, Mil95, PRB07]. type-checking [Mil95]. Type-Safe [Loc18, Gro03]. typed [DMBM16]. Types [AFF06, FFLQ08, Ten98, BAM07, KS93, RMV22, VGR06]. typings [Smi06].

UCITA [Gar01]. UK [AOV^{+}99]. ULT [PG03]. Ultra [PWL^{+}11]. Ultra-Scale [PWL^{+}11]. UML [SK12]. Unbounded [CNV^{+}06, FKP15, BDLM07]. uncommon [BDLM07]. Unconventional [DSAD^{+}18]. Uncover [WS08]. underdetermined [Kub15]. Undergraduate [BLPV04]. Understandable [MSM^{+}16]. Understanding [BZ07, TLA^{+}02, EPAG16, JGS^{+}19, RRP06]. Undocumented [SW97]. Unfoldings [KH18a, SPDLK^{+}17]. Unicode [Swi09]. Unified [Wei98b, ABG^{+}08, GKZ12]. Uniform [BDN02, SKG^{+}11]. unifying [MS03]. unimodular [D’H92]. unintrusive [HDT^{+}13]. uniprocessor [GL98a, Yan97]. uniprocessors [BRE92, EJK^{+}96]. Uniscape [Ano98b]. UNISIM [LS11]. UNISIM-Based [LS11]. unit [CBM10, Par91, PAB^{+}14]. United [ACM94c]. Unithreaded [RLJ^{+}09]. Units [RKK15, Gun97]. univariate [CMX10]. University [IEE99]. UNIX [Ano00b, FG91, JJ91, Kor89, MS87, MS89, Nor96, RR96, RR03, Yoo96a, Ano98b, Ric91]. Unix-to-NT [Ano98b]. UnixWare [Rod94, Rod95b]. unlocking [XSaJ08]. unravel [But14]. Unraveling [Bec00]. Unsynchronized [DSR15]. unveiled [Ano95a, Ano95b]. Unveiling [AAC^{+}15]. up-and-downdating [VV11]. UPC [EGC02, FA19]. updates [NH09]. Updating [HSS^{+}14, HSD^{+}12, NHFP08]. Ur [Chl15b]. Ur/Web [Chl15b]. URL [TLA^{+}02]. USA [ACM94a, ACM94d, Cha05, Hol12, ACM96, ACM98d, ACM00, Ano90, EV01, IEE89, IEE94a, IEE96, IEE02, SS96, USE89, USE91a, USE91b, USE92a, USE93a, USE93b, USE00b, USE00a, USE01]. Usage [BS96, Kor89, VS11b]. usages [BFSK20]. Use [Bak95a, FJ08, HW92, WWW^{+}02]. Use-once [Bak95a]. Useful [Pet03]. USENIX [Ano90, Ano94f]. User [ABLL92, DLM99, Eng00, GRS97, MQW95, SLT03, BF08, GP05, GRR06, HF96, Li05, MSLM91, OT95, SLT02, TNB^{+}95, YZYL07]. User-Level [ABLL92, SLT03, MQW95, GRR06, MSLM91, OT95, SLT02, YZYL07]. User-Space [Eng00, GRS97]. Using [Ano99, ABH^{+}00, AZG17, BDN02, BTL^{+}19, BBC^{+}00, BLG01, BTE98, BAZ^{+}19, CRE99, Cor00, DS16, DTLW16, DBRD91, GH03, HBG01, HJT^{+}93, HBTG98, Hei03, How00, KMjC02, KH18b, Kwo03, KET06b, LFA96, MPD04, McM98a, McM98b, Mix94, MM07, PF01, PBR^{+}15, PO03, SW08, SCD^{+}15, SEP96, SLT02, WKG17, WJ12, Whi03, ZLJ16, Ano96, Bar09, BCM^{+}07, CML00, Cat94, CTYP02, CDD^{+}10, CVJL08, CKZ12, DESE13, GCC15, GMB93, GEG07, Hig97, HH97, JWTG11, JJY^{+}03, JCP17, KASD07, KBF^{+}12, LK15, MM14, NPC06, NWT^{+}07, Nik94, PT03, RKM^{+}10a, RKM^{+}10b, RM99, RPNT05, SLGZ99, SLP^{+}09, TP18, TFG10, Tod95, TAN04, VPC02, VD08, ZJS^{+}11, KSB^{+}08]. UT [Hol12]. Utility [FHM95a, JSMP13, WLT19, FHM95b]. Utility-based [JSMP13]. Utilization [PFP^{+}22, Squ94]. Utilizing [ES97, WZSK19]. UX [Ano95a, Ano95b, Yam96].

V [EKB^{+}92, Pea92, FG91, PST^{+}92]. v1.0 [Ano00b]. Validating [LB17]. Validation [BMV03, DSR17, LB17, SCB15]. Valley [GBK^{+}09]. value [DG99, TFG10, ZCSM02a, ZCSM02b]. Values [EUVG06]. Variable [CA20, Evr01, dB09]. Variables [Hol98c, Whi03, Bak95a]. variants [RB22]. variation [PGB12]. variety [CML00]. VAX [Gil88]. VDOM [SSB^{+}22]. Vector [CSM^{+}21, Goo97, HHOM91, HHOM92, KBH^{+}04a, KBH^{+}04b, KKS^{+}08, LRZ16, SSB^{+}22, VD08, CS95a, CS95b, CSV10, KS21, KBA08]. Vector-Processor [HHOM91, HHOM92]. Vector-Thread [KBH^{+}04a, KBH^{+}04b, KBA08]. vectorization [cC91, JMS^{+}10, RKHT17]. vectorized [TP18]. vectors [KTK12]. Velodrome [FFY08]. Verification [ÁMdBdRS02, BCR01, Chl15a, DRV02, EGP14, FK12, ITF^{+}22, JBK18, KKW14, YDLW20, BMTZ21, BK13, CASA14, DCK07, EG11, FFQS05, HLGD19, NSH14, Stä05]. Verified [Loc18]. verifiers [GLPR12]. Verifying [GMR09, RKCW98, GPR11]. version [NHFP08, TV10]. version-consistent [NHFP08]. versions [BD06]. Versus [NSP^{+}14, Ann96, Yam96, dlPRGB99]. Vertex [CNZS17]. Vertex- [CNZS17]. Very [AOV^{+}99, Pet03]. VI [ACM94d, Ano03]. via [BCZY16, CCWY17, FBF01, Hig97, ITF^{+}22, KRBJ12, KGPH12, Kim14, LWV^{+}10, LZTZ15, LEL^{+}97a, LEL^{+}97b, RM00, SCCP13, SMD^{+}10, Ten98, VV11, WCW^{+}04b, WCW^{+}04c, WCW^{+}04d, WCW^{+}04a, WLK^{+}09]. Viability [KLH97].

Video [BC00]. view [KTLK13, PT91]. Vina [TO10]. Violations [MWP^{+}21]. Virtual [BSSS14, BBM09, KG05, KKDV03, PRB07, PHBC18, USE01, WCW^{+}04a, DLM99, DPZ97, DC99, DC00, MN03, MRG17, Ven97, WCW^{+}04b, WCW^{+}04c, WCW^{+}04d, WK08a, WK08b, WK08c]. Virtualization [LRZ16, YSS^{+}17, ABB^{+}15]. Virtualized [FNS^{+}22]. Virtually [LB92]. virtues [NJK16]. virus [GJ11]. viscous [RM99]. Visual [PTMB09, Dil93, McM96c, Esp96, Nag01]. Visualization [Ano97a, ACR01, Cal02, Caz02, MSU^{+}16, BCHS00, CSB00, MKK99, NCA93]. Visualizing [CdOS01, WT10, ACD^{+}18, DSEE13]. Visually [Dru95]. VLIW [For97, GSL10, ÖCS01]. VLSI [ABC^{+}93]. VM [FGG14]. VMs [KKJ^{+}13]. VMT [FNS^{+}22]. voltage [MTPT12]. volumes [Koo93]. VRSync [MTPT12]. vs [EHP^{+}07, GBK^{+}09, MMTW10, MCFT99, SSkP^{+}07, SKP^{+}02]. VTK [MSU^{+}16]. VTK-m [MSU^{+}16]. vulnerability [SSN10, WHG07].

WA [LCK11, ACM93c, IEE94a, IEE94d]. Wabi [Ano97a]. wait [PMCP22]. wait-free [PMCP22]. Waiting [LA93]. Waits [How00]. WAN [Yas95]. Wanted [Ano94g]. Warnings [CJW^{+}15]. warp [FSYA09, MTS10, Rei95, Tam95]. was [San04]. Washington [ACM92, Ano90, IEE94c, USE98a]. Watch [Ano97b]. water [LVA^{+}13]. Wave [Ano00b, BBC^{+}00, LS07, WQLJ18]. wave-based [WQLJ18]. wavelet [TKHG04]. Way [KAO05, MTN^{+}00, Rin99, ZJFA09, FGT96]. Ways [Wei97]. Weak [KZC15, TVD14]. Weakening [WRPP19]. Weaving [Pra95b]. Web [Ano94d, Swi09, Chl15a, Chl15b, Hig97, MGI14, PCM16, VP16].

WebAssembly [WRPP19]. Webrelay [Zha00]. WebThreads [Ano97a]. week [Ano95a, Ano95b]. weeks [But14]. Weight [IXS19, Way95]. Weighted [CNZS17, EE14, IMMP20, HFV^{+}12]. Weighted-IPC [EE14]. weighting [VS11b]. Weightless [SPY^{+}93]. Weld [ÖCS01]. well [Kub15]. well-determined [Kub15]. West [EV01]. WG2.5 [BT01]. Wheeler [BVP^{+}19, LHS16, NTR16]. Where [EHP^{+}07]. Whole [GN96, BBM09]. Whole-Program [GN96]. Wide [Ano94d, Ano96, FGT96]. wide-area [FGT96]. Widening [KKW14]. Will [BVM19, Ano95a, Ano95b]. WiMAX [CDD^{+}10]. Win32 [Bec01, BW97, CW98, Har99, How00, Lar97, PG99]. window [VS11b]. Windows [USE98a, HKT93, YZYL07, Hig97, Lee93, PG96, Pra95c, Pra95b, TCI98, Tim03, Yam96]. Winter [Ano90, USE89, USE91b, USE93b]. Wired [DHR^{+}01]. Within [BP05]. without [Gus05, LZBW14, Pla02]. woes [Ver97]. WOMPAT [Cha05, EV01]. Work [Ber96b, SKV21, Wal95, WYT^{+}20, ALHH08, Ber96a, BL94, BL99, Lep95, OdSSP12, RL14]. work-optimal [Lep95]. Work-stealing [WYT^{+}20, ALHH08, RL14]. worker [SCM05]. Workers [VP16]. Workflow [KD22]. workflows [FGG14]. Working [BT01]. Workload [EE14, KTR^{+}04, SSYG97, LBE^{+}98]. Workloads [DS09, FNS^{+}22, GVT^{+}17, KML04, LYH16, RCC12, SLJ^{+}18, CML00, PPGS20, SQP08a, SQP08b, SQP08c, WA08]. WorkPlace [Ano97c, Bra97]. works [Hig97, San04]. Workshop [ACM98a, RM03, Ano94e, Cha05, EV01, IEE89, IEE94a, IEE94d, Ass96, USE96, FR95]. Workstation [Ano00b, HN91, IEE89]. Workstations [KLH97, Lu98, LGH94, RGK99, PH97]. World [Ano92a, Ano92b, Ano94d, Ano96, Sut99, BBM09, Hol98d, Hol98a, Hol98b, Hol98c, Hol99a, Hol99b, WLG^{+}14]. World-wide [Ano96]. Wrapper [AS14]. Wrappers [Hub01]. Write [Sho97a, Sho97b]. Writer [Ano97a]. written [ND13]. WWOS [IEE89]. WWOS-II [IEE89].

X [Ano00b, Smi92, Sri95, TCS20, MSM^{+}16]. x86 [RMV22]. Xeon [SCD^{+}15]. Xlib [Gil93, STW93]. XML [BVL09, DWYB10]. XML-Based [BVL09]. XMT [DV99, VV00, BÇG14, VTSM12, VDBN98]. XMT-2 [BÇG14]. XPS [Ger95].

y-cruncher [Yee20]. Year [Ano99]. Yokohama [Ano03]. York [IEE90]. Yosemite [Ano00b].

z13 [ABB^{+}15, CJB^{+}15]. Zurich [Lak96].

References

Antoniu:2001:HSC

[A^{+}01] Gabriel Antoniu et al. The Hyperion system: Compiling multithreaded Java bytecode for distributed execution. Parallel Computing, 27(10):1279– 1297, September 2001. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                                     tronic). URL http://www.
                                                     elsevier.com/gej-ng/10/
                                                     35/21/47/40/27/abstract.
                                                     html;   http://www.elsevier.
                                                     nl/gej-ng/10/35/21/47/40/
                                                     27/article.pdf.

Aliaga:2015:UPE

[AAC^{+}15] José I. Aliaga, Hartwig Anzt, Maribel Castillo, Juan C. Fernández, Germán León, Joaquín Pérez, and Enrique S.

Quintana-Ortí. Unveiling the performance-energy tradeoff in iterative linear system solvers for multithreaded processors. Concurrency and Computation: Practice and Experience, 27(4):885– 904, March 25, 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alverson:1992:EHP

[AACK92] G. A. Alverson, R. Alverson, D. Callahan, and B. Koblenz. Exploiting heterogeneous parallelism on a multi-threaded multiprocessor. In ACM [ACM92], pages 188–197. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Amamiya:2009:CBN

[AAHF09] Satoshi Amamiya, Makoto Amamiya, Ryuzo Hasegawa, and Hiroshi Fujita. A continuation-based noninterruptible multithreading processor architecture. The Journal of Supercomputing, 47 (2):228–252, February 2009. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=47&
issue=2&spage=228.

Abdulla:2019:OSM

[AAJ^{+}19] Parosh Aziz Abdulla, Mohamed Faouzi Atig, Bengt Jonsson, Magnus Lång, Tuan Phong Ngo, and Konstantinos Sagonas. Optimal stateless model checking for reads-from equivalence under sequential consistency. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):150:1–150:29, October 2019. URL https:

//dl.acm.org/doi/abs/10.
1145/3360576.

Athanasaki:2008:EPL

[AAKK08] Evangelia Athanasaki, Nikos Anastopoulos, Kornilios Kourtis, and Nectarios Koziris. Exploring the performance limits of simultaneous multithreading for memory intensive applications. The Journal of Supercomputing, 44 (1):64–97, April 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=44&
issue=1&spage=64.

Antoniu:2001:DPP

[AB01] Gabriel Antoniu and Luc Bougé. DSM-PM2: a portable implementation platform for multithreaded DSM consistency protocols. Lecture Notes in Computer Science, 2026:55–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2026/20260055.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2026/20260055.
pdf.

Antoniu:2002:IMP

[AB02] Gabriel Antoniu and Luc Bougé. Implementing multithreaded protocols for release consistency on top of the generic DSM-PM platform. Lecture Notes in Computer Science, 2326:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2326/23260179.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2326/23260179.
pdf.

Axnix:2015:IZF

[ABB^{+}15] C. Axnix, G. Bayer, H. Bohm, J. von Buttlar, M. S. Farrell, L. C. Heller, J. P. Kubala, S. E. Lederer, R. Mansell, A. Nunez Mencias, and S. Usenbinz. IBM z13 firmware innovations for simultaneous multithreading and I/O virtualization. IBM Journal of Research and Development, 59(??):11:1–11:11, ???? 2015. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Agarwal:1993:SMV

[ABC^{+}93] Anant Agarwal, Jonathan Babb, David Chaiken, Godfrey D’Souza, Kirk Johnson, David Kranz, John Kubiatowicz, Beng-Hong Lim, Gino Maa, and Ken Mackenzie. Sparcle: a multithreaded VLSI processor for parallel computing. Lecture Notes in Computer Science, 748:359– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antonopoulos:2009:ASH

[ABC^{+}09] Christos D. Antonopoulos, Filip Blagojevic, Andrey N. Chernikov, Nikos P. Chrisochoides, and Dimitrios S. Nikolopoulos. Algorithm, software, and hardware optimizations for Delaunay mesh generation on simultaneous multithreaded architectures. Journal of Parallel and Distributed Computing, 69(7):601–612, July 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Aliaga:2015:CMS

[ABC^{+}15] José I. Aliaga, José M. Badía, Maribel Castillo, Davor Davidović, Rafael Mayo, and Enrique S. Quintana-Ortí. Outof-core macromolecular simulations on multithreaded architectures. Concurrency and Computation: Practice and Experience, 27(6):1540– 1550, April 25, 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aliaga:2012:SDG

[ABD^{+}12] José I. Aliaga, Paolo Bientinesi, Davor Davidović, Edoardo Di Napoli, Francisco D. Igual, and Enrique S. Quintana-Ortí. Solving dense generalized eigenproblems on multi-threaded architectures. Applied Mathematics and Computation, 218 (22):11279–11289, July 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S009630031200505X.

Agarwal:2010:DDP

[ABF^{+}10] R. Agarwal, S. Bensalem, E. Farchi, K. Havelund, Y. Nir-Buchbinder, S. Stoller, S. Ur, and L. Wang. Detection of deadlock potentials in multithreaded programs. IBM Journal of Research and Development, 54(5):3:1–3:15, ???? 2010. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Auerbach:2008:FTG

[ABG^{+}08] Joshua Auerbach, David F.

Bacon, Rachid Guerraoui, Jesper Honig Spring, and Jan Vitek. Flexible task graphs: a unified restricted thread programming model for Java. ACM SIGPLAN Notices, 43 (7):1–11, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Antoniu:2000:IJC

[ABH^{+}00] Gabriel Antoniu, Luc Bougé, Philip Hatcher, Mark Mac-Beth, Keith McGuigan, and Raymond Namyst. Implementing Java consistency using a generic, multithreaded DSM runtime system. Lecture Notes in Computer Science, 1800:560–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1800/18000560.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1800/18000560.
pdf.

Antoniu:2001:CMJ

[ABH^{+}01] Gabriel Antoniu, Luc Bougé, Philip Hatcher, Mark Mac-Beth, Keith McGuigan, and Raymond Namyst. Compiling multithreaded Java bytecode for distributed execution (distinguished paper). Lecture Notes in Computer Science, 1900:1039–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1900/19001039.
htm;   http://link.springer-
ny.com/link/service/series/

0558/papers/1900/19001039.
pdf.

Anderson:1992:SAE

[ABLL92] Thomas E. Anderson, Brian N.

Bershad, Edward D. Lazowska, and Henry M. Levy. Scheduler activations: Effective kernel support for the user-level management of parallelism. ACM Transactions on Computer Systems, 10(1): 53–79, February 1992. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tocs/1992-10-1/
p53-anderson/.

Amestoy:2019:PSB

[ABLM19] Patrick R. Amestoy, Alfredo Buttari, Jean-Yves L’Excellent, and Theo Mary. Performance and scalability of the block low-rank multifrontal factorization on multicore architectures. ACM Transactions on Mathematical Software, 45(1):2:1–2:26, March 2019. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3242094.

Antoniu:1999:ETT

[ABN99] G. Antoniu, L. Bouge, and R. Namyst. An efficient and transparent thread migration scheme in the PM2 runtime system. Lecture Notes in Computer Science, 1586:496– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aumage:2000:PAM

[ABN00] Olivier Aumage, Luc Bougé, and Raymond Namyst. A portable and adaptative multiprotocol communication library for multithreaded runtime systems. Lecture Notes in Computer Science, 1800:1136–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1800/18001136.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1800/18001136.
pdf.

Antoniu:2000:CDP

[ABNP00] G. Antoniu, L. Bougé, R. Namyst, and C. Pérez. Compiling data-parallel programs to a distributed runtime environment with thread isomigration. Parallel Processing Letters, 10(2/3):201– ??, September 2000. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL http:

//ejournals.wspc.com.sg/
ppl/10/1002_03/S0129626400000202.
html.

Aleen:2009:CAS

[AC09] Farhana Aleen and Nathan Clark. Commutativity analysis for software parallelization: letting program transformations see the big picture. ACM SIGPLAN Notices, 44(3):241–252, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Almasi:2003:DCD

[ACC^{+}03] George Almási, Călin Caşcaval, José G. Castaños, Monty Denneau, Derek Lieber, José E. Moreira, and Henry S. Warren, Jr. Dissecting Cyclops: a detailed analysis of a multithreaded architecture. ACM SIGARCH Computer Architecture News, 31 (1):26–38, March 2003. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Adams:2018:TTV

[ACD^{+}18] Joel C. Adams, Patrick A.

Crain, Christopher P. Dilley, Christiaan D. Hazlett, Elizabeth R. Koning, Serita M. Nelesen, Javin B. Unger, and Mark B. Vande Stel. TSGL: A tool for visualizing multithreaded behavior. Journal of Parallel and Distributed Computing, 118 (part 1)(??):233–246, August 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S0743731518301035.

ACM:1992:CPI

[ACM92] ACM, editor. Conference proceedings / 1992 International Conference on Supercomputing, July 19–23, 1992, Washington, DC. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

ACM:1993:CRT

[ACM93a]  ACM, editor.  Confer-

ence record of the Twentieth Annual ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the symposium, Charleston, South Carolina, January 10– 13, 1993. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15 1993. URL

                                        http://www.acm.org/pubs/
                                        contents/proceedings/plan/
                                        158511/index.html. ACM

order number 549930.

ACM:1993:PTF

[ACM93b] ACM, editor. Proceedings of the twenty-fifth annual ACM Symposium on the Theory of Computing, San Diego, California, May 16–18, 1993. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

ACM:1993:TCS

[ACM93c]  ACM, editor.  TRI-Ada

’93: Conference — September 1993, Seattle, WA, TRIADA -proceedings- 1993. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-621-2. LCCN ???? ACM Order No. 825930.

ACM:1994:ASC

[ACM94a] ACM, editor. ACM SIGPLAN ’94 Conference on Programming Language Design and Implementation (PLDI). Orlando, FL, USA, 20–24 June, 1994, volume 29(6) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, June 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1994:CRP

[ACM94b] ACM, editor. Conference record of POPL ’94, 21st ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the Symposium: Portland, Oregon, January 17–21, 1994. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-636-0. LCCN QA76.7 .A15 1994. URL

http://www.acm.org/pubs/
contents/proceedings/plan/
174675/index.html.

ACM:1994:IPI

[ACM94c] ACM, editor. ISSAC ’94:

Proceedings of the 1994 International Symposium on Symbolic and Algebraic Computation: July 20–22, 1994, Oxford, England, United Kingdom. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

ACM:1994:SIC

[ACM94d] ACM, editor. Sixth International Conference on Architectural Support for Programming Languages and Operating Systems (ASPLOS-VI). San Jose, CA, USA, 4–7 October, 1994, volume 29(11) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1995:CPI

[ACM95a] ACM, editor. Conference proceedings of the 1995 International Conference on Supercomputing, Barcelona, Spain, July 3–7, 1995, Conference Proceedings of the International Conference on Supercomputing. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995.

ACM:1995:CRP

[ACM95b] ACM, editor. Conference record of POPL ’95, 22nd ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the Symposium: San Francisco, California, January 22–25, 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-692-1. LCCN QA 76.7 A11 1995. URL

http://www.acm.org/pubs/
contents/proceedings/plan/
199448/index.html. ACM

order number: 549950.

ACM:1996:FCP

[ACM96] ACM, editor. FCRC ’96:

Conference proceedings of the 1996 International Conference on Supercomputing: Philadelphia, Pennsylvania, USA, May 25–28, 1996. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

ACM:1998:AWJ

[ACM98a] ACM, editor. ACM 1998 Workshop on Java for High-Performance Network Computing. ACM Press, New York, NY 10036, USA, 1998. ISBN ???? LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/
program.html. Possibly un-

published, except electronically.

ACM:1998:CRP

[ACM98b]  ACM, editor.  Confer-

ence record of POPL ’98: the 25th ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the Symposium, San Diego, California, 19–21 January 1998. ACM Press, New York, NY 10036, USA, 1998. ISBN 0-89791-979-3. LCCN QA76.7 .A15 1998. URL

http://www.acm.org/pubs/
contents/proceedings/plan/
268946/index.html. ACM

order number: 549981.

ACM:1998:PAI

[ACM98c] ACM, editor. Proceedings: the 25th Annual International Symposium on Computer Architecture, June 27–July 1, 1998, Barcelona, Spain, volume 26(3) of Computer architecture news. ACM Press, New York, NY 10036, USA, 1998. ISBN 0-8186-8491-7, 0-8186-8492-5, 0-8186-8493-3. LCCN QA76.9.A73 S97 1998.

URL http://portal.acm.
org/toc.cfm?id=279358;
http://portal.acm.org/toc.
cfm?id=285930. ACM Order

Number 414984. IEEE Computer Society Order Number PR08491; IEEE Order Plan Catalog Number 98CB36235.

ACM:1998:SHP

[ACM98d] ACM, editor. SC’98: High Performance Networking and Computing: Proceedings of the 1998 ACM/IEEE SC98 Conference: Orange County Convention Center, Orlando, Florida, USA, November 7–13, 1998. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN ???? LCCN ???? URL http:/

/www.supercomp.org/sc98/
papers/.

ACM:1999:PASa

[ACM99a] ACM, editor. Proceedings of the ACM SIGPLAN ’99 Conference on Programming Language Design and Implementation (PLDI ’99), Atlanta, Georgia, 2–4 May 1999. ACM Press, New York, NY 10036, USA, 1999. ISBN ???? LCCN ????

ACM:1999:SPO

[ACM99b] ACM, editor. SC’99: Oregon Convention Center 777 NE Martin Luther King Jr. Boulevard, Portland, Oregon, November 11–18, 1999. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999.

ACM:2000:SHP

[ACM00]  ACM, editor.  SC2000:

High Performance Networking and Computing. Dallas Convention Center, Dallas, TX, USA, November 4–10, 2000. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA,

                                        2000.  URL http://www.
                                        sc2000.org/proceedings/
                                        info/fp.pdf.

ACM:2001:PAJ

[ACM01]  ACM, editor.  Proceed-

ings of the ACM 2001 Java Grande/ISCOPE Conference: Palo Alto, Calif., June 2– 4, 2001. ACM Press, New York, NY 10036, USA, 2001. ISBN 1-58113-359-6. LCCN QA76.9.O35 A26 2001.

ACM:2003:SII

[ACM03] ACM, editor. SC2003: Igniting Innovation. Phoenix, AZ, November 15–21, 2003. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 1-58113-695-1. LCCN ????

Arvind:1997:MSC

[ACMA97] Arvind, A. Caro, J.-W. Maessen, and S. Aditya. A multithreaded substrate and compilation model for the implicitly parallel language pH. Lecture Notes in Computer Science, 1239:519–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attali:2001:GVJ

[ACR01] Isabelle Attali, Denis Caromel, and Marjorie Russo.

Graphical visualization of Java objects, threads, and locks. IEEE Distributed Systems Online, 2(1), 2001. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0101/features/att0101_print.
htm.

Adams:2008:ENE

[AD08] Michael D. Adams and R. Kent Dybvig. Efficient nondestructive equality checking for trees and graphs. ACM SIGPLAN Notices, 43(9):179–188, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adamo:1998:MTO

[Ada98] Jean-Marc Adamo. Multithreaded object-oriented MPIbased message passing interface: the ARCH library, volume SECS 446 of The Kluwer international series in engineering and computer science. Kluwer Academic Publishers, Dordrecht, The Netherlands; Boston, MA, USA, 1998. ISBN 0-7923-8165-3. xiv + 185 pp. LCCN TK5102.5.A293 1998. US$120.00.

Abraham:2005:ABP

[ÁdBdRS05] Erika Ábrahám, Frank S.

de Boer, Willem-Paul de Roever, and Martin Steffen. An assertion-based proof system for multithreaded Java. Theoretical Computer Science, 331 (2–3):251–290, February 25, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abraham:2008:DPS

[ÁdBdRS08] Erika Ábrahám, Frank S. de Boer, Willem-Paul de Roever, and Martin Steffen. A deductive proof system for multithreaded Java with exceptions. Fundamenta Informaticae, 82(4):391–463, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abraham:2003:TSP

[AddS03] E. Abraham, F. S. deBoer, W. P. deRoever, and M. Steffen. A tool-supported proof system for multithreaded Java. Lecture Notes in Computer Science, 2852:1– 32, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abadi:2006:TSL

[AFF06] Martin Abadi, Cormac Flanagan, and Stephen N. Freund. Types for safe locking: Static race detection for Java. ACM Transactions on Programming Languages and Systems, 28 (2):207–255, March 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Arnold:1996:MPJ

[AG96] K. Arnold and J. Gosling.

Multithreaded programming in Java. Web Techniques, 1(7): 34–40, 42–43, October 1996. CODEN WETEFA. ISSN 1086-556X.

Agerwala:2006:SRC

[AG06] T. Agerwala and M. Gupta.

Systems research challenges: a scale-out perspective. IBM Journal of Research and Development, 50(2/3):173–??, March /May 2006. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/502/agerwala.
html.

Akbari:2021:EMT

[AG21] Amir Akbari and Dennis Giannacopoulos. An efficient multi-threaded Newton– Raphson algorithm for strong coupling modeling of multiphysics problems. Computer Physics Communications, 258 (??):Article 107563, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465520302708.

Agarwal:1989:PTM

[Aga89] Anant Agarwal. Performance tradeoffs in multithreaded processors. Technical Report 89-566, Massachusetts Institute of Technology, Microsystems Program Office, Cambridge, MA, USA, 1989. 30 pp.

Agarwal:1991:PTM

[Aga91] Anant Agarwal. Performance tradeoffs in multithreaded processors. Technical report MIT/LCS/TR 501; VLSI memo no. 89-566, Laboratory for Computer Science, Massachusetts Institute of Technology, Cambridge, MA, USA, 1991. 39 pp.

Agarwal:1992:PTM

[Aga92] Anant Agarwal. Performance tradeoffs in multithreaded processors. IEEE Transactions on Parallel and Distributed Systems, 3(5):525– 539, September 1992. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anderson:2008:SCD

[AGEB08] Zachary Anderson, David Gay, Rob Ennals, and Eric Brewer. SharC: checking data sharing strategies for multithreaded C. ACM SIGPLAN Notices, 43(6):149–158, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

AlBarakat:2018:MFM

[AGJ18] Laith M. AlBarakat, Paul Gratz, V, and Daniel A. Jimenez. MTB-Fetch: Multithreading aware hardware prefetching for chip multiprocessors. IEEE Computer Architecture Letters, 17(2):175– 178, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Amrhein:1996:CSM

[AGK96] Beatrice Amrhein, Oliver Gloor, and Wolfgang Küchlin. A case study of multithreaded Gröbner basis completion. In Lakshman Y. N. [Lak96], pages 95–

102. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996.

URL http://www.acm.org:
80/pubs/citations/proceedings/
issac/236869/p95-amrhein/

.

Anderson:2009:LAC

[AGN09] Zachary R. Anderson, David Gay, and Mayur Naik. Lightweight annotations for controlling sharing in concurrent data structures. ACM SIGPLAN Notices, 44(6):98– 109, June 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Akkary:2000:CSM

[AH00] Haitham Akkary and Sébastien Hily. The case for speculative multithreading on SMT processors. Lecture Notes in Computer Science, 1940:59– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/1940/19400059.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1940/19400059.
                                        pdf.

Abdulla:2008:MCR

[AHK08] Parosh Aziz Abdulla, Frédéric Haziza, and Mats Kindahl. Model checking race-freeness. ACM SIGARCH Computer Architecture News, 36(5):72– 79, December 2008. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Adiletta:2002:PSA

[AHW02] Matthew Adiletta, Donald Hooper, and Myles Wilde. Packet over SONET: Achieving 10 Gigabit/sec packet processing with an IXP2800. Intel Technology Journal, 6 (3):29–39, August 15, 2002. ISSN 1535-766X. URL

                                        http://developer.intel.
                                        com/technology/itj/2002/
                                        volume06issue03/art05_packetoversonet/
                                        p01_abstract.htm;   http:
                                        //developer.intel.com/technology/
                                        itj/2002/volume06issue03/
                                        art05_packetoversonet/vol6iss3_
                                        art05.pdf.

Aitken:1996:MCJ

[Ait96] Gary Aitken. Moving from C++ to Java. Dr. Dobb’s Journal of Software Tools, 21 (3):52, 54–56, March 1996.

CODEN DDJOEB. ISSN 1044-789X.

Ahn:2012:ISE

[AJK^{+}12] Jung Ho Ahn, Norman P.

Jouppi, Christos Kozyrakis, Jacob Leverich, and Robert S. Schreiber. Improving system energy efficiency with memory rank subsetting. ACM Transactions on Architecture and Code Optimization, 9(1):4:1– 4:??, March 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Azagury:1999:NIR

[AKP99] Alain Azagury, Elliot K.

Kolodner, and Erez Petrank. A note on the implementation of replication-based garbage collection for multithreaded applications and multiprocessor environments. Parallel Processing Letters, 9(3):391– ??, September 1999. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Aciicmez:2006:PSB

[AKS06] Onur Aciiçmez, Çetin Kaya Koç, and Jean-Pierre Seifert. On the power of simple branch prediction analysis. Technical report, School of EECS, Oregon State University, Corvallis, OR 97331, USA, October 2006.

URL http://eprint.iacr.
org/2006/351;   http://
eprint.iacr.org/2006/351.
pdf.

Arjomand:2016:BAP

[AKSD16] Mohammad Arjomand, Mahmut T. Kandemir, Anand Sivasubramaniam, and Chita R. Das. Boosting access parallelism to PCM-based main memory. ACM SIGARCH Computer Architecture News, 44(3):695–706, June 2016. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Arman:2021:OHP

[AL21] Arif Arman and Dmitri Loguinov. Origami: a highperformance mergesort framework. Proceedings of the VLDB Endowment, 15(2): 259–271, October 2021. CO-DEN ???? ISSN 2150-

                                        8097.  URL https://
                                        dl.acm.org/doi/10.14778/
                                        3489496.3489507.

Amer:2018:LCM

[ALB^{+}18] Abdelhalim Amer, Huiwei Lu, Pavan Balaji, Milind Chabbi, Yanjie Wei, Jeff Hammond, and Satoshi Matsuoka. Lock contention management in multithreaded MPI. ACM Transactions on Parallel Computing (TOPC), 5 (3):12:1–12:??, January 2018. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3275443.

Alfieri:1994:EKI

[Alf94] R. A. Alfieri. An efficient kernel-based implementation of POSIX threads. In Anonymous [Ano94f], pages 59–72. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

Agrawal:2008:AWS

[ALHH08] Kunal Agrawal, Charles E.

Leiserson, Yuxiong He, and Wen Jing Hsu. Adaptive work-stealing with parallelism feedback. ACM Transactions on Computer Systems, 26 (3):7:1–7:32, September 2008. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Agrawal:2010:HLF

[ALS10] Kunal Agrawal, Charles E.

Leiserson, and Jim Sukha. Helper locks for fork-join parallel programming. ACM SIG-PLAN Notices, 45(5):245–256, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ahn:2009:MDE

[ALSJ09] Jung Ho Ahn, Jacob Leverich, Robert S. Schreiber, and Norman P. Jouppi. Multicore DIMM: an energy efficient memory module with independently controlled DRAMs. IEEE Computer Architecture Letters, 8(1):5–8, January/ June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Amer:2015:MRC

[ALW^{+}15] Abdelhalim Amer, Huiwei Lu, Yanjie Wei, Pavan Balaji, and Satoshi Matsuoka. MPI+Threads: runtime contention and remedies. ACM SIGPLAN Notices, 50(8):239– 248, August 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Amamiya:1989:DFC

[Ama89] M. Amamiya. Data flow computing and parallel reduction machine. Future Generation Computer Systems, 4 (??):53–67, ???? 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Amaranth:1998:TBM

[Ama98]  Paul Amaranth.  A Tcl-

based multithreaded test harness. In USENIX [USE98b], page ?? ISBN 1-880446-98-7. LCCN QA76.73.T44 T34 1998. URL http://db.

                                        usenix.org/publications/
                                        library/proceedings/tcl98/
                                        amaranth.html.

Aamodt:2003:FMO

[AMC^{+}03] Tor M. Aamodt, Pedro Marcuello, Paul Chow, Antonio González, Per Hammarlund, Hong Wang, and John P. Shen. A framework for modeling and optimization of prescient instruction prefetch. ACM SIGMETRICS Performance Evaluation Review, 31 (1):13–24, June 2003. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Abraham-Mumm:2002:VJR

[ÁMdBdRS02] Erika Ábrahám-Mumm, Frank S.

de Boer, Willem-Paul de Roever, and Martin Steffen. Verification for Java’s reentrant multithreading concept. Lecture Notes in Computer Science, 2303:5– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2303/23030005.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2303/23030005.
pdf.

Azizi:2009:AEC

[AMPH09] Omid Azizi, Aqeel Mahesri, Sanjay J. Patel, and Mark Horowitz. Area-efficiency in CMP core design: cooptimization of microarchitecture and physical design. ACM SIGARCH Computer Architecture News, 37 (2):56–65, May 2009. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Aiex:1998:CMT

[AMRR98] R. M. Aiex, S. L. Martins, C. C. Ribeiro, and N. D. L. R. Rodriguez. Cooperative multi-thread parallel tabu search with an application to circuit partitioning. Lecture Notes in Computer Science, 1457:310–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anju:2022:MID

[AN22] M. A. Anju and Rupesh Nasre. Multi-interval Dom-Lock: Toward improving concurrency in hierarchies. ACM Transactions on Parallel Computing (TOPC), 9(3):12:1–12:27, September 2022. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3543543.

Annavaram:1996:BVN

[Ann96] Murali Annavaram. Blocking versus non-blocking: issues and tradeoffs in multithreaded code execution. Thesis (M.S.), Colorado State University, Fort Collins, CO, USA, 1996. viii + 57 pp.

Anonymous:1990:PWU

[Ano90] Anonymous, editor. Proceedings of the Winter 1990 USENIX Conference, Washington, DC, USA, January 22–26, 1990. USENIX Association, Berkeley, CA, USA, 1990.

Anonymous:1991:PIS

[Ano91] Anonymous, editor. Proceedings of the International Symposium on Supercomputing: Fukuoka, Japan, November 6–8, 1991. Kyushu University Press, Fukuoka, Japan, 1991. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Anonymous:1992:MWPa

[Ano92a] Anonymous. It’s a multithreaded world, part 1: Multithreaded operating systems are becoming the norm. Here’s how your applications can exploit them. Byte Magazine, 17 (5):289–??, May 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:MWPb

[Ano92b] Anonymous. It’s a multithreaded world, part 2: Multithreaded operating systems are taking over. Are your applications ready? Byte Magazine, 17(6):351–??, June 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:ICS

[Ano94a] Anonymous, editor. 1994 International Computer Symposium Conference Proceedings. Nat. Chiao Tung Univ, Hsinchu, Taiwan, 1994. ISBN ???? LCCN ???? 2 vol.

Anonymous:1994:MDP

[Ano94b]  Anonymous.  Multiproces-

sor desktops are proliferating, even though there remains a shortage of multithreaded applications for them. Open Systems Today, 165:60–??, December 1994. ISSN 1061-0839.

Anonymous:1994:DCT

[Ano94c] Anonymous. On the design of Chant: a talking threads package. In IEEE [IEE94c], pages 350–359. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PIW

[Ano94d] Anonymous, editor. Proceedings of the 2nd International World Wide Web conference, Mosaic and the Web, October 1994, Ramada-Congress Hotel, 520 South Michigan Avenue, Chicago, IL, volume 18(6) of Online & CDROM review: the international journal of online & optical information systems. Learned Information, Medford, NJ, USA, 1994. CO-DEN ONCDEW. ISSN 0309-

314X. URL http://www.
ncsa.uiuc.edu/SDG/IT94/
Proceedings/WWW2_Proceedings.
html.

Anonymous:1994:SIP

[Ano94e] Anonymous. Special issue: panel sessions of the 1991 Workshop on Multithreaded Computers, November 22, 1991, Albuquerque, New Mexico, in conjunction with Supercomputing ’91. Computer architecture news, 22(1):2–33, 1994.

Anonymous:1994:USC

[Ano94f] Anonymous, editor. USENIX Summer conference: – June 1994, Boston, MA, USENIX Conference Proceedings 1994. USENIX Association, Berkeley, CA, USA, 1994. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

Anonymous:1994:WMC

[Ano94g] Anonymous. Wanted: The Multithreaded CIO. Datamation, 40(8):34–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:HUW

[Ano95a] Anonymous. HP-UX 10.0 will be unveiled this week, with newly tuned kernel and I/O paths, plus a multithreaded NFS implementation. Open Systems Today, 168:34–??, February 1995. ISSN 1061-0839.

Anonymous:1995:HWB

[Ano95b] Anonymous. HP-UX 10.0 will be unveiled this week, with newly tuned kernel and I/O paths, plus a multithreaded NFS implementation. Open Systems Today, 168:34–??, February 1995. ISSN 1061-0839.

Anonymous:1996:WWD

[Ano96] Anonymous. World-wide distributed system using Java and the Internet. IEEE International Symposium on High Performance Distributed Computing, Proceedings, pages 11–18, 1996. CODEN PID-CFB. ISSN 1082-8907. IEEE catalog number 96TB100069.

Anonymous:1997:NPW

[Ano97a]  Anonymous.  New prod-

ucts: WebThreads 1.0.1; QUERYFLEX Report Writer; Linux Pro Desktop 1.0; NDP Fortran for Linux; Numerics and Visualization for Java; Craftworks Linux/AXP 2.2; InfoDock Linux Software Development Toolset; Caldera Wabi 2.2 for Linux. Linux Journal, 34:??, February 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1997:TWP

[Ano97b]  Anonymous.  Tech watch

— pattern-recognition system. Piecing together history. 3D semiconductor simulation. Multi-threaded architecture. Computer Graphics World, 20 (9):15–??, September 1997. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1997:TNR

[Ano97c] Anonymous. Technology news & reviews: Chemkin software; OpenMP Fortran Standard; ODE toolbox for Matlab; Java products; Scientific WorkPlace 3.0. IEEE Computational Science & Engineering, 4(4):75–??, October/December 1997. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1997/pdf/c4075.
pdf.

Anonymous:1998:MS

[Ano98a] Anonymous. Multithreaded system. IEEE Micro, 18(3): 76, May/June 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1998:NTS

[Ano98b] Anonymous. New tools: Software development: Uniscape’s internationalization library; Global Technologies’ Unixto-NT solution; KAI’s multithreaded Java debugging tool; Price Systems’ parametric forecasting tool. Computer, 31(6):98, 102, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co1998/pdf/r6098.
pdf.

Anonymous:1999:BST

[Ano99] Anonymous. Bookshelf: Surviving the top ten challenges of software development; the Year 2000 crisis; the continuing challenge; software project survival guide; objectoriented multithreading using C++. IEEE Software, 16 (1):114–??, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:

//dlib.computer.org/so/
books/so1999/pdf/s1114.
pdf.

Anonymous:2000:CCI

[Ano00a] Anonymous, editor. Cool Chips III: An International Symposium on Low-Power and High-Speed Chips, Kikai-Shinko-Kaikan, Tokyo, Japan April 24–25, 2000. ????, ????, 2000.

Anonymous:2000:NPAa

[Ano00b]  Anonymous.  New prod-

ucts: AVP for Linux/FreeBSD UNIX, Kaspersky Lab Ltd.; API PowerRAC Chassis 320, Alpha Processor Inc.; ODBC-ODBC Bridge, Easysoft Ltd.; LinkScan 6.1, Electronic Software Publishing Corporation; Metro-X Enhanced Server CD, Metro Link, Inc.; P-STAT Statistical Software, P-STAT, Inc.; System Manager in a Box v1.0, Pega-Soft Canada; PGI Workstation 3.1, PGI; Quick Restore 2.6, Workstation Solutions, Inc.; Threads.h++ and Tools.h++ Professional, Rogue Wave Software; Scriptics Connect 1.0, 1.1, Scriptics Corporation; TapeWare 6.2 Backup Software, Yosemite Technologies, Inc.; DoubleVision for Linux Systems, Tridia Corporation. Linux Journal, 71:??, March 2000. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2000:SLT

[Ano00c] Anonymous. Strictly on-line:

T/TCP: TCP for Transactions by Mark Stacey, Ivan Griffin and John Nelson; POSIX Thread Libraries by Felix Garcia and Javier Fernandez; Linux and Open-Source Applications by Peter Jones and M. B. Jorgenson; Laptops for Linux! by Jason Kroll. Linux Journal, 70:??, February 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue70/3075.
html;   http://noframes.
linuxjournal.com/lj-issues/
issue70/3184.html;   http:
//noframes.linuxjournal.
com/lj-issues/issue70/3683.
html;   http://noframes.
linuxjournal.com/lj-issues/
issue70/3766.html.

Anonymous:2001:ESM

[Ano01] Anonymous. Errata: “Speculative Multithreaded Processors”. Computer, 34(5):7, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2001/pdf/r5004.

pdf. See [SR01b].

Anonymous:2002:ST

[Ano02]  Anonymous.  Speculative

threads. ACM SIGARCH Computer Architecture News, 30(5):??, December 2002. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Anonymous:2003:CCV

[Ano03] Anonymous, editor. Cool Chips VI:An International Symposium on Low-Power and High-Speed Chips, Yokohama Joho Bunka Center, Yokohama, Japan (Yokohama Media & Communications Center, Yokohama, Japan) April 16–18, 2003. ????, ????, 2003. ISBN ???? LCCN ????

Anonymous:2005:ECS

[Ano05] Anonymous. Errata: Characterization of Simultaneous Multithreading (SMT) Efficiency in POWER5. IBM Journal of Research and Development, 49(6):1003– ??, November 2005. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/496/errata.html.

See [MMM^{+}05].

Atkinson:1999:PTF

[AOV^{+}99] Malcolm P. Atkinson, Maria E. Orlowska, Patrick Valduriez, Stanley B. Zdonik, and Michael L. Brodie, editors. Proceedings of the Twentyfifth International Conference on Very Large Databases, Edinburgh, Scotland, UK, 7– 10 September, 1999. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. Also known as VLDB’99.

Arnau:2012:BMG

[APX12] José-María Arnau, Joan-Manuel Parcerisa, and Polychronis Xekalakis. Boosting mobile GPU performance with a decoupled access/execute fragment processor. ACM SIGARCH Computer Architecture News, 40 (3):84–93, June 2012. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ISCA ’12 conference proceedings.

Areias:2017:SDP

[AR17] Miguel Areias and Ricardo Rocha. On scaling dynamic programming problems with a multithreaded tabling Prolog system. The Journal of Systems and Software, 125(??):417–426, March 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

//www.sciencedirect.com/
science/article/pii/S0164121216300929.

Areias:2019:MDL

[AR19] Miguel Areias and Ricardo Rocha. Multi-dimensional lock-free arrays for multithreaded mode-directed tabling in Prolog. Concurrency and Computation: Practice and Experience, 31(5):e4491:1– e4491:??, March 10, 2019.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Algosaibi:2020:PBT

[ARA20] Abdulelah Algosaibi, Khaled Ragab, and Saleh Albahli. Parallel-based techniques for managing and analyzing the performance on semantic graph. Parallel Processing Letters, 30(02):??, June 2020. ISSN 0129-6264 (print), 1793-642X (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0129626420500073.

Adiletta:2002:NGI

[ARB^{+}02] Matthew Adiletta, Mark Rosenbluth, Debra Bernstein, Gilbert Wolrich, and Hugh Wilkinson. The next generation of Intel IXP network processors. Intel Technology Journal, 6(3):6–18, August 15, 2002. ISSN 1535-766X. URL

                                        http://developer.intel.
                                        com/technology/itj/2002/
                                        volume06issue03/art01_nextgenixp/
                                        p01_abstract.htm;   http:
                                        //developer.intel.com/technology/
                                        itj/2002/volume06issue03/
                                        art01_nextgenixp/vol6iss3_
                                        art01.pdf.

Arunachalam:1992:EMM

[Aru92] Prakash Arunachalam. Evaluation of a multithreaded microprocessor with MIPS R3000 instruction set. Thesis (M.S. in Engineering), University of Texas at Austin, Austin, TX, USA, 1992. vii + 45 pp.

Addison:2003:OIA

[ARvW03] C. Addison, Y. Ren, and M. van Waveren. OpenMP issues arising in the development of parallel BLAS and LAPACK libraries. Scientific Programming, 11(2):95– 104, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Awile:2014:PWF

[AS14] Omar Awile and Ivo F.

Sbalzarini. A Pthreads wrapper for Fortran 2003. ACM Transactions on Mathematical Software, 40(3):19:1–19:15, April 2014. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

USENIX:1996:ATT

[Ass96] USENIX Association, editor.

4th Annual Tcl/Tk Workshop ’96, July 10–13, 1996. Monterey, CA. USENIX Association, Berkeley, CA, USA, July 10–13, 1996. ISBN 1-880446-78-2. LCCN QA76.73.T44 T44 1996.

Asyabi:2019:COS

[ASSS19] Esmail Asyabi, Erfan Sharafzadeh, SeyedAlireza SanaeeKohroudi, and Mohsen Sharifi. CTS: an operating system CPU scheduler to mitigate tail latency for latency-sensitive multi-threaded applications. Journal of Parallel and Distributed Computing, 133(??):

232–243, November 2019. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0743731518302387.

Altiparmak:2016:MMF

[AT16] N. Altiparmak and A. S. Tosun. Multithreaded maximum flow based optimal replica selection algorithm for heterogeneous storage architectures. IEEE Transactions on Computers, 65(5):1543–1557, May 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Adl-Tabatabai:2006:CRS

[ATLM^{+}06] Ali-Reza Adl-Tabatabai, Brian T. Lewis, Vijay Menon, Brian R. Murphy, Bratin Saha, and Tatiana Shpeisman. Compiler and runtime support for efficient software transactional memory. ACM SIG-PLAN Notices, 41(6):26–37, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Arslan:2021:ESR

[AU21] Sanem Arslan and Osman Unsal. Efficient selective replication of critical code regions for SDC mitigation leveraging redundant multithreading. The Journal of Supercomputing, 77(12):14130– 14160, December 2021. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s11227-021-03804-6.

Arteaga:2017:GFG

[AZG17] Jaime Arteaga, Stéphane Zuckerman, and Guang R. Gao. Generating fine-grain multithreaded applications using a multigrain approach. ACM Transactions on Architecture and Code Optimization, 14(4):47:1–47:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Boehm:2008:FCC

[BA08] Hans-J. Boehm and Sarita V.

Adve. Foundations of the C++ concurrency memory model. ACM SIGPLAN Notices, 43(6):68–78, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bocchino:2009:TES

[BAD^{+}09] Robert L. Bocchino, Jr., Vikram S. Adve, Danny Dig, Sarita V. Adve, Stephen Heumann, Rakesh Komuravelli, Jeffrey Overbey, Patrick Simmons, Hyojin Sung, and Mohsen Vakilian. A type and effect system for deterministic parallel Java. ACM SIGPLAN Notices, 44(10):97–116, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergan:2010:CCRa

[BAD^{+}10a] Tom Bergan, Owen Anderson, Joseph Devietti, Luis Ceze, and Dan Grossman. CoreDet: a compiler and runtime system for deterministic multithreaded execution. ACM SIGARCH Computer Architecture News, 38 (1):53–64, March 2010. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Bergan:2010:CCRb

[BAD^{+}10b] Tom Bergan, Owen Anderson, Joseph Devietti, Luis Ceze, and Dan Grossman. Core-Det: a compiler and runtime system for deterministic multithreaded execution. ACM SIGPLAN Notices, 45(3):53– 64, March 2010. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1995:UOV

[Bak95a]  Henry G. Baker.  “use-

once” variables and linear objects: storage management, reflection and multithreading. ACM SIGPLAN Notices, 30(1):45–52, January 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1995:GTP

[Bak95b] Mary Baker. Going threadbare (panel session): sense or sedition? a debate on the threads abstraction. Operating Systems Review, 29(5): 227, December 1995. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Baldwin:2002:LMF

[Bal02] John H. Baldwin. Locking in the multithreaded FreeBSD kernel. In USENIX [USE02], pages 27–35. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736 2002. URL http://

www.usenix.org/publications/
library/proceedings/bsdcon02/
baldwin.html.

Bic:1993:EUI

[BAM93] Lubomir Bic and Mayez Al-Mouhamed. The EM-4 under implicit parallelism. Journal of Parallel and Distributed Computing, 19(3): 255–261, November 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1109/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1109/production/pdf.

Burckhardt:2007:CCC

[BAM07] Sebastian Burckhardt, Rajeev Alur, and Milo M. K. Martin. CheckFence: checking consistency of concurrent data types on relaxed memory models. ACM SIG-PLAN Notices, 42(6):12–21, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barkstrom:2009:UAS

[Bar09] Bruce R. Barkstrom. On using Ada to solve problems in computational economics and related disciplines with concurrent, multiagent algorithms. ACM SIGADA Ada Letters, 29(3):61–72, December 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bauer:1992:PCE

[Bau92] Barr E. Bauer. Parallel C extensions. Dr. Dobb’s Journal of Software Tools, 17(8): 110, 112–114, 124, 127, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Budhkar:2019:AMD

[BAZ^{+}19] Prerna Budhkar, Ildar Absalyamov, Vasileios Zois, Skyler Windh, Walid A. Najjar, and Vassilis J. Tsotras. Accelerating in-memory database selections using latency masking hardware threads. ACM Transactions on Architecture and Code Optimization, 16(2):13:1–13:??, May 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Bolding:2000:MSM

[BB00] Barry Bolding and Kim Baldridge. Multithreaded shared memory parallel implementation of the electronic structure code GAMESS. Computer Physics Communications, 128(1–2):55–66, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465500000679.

Barros:2020:ALS

[BB20] D. A. Barros and C. Bentes.

Analyzing the loop scheduling mechanisms on Julia multithreading. In 2020 IEEE 32nd International Symposium on Computer Architecture and High Performance Computing (SBAC-PAD), pages 257–

264. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2020.

Bova:2000:DLP

[BBC^{+}00] Steve W. Bova, Clay P. Breshears, Christine E. Cuicchi, Zeki Demirbilek, and Henry A. Gabb. Dual-level parallel analysis of harbor wave response using MPI and OpenMP. The International Journal of High Performance Computing Applications, 14 (1):49–64, Spring 2000. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Balter:1991:AIG

[BBD^{+}91] R. Balter, J. Bernadat, D. Decouchant, A. Duda, A. Freyssinet, S. Krakowiak, M. Meysembourg, P. Le Dot, H. Nguyen Van, E. Paire, M. Riveill, C. Roison, X. Rousset de Pina, R. Scioville, and G. Vandôme. Architecture and implementation of guide, an object-oriented distributed system. Computing Systems, 4(1):31–67, Winter 1991. CO-DEN CMSYE2. ISSN 0895-6340.

Ball:2011:PPT

[BBdH^{+}11] Thomas Ball, Sebastian Burckhardt, Peli de Halleux, Madan Musuvathi, and Shaz Qadeer. Predictable and progressive testing of multithreaded code. IEEE Software, 28(3):75–83, May/June 2011. CODEN IESEDJ. ISSN 0740-7459 (print), 0740-7459 (electronic).

Balis:2002:CPM

[BBFW02] B. Balis, M. Bubak, W. Funika, and R. Wismüller. A concept of portable monitoring of multithreaded programs. Lecture Notes in Computer Science, 2330:884–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2330/23300884.
htm;   http://link.springer-
ny.com/link/service/series/

0558/papers/2330/23300884.
pdf.

Balis:2003:MSM

[BBFW03] Bartosz Baliś, Marian Bubak, Włodzimierz Funika, and Roland Wismüller. A monitoring system for multithreaded applications. Future Generation Computer Systems, 19(5):641–650, July 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Balaji:2010:FGM

[BBG^{+}10] Pavan Balaji, Darius Buntinas, David Goodell, William Gropp, and Rajeev Thakur. Fine-grained multithreading support for hybrid threaded MPI programming. The International Journal of High Performance Computing Applications, 24(1):49–57, February 2010. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://

hpc.sagepub.com/content/
24/1/49.full.pdf+html.

Bender:2017:TLM

[BBH^{+}17] Michael A. Bender, Jonathan W.

Berry, Simon D. Hammond, K. Scott Hemmert, Samuel McCauley, Branden Moore, Benjamin Moseley, Cynthia A. Phillips, David Resnick, and Arun Rodrigues. Two-level main memory codesign: Multi-threaded algorithmic primitives, analysis, and simulation. Journal of Parallel and Distributed Computing, 102(??):213–228, April 2017. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S074373151630185X.

Bratanov:2009:VMW

[BBM09] Stanislav Bratanov, Roman Belenov, and Nikita Manovich. Virtual machines: a whole new world for performance analysis. Operating Systems Review, 43(2): 46–55, April 2009. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Butler:2011:BAM

[BBSG11] Michael Butler, Leslie Barnes, Debjit Das Sarma, and Bob Gelinas. Bulldozer: An approach to multithreaded compute performance. IEEE Micro, 31(2):6–15, March/April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Barabash:2005:PIM

[BBYG^{+}05] Katherine Barabash, Ori Ben-Yitzhak, Irit Goft, Elliot K. Kolodner, Victor Leikehman, Yoav Ossia, Avi Owshanko, and Erez Petrank. A parallel, incremental, mostly concurrent garbage collector for servers. ACM Transactions on Programming Languages and Systems, 27(6): 1097–1146, November 2005.

CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Buhr:1994:TRM

[BC94] R. J. A. Buhr and R. S.

Casselman. Timethread-role maps for object-oriented design of real-time-and-distributed systems. ACM SIGPLAN Notices, 29(10):301, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ball:1998:MTA

[BC98] Steve Ball and John Miller Crawford. Multi-threaded assignment surprises. Java Report: The Source for Java Development, 3(??):??, September 1998. CODEN JREPFI. ISSN 1086-4660. URL http:/

/archive.javareport.com/
9809/html/from_pages/ftp_
col1.shtml.

Bhandarkar:2000:PPM

[BC00] Suchendra M. Bhandarkar and Shankar R. Chandrasekaran. Parallel parsing of MPEG video in a multi-threaded multiprocessor environment. Lecture Notes in Computer Science, 1800:194–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1800/18000194.
htm;   http://link.springer-

                                        ny.com/link/service/series/
                                        0558/papers/1800/18000194.
                                        pdf.

Boudol:2002:NCP

[BC02] Gérard Boudol and Ilaria Castellani. Noninterference for concurrent programs and thread systems. Theoretical Computer Science, 281(1-2): 109–130, May 2002. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bronson:2010:PCB

[BCCO10] Nathan G. Bronson, Jared Casper, Hassan Chafi, and Kunle Olukotun. A practical concurrent binary search tree. ACM SIGPLAN Notices, 45(5):257–268, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Banerjee:1995:PCD

[BCG^{+}95] Prithviraj Banerjee, John A. Chandy, Manish Gupta, Eugene W. Hodges IV, John G. Holm, Antonio Lain, Daniel J. Palermo, Shankar Ramaswamy, and Ernesto Su. The Paradigm compiler for distributed-memory multicomputers. Computer, 28(10): 37–47, October 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boneti:2008:SCP

[BCG^{+}08] Carlos Boneti, Francisco J.

Cazorla, Roberto Gioiosa, Alper Buyuktosunoglu, Chen-Yong Cher, and Mateo Valero. Software-controlled priority characterization of POWER5 processor. ACM SIGARCH Computer Architecture News, 36(3):415–426, June 2008. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Bergan:2013:ICS

[BCG13] Tom Bergan, Luis Ceze, and Dan Grossman. Inputcovering schedules for multithreaded programs. ACM SIGPLAN Notices, 48(10): 677–692, October 2013. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Bokhari:2014:MMM

[BÇG14] Shahid H. Bokhari, Ümit V.

Çatalyürek, and Metin N. Gurcan. Massively multithreaded maxflow for image segmentation on the Cray XMT-2. Concurrency and Computation: Practice and Experience, 26(18):2836–2855, December 25, 2014. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bedy:2000:VSM

[BCHS00] Michael Bedy, Steve Carr, Xianlong Huang, and Ching-Kuang Shene. A visualization system for multithreaded programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):1–5, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Biagioni:1998:SST

[BCL^{+}98] Edoardo Biagioni, Ken Cline, Peter Lee, Chris Okasaki, and Chris Stone. Safefor-space threads in Standard ML. Higher-Order and Symbolic Computation, 11 (2):209–225, December 1998. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=1388-3690&volume=11&
issue=2&spage=209;   http:
//www.wkap.nl/oasis.htm/
187569.

Benner:2007:SLS

[BCM^{+}07] Peter Benner, Maribel Castillo, Rafael Mayo, Enrique S. Quintana-Ortí, and Gregorio Quintana-Ortí. Stabilizing large-scale generalized systems on parallel computers using multithreading and message-passing. Concurrency and Computation: Practice and Experience, 19 (4):531–542, March 25, 2007.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ball:2001:PVM

[BCR01] Thomas Ball, Sagar Chaki, and Sriram K. Rajamani. Parameterized verification of multithreaded software libraries. Lecture Notes in Computer Science, 2031:158– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2031/20310158.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2031/20310158.
pdf.

Bajaj:2011:FFP

[BCS11] Chandrajit L. Bajaj, Rezaul Chowdhury, and Vinay Siddahanavalli. F^{2}Dock: Fast Fourier protein-protein docking. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 8(1): 45–58, January 2011. CO-DEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Badamo:2016:IPE

[BCZY16] Michael Badamo, Jeff Casarona, Minshu Zhao, and Donald Yeung. Identifying powerefficient multicore cache hierarchies via reuse distance analysis. ACM Transactions on Computer Systems, 34(1):

3:1–3:??, April 2016. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Beyls:2000:CGM

[BD00] K. E. Beyls and E. H. D’Hollander. Compiler generated multithreading to alleviate memory latency. J.UCS: Journal of Universal Computer Science, 6(10):968–993, October 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_6_10/
compiler_generated_multithreading_
to.

Brzuszek:2006:MTS

[BD06] Marcin Brzuszek and Andrzej Daniluk. Multithreaded transactions in scientific computing: New versions of a computer program for kinematical calculations of RHEED intensity oscillations. Computer Physics Communications, 175 (10):678–681, November 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465506002979.

Bic:1998:MAD

[BDF98] Lubomir Bic, Michael B. Dillencourt, and Munehiro Fukuda. Mobile agents, DSM, coordination, and selfmigrating threads: a common framework. UCI-ICS technical report 98-33, Information and Computer Science, University of California, Irvine, Irvine, CA, October 8, 1998. 11 pp.

Bracy:2006:DAC

[BDJ06] A. Bracy, K. Doshi, and Q. Jacobson. Disintermediated active communication. IEEE Computer Architecture Letters, 5(2):15, February 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Blundell:2007:MFC

[BDLM07] Colin Blundell, Joe Devietti, E. Christopher Lewis, and Milo M. K. Martin. Making the fast case common and the uncommon case simple in unbounded transactional memory. ACM SIGARCH Computer Architecture News, 35 (2):24–34, May 2007. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Bangs:1998:BOS

[BDM98] Gaurav Bangs, Peter Druschel, and Jeffrey C. Mogul. Better operating system features for faster network servers. ACM SIGMETRICS Performance Evaluation Review, 26(3):23–30, December 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bouge:2002:IRE

[BDN02] L. Bougé, V. Danjean, and R. Namyst. Improving reactivity to I/O events in multithreaded environments using a uniform, scheduler-centric API. Lecture Notes in Computer Science, 2400:605–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2400/24000605.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2400/24000605.
pdf.

Bouajjani:2012:ARP

[BE12] Ahmed Bouajjani and Michael Emmi. Analysis of recursively parallel programs. ACM SIG-PLAN Notices, 47(1):203–214, January 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bouajjani:2013:ARP

[BE13] Ahmed Bouajjani and Michael Emmi. Analysis of recursively parallel programs. ACM Transactions on Programming Languages and Systems, 35(3):10:1–10:??, November 2013. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Becker:2000:JSU

[Bec00] Pete Becker. The journeyman’s shop: Unraveling multithreading. C/C++ Users Journal, 18(8):71–??, August 2000. CODEN CCUJEX. ISSN 1075-2838.

Becker:2001:SMW

[Bec01] Thomas Becker. Synchronization monitors for Win32. Dr. Dobb’s Journal of Software Tools, 26(12):46, 48, 50–52, 54, December 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_12/monitor.txt;   http:
//www.ddj.com/ftp/2001/
2001_12/monitor.zip.

Beddow:1991:MTC

[Bed91] A. J. M. Beddow. Multithreaded C functions. C Users Journal, 9(1):57–??, January 1991. ISSN 0898-9788.

Beebe:1998:BPA

[Bee98] Nelson H. F. Beebe. A bibliography of publications about multithreading. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 7, 1998. 15 pp. URL https:

//www.math.utah.edu/pub/
tex/bib/index-table-m.html#
multithreading. This report

is updated frequently.

Borkenhagen:2000:MPP

[BEKK00] J. M. Borkenhagen, R. J.

Eickemeyer, R. N. Kalla, and S. R. Kunkel. A multithreaded PowerPC processor for commercial servers. IBM Journal of Research and Development, 44(6):885–898, November 2000. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/446/borkenhagen.
html.

Berg:1996:HDT

[Ber96a] C. Berg. How do threads work and how can I create a general-purpose event? Dr. Dobb’s Journal of Software Tools, 21(11):111–115, 126– 127, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:JQH

[Ber96b] Cliff Berg. Java Q and A: How do threads work and how can I create a general-purpose event? Dr. Dobb’s Journal of Software Tools, 21(11):111–??, November 1, 1996. CODEN DDJOEB. ISSN 1044-789X.

Bettcher:1973:TSR

[Bet73] C. W. Bettcher. Thread standardization and relative cost. ACM SIGARCH Computer Architecture News, 2 (1):9, January 1973. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Bhowmik:2004:GCF

[BF04] Anasua Bhowmik and Manoj Franklin. A general compiler framework for speculative multithreaded processors. IEEE Transactions on Parallel and Distributed Systems, 15(8):713–724, August 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/dl/
trans/td/2004/08/l0713.
htm;   http://csdl.computer.
org/dl/trans/td/2004/08/
l0713.pdf.

Bahmann:2008:EFK

[BF08] Helge Bahmann and Konrad Froitzheim. Extending futex for kernel to user notification. Operating Systems Review, 42 (5):18–26, July 2008. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bhatotia:2015:ITL

[BFA^{+}15] Pramod Bhatotia, Pedro Fonseca, Umut A. Acar, Björn B. Brandenburg, and Rodrigo Rodrigues. iThreads: a threading library for parallel incremental computation. ACM SIGPLAN Notices, 50(4):645–659, April 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bagherzadeh:2020:ACB

[BFSK20] Mehdi Bagherzadeh, Nicholas Fireman, Anas Shawesh, and Raffi Khatchadourian. Actor concurrency bugs: a comprehensive study on symptoms, root causes, API usages, and differences. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):214:1–214:32, November 2020. URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3428282.

Bergan:2014:SEM

[BGC14] Tom Bergan, Dan Grossman, and Luis Ceze. Symbolic execution of multithreaded programs from arbitrary program contexts. ACM SIGPLAN Notices, 49(10):491–506, October 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baghsorkhi:2012:EPE

[BGDmWH12] Sara S. Baghsorkhi, Isaac Gelado, Matthieu Delahaye, and Wen mei W. Hwu. Efficient performance evaluation of memory hierarchy for highly multithreaded graphics processors. ACM SIG-PLAN Notices, 47(8):23–34, August 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Bic:1995:ATD

[BGG95] Lubomir Bic, Guang R. Gao, and Jean-Luc Gaudiot. Advanced topics in dataflow computing and multithreading. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-6541-6, 0-8186-6540-8 (paperback). x + 450 pp. LCCN QA76.9.A73A356 1994.

Burgess:2012:EFL

[BGH^{+}12] David Burgess, Edmund Gieske, James Holt, Thomas Hoy, and Gary Whisenhunt. e6500: Freescale’s lowpower, high-performance multithreaded embedded processor. IEEE Micro, 32(5):26– 36, September/October 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Buendgen:1994:MAT

[BGK94a] R. Buendgen, M. Goebel, and W. Kuechlin. Multithreaded AC term rewriting. In Hong [Hon94], pages 84–93. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Buendgen:1994:MTA

[BGK94b] R. Buendgen, M. Goebel, and W. Kuechlin. Multithreaded AC term rewriting. In Hong [Hon94], pages 84–93. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Bundgen:1994:FPC

[BGK94c] Reinhard Bündgen, Manfred Göbel, and Wolfgang Küchlin. A fine-grained parallel completion procedure. In ACM [ACM94c], pages 269–277. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

URL http://www.acm.org:
80/pubs/citations/proceedings/
issac/190347/p269-bundgen/

.

Bundgen:1996:SCM

[BGK96] Reinhard Bündgen, Manfred Göbel, and Wolfgang Küchlin. Strategy compliant multithreaded term completion. Journal of Symbolic Computation, 21(4/5/6):475–506 (or 475–505??), April, May & June 1996. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Parallel symbolic computation.

Blundell:2006:AGT

[BGP06] Colin Blundell, Dimitra Giannakopoulou, and Corina S. Păsăreanu. Assume-guarantee testing. ACM SIGSOFT Software Engineering Notes, 31 (2):1:1–1:??, March 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bednorz:1997:CDA

[BGZ97] M. Bednorz, A. Gwozdowski, and K. Zieliński. Contextual debugging and analysis of multithreaded applications. Concurrency: Practice and Experience, 9 (2):123–139, February 1997. CODEN CPEXEI. ISSN 1040-3108. URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?ID=
13852;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13852&PLACEBO=IE.pdf.

Byrd:1995:MPA

[BH95] G. T. Byrd and M. A. Holliday. Multithreaded processor architectures. IEEE Spectrum, 32(8):38–46, August 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bouchenak:2004:EIE

[BHK^{+}04] S. Bouchenak, D. Hagimont, S. Krakowiak, N. De Palma, and F. Boyer. Experiences implementing efficient Java thread serialization, mobility and persistence. Software— Practice and Experience, 34 (4):355–393, April 10, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bubeck:1995:DSC

[BHKR95] T. Bubeck, M. Hiller, W. Kuchlin, and W. Rosenstiel. Distributed symbolic computation with DTS. In Ferreira and Rolim [FR95], pages 231–248. ISBN 3-540-60321-2. LCCN QA76.642.I59 1995.

Barekas:2003:MAO

[BHP^{+}03] Vasileios K. Barekas, Panagiotis E. Hadjidoukas, Eleftherios D. Polychronopoulos, et al. A multiprogramming aware OpenMP implementation. Scientific Programming, 11(2):133–141, 2003. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bientinesi:2011:CFS

[BIK^{+}11] Paolo Bientinesi, Francisco D. Igual, Daniel Kressner, Matthias Petschow, and Enrique S. Quintana-Ortí. Condensed forms for the symmetric eigenvalue problem on multithreaded architectures. Concurrency and Computation: Practice and Experience, 23 (7):694–707, May 2011. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Birrell:1989:IPT

[Bir89] Andrew D. Birrell. An introduction to programming with threads. SRC reports 35, Digital Systems Research Center, Palo Alto, CA, USA, January 6, 1989. 35 pp.

Blumofe:1995:CEM

[BJK^{+}95] Robert D. Blumofe, Christopher F. Joerg, Bradley C. Kuszmaul, Charles E. Leiserson, Keith H. Randall, and Yuli Zhou. Cilk: an efficient multithreaded runtime system. ACM SIGPLAN Notices, 30(8):207–216, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blumofe:1996:CEM

[BJK^{+}96] Robert D. Blumofe, Christopher F. Joerg, Bradley C. Kuszmaul, Charles E. Leiserson, Keith H. Randall, and Yuli Zhou. Cilk: An efficient multithreaded runtime system. Journal of Parallel and Distributed Computing, 37(1):55–69, August 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0107/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0107/production/pdf.

Bhandarkar:1996:MPM

[BK96] M. A. Bhandarkar and L. V.

Kale. MICE: a prototype MPI implementation in Converse environment. In IEEE [IEE96], pages 26–31. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Beckert:2013:DLD

[BK13] Bernhard Beckert and Vladimir Klebanov. A Dynamic Logic for deductive verification of multi-threaded programs. Formal Aspects of Computing, 25(3): 405–437, May 2013. CO-DEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00165-012-0261-4.

Bond:2013:OCC

[BKC^{+}13] Michael D. Bond, Milind Kulkarni, Man Cao, Minjia Zhang, Meisam Fathi Salmi, Swarnendu Biswas, Aritra Sengupta, and Jipeng Huang.

OCTET: capturing and controlling cross-thread dependences efficiently. ACM SIG-PLAN Notices, 48(10):693– 712, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOP-SLA ’13 conference proceedings.

Basile:2006:ARM

[BKI06] Claudio Basile, Zbigniew Kalbarczyk, and Ravishankar K. Iyer. Active replication of multithreaded applications. IEEE Transactions on Parallel and Distributed Systems, 17(5):448–465, May 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /csdl.computer.org/comp/
                                        trans/td/2006/05/l0448s.
                                        pdf.

Bujanovic:2017:HBA

[BKK17] Zvonimir Bujanović, Lars Karlsson, and Daniel Kressner. A Householder-based algorithm for Hessenbergtriangular reduction. arxiv.org, ??(??):??, October 23, 2017.

                                        URL https://arxiv.org/
                                        abs/1710.08538.

Blumofe:1993:SES

[BL93] Robert D. Blumofe and Charles E. Leiserson. Spaceefficient scheduling of multithreaded computations. In ACM [ACM93b], pages 362– 371. ISBN 0-89791-591-7.

LCCN QA 76.6 A13 1993.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/167088/p362-blumofe/
p362-blumofe.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/167088/
p362-blumofe/. ACM order

no. 508930.

Blumofe:1994:SMC

[BL94] R. D. Blumofe and C. E.

Leiserson. Scheduling multithreaded computations by work stealing. In Goldwasser [Gol94], pages 356–368. CO-DEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Bianchini:1996:EPM

[BL96] Ricardo Bianchini and Beng-Hong Lim. Evaluating the performance of multithreading and prefetching in multiprocessors. Journal of Parallel and Distributed Computing, 37(1):83–97, August 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0109/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0109/production/pdf.

Blumofe:1998:SES

[BL98] Robert D. Blumofe and Charles E. Leiserson. Spaceefficient scheduling of multithreaded computations. SIAM Journal on Computing, 27 (1):202–229, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25947.

Blumofe:1999:SMC

[BL99] Robert D. Blumofe and Charles E. Leiserson. Scheduling multithreaded computations by work stealing. Journal of the ACM, 46(5):720– 748, September 1999. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/jacm/1999-46-5/
                                        p720-blumofe/.

Bordawekar:1997:EEH

[BLCD97] Rajesh Bordawekar, Steven Landherr, Don Capps, and Mark Davis. Experimental evaluation of the Hewlett– Packard Exemplar file system. ACM SIGMETRICS Performance Evaluation Review, 25(3):21–28, December 1997. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Broberg:2001:POU

[BLG01] Magnus Broberg, Lars Lundberg, and Håkan Grahn. Performance optimization using extended critical path analysis in multithreaded programs on multiprocessors. Journal of Parallel and Distributed Computing, 61(1):115–136, January 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1667;   http://www.
idealibrary.com/links/doi/
10.1006/jpdc.2000.1667/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1667/ref.

Blundell:2006:STM

[BLM06] C. Blundell, E. C. Lewis, and M. M. K. Martin. Subtleties of transactional memory atomicity semantics. IEEE Computer Architecture Letters, 5 (2):17, February 2006. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Bucker:2004:TUC

[BLPV04] H. M. Bucker, B. Lang, H. J.

Pflug, and A. Vehreschild. Threads in an undergraduate course: a Java example illuminating different multithreading approaches. Lecture Notes in Computer Science, 3044:882–891, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blumofe:1992:MSM

[Blu92] Robert D. (Robert David)

Blumofe. Managing storage for multithreaded computations. Thesis (M.S.), Massachusetts Institute of Technology, Laboratory for Computer Science, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1992. 83 pp. Also available as Report MIT/LCS/TR 552.

Blumofe:1995:EMP

[Blu95] Robert D. (Robert David) Blumofe. Executing multithreaded programs efficiently. Thesis (Ph.D.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1995. 145 pp.

Bolinger:1991:PSH

[BM91] D. Bolinger and S. Mangalat. Parallelizing signal handling and process management in OSF/1. In USENIX [USE91a], pages 105–122. LCCN QAX 27.

Baquero:1994:CAC

[BM94] Carlos Baquero and Francisco Moura. Concurrency annotations in C++. ACM SIG-PLAN Notices, 29(7):61–67, July 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergstra:2007:SCE

[BM07] J. A. Bergstra and C. A. Middelburg. Synchronous cooperation for explicit multithreading. Acta Informatica, 44(7–8):525–569, December 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0001-5903&volume=44&
issue=7&spage=525.

Berger:2000:HSMa

[BMBW00a] Emery D. Berger, Kathryn S.

McKinley, Robert D. Blumofe, and Paul R. Wilson. Hoard: a scalable memory allocator for multithreaded applications. ACM SIGARCH Computer Architecture News, 28(5):117– 128, December 2000. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Berger:2000:HSMb

[BMBW00b] Emery D. Berger, Kathryn S.

McKinley, Robert D. Blumofe, and Paul R. Wilson. Hoard: a scalable memory allocator for multithreaded applications. ACM SIG-PLAN Notices, 35(11):117– 128, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berger:2000:HSMc

[BMBW00c] Emery D. Berger, Kathryn S.

McKinley, Robert D. Blumofe, and Paul R. Wilson. Hoard: a scalable memory allocator for multithreaded applications. Operating Systems Review, 34(5): 117–128, December 2000. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Balkind:2016:OOS

[BMF^{+}16] Jonathan Balkind, Michael McKeown, Yaosheng Fu, Tri Nguyen, Yanqi Zhou, Alexey Lavrov, Mohammad Shahrad, Adi Fuchs, Samuel Payne, Xiaohua Liang, Matthew Matl, and David Wentzlaff. Open-Piton: an open source manycore research framework. Operating Systems Review, 50 (2):217–232, June 2016. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Balkind:2019:OOS

[BMF^{+}19] Jonathan Balkind, Michael McKeown, Yaosheng Fu, Tri Nguyen, Yanqi Zhou, Alexey Lavrov, Mohammad Shahrad, Adi Fuchs, Samuel Payne, Xiaohua Liang, Matthew Matl, and David Wentzlaff. Open-Piton: an open source hardware platform for your research. Communications of the ACM, 62(12):79–87, December 2019. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://cacm.
acm.org/magazines/2019/
12/241058/fulltext.

Bouge:1999:ECM

[BMN99] L. Bouge, J.-F. Mehaut, and R. Namyst. Efficient communications in multithreaded runtime systems. Lecture Notes in Computer Science, 1586:468–482, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baker:1994:EPP

[BMR94] T. P. Baker, Frank Mueller, and Viresh Rustagi. Experience with a prototype of the POSIX “minimal realtime system profile”. In IEEE [IEE94d], pages 12–17. ISBN 0-8186-5710-3. LCCN QA76.54.I173 1994.

Baumann:2021:CBV

[BMTZ21] Pascal Baumann, Rupak Majumdar, Ramanathan S. Thinniyam, and Georg Zetzsche. Context-bounded verification of liveness properties for multithreaded shared-memory programs. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 44:1–44:31, January 2021.

URL https://dl.acm.org/
doi/10.1145/3434325.

Briguglio:2003:PPM

[BMV03] Sergio Briguglio, Beniamino Di Martino, and Gregorio Vlad. A performance-prediction model for PIC applications on clusters of symmetric multiprocessors: Validation with hierarchical HPF + OpenMP implementation. Scientific Programming, 11(2):159–176, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Brunst:2001:GBP

[BNH01] Holger Brunst, Wolfgang E. Nagel, and Hans-Christian Hoppe. Group-based performance analysis for multithreaded SMP cluster applications. Lecture Notes in Computer Science, 2150:148– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2150/21500148.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2150/21500148.
pdf.

Burnim:2011:SCSa

[BNS11a] Jacob Burnim, George Necula, and Koushik Sen. Specifying and checking semantic atomicity for multithreaded programs. ACM SIGARCH Computer Architecture News, 39(1):79–90, March 2011. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Burnim:2011:SCSb

[BNS11b] Jacob Burnim, George Necula, and Koushik Sen. Specifying and checking semantic atomicity for multithreaded programs. ACM SIGPLAN Notices, 46(3):79–90, March 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burnim:2012:SCS

[BNS12] Jacob Burnim, George Necula, and Koushik Sen. Specifying and checking semantic atomicity for multithreaded programs. ACM SIGPLAN Notices, 47(4):79–90, April 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benson:1996:DMS

[BO96] G. D. Benson and R. A. Olsson. The design of microkernel support for the SR concurrent programming language. In Szymanski and Sinharoy [SS96], pages 227–240. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Bull:2001:MSO

[BO01] J. Mark Bull and Darragh O’Neill. A microbenchmark suite for OpenMP 2.0. ACM SIGARCH Computer Architecture News, 29(5):41– 48, December 2001. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Blandy:2017:PR

[BO17] Jim Blandy and Jason Orendorff. Programming Rust. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2017. ISBN 1-4919-2728-3 (paperback), 1-4919-2727-5, 1-4919-2723-2 (e-book), 1-4919-2725-9 (e-book). xx + 598 pp. LCCN QA76.73.R88 B53 2017. URL http://

proquest.safaribooksonline.
com/9781491927274.

Boehm:2005:TCI

[Boe05] Hans-J. Boehm. Threads cannot be implemented as a library. ACM SIGPLAN Notices, 40(6):261–268, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bond:2013:GDG

[Bon13] Michael Bond. GPUDet: a deterministic GPU architecture. ACM SIGPLAN Notices, 48 (4):1–12, April 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boothe:1993:EMC

[Boo93] Bob Boothe. Evaluation of multithreading and caching in large shared memory parallel computers. Thesis (Ph.D.), University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, July 1993. ix + 169 pp. Also available as Report UCB/CSD 93/766.

Brinkschulte:2005:ICA

[BP05] U. Brinkschulte and M. Pacher.

Implementing control algorithms within a multithreaded Java microcontroller. Lecture Notes in Computer Science, 3432:33–49, 2005. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brais:2019:AAM

[BP19] Hadi Brais and Preeti Ranjan Panda. Alleria: an advanced memory access profiling framework. ACM Transactions on Embedded Computing Systems, 18(5s):81:1– 81:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358193.

Boehm:2007:MCC

[BPL07] Hans Boehm, Bill Pugh, and Doug Lea. Multithreading in C and C++. ;login: the USENIX Association newsletter, 32(1):??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2007-volume-32-
number-1/multithreading-
c-and-c.

Boroday:2005:DAJ

[BPSH05] S. Boroday, A. Petrenko, J. Singh, and H. Hallal. Dynamic analysis of Java applications for multithreaded antipatterns. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Boothe:1992:IMT

[BR92] Bob Boothe and Abhiram Ranade. Improved multithreading techniques for hiding communication latency in multiprocessors. ACM SIGARCH Computer Architecture News, 20(2):214–223, May 1992. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Bogdanas:2015:KJC

[BR15] Denis Bogdanas and Grigore Rosu. K-Java: a complete semantics of Java. ACM SIG-PLAN Notices, 50(1):445–456, January 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bramley:1997:TNRb

[Bra97]  Randall Bramley.  Tech-

nology news & reviews: Chemkin software; OpenMP Fortran Standard; ODE toolbox for Matlab; Java products; Scientific WorkPlace 3.0. IEEE Computational Science & Engineering, 4 (4):75–78, October/December 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1997/pdf/c4075.
pdf.

Bershad:1992:FME

[BRE92] Brian N. Bershad, David D.

Redell, and John R. Ellis. Fast mutual exclusion for uniprocessors. ACM SIGPLAN Notices, 27(9):223–233, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/asplos/143365/
p223-bershad/.

Brebner:2002:MLC

[Bre02]  Gordon Brebner.  Multi-

threading for logic-centric systems. Lecture Notes in Computer Science, 2438:5– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2438/24380005.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2438/24380005.
pdf.

Briot:1989:OAS

[Bri89] Jean-Pierre Briot. From objects to actors: study of a limited symbiosis in Smalltalk-80. ACM SIGPLAN Notices, 24(4):69–72, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/plan/67386/
                                        p69-briot/.

Brightwell:2003:DIP

[BRM03] Ron Brightwell, Rolf Riesen, and Arthur B. Maccabe. Design, implementation, and performance of MPI on Portals 3.0. The International Journal of High Performance Computing Applications, 17 (1):7–20, Spring 2003. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Barthe:2010:SMP

[BRRS10] Gilles Barthe, Tamara Rezk, Alejandro Russo, and Andrei Sabelfeld. Security of multithreaded programs by compilation. ACM Transactions on Information and System Security, 13(3):21:1–21:??, July 2010. CODEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

Bellosa:1996:PIL

[BS96] Frank Bellosa and Martin Steckermeier. The performance implications of locality information usage in shared-memory multiprocessors. Journal of Parallel and Distributed Computing, 37(1):113–121, August 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

                                        http://www.idealibrary.

com/links/doi/10.1006/jpdc.
1996.0112/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0112/production/pdf.

Broadman:1999:ECM

[BS99] Allen Broadman and Eric Shaw. Executing a class member in its own thread. C/C++ Users Journal, 17(12):??, December 1999. CODEN CCU-JEX. ISSN 1075-2838.

Boussinot:2000:JTS

[BS00] Frédéric Boussinot and Jean-Ferdy Susini. Java threads and SugarCubes. Software— Practice and Experience, 30 (5):545–566, April 25, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/71004433/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=71004433&PLACEBO=IE.
pdf.

Bacon:2006:BFL

[BS06] D. F. Bacon and X. Shen.

Braids and fibers: Language constructs with architectural support for adaptive responses to memory latencies. IBM Journal of Research and Development, 50 (2/3):209–??, March /May 2006. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/502/bacon.html.

Bokhari:2010:EPM

[BS10a] Shahid Bokhari and Joel Saltz. Exploring the performance of massively multithreaded architectures. Concurrency and Computation: Practice and Experience, 22 (5):588–616, April 10, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Burnim:2010:ACD

[BS10b] Jacob Burnim and Koushik Sen. Asserting and checking determinism for multithreaded programs. Communications of the ACM, 53 (6):97–105, June 2010. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bartolini:2014:AFG

[BSSS14] Davide B. Bartolini, Filippo Sironi, Donatella Sciuto, and Marco D. Santambrogio. Automated fine-grained CPU provisioning for virtual machines. ACM Transactions on Architecture and Code Optimization, 11(3):27:1–27:??, October 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Boisvert:2001:ASS

[BT01] Ronald F. Boisvert and Ping Tak Peter Tang, editors. The architecture of scientific software: IFIP TC2/WG2.5 Working Conference on the Architecture of Scientific Software, October 2–4, 2000, Ottawa, Canada, volume 60 of IFIP. Kluwer Academic Publishers, Dordrecht, The Netherlands; Boston, MA, USA, 2001. ISBN 0-7923-7339-1. LCCN QA76.758 .I345 2000.

Brunett:1998:IET

[BTE98] Sharon M. Brunett, John Thornley, and Marrq Ellenbecker. An initial evaluation of the Tera multithreaded architecture and programming system using the C3I parallel benchmark suite. In ACM [ACM98d], page ?? ISBN ???? LCCN ???? URL http:/

/www.supercomp.org/sc98/
TechPapers/sc98_FullAbstracts/
Brunett1063/Index.htm.

Bouksiaa:2019:UDE

[BTL^{+}19] M. S. M. Bouksiaa, F. Trahay, A. Lescouet, G. Voron, R. Dulong, A. Guermouche, É. Brunet, and G. Thomas. Using differential execution analysis to identify thread interference. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2866–2878, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Butenhof:1997:PPT

[But97] David R. Butenhof. Programming with POSIX threads. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-63392-2. xviii + 381 pp. LCCN QA76.76.T55B88 1997. US$31.95. URL http://www.

                                        amazon.com/exec/obidos/
                                        ASIN/0201633922/ref=sim_
                                        books/002-4892305-5599452.

Buttari:2013:FGM

[But13] Alfredo Buttari. Fine-grained multithreading for the multifrontal QR factorization of sparse matrices. SIAM Journal on Scientific Computing, 35(4):C323–C345, ???? 2013. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Butcher:2014:SCM

[But14]  Paul N. Butcher.  Seven

concurrency models in seven weeks: when threads unravel. The Pragmatic Programmers. The Pragmatic Bookshelf, Dallas, TX, USA, 2014. ISBN 1-937785-65-3 (paperback), 1-941222-27-7 (e-book). xiii + 275 pp. LCCN QA76.642 .B88 2014. URL http://

                                        proquest.safaribooksonline.
                                        com/?fpi=9781941222737.

Bik:1997:JPJ

[BVG97] Aart J. C. Bik, Juan E. Villacis, and Dennis B. Gannon. javar: a prototype Java restructuring compiler. Concurrency: Practice and Experience, 9(11): 1181–1191, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?ID=
13819;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13819&PLACEBO=IE.pdf.

Special Issue: Java for computational science and engineering — simulation and modeling II.

Barnes:2009:XBA

[BVL09] Christopher Barnes, Pranav Vaidya, and Jaehwan John Lee. An XML-based ADL framework for automatic generation of multithreaded computer architecture simulators. IEEE Computer Architecture Letters, 8(1):13–16, January/ June 2009. ISSN 1556-6056 (print), 1556-6064 (electronic).

Bajczi:2019:WMP

[BVM19] Levente Bajczi, András Vörös, and Vince Molnár. Will my program break on this faulty processor?: Formal analysis of hardware fault activations in concurrent embedded software. ACM Transactions on Embedded Computing Systems, 18(5s):89:1–89:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3358238.

Bonizzoni:2019:MMB

[BVP^{+}19] Paola Bonizzoni, Gianluca Della Vedova, Yuri Pirola, Marco Previtali, and Raffaella Rizzi.

Multithread multistring Burrows– Wheeler transform and longest common prefix array. Journal of Computational Biology, 26(9):948–961, September 2019. CODEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (electronic). URL

https://www.liebertpub.
com/doi/abs/10.1089/cmb.
2018.0230;   https://www.
liebertpub.com/doi/pdf/
10.1089/cmb.2018.0230.

Beveridge:1997:MAW

[BW97] Jim Beveridge and Robert Wiener. Multithreading applications in Win32: the complete guide to threads. Addison-Wesley Developers Press, Reading, MA, USA, 1997. ISBN 0-201-44234-5 (pb) 0-201-18385-4 (CD-ROM). xviii + 368 pp. LCCN QA76.76.O63 B478 1997.

Bai:2015:SPA

[BWDZ15] Xiuxiu Bai, Endong Wang, Xiaoshe Dong, and Xingjun Zhang. A scalability prediction approach for multithreaded applications on manycore processors. The Journal of Supercomputing, 71(11):4072–4094, November 2015. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11227-015-1505-x.

Basharahil:2005:DSA

[BWXF05] Ramzi Basharahil, Brian Wims, Cheng-Zhong Xu, and Song Fu. Distributed shared arrays: An integration of message passing and multithreading on SMP clusters. The Journal of Supercomputing, 31(2):161–184, February 2005. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=31&
issue=2&spage=161.

Berger:2009:GSM

[BYLN09] Emery D. Berger, Ting Yang, Tongping Liu, and Gene Novark. Grace: safe multithreaded programming for C/ C++. ACM SIGPLAN Notices, 44(10):81–96, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benaya:2007:UTA

[BZ07] Tamar Benaya and Ela Zur.

Understanding threads in an advanced Java course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):323, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Criswell:2020:SPC

[CA20] K. Criswell and T. Adegbija. A survey of phase classification techniques for characterizing variable application behavior. IEEE Transactions on Parallel and Distributed Systems, 31(1):224–236, January 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Calcote:1997:TPS

[Cal97] John Calcote. Thread pools and server performance. Dr. Dobb’s Journal of Software Tools, 22(7):60–??, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Calkins:2000:ITT

[Cal00] Charles Calkins. Integrating threads with template classes. C/C++ Users Journal, 18(5): 32–??, May 2000. CODEN CCUJEX. ISSN 1075-2838.

Callaway:2002:VTR

[Cal02] John Callaway. Visualization of threads in a running Java program. Thesis (M.S.), University of California, Santa Cruz, Santa Cruz, CA, USA, 2002.

Caromel:1989:GMC

[Car89a] Denis Caromel. A general model for concurrent and distributed object-oriented programming. ACM SIGPLAN Notices, 24(4):102–104, April 1989. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/plan/67386/
p102-caromel/.

CarrerasVaquer:1989:APE

[Car89b] Carlos Carreras Vaquer. Architecture and performance evaluation of a multithreaded cache design. Thesis (M.S. in Engineering), University of Texas at Austin, Austin, TX, USA, 1989. xii + 108 pp.

Campanoni:2008:PDC

[CAR08] Simone Campanoni, Giovanni Agosta, and Stefano Crespi Reghizzi. A parallel dynamic compiler for CIL bytecode. ACM SIGPLAN Notices, 43(4):11–20, April 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Catano:2014:CSL

[CASA14] Néstor Cataño, Ijaz Ahmed, Radu I. Siminiceanu, and Jonathan Aldrich. A case study on the lightweight verification of a multi-threaded task server. Science of Computer Programming, 80(??): 169–187, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167642313000178.

Catanzaro:1994:MSA

[Cat94] Ben J. Catanzaro. Multiprocessor system architectures: a technical survey of multiprocessor/multithreaded systems using SPARC, multilevel bus architectures and Solaris (SunOS). P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-089137-1. xxxii + 493 pp. LCCN QA76.5.C3864 1994.

Cazals:2002:NID

[Caz02] Frédéric Cazals. Non-intrusive debugging and incremental visualization with the geometric stethoscope. Journal of Graphics Tools: JGT, 7(2):27–40, 2002. CODEN JGTOFD. ISSN 1086-7651.

                                        URL http://www.acm.org/
                                        jgt/papers/Cazals02/.

Caswell:1989:IMD

[CB89] Deborah L. Caswell and David L. Black. Implementing a Mach debugger for multithreaded applications. Research paper CMU-CS-89-154, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, November 1989. 13 pp. To appear in the Conference Proceedings of Winter 1990 USENIX Technical Conference and Exhibition, Washington, DC, January, 1990.

Caswell:1990:IMD

[CB90] D. Caswell and D. Black. Implementing a Mach debugger for multithreaded applications. In Anonymous [Ano90], pages 25–39.

Creech:2016:TSS

[CB16] Timothy Creech and Rajeev Barua. Transparently space sharing a multicore among multiple processes. ACM Transactions on Parallel Computing (TOPC), 3(3): 17:1–17:??, December 2016. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Coons:2010:GEU

[CBM10] Katherine E. Coons, Sebastian Burckhardt, and Madanlal Musuvathi. GAMBIT: effective unit testing for concurrency libraries. ACM SIG-PLAN Notices, 45(5):15–24, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cui:2000:MPC

[CBN^{+}00] J. Cui, J. L. Bordim, K. Nakano, T. Hayashi, and N. Ishii. Multithreaded parallel computer model with performance evaluation. Lecture Notes in Computer Science, 1800:155–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1800/18000155.
htm;   http://link.springer-

                                        ny.com/link/service/series/
                                        0558/papers/1800/18000155.
                                        pdf.

Chiueh:1991:MTV

[cC91]  Tzi cker Chiueh.  Multi-

threaded vectorization. ACM SIGARCH Computer Architecture News, 19(3):352–361, May 1991. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Chang:2004:TSP

[CC04] B. M. Chang and J. D. Choi. Thread-sensitive pointsto analysis for multithreaded Java programs. Lecture Notes in Computer Science, 3280:945–954, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:2014:MSD

[CC14] Y. Cai and W. K. Chan. Magiclock: Scalable detection of potential deadlocks in large-scale multithreaded programs. IEEE Transactions on Software Engineering, 40(3):266–281, March 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6718069.

Chen:2018:ESE

[CC18] Kuan-Chung Chen and Chung-Ho Chen. Enabling SIMT execution model on homogeneous multi-core system. ACM Transactions on Architecture and Code Optimization, 15(1): 6:1–6:??, April 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Chen:2012:MLS

[CCC12] Chih-Yuan Chen, Jhong-Yi Ciou, and Rong-Guey Chang. Multi-level simultaneous multithreading scheduling to reduce the temperature of register files. Concurrency and Computation: Practice and Experience, 24(12):1296–1316, August 25, 2012. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2011:MJP

[CCH11] Kuo-Yi Chen, J. Morris Chang, and Ting-Wei Hou. Multithreading in Java: Performance and scalability on multicore systems. IEEE Transactions on Computers, 60(11):1521–1534, November 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=5661769.

Chen:2016:TMR

[CCK^{+}16] Kuan-Hsun Chen, Jian-Jia Chen, Florian Kriebel, Semeen Rehman, Muhammad Shafique, and Jörg Henkel. Task mapping for redundant multithreading in multi-cores with reliability and performance heterogeneity. IEEE Transactions on Computers, 65(11):3441–3455, November 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chinya:2011:BDP

[CCW^{+}11] Gautham N. Chinya, Jamison D. Collins, Perry H. Wang, Hong Jiang, Guei-Yuan Lueh, Thomas A. Piazza, and Hong Wang. Bothnia: a dual-personality extension to the Intel integrated graphics driver. Operating Systems Review, 45(1): 11–20, January 2011. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:2017:IGP

[CCWY17] Li-Jhan Chen, Hsiang-Yun Cheng, Po-Han Wang, and Chia-Lin Yang. Improving GPGPU performance via cache locality aware thread block scheduling. IEEE Computer Architecture Letters, 16 (2):127–131, July/December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Chetlur:2010:SWM

[CDD^{+}10] M. Chetlur, U. Devi, P. Dutta, P. Gupta, L. Chen, Z. Zhu, S. Kalyanaraman, and Y. Lin. A software WiMAX medium access control layer using massively multithreaded processors. IBM Journal of Research and Development, 54(1):??, ???? 2010. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/abstracts/rd/541/
chetlur-dutta.html.

Chandra:2001:PPO

[CDK^{+}01] Rohit Chandra, Leonardo Dagum, David Kohr, Dror Maydan, Jeff McDonald, and Ramesh Menon. Parallel Programming in OpenMP. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2001. ISBN 1-55860-671-8. xvi + 230 pp. LCCN QA76.642 .P38 2001. US$39.95.

URL http://www.mkp.com/
books_catalog/catalog.asp?
ISBN=1-55860-671-8.

Chung:2013:LBD

[CDL13] Eric S. Chung, John D. Davis, and Jaewon Lee. LINQits: big data on little clients. ACM SIGARCH Computer Architecture News, 41(3): 261–272, June 2013. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ICSA ’13 conference proceedings.

ChassindeKergommeaux:2001:PEE

[CdOS01] Jacques Chassin de Kergommeaux and Benhur de Oliveira Stein. Pajé: An extensible environment for visualizing multithreaded programs executions. Lecture Notes in Computer Science, 1900:133–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/1900/19000133.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1900/19000133.
                                        pdf.

Conoci:2021:PCP

[CDP^{+}21] Stefano Conoci, Pierangelo Di Sanzo, Alessandro Pellegrini, Bruno Ciciani, and Francesco Quaglia. On power capping and performance optimization of multithreaded applications. Concurrency and Computation: Practice and Experience, 33(13):e6205:1–e6205:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2022:EMA

[CFC^{+}22] Jianyi Cheng, Shane T. Fleming, Yu Ting Chen, Jason Anderson, John Wickerson, and George A. Constantinides. Efficient memory arbitration in high-level synthesis from multi-threaded code. IEEE Transactions on Computers, 71(4):933–946, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Catalyurek:2012:GCA

[ÇFG^{+}12] Ümit V. Çatalyürek, John Feo, Assefaw H. Gebremedhin, Mahantesh Halappanavar, and Alex Pothen. Graph coloring algorithms for multicore and massively multithreaded architectures. Parallel Computing, 38(10–11): 576–594, October/November 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167819112000592.

Canetti:1991:PCP

[CFK^{+}91] R. Canetti, L. P. Fertig, S. A.

Kravitz, D. Malki, R. Y. Pinter, S. Porat, and A. Teperman. The parallel C (pC) programming language. IBM Journal of Research and Development, 35(5/6):727–741, September/November 1991. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Cerin:2006:MSS

[CGK06] Christophe Cérin, Jean-Luc Gaudiot, and Michel Koskas. A multithreaded SQL service. Parallel Processing Letters, 16 (2):245–259, June 2006. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Culler:1992:AMMa

[CGL92a] David E. Culler, Michial Gunter, and James C. Lee. Analysis of multithreaded microprocessors under multiprogramming. Report UCB/CSD 92/687, University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, May 1992. 17 pp.

Culler:1992:AMMb

[CGL92b] David E. Culler, Michial Gunter, and James C. Lee. Analysis of multithreaded microprocessors under multiprogramming. ACM SIGARCH Computer Architecture News, 20(2):438, May 1992. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Cattaneo:1992:ACT

[CGR92] G. Cattaneo, G. Di Giore, and M. Ruotolo. Another C threads library. ACM SIG-PLAN Notices, 27(12):81–90, December 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Castello:2020:ATL

[CGS^{+}20] A. Castelló, R. M. Gual, S. Seo, P. Balaji, E. S. Quintana-Ortí, and A. J. Peña. Analysis of threading libraries for high performance computing. IEEE Transactions on Computers, 69(9): 1279–1292, September 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Culler:1993:TCC

[CGSV93] David E. Culler, Seth Copen Goldstein, Klaus Erik Schauser, and Thorsten Von Eicken.

TAM – a compiler controlled Threaded Abstract Machine. Journal of Parallel and Distributed Computing, 18(3):347–370, July 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1070/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1070/production/pdf.

Chong:1995:PAF

[CH95] Yong-Kim Chong and Kai Hwang. Performance analysis of four memory consistency models for multithreaded multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 6(10):1085–1099, October 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//www.computer.org/tpds/
td1995/l1085abs.htm.

Chaudhuri:2004:SAN

[CH04] Mainak Chaudhuri and Mark Heinrich. SMTp: An Architecture for Next-generation Scalable Multi-threading. ACM SIGARCH Computer Architecture News, 32(2):124, March 2004. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Chaudhry:2002:PTS

[Cha02]  Puneesh Chaudhry.  A

per-thread singleton class. C/C++ Users Journal, 20(5): 14–??, May 2002. CODEN CCUJEX. ISSN 1075-2838.

Chapman:2005:SMP

[Cha05] Barbara M. Chapman, editor. Shared memory parallel programming with OpenMP: 5th International Workshop on OpenMP Applications and Tools, WOMPAT 2004, Houston, TX, USA, May 17–18, 2004: Revised selected papers, volume 3349 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CO-DEN LNCSD9. ISBN 3-540-24560-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 .A1 L42 NO.3349. URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=issue&
                                        issn=0302-9743&volume=3349;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=volume&
                                        id=doi:10.1007/b105895.

Chen:2003:CSS

[CHH^{+}03] Peng-Sheng Chen, Ming-Yu Hung, Yuan-Shin Hwang, Roy Dz-Ching Ju, and Jenq Kuen Lee. Compiler support for speculative multithreading architecture with probabilistic points-to analysis. ACM SIG-PLAN Notices, pages 25–36, 2003. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chlipala:2015:NIM

[Chl15a] Adam Chlipala. From network interface to multithreaded Web applications: a case study in modular program verification. ACM SIGPLAN Notices, 50(1):609–622, January 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chlipala:2015:UWS

[Chl15b] Adam Chlipala. Ur/Web:

a simple model for programming the Web. ACM SIG-PLAN Notices, 50(1):153–165, January 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chowdhury:1992:PEA

[Cho92] Indranil Chowdhury. Performance evaluation and architecture of an instruction cache for multithreaded RISC processor. Thesis (M.S. in Engineering), University of Texas at Austin, Austin, TX, USA, 1992. x + 93 pp.

Chong:1993:EMC

[Cho93] Yong-Kim Chong. Effects of memory consistency models on multithreaded multiprocessor performance. Thesis (M.S.), University of Southern California, Los Angeles, CA, USA, 1993. viii + 89 pp.

Chrisochoides:1995:MMDa

[Chr95a] Nikos Chrisochoides. Multithreaded model for dynamic load balancing parallel adaptive PDE computations. Technical report CTC95, TR221, Cornell Theory Center, Cornell University, Ithaca, NY, USA, 1995. 23 pp.

Chrisochoides:1995:MMDb

[Chr95b] Nikos Chrisochoides. Multithreaded model for dynamic load balancing parallel adaptive PDE computations. NASA contractor report 198244; ICASE report 95-83., Institute for Computer Applications in Science and Engineering NASA Langley Research Center, Hampton, VA, USA, November 1995. i + 23 + i pp. To appear in Applied Numerical Mathematics Journal.

Chrisochoides:1996:MMD

[Chr96] Nikos Chrisochoides. Multithreaded model for the dynamic load-balancing of parallel adaptive PDE computations. Applied Numerical Mathematics: Transactions of IMACS, 20(4):349–365, June 3, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
apnum/cas_sub/browse/browse.

cgi?year=1996&volume=20&
issue=4&aid=652.

Christiaens:2001:JRR

[Chr01] Mark Christiaens. JaRec:

Record/replay for multithreaded Java programs. In USENIX [USE01], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/
JVM_wips/S07.pdf.

Catalan:2017:TEM

[CIM^{+}17] Sandra Catalán, Francisco D.

Igual, Rafael Mayo, Rafael Rodríguez-Sánchez, and Enrique S. Quintana-Ortí. Time and energy modeling of a high-performance multithreaded Cholesky factorization. The Journal of Supercomputing, 73(1):139– 151, January 2017. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ching:1991:EAP

[CJ91] W.-M. Ching and D. Ju. Execution of automatically parallelized APL programs on RP3. IBM Journal of Research and Development, 35(5/6):767– 777, September/November 1991. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Curran:2015:IZM

[CJB^{+}15] B. W. Curran, C. Jacobi, J. J. Bonanno, D. A. Schroter, K. J. Alexander, A. Puranik, and M. M. Helms. The IBM z13 multithreaded microprocessor. IBM Journal of Research and Development, 59(4–5):1:1–1:13, July/September 2015. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Cejtin:1995:HOD

[CJK95] Henry Cejtin, Suresh Jagannathan, and Richard Kelsey. Higher-order distributed objects. ACM Transactions on Programming Languages and Systems, 17(5):704–739, September 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0164-0925/
213986.html.

Cai:2015:ADB

[CJW^{+}15] Yan Cai, Changjiang Jia, Shangru Wu, Ke Zhai, and Wing Kwong Chan. ASN: A dynamic barrier-based approach to confirmation of deadlocks from warnings for large-scale multithreaded programs. IEEE Transactions on Parallel and Distributed Systems, 26(1):13– 23, January 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://www.
computer.org/csdl/trans/
td/2015/01/06747310-abs.
html.

Carter:1994:HSF

[CKD94] Nicholas P. Carter, Stephen W.

Keckler, and William J. Dally. Hardware support for fast capability-based addressing. ACM SIGPLAN Notices, 29(11):319–327, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/asplos/195473/
p319-carter/.

Cenciarelli:1997:SMJ

[CKRW97a] P. Cenciarelli, A. Knapp, B. Reus, and M. Wirsing. From sequential to multithreaded Java: An eventbased operational semantics. Lecture Notes in Computer Science, 1349:75–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cenciarelli:1997:SMT

[CKRW97b] P. Cenciarelli, A. Knapp, B. Reus, and M. Wirsing. From sequential to multithreaded Java: An eventbased operational semantics. Lecture Notes in Computer Science, 1349:75–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cenciarelli:1999:EBS

[CKRW99] P. Cenciarelli, A. Knapp, B. Reus, and M. Wirsing. An event-based structural operational semantics of multi-threaded Java. Lecture Notes in Computer Science, 1523:157–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clements:2012:SAS

[CKZ12] Austin T. Clements, M. Frans Kaashoek, and Nickolai Zeldovich. Scalable address spaces using RCU balanced trees. ACM SIGARCH Computer Architecture News, 40 (1):199–210, March 2012. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ASPLOS ’12 conference proceedings.

Chaudhry:1994:CMP

[CL94] Ghulam Chaudhry and Xuechang Li. A case for the multithreaded processor architecture. ACM SIGARCH Computer Architecture News, 22 (4):55–59, September 1994. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Caudal:1995:DEM

[CL95] F. Caudal and B. Lecussan. Design and evaluation of a multi-threaded architecture for parallel graph reduction. Lecture Notes in Computer Science, 964:411– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:2000:SCP

[CL00] Sung-Eun Choi and E. Christopher Lewis. A study of common pitfalls in simple multithreaded programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1): 325–329, March 2000. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chase:1994:SPS

[CLFL94] Jeffrey S. Chase, Henry M.

Levy, Michael J. Feeley, and Edward D. Lazowska. Sharing and protection in a singleaddress-space operating system. ACM Transactions on Computer Systems, 12(4): 271–307, November 1994. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tocs/1994-12-4/
p271-chase/.

Choi:2002:EPD

[CLL^{+}02] Jong-Deok Choi, Keunwoo Lee, Alexey Loginov, Robert O’Callahan, Vivek Sarkar, and Manu Sridharan. Efficient and precise datarace detection for multithreaded objectoriented programs. ACM SIG-PLAN Notices, 37(5):258–269, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cormen:2009:IA

[CLRS09] Thomas H. Cormen, Charles Eric Leiserson, Ronald L. Rivest, and Clifford Stein, editors. Introduction to algorithms. MIT Press, Cambridge, MA, USA, third edition, 2009. ISBN 0-262-03384-4 (hardcover), 0-262-53305-7 (paperback). xix + 1292 pp. LCCN QA76.6 .C662 2009.

Chapman:1998:OHI

[CM98] B. Chapman and P. Mehrotra. OpenMP and HPF: Integrating two paradigms. Lecture Notes in Computer Science, 1470:650–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cugu:2020:PMS

[ÇM20] Ilke Çugu and Murat Manguoglu. A parallel multithreaded sparse triangular linear system solver. Computers and Mathematics with Applications, 80(2):371–385, July 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122119304602.

Curtis-Maury:2008:PBP

[CMBAN08] Matthew Curtis-Maury, Filip Blagojevic, Christos D. Antonopoulos, and Dimitrios S. Nikolopoulos. Prediction-based powerperformance adaptation of multithreaded scientific codes.

IEEE Transactions on Parallel and Distributed Systems, 19(10):1396–1410, October 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cain:2013:RAS

[CMF^{+}13] Harold W. Cain, Maged M.

Michael, Brad Frey, Cathy May, Derek Williams, and Hung Le. Robust architectural support for transactional memory in the Power architecture. ACM SIGARCH Computer Architecture News, 41 (3):225–236, June 2013. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ICSA ’13 conference proceedings.

Cahir:2000:PMM

[CMK00] Margaret Cahir, Robert Moench, and Alice E. Koniges. Programming models and methods. In Koniges [Kon00], chapter 3, pages 27–54. ISBN 1-55860-540-1. LCCN QA76.58 .I483 2000. Discusses PVM, MPI, SHMEM, High-Performance Fortran, and POSIX threads.

Cahoon:2000:EPD

[CML00] Brendon Cahoon, Kathryn S.

McKinley, and Zhihong Lu. Evaluating the performance of distributed architectures for information retrieval using a variety of workloads. ACM Transactions on Information Systems, 18(1):1–43, January 2000. CODEN ATISET. ISSN 1046-8188. URL http://www.

                                        acm.org/pubs/citations/
                                        journals/tois/2000-18-1/
                                        p1-cahoon/.

Carroll:2019:ACM

[CmL19a] Shane Carroll and Wei ming Lin. Applied on-chip machine learning for dynamic resource control in multithreaded processors. Parallel Processing Letters, 29(03):??, September 2019. ISSN 0129-6264 (print), 1793-642X (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0129626419500130.

Carroll:2019:RRT

[CmL19b] Shane Carroll and Wei ming Lin. Round robin thread selection optimization in multithreaded processors. Parallel Processing Letters, 29(01): ??, March 2019. ISSN 0129-6264 (print), 1793-642X (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0129626419500038.

Carroll:2021:ELT

[CmL21] Shane Carroll and Wei ming Lin. Exploiting long-term temporal cache access patterns for LRU insertion prioritization. Parallel Processing Letters, 31(02):??, June 2021. ISSN 0129-6264 (print), 1793-642X (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0129626421500109.

Carr:2003:TPT

[CMS03] Steve Carr, Jean Mayo, and Ching-Kuang Shene. Thread-Mentor: a pedagogical tool for multithreaded programming. ACM Journal on Educational Resources in Computing (JERIC), 3(1):1–30, March 2003. CODEN ???? ISSN 1531-4278.

Chen:2010:CCM

[CMX10] Changno Chen, Marc Moreno Maza, and Yuzhen Xie. Cache complexity and multicore implementation for univariate real root isolation. ACM Communications in Computer Algebra, 44(3):97–98, September 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Che:2014:ALM

[CN14] Hao Che and Minh Nguyen.

Amdahl’s Law for multithreaded multicore processors. Journal of Parallel and Distributed Computing, 74(10):3056–3069, October 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0743731514001142.

Cabodi:2013:TBM

[CNQ13] Gianpiero Cabodi, Sergio Nocco, and Stefano Quer. Thread-based multi-engine model checking for multicore platforms. ACM Transactions on Design Automation of Electronic Systems., 18(3): 36:1–36:??, July 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chuang:2006:UPB

[CNV^{+}06] Weihaw Chuang, Satish Narayanasamy, Ganesh Venkatesh, Jack Sampson, Michael Van Biesbrouck, Gilles Pokam, Brad Calder, and Osvaldo Colavin. Unbounded page-based transactional memory. ACM SIG-PLAN Notices, 41(11):347– 358, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cui:2017:MTA

[CNZS17] Huanqing Cui, Jian Niu, Chuanai Zhou, and Minglei Shu. A multi-threading algorithm to detect and remove cycles in vertex- and arc-weighted digraph. Algorithms (Basel), 10(4), December 2017. CODEN AL-GOCH. ISSN 1999-4893 (elec-

                                        tronic). URL https://www.
                                        mdpi.com/1999-4893/10/4/
                                        115.

Colvin:1990:CTS

[Col90a] Gregory Colvin. CUG306 thread and synapsys. C Users Journal, 8(3):131–??, March 1990. ISSN 0898-9788.

Colvin:1990:MLT

[Col90b] Gregory Colvin. Multitasking with lightweight threads.

C Users Journal, 8(3):55–??, March 1990. ISSN 0898-9788.

Coorg:1995:PNS

[Coo95] S. R. Coorg. Partitioning nonstrict functional languages for multi-threaded code generation. Lecture Notes in Computer Science, 983:82–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cook:2002:REJ

[Coo02]  Jonathan J. Cook.  Re-

verse execution of Java bytecode. The Computer Journal, 45(6):608–619, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_
journal/hdb/Volume_45/Issue_
06/450608.sgm.abs.html;
http://www3.oup.co.uk/
computer_journal/hdb/Volume_
45/Issue_06/pdf/450608.
pdf.

Corbett:2000:USA

[Cor00]  James C. Corbett.  Us-

ing shape analysis to reduce finite-state models of concurrent Java programs. ACM Transactions on Software Engineering and Methodology, 9(1):51–93, January 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tosem/
2000-9-1/p51-corbett/p51-

corbett.pdf;   http://www.
acm.org/pubs/citations/
journals/tosem/2000-9-1/
p51-corbett/.

Choi:2008:ABP

[CPT08] Bumyong Choi, Leo Porter, and Dean M. Tullsen. Accurate branch prediction for short threads. Operating Systems Review, 42(2):125– 134, March 2008. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Clark:2002:AMT

[CR02] Keith Clark and Peter J. Robinson. Agents as multithreaded logical objects. Lecture Notes in Computer Science, 2407:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2407/24070033.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2407/24070033.
pdf.

Cappello:1999:PNB

[CRE99] F. Cappello, O. Richard, and D. Etiemble. Performance of the NAS benchmarks on a cluster of SMP PCs using a parallelization of the MPI programs with OpenMP. Lecture Notes in Computer Science, 1662:339–350, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Criscolo:1998:JQH

[Cri98a]  Mike Criscolo.  Java Q

and A: How do I queue Java threads? Dr. Dobb’s Journal of Software Tools, 23(10):127–129, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
1998/1998_10/jqa108.txt;
http://www.ddj.com/ftp/
1998/1998_10/jqa108.zip.

Criscolo:1998:JQ

[Cri98b] Mike Criscolo. Java Q&A:

How do I queue Java threads? Dr. Dobb’s Journal of Software Tools, 23(10):127–129, October 1998. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/1998/1998_10/jqa108.
txt;   http://www.ddj.com/
ftp/1998/1998_10/jqa108.
zip.

Cromwell:1998:PBD

[Cro98] Jeff Cromwell. Programmer’s bookshelf: The dawning of the age of multithreading. Dr. Dobb’s Journal of Software Tools, 23(9):127, 129, September 1998. CODEN DDJOEB. ISSN 1044-789X.

Chang:1995:CSM

[CS95a] C.-Y. Chang and J.-P. Sheu.

Compile-time scheduling of multithread with data localities on multiple vector processors. Concurrency: Practice and Experience, 7(5):349– 369, August 1995. CODEN CPEXEI. ISSN 1040-3108.

Chang:1995:CTS

[CS95b] C.-Y. Chang and J.-P. Sheu. Compile-time scheduling of multithread with data localities on multiple vector processors. Concurrency: Practice and Experience, 7(5):349– 369, August 1995. CODEN CPEXEI. ISSN 1040-3108.

Carr:2000:PCL

[CS00] Steve Carr and Ching-Kuang Shene. A portable class library for teaching multithreaded programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):124– 127, September 2000. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carothers:2002:CMP

[CS02] Christopher D. Carothers and Boleslaw K. Szymanski. Checkpointing multithreaded programs. Dr. Dobb’s Journal of Software Tools, 27(8):??, August 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2002/
2002_08/checkpt.txt.

Chen:2012:CLA

[CS12] Guancheng Chen and Per Stenstrom. Critical lock analysis: diagnosing critical section bottlenecks in multithreaded applications. In Hollingsworth [Hol12], pages 71:1–71:11. ISBN 1-4673-0804-8. URL http:

//conferences.computer.
org/sc/2012/papers/1000a099.
pdf.

ChassindeKergommeaux:2000:PIV

[CSB00] J. Chassin de Kergommeaux, B. Stein, and P. E. Bernard. Pajé, an interactive visualization tool for tuning multithreaded parallel applications. Parallel Computing, 26(10): 1253–1274, August 15, 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-ng/
10/35/21/42/31/24/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/35/21/42/31/
24/article.pdf.

Chappell:1999:SSM

[CSK^{+}99] Robert S. Chappell, Jared Stark, Sangwook P. Kim, Steven K. Reinhardt, and Yale N. Patt. Simultaneous subordinate microthreading (SSMT). ACM SIGARCH Computer Architecture News, 27(2):186–195, May 1999. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Constantinou:2005:PIS

[CSM^{+}05] Theofanis Constantinou, Yiannakis Sazeides, Pierre Michaud, Damien Fetis, and Andre Seznec. Performance implications of single thread migration on a chip multi-core. ACM SIGARCH Computer Architecture News, 33(4):80– 91, November 2005. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Cheikh:2021:KDV

[CSM^{+}21] A. Cheikh, S. Sordillo, A. Mastrandrea, F. Menichelli, G. Scotti, and M. Olivieri. Klessydra-T: Designing vector coprocessors for multithreaded edge-computing cores. IEEE Micro, 41(2):64– 71, March/April 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Culler:1991:FGPa

[CSS^{+}91a] David E. Culler, Anurag Sah, Klaus E. Schauser, Thorsten von Eicken, and John Wawrzynek. Fine-grain parallelism with minimal hardware support: a compilercontrolled threaded abstract machine. ACM SIGARCH Computer Architecture News, 19(2):164–175, April 1991. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Culler:1991:FGPb

[CSS^{+}91b] David E. Culler, Anurag Sah, Klaus E. Schauser, Thorsten von Eicken, and John Wawrzynek. Finegrain parallelism with minimal hardware support: a compiler-controlled threaded abstract machine. ACM SIG-PLAN Notices, 26(4):164–175, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Culler:1991:FGPc

[CSS^{+}91c] David E. Culler, Anurag Sah, Klaus E. Schauser, Thorsten von Eicken, and John Wawrzynek. Finegrain parallelism with minimal hardware support: a compiler-controlled threaded abstract machine. Operating Systems Review, 25(3S): 164–175, April 1991. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Choi:2010:MDA

[CSV10] Jee W. Choi, Amik Singh, and Richard W. Vuduc. Modeldriven autotuning of sparse matrix-vector multiply on GPUs. ACM SIGPLAN Notices, 45(5):115–126, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Christopher:2000:HPJ

[CT00] Thomas Christopher and George Thiruvathukal. High Performance Java Platform Computing: Multithreaded and Networked Programming. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-016164-0. xxii + 409 pp. LCCN ???? US$49.99.

URL http://www.sun.com/
books/catalog/christopher/

.

Chappell:2002:DPB

[CTYP02] Robert S. Chappell, Francis Tseng, Adi Yoaz, and Yale N. Patt. Difficult-path branch prediction using subordinate microthreads. ACM SIGARCH Computer Architecture News, 30(2):307–317, May 2002. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Caromel:1998:JFS

[CV98] Denis Caromel and Julien Vayssiere. A Java framework for seamless sequential, multi-threaded, and distributed programming. In ACM [ACM98a], page ?? ISBN ???? LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/
papers/javapp.pdf;   http:
//www.cs.ucsb.edu/conferences/
java98/papers/javapp.ps.

Possibly unpublished, except electronically.

Chen:2018:ROM

[CvdBC18] Kuan-Hsun Chen, Georg von der Brüggen, and Jian-Jia Chen. Reliability optimization on multi-core systems with multi-tasking and redundant multi-threading. IEEE Transactions on Computers, 67(4):484–497, ???? 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/document/8094023/.

Chugh:2008:DAC

[CVJL08] Ravi Chugh, Jan W. Voung, Ranjit Jhala, and Sorin Lerner. Dataflow analysis for concurrent programs using datarace detection. ACM SIGPLAN Notices, 43(6):316– 326, June 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:1998:WMP

[CW98] Aaron Cohen and Mike Woodring. Win32 Multithreaded Programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 1-56592-296-4. xv + 705 pp. LCCN QA76.76.O63 C633 1998. US$39.95.

URL http://www.ora.com/
catalog/multithread/;
http://www.oreilly.com/
catalog/multithread.

Chakravarti:2003:ISM

[CWHB03] A. Chakravarti, X. Wang, J. Hallstrom, and G. Baumgartner. Implementation of strong mobility for multithreaded agents in Java. In Proceedings of the International Conference on Parallel Processing, pages 321–332. ????, ????, 2003. CODEN ???? ISSN 0190-3918.

Chakraborty:2006:CSE

[CWS06] Koushik Chakraborty, Philip M. Wells, and Gurindar S. Sohi. Computation spreading: employing hardware migration to specialize CMP cores onthe-fly. ACM SIGPLAN Notices, 41(11):283–292, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Choi:2009:HCS

[CY09] Seungryul Choi and Donald Yeung. Hill-climbing SMT processor resource distribution. ACM Transactions on Computer Systems, 27(1):1:1– 1:??, February 2009. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Chin:2018:EAN

[CYYL18] Wei-Sheng Chin, Bo-Wen Yuan, Meng-Yuan Yang, and Chih-Jen Lin. An efficient alternating Newton method for learning factorization machines. ACM Transactions on Intelligent Systems and Technology (TIST), 9(6):72:1– 72:??, November 2018. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3230710.

Chen:1998:MTO

[CYZ98] Jiajun Chen, Xiaodong Yuan, and Guolian Zhengp. A multithreaded object-oriented programming model. ACM SIG-SOFT Software Engineering Notes, 23(3):83–86, May 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Choi:2002:IFI

[CZ02] Jong-Deok Choi and Andreas Zeller. Isolating failureinducing thread schedules. ACM SIGSOFT Software Engineering Notes, 27(4): 210–220, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cao:2017:HRD

[CZS^{+}17] Man Cao, Minjia Zhang, Aritra Sengupta, Swarnendu Biswas, and Michael D. Bond. Hybridizing and relaxing dependence tracking for efficient parallel runtime support. ACM Transactions on Parallel Computing (TOPC), 4(2):9:1–9:??, October 2017. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Cao:2016:DBG

[CZSB16] Man Cao, Minjia Zhang, Aritra Sengupta, and Michael D. Bond. Drinking from both glasses: combining pessimistic and optimistic tracking of cross-thread dependences. ACM SIGPLAN Notices, 51(8):20:1–20:??, August 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cai:2013:TST

[CZWC13] Yan Cai, Ke Zhai, Shangru Wu, and W. K. Chan. Team-Work: synchronizing threads globally to detect real deadlocks for multithreaded programs. ACM SIGPLAN Notices, 48(8):311–312, August 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’13 Conference proceedings.

Daniluk:2009:MTS

[Dan09]  Andrzej Daniluk.  Mul-

tithreaded transactions in scientific computing. The Growth06 v2 program. Computer Physics Communications, 180(7):1219–1220, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465509000393.

Davis:2011:ASM

[Dav11] Timothy A. Davis. Algorithm 915, SuiteSparseQR: Multifrontal multithreaded rankrevealing sparse QR factorization. ACM Transactions on Mathematical Software, 38(1): 8:1–8:22, November 2011. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Day:1992:INB

[Day92a] Michael Day. Implementing NLM-Based client/server architectures. Dr. Dobb’s Journal of Software Tools, 17(10): 78–84, October 1992. CODEN DDJOEB. ISSN 1044-789X.

Day:1992:INC

[Day92b] Michael Day. Implementing NLM-Based client/server architectures. Dr. Dobb’s Journal of Software Tools, 17(10): 78–84, October 1992. CODEN DDJOEB. ISSN 1044-789X.

deBoer:2009:SVC

[dB09] F. S. de Boer. A sharedvariable concurrency analysis of multi-threaded objectoriented programs. Theoretical Computer Science, 410(2– 3):128–141, February 6, 2009. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Draves:1991:UCI

[DBRD91] Richard P. Draves, Brian N.

Bershad, Richard F. Rashid, and Randall W. Dean. Using continuations to implement thread management and communication in operating systems. Operating Systems Review, 25(5):122–136, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Duda:1999:BVT

[DC99] Kenneth J. Duda and David R. Cheriton. Borrowed-virtual-time (BVT) scheduling: supporting latency-sensitive threads in a general-purpose scheduler. Operating Systems Review, 33(5):261–276, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Duda:2000:BVT

[DC00] Kenneth J. Duda and David R. Cheriton. Borrowed-virtual-time (BVT) scheduling: supporting latency-sensitive threads in a general-purpose scheduler. Operating Systems Review, 34(2):27–28, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dou:2007:CCM

[DC07] Jialin Dou and Marcelo Cintra. A compiler cost model for speculative parallelization. ACM Transactions on Architecture and Code Optimization, 4(2):12:1–12:??, June 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Das:2007:FVT

[DCK07] Dipankar Das, P. P. Chakrabarti, and Rajeev Kumar. Functional verification of task partitioning for multiprocessor embedded systems. ACM Transactions on Design Automation of Electronic Systems., 12(4):44:1– 44:??, September 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

DeLozier:2018:SSO

[DELD18] Christian DeLozier, Ariel Eizenberg, Brandon Lucia, and Joseph Devietti. SOFRITAS: Serializable orderingfree regions for increasing thread atomicity scalably. ACM SIGPLAN Notices, 53(2):286–300, February 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dennis:1994:MMP

[Den94] Jack B. Dennis. Machines and models for parallel computing. International Journal of Parallel Programming, 22(1): 47–77, February 1994. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

DuBois:2013:CSI

[DESE13] Kristof Du Bois, Stijn Eyerman, Jennifer B. Sartor, and Lieven Eeckhout. Criticality stacks: identifying critical threads in parallel programs using synchronization behavior. ACM SIGARCH Computer Architecture News, 41 (3):511–522, June 2013. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ICSA ’13 conference proceedings.

Silva:2019:RFG

[DFC^{+}19] Lucas Bragança Da Silva, Ricardo Ferreira, Michael Canesche, Marcelo M. Menezes, Maria D. Vieira, Jeronimo Penha, Peter Jamieson, and José Augusto M. Nacif. READY: a fine-grained multithreading overlay framework for modern CPU–FPGA dataflow applications. ACM Transactions on Embedded Computing Systems, 18(5s): 56:1–56:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3358187.

DeWitt:1999:PTL

[DG99] Anthony DeWitt and Thomas Gross. The potential of thread-level speculation based on value profiling. ACM SIGARCH Computer Architecture News, 27(1):22, March 1999. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Domani:2003:TLH

[DGK^{+}03] Tamar Domani, Gal Goldshtein, Elliot K. Kolodner, Ethan Lewis, Erez Petrank, and Dafna Sheinwald. Thread-local heaps for Java. ACM SIGPLAN Notices, 38(2s):183–194, February 2003. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dosanjh:2020:TQM

[DGSB20] Matthew G. F. Dosanjh, Ryan E. Grant, Whit Schonbein, and Patrick G. Bridges. Tail queues: a multi-threaded matching architecture. Concurrency and Computation: Practice and Experience, 32 (3):e5158:1–e5158:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DHollander:1992:PLL

[D’H92] Erik H. D’Hollander. Partitioning and labeling of loops by unimodular transformations. IEEE Transactions on Parallel and Distributed Systems, 3(4):465–476, July 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

DeRusso:1998:MEH

[DH98] Joe DeRusso, III and Peter Haggar. Multithreaded exception handling in Java. Java Report: The Source for Java Development, 3(??):??, August 1998. CODEN JREPFI. ISSN 1086-4660. URL http:/

/archive.javareport.com/
9808/html/from_pages/ftp_
feature.shtml.

Dolby:2012:DCA

[DHM^{+}12] Julian Dolby, Christian Hammer, Daniel Marino, Frank Tip, Mandana Vaziri, and Jan Vitek. A data-centric approach to synchronization. ACM Transactions on Programming Languages and Systems, 34(1):4:1–4:48, April 2012. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Duncan:2001:LPD

[DHR^{+}01] Ray Duncan, Duncan Harris, Douglas Reilly, Craig Rodrigues, Michael Birken, and Paul S. Person. Letters: Plugin desupport; threading and the .Net framework; CORBA interoperability; game over for Java; totally wired. Dr. Dobb’s Journal of Software Tools, 26(11):10, 12, November 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/.

Dillon:1993:VEM

[Dil93] Laura K. Dillon. A visual execution model for Ada tasking. ACM Transactions on Software Engineering and Methodology, 2(4):311–345, October 1993. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tosem/
1993-2-4/p311-dillon/p311-
dillon.pdf;   http://www.
acm.org/pubs/citations/

journals/tosem/1993-2-4/
p311-dillon/.

Dill:2000:MCJ

[Dil00] David Dill. Model checking Java programs (abstract only). ACM SIGSOFT Software Engineering Notes, 25 (5):179, September 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Divekar:1995:IMP

[Div95] Ravindra Divekar. The impact of multithreading on the performance of superscalar processors. Thesis (M.A.), State University of New York at Binghamton, Thomas J. Watson School of Engineering and Applied Science, Binghamton, NY, USA, 1995. vi + 73 pp.

Dam:2010:PCI

[DJLP10] Mads Dam, Bart Jacobs, Andreas Lundblad, and Frank Piessens. Provably correct inline monitoring for multithreaded Java-like programs. Journal of Computer Security, 18(1):37–59, ???? 2010. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Karniadakis:2002:DLP

[DK02] Suchuan Dong and George Em.

Karniadakis. Dual-level parallelism for deterministic and stochastic CFD problems. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN

???? URL http://www.sc-
2002.org/paperpdfs/pap.
pap137.pdf.

Denniston:2016:DH

[DKA16] Tyler Denniston, Shoaib Kamil, and Saman Amarasinghe. Distributed Halide. ACM SIGPLAN Notices, 51 (8):5:1–5:??, August 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dubey:1994:APM

[DKF94] Pradeep Dubey, Arvind Krishna, and M. J. (Michael J.) Flynn. Analytical performance modeling for a spectrum of multithreaded machines. Research report RC 19549 (85007), IBM T. J. Watson Research Center, Yorktown Heights, NY, USA, May 3, 1994. 27 pp.

Ding:2018:IOC

[DKG18] Bailu Ding, Lucja Kot, and Johannes Gehrke. Improving optimistic concurrency control through transaction batching and operation reordering. Proceedings of the VLDB Endowment, 12(2):169–182, October 2018. CODEN ???? ISSN 2150-8097.

Doligez:1993:CGG

[DL93] Damien Doligez and Xavier Leroy. A concurrent, generational garbage collector for a multithreaded implementation of ML. In ACM [ACM93a], pages 113–

123. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15 1993.

URL http://www.acm.org:
80/pubs/citations/proceedings/
plan/158511/p113-doligez/

. ACM order number 549930.

Devietti:2009:DDS

[DLCO09] Joseph Devietti, Brandon Lucia, Luis Ceze, and Mark Oskin. DMP: deterministic shared memory multiprocessing. ACM SIGPLAN Notices, 44(3):85–96, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dongarra:1999:RAP

[DLM99] J. J. Dongarra, E. Luque, and Tomas Margalef, editors. Recent advances in parallel virtual machine and message passing interface: 6th European PVM/ MPI Users’ Group Meeting, Barcelona, Spain, September 26–29, 1999: Proceedings, volume 1697 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CO-DEN LNCSD9. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999. URL

http://link.springer-ny.
com/link/service/series/

                                        0558/tocs/t1697.htm;   http:
                                        //www.springerlink.com/
                                        openurl.asp?genre=issue&
                                        issn=0302-9743&volume=1697.

delaPuente:1999:RTP

[dlPRGB99] Juan A. de la Puente, José F. Ruiz, and Jesús M. González-Barahona. Real-time programming with GNAT: Specialized kernels versus POSIX threads. ACM SIGADA Ada Letters, 19(2):73–77, June 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Demange:2013:PBB

[DLZ^{+}13] Delphine Demange, Vincent Laporte, Lei Zhao, Suresh Jagannathan, David Pichardie, and Jan Vitek. Plan B: a buffered memory model for Java. ACM SIGPLAN Notices, 48(1):329–342, January 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dagum:1998:OIS

[DM98] Leonardo Dagum and Ramesh Menon. OpenMP: An industry-standard API for shared-memory programming. IEEE Computational Science & Engineering, 5(1): 46–55, January/March 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1998/pdf/c1046.

pdf;   http://www.computer.
org/cse/cs1998/c1046abs.
htm.

Daloze:2016:ETS

[DMBM16] Benoit Daloze, Stefan Marr, Daniele Bonetta, and Hanspeter Mössenböck. Efficient and thread-safe objects for dynamicallytyped languages. ACM SIG-PLAN Notices, 51(10):642– 659, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Drechsler:2018:TSR

[DMSM18] Joscha Drechsler, Ragnar Mogk, Guido Salvaneschi, and Mira Mezini. Threadsafe reactive programming. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA): 107:1–107:30, October 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3276477.

Dorfman:1994:EMO

[DN94] Len Dorfman and Marc J.

Neuberger. Effective multithreading in OS/2. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-017841-0 (paperback). xii + 288 pp. LCCN QA76.76.O63D6694 1994. US$34.95.

Devietti:2012:RRC

[DNB^{+}12] Joseph Devietti, Jacob Nelson, Tom Bergan, Luis Ceze, and Dan Grossman. RCDC: a relaxed consistency deterministic computer. ACM SIG-PLAN Notices, 47(4):67–78, April 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Danjean:2000:IKA

[DNR00] Vincent Danjean, Raymond Namyst, and Robert D. Russell. Integrating kernel activations in a multithreaded runtime system on top of L INUX. Lecture Notes in Computer Science, 1800:1160–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1800/18001160.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1800/18001160.
pdf.

Dublish:2016:CCG

[DNT16] Saumay Dublish, Vijay Nagarajan, and Nigel Topham. Cooperative caching for GPUs. ACM Transactions on Architecture and Code Optimization, 13(4):39:1–39:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Dorojevets:1995:MDA

[DO95] M. N. Dorojevets and V. G. Oklobdzija. Multithreaded decoupled architecture. International Journal of High Speed Computing (IJHSC), 7 (3):465–??, 1995. CODEN IH-SCEZ. ISSN 0129-0533.

Donalson:1992:DDP

[Don92] Douglas Dale Donalson. DISC:

a dynamic performance evaluation of a multi-thread architecture. Thesis (M.S.), Electrical and Computer Engineering Department, University of California, Santa Barbara, Santa Barbara, CA, USA, 1992. ix + 88 pp.

Donnelly:2002:LTT

[Don02] Austin Donnelly. Lightweight thread tunnelling in network applications. Lecture Notes in Computer Science, 2546:48– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2546/25460048.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2546/25460048.pdf.

Dou:1997:ISV

[DPZ97] Yong Dou, Zhengbing Pang, and Xingming Zhou. Implementing a software virtual shared memory on PVM. In IEEE [IEE97], page ?? ISBN 0-8186-7876-3 (paperback and case), 0-8186-7878-X (microfiche). LCCN QA76.58 .A4 1997.

Drake:1996:IJT

[Dra96] Donald G. Drake. Introduction to Java threads. JavaWorld: IDG’s magazine for the Java community, 1 (2):??, April 1996. CO-DEN ???? ISSN 1091-

8906.  URL http://www.
javaworld.com/javaworld/
jw-04-1996/jw-04-threads.
htm.

Drusinsky:1995:VDE

[Dru95] Doron Drusinsky. Visually designing embedded-systems applications. Dr. Dobb’s Journal of Software Tools, 20(6): 62, 64, 66, 68, 104–106, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Delzanno:2002:TAV

[DRV02] Giorgio Delzanno, Jean-François Raskin, and Laurent Van Begin. Towards the automated verification of multithreaded Java programs. Lecture Notes in Computer Science, 2280:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2280/22800173.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2280/22800173.
pdf.

Desai:2009:AIC

[DS09] Aniruddha Desai and Jugdutt Singh. Architecture independent characterization of embedded Java workloads. IEEE Computer Architecture Letters, 8(1):29–32, January/ June 2009. CODEN ????

ISSN 1556-6056 (print), 1556-6064 (electronic).

Deniz:2016:UML

[DS16] Etem Deniz and Alper Sen.

Using machine learning techniques to detect parallel patterns of multi-threaded applications. International Journal of Parallel Programming, 44(4):867–900, August 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10766-015-0396-z.

Deiana:2018:UPN

[DSAD^{+}18] Enrico A. Deiana, Vincent St-Amour, Peter A. Dinda, Nikos Hardavellas, and Simone Campanoni. Unconventional parallelization of nondeterministic applications. ACM SIGPLAN Notices, 53(2):432–447, February 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bois:2013:BGV

[DSEE13] Kristof Du Bois, Jennifer B.

Sartor, Stijn Eyerman, and Lieven Eeckhout. Bottle graphs: visualizing scalability bottlenecks in multi-threaded applications. ACM SIGPLAN Notices, 48(10):355–372, October 2013. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Dang:2017:ECB

[DSG17] Hoang-Vu Dang, Marc Snir, and William Gropp. Eliminating contention bottlenecks in multithreaded MPI. Parallel Computing, 69(??):1– 23, November 2017. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0167819117301187.

Dohi:2010:IPE

[DSH^{+}10] Keisuke Dohi, Yuichiro Shibata, Tsuyoshi Hamada, Tomonari Masada, Kiyoshi Oguri, and Duncan A. Buell. Implementation of a programming environment with a multithread model for reconfigurable systems. ACM SIGARCH Computer Architecture News, 38(4):40– 45, September 2010. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Das:2015:SBP

[DSR15] Madan Das, Gabriel Southern, and Jose Renau. Sectionbased program analysis to reduce overhead of detecting unsynchronized thread communication. ACM Transactions on Architecture and Code Optimization, 12(2):23:1–23:??, July 2015. CODEN ????

ISSN 1544-3566 (print), 1544-3973 (electronic).

Dutta:2017:SVC

[DSR17] Sudakshina Dutta, Dipankar Sarkar, and Arvind Rawat. Synchronization validation for cross-thread dependences in parallel programs. International Journal of Parallel Programming, 45(6):1326– 1365, December 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Ding:2015:OCA

[DTK^{+}15] Wei Ding, Xulong Tang, Mahmut Kandemir, Yuanrui Zhang, and Emre Kultursay. Optimizing off-chip accesses in multicores. ACM SIG-PLAN Notices, 50(6):131–142, June 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

David:2014:CMC

[DTLM14] Florian David, Gael Thomas, Julia Lawall, and Gilles Muller. Continuously measuring critical section pressure with the free-lunch profiler. ACM SIGPLAN Notices, 49(10):291–307, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Diavastos:2016:ITD

[DTLW16] Andreas Diavastos, Pedro Trancoso, Mikel Luján, and Ian Watson. Integrating transactions into the datadriven multi-threading model using the TFlux platform. International Journal of Parallel Programming, 44(2): 257–277, April 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10766-015-0369-2.

Deveci:2018:MSM

[DTR18] Mehmet Deveci, Christian Trott, and Sivasankaran Rajamanickam. Multithreaded sparse matrix–matrix multiplication for many-core and GPU architectures. Parallel Computing, 78(??):33– 46, October 2018. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167819118301923.

Dubey:1995:SSM

[Dub95]  Pradeep Dubey.  Single-

program speculative multithreading (SPSM) architecture: compiler-assisted finegrained multithreading. Research report RC 19928 (88233), IBM T. J. Watson Research Center, Yorktown Heights, NY, USA, February 6, 1995. 25 pp.

Dugger:1995:MC

[Dug95] Jim Dugger. Multithreading in C++. C/C++ Users Journal, 13(11):23–??, November 1995. CODEN CCUJEX. ISSN 1075-2838.

Dascal:1999:ELR

[DV99] Shlomit Dascal and Uzi Vishkin. Experiments with list ranking for explicit multithreaded (XMT) instruction parallelism (extended abstract). Lecture Notes in Computer Science, 1668:43– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1668/16680043.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1668/16680043.
pdf.

DePestel:2018:RRP

[DVAE18] Sander De Pestel, Sam Van den Steen, Shoaib Akram, and Lieven Eeckhout. RPPM: Rapid performance prediction of multithreaded applications on multicore hardware. IEEE Computer Architecture Letters, 17(2):183–186, July/ December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Devietti:2012:RAS

[DWS^{+}12] Joseph Devietti, Benjamin P.

Wood, Karin Strauss, Luis Ceze, Dan Grossman, and Shaz Qadeer. RADISH: always-on sound and complete \underline{Ra}ce \underline{D}etection \underline{i}n \underline{S}oftware and \underline{H}ardware. ACM SIGARCH Computer Architecture News, 40(3):201– 212, June 2012. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ISCA ’12 conference proceedings.

Ding:2010:PCM

[DWYB10] Jason Jianxun Ding, Abdul Waheed, Jingnan Yao, and Laxmi N. Bhuyan. Performance characterization of multi-thread and multi-core processors based XML application oriented networking systems. Journal of Parallel and Distributed Computing, 70(5):584–597, May 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dyer:1998:CAS

[Dye98] Dave Dyer. Can Assure save Java from the perils of multithreading? JavaWorld: IDG’s magazine for the Java community, 3(10):??, 1998. CODEN ???? ISSN 1091-

8906.  URL http://www.
javaworld.com/javaworld/
jw-10-1998/jw-10-assure.
htm.

Ding:2012:CDF

[DZKS12] Wei Ding, Yuanrui Zhang, Mahmut Kandemir, and Seung Woo Son. Compilerdirected file layout optimization for hierarchical storage systems. In Hollingsworth [Hol12], pages 41:1–41:11. ISBN 1-4673-0804-8. URL

http://conferences.computer.
org/sc/2012/papers/1000a030.
pdf.

Elwasif:2001:AMT

[EBKG01] Wael R. Elwasif, David E.

Bernholdt, James A. Kohl, and G. A. Geist. An architecture for a multi-threaded harness kernel. Lecture Notes in Computer Science, 2131:126– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2131/21310126.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310126.
pdf.

Eskilson:1998:SMM

[EC98] Jesper Eskilson and Mats Carlsson. SICStus MT — a multithreaded execution environment for SICStus Prolog. Lecture Notes in Computer Science, 1490:36–53, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1490/14900036.
htm;   http://link.springer-
ny.com/link/service/series/

                                        0558/papers/1490/14900036.
                                        pdf.

Esmaeilzadeh:2012:LBL

[ECX^{+}12] Hadi Esmaeilzadeh, Ting Cao, Yang Xi, Stephen M. Blackburn, and Kathryn S. McKinley. Looking back on the language and hardware revolutions: measured power, performance, and scaling. ACM SIGPLAN Notices, 47(4):319– 332, April 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eyerman:2009:MLP

[EE09a] Stijn Eyerman and Lieven Eeckhout. Memory-level parallelism aware fetch policies for simultaneous multithreading processors. ACM Transactions on Architecture and Code Optimization, 6(1):3:1– 3:??, March 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Eyerman:2009:PTC

[EE09b] Stijn Eyerman and Lieven Eeckhout. Per-thread cycle accounting in SMT processors. ACM SIGPLAN Notices, 44(3):133–144, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eyerman:2010:PJS

[EE10] Stijn Eyerman and Lieven Eeckhout. Probabilistic job symbiosis modeling for SMT processor scheduling. ACM SIGPLAN Notices, 45(3):91– 102, March 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eyerman:2012:PMJ

[EE12] Stijn Eyerman and Lieven Eeckhout. Probabilistic modeling for job symbiosis scheduling on SMT processors. ACM Transactions on Architecture and Code Optimization, 9(2):7:1–7:??, June 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Eyerman:2014:RCW

[EE14] Stijn Eyerman and Lieven Eeckhout. Restating the case for weighted-IPC metrics to evaluate multiprogram workload performance. IEEE Computer Architecture Letters, 13(2):93–96, July/December 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Eggers:1997:SMP

[EEL^{+}97] Susan J. Eggers, Joel S.

Emer, Henry M. Levy, Jack L. Lo, Rebecca L. Stamm, and Dean M. Tullsen. Simultaneous multithreading: a platform for next-generation processors. IEEE Micro, 17(5): 12–19, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1997/pdf/m5012.
                                        pdf;   http://www.computer.
                                        org/micro/mi1997/m5012abs.
                                        htm.

Edelstein:2003:FTM

[EFG^{+}03] Orit Edelstein, Eitan Farchi, Evgeny Goldin, Yarden Nir, Gil Ratsaby, and Shmuel Ur. Framework for testing multithreaded Java programs. Concurrency and Computation: Practice and Experience, 15 (3–5):485–499, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Emmi:2007:LA

[EFJM07] Michael Emmi, Jeffrey S. Fischer, Ranjit Jhala, and Rupak Majumdar. Lock allocation. ACM SIGPLAN Notices, 42(1):291–296, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Edelstein:2001:MJP

[EFN^{+}01] Orit Edelstein, Eitan Farchi, Yarden Nir, Gil Ratsaby, and Shmuel Ur. Multithreaded Java program test generation. In ACM [ACM01], page 181. ISBN 1-58113-359-6. LCCN QA76.9.O35

                                        A26 2001. URL http://www.
                                        philippsen.com/JGI2001/
                                        camerareadyabstracts/18.
                                        html;   http://www.philippsen.

com/JGI2001/finalpapers/
18500181.ps.

Edelstein:2002:MJP

[EFN^{+}02] O. Edelstein, E. Farchi, Y. Nir, G. Ratsaby, and S. Ur. Multithreaded Java program test generation. IBM Systems Journal, 41(1):111– 125, ???? 2002. CODEN IBMSA7. ISSN 0018-8670.

URL http://www.research.
ibm.com/journal/sj/411/
edelstein.html;   http:
//www.research.ibm.com/
journal/sj/411/edelstein.
pdf.

Esparza:2011:CPB

[EG11] Javier Esparza and Pierre Ganty. Complexity of patternbased verification for multithreaded programs. ACM SIGPLAN Notices, 46(1):499– 510, January 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

El-Ghazawi:2002:UPP

[EGC02] Tarek El-Ghazawi and François Cantonnet. UPC performance and potential: a NPB experimental study. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN

???? URL http://www.sc-
2002.org/paperpdfs/pap.
pap316.pdf.

Eggers:2010:AL

[Egg10] Susan Eggers. 2010 Athena lecture. ACM SIGPLAN Notices, 45(6):98, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Esparza:2014:PBV

[EGP14] Javier Esparza, Pierre Ganty, and Tomás Poch. Patternbased verification for multithreaded programs. ACM Transactions on Programming Languages and Systems, 36 (3):9:1–9:??, September 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Elmasri:1995:TCL

[EHG95] N. Elmasri, H. H. J. Hum, and G. R. Gao. The threaded communication library: Preliminary experiences on a multiprocessor with dual-processor nodes. In ACM [ACM95a], pages 195–199. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995.

Emer:2007:STV

[EHP^{+}07] Joel Emer, Mark D. Hill, Yale N. Patt, Joshua J. Yi, Derek Chiou, and Resit Sendag. Single-threaded vs. multithreaded: Where should we focus? IEEE Micro, 27(6): 14–24, November/December 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Eytani:2007:TFB

[EHSU07] Yaniv Eytani, Klaus Havelund, Scott D. Stoller, and Shmuel Ur. Towards a framework and a benchmark for testing tools for multi-threaded programs. Concurrency and Computation: Practice and Experience, 19(3):267–279, March 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Eickemeyer:1997:EMP

[Eic97] Richard J. Eickemeyer. Evaluation of multithreaded processors and thread-switch policies. Research report RC 20956 (92759), IBM T. J. Watson Research Center, Yorktown Heights, NY, USA, August 18, 1997. 16 pp.

Eager:1993:CER

[EJ93] Derek L. Eager and John Jahorjan. Chores: Enhanced run-time support for shared-memory parallel computing. ACM Transactions on Computer Systems, 11(1): 1–32, February 1993. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tocs/1993-11-1/
p1-eager/.

Eickemeyer:1996:EMU

[EJK^{+}96] Richard J. Eickemeyer, Ross E.

Johnson, Steven R. Kunkel, Mark S. Squillante, and Shiafun Liu. Evaluation of multithreaded uniprocessors for commercial application environments. ACM SIGARCH Computer Architecture News, 24(2):203–212, May 1996. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Ediger:2013:GMA

[EJRB13] David Ediger, Karl Jiang, E. Jason Riedy, and David A. Bader. GraphCT: Multithreaded algorithms for massive graph analysis. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2220–2229, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Eykholt:1992:BMM

[EKB^{+}92] J. R. Eykholt, S. R. Kleiman, S. Barton, R. Faulkner, D. Stein, M. Smith, A. Shivalingiah, J. Voll, M. Weeks, and D. Williams. Beyond multiprocessing: Multithreading the System V Release 4 kernel. In USENIX [USE92a], pages 11–18. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Eggers:1990:TEI

[EKKL90] S. J. Eggers, David R. Keppel, Eric J. Koldinger, and Henry M. Levy. Techniques for efficient inline tracing on a shared-memory multiprocessor. ACM SIGMETRICS Performance Evaluation Review, 18(1):37–47, May 1990. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

English:1995:MC

[Eng95] John English. Multithreading in C++. ACM SIG-PLAN Notices, 30(4):21–28, April 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Engelschall:2000:PMS

[Eng00] Ralf S. Engelschall. Portable multithreading — the signal stack trick for userspace thread creation. In USENIX [USE00a], pages 239–249. ISBN 1-880446-22-

7. LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/engelschall.
html.

Evtyushkin:2016:UMC

[EPAG16] Dmitry Evtyushkin, Dmitry Ponomarev, and Nael Abu-Ghazaleh. Understanding and mitigating covert channels through branch predictors. ACM Transactions on Architecture and Code Optimization, 13(1):10:1–10:??, April 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Elmas:2007:GRT

[EQT07] Tayfun Elmas, Shaz Qadeer, and Serdar Tasiran. Goldilocks: a race and transaction-aware Java runtime. ACM SIG-PLAN Notices, 42(6):245–255, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Emerson:1997:USW

[ES97] E. A. Emerson and A. P. Sistla. Utilizing symmetry when model-checking under fairness assumptions: An automata-theoretic approach. ACM Transactions on Programming Languages and Systems, 19(4): 617–638, July 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/toplas/1997-19-
                                        4/p617-emerson/.

Esposito:1996:MVB

[Esp96] Dino Esposito. Multithreading and Visual Basic. Dr. Dobb’s Journal of Software Tools, 21(12):46–??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Estep:1993:LMM

[Est93] James L. Estep. Lightweight multithreaded multimedia conference server. Thesis (M.S.), West Virginia University, Morgantown, WV, USA, 1993. vi + 57 pp.

Ergin:2006:ENV

[EUVG06] O. Ergin, O. Unsal, X. Vera, and A. Gonzalez. Exploiting narrow values for soft error tolerance. IEEE Computer Architecture Letters, 5(2):12, February 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Eigenmann:2001:OSM

[EV01] Rudolf Eigenmann and Michael J.

Voss, editors. OpenMP shared memory parallel programming: International Workshop on OpenMP Applications and Tools, WOMPAT 2001, West Lafayette, IN, USA, July 30–31, 2001: proceedings, volume 2104 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42346-X (paperback). LCCN QA76.642 .I589 2001; QA267.A1 L43 no.2104. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t2104.htm.

Evripidou:2001:MDD

[Evr01] Paraskevas Evripidou. D^{3}- Machine: a decoupled datadriven multithreaded architecture with variable resolution support. Parallel Computing, 27(9):1197– 1225, August 2001. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-ng/
10/35/21/47/35/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/35/21/47/35/
25/article.pdf.

Engelhardt:1996:PIP

[EW96] Dean Engelhardt and Andrew Wendelborn. A partitioningindependent paradigm for nested data parallelism. International Journal of Parallel Programming, 24(4):291– 317, August 1996. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Fraguela:2019:EDP

[FA19] B. B. Fraguela and D. Andrade. Easy dataflow programming in clusters with UPC++ DepSpawn. IEEE Transactions on Parallel and Distributed Systems, 30(6): 1267–1282, June 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fan:1993:LMC

[Fan93]  Xiaoming Fan.  Latency-

directed multithreaded computation and its architectural support. Thesis (Ph.D.), Universität Hamburg, Aachen, Germany, 1993. xi + 174 + 22 + 11 pp. Summary in German.

Farber:1996:EAM

[Far96] Philipp Farber. Execution architecture of the multithreaded ADAM prototype. Thesis (doctoral), Swiss Federal Institute of Technology, Zurich, Switzerland, 1996. iv + 127 pp.

Figueiredo:2001:IPH

[FBF01] Renato J. O. Figueiredo, Jeffrey P. Bradford, and José A. B. Fortes. Improving the performance of heterogeneous DSMs via multithreading. Lecture Notes in Computer Science, 1981:168–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1981/19810168.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1981/19810168.
pdf.

Fiske:1995:TPT

[FD95] Stuart Fiske and William J.

Dally. Thread prioritization: a thread scheduling mechanism for multiple-context parallel processors. Future Generation Computer Systems, 11 (6):503–518, October 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Feuerstein:1996:MTP

[FD96] E. Feuerstein and A. S. De Loma. On multi-threaded paging. Lecture Notes in Computer Science, 1178:417– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feuerstein:2002:LMT

[FdL02] E. Feuerstein and A. Strejilevich de Loma. On-line multi-threaded paging. Algorithmica, 32(1):36–60, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0178-4617&volume=32&
issue=1&spage=36.

Fekete:2008:TSD

[Fek08] Alan D. Fekete. Teaching students to develop thread-safe Java classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):119–123, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Ferrara:2013:GSA

[Fer13] P. Ferrara. A generic static analyzer for multithreaded Java programs. Software— Practice and Experience, 43 (6):663–684, June 2013. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Flanagan:2004:ADA

[FF04] Cormac Flanagan and Stephen N. Freund. Atomizer: a dynamic atomicity checker for multithreaded programs. ACM SIGPLAN Notices, 39(1):256– 267, January 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flanagan:2008:ADA

[FF08] Cormac Flanagan and Stephen N.

Freund. Atomizer: a dynamic atomicity checker for multithreaded programs. Science of Computer Programming, 71 (2):89–109, April 1, 2008. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Flanagan:2009:FEP

[FF09] Cormac Flanagan and Stephen N.

Freund. FastTrack: efficient and precise dynamic race detection. ACM SIGPLAN Notices, 44(6):121–133, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flanagan:2010:AMD

[FF10] Cormac Flanagan and Stephen N.

Freund. Adversarial memory for detecting destructive races. ACM SIGPLAN Notices, 45(6):244–254, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fezzardi:2020:ABD

[FF20] Pietro Fezzardi and Fabrizio Ferrandi. Automated bug detection for high-level synthesis of multi-threaded irregular applications. ACM Transactions on Parallel Computing (TOPC), 7(4):27:1–27:26, December 2020. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3418086.

Flanagan:2008:TAS

[FFLQ08] Cormac Flanagan, Stephen N. Freund, Marina Lifshin, and Shaz Qadeer. Types for atomicity: Static checking and inference for Java. ACM Transactions on Programming Languages and Systems, 30(4): 20:1–20:52, July 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Flanagan:2004:EPA

[FFQ04] Cormac Flanagan, Stephen N. Freund, and Shaz Qadeer. Exploiting purity for atomicity. ACM SIGSOFT Software Engineering Notes, 29 (4):221–231, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Flanagan:2005:MVM

[FFQS05] Cormac Flanagan, Stephen N. Freund, Shaz Qadeer, and Sanjit A. Seshia. Modular verification of multithreaded programs. Theoretical Computer Science, 338(1–3):153– 183, June 10, 2005. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flanagan:2008:VSC

[FFY08] Cormac Flanagan, Stephen N.

Freund, and Jaeheon Yi. Velodrome: a sound and complete dynamic atomicity checker for multithreaded programs. ACM SIGPLAN Notices, 43(6):293–303, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Faulkner:1991:PFS

[FG91] Roger Faulkner and Ron Gomes. The process file system and process model in UNIX System V. In USENIX [USE91b], pages 243–252. LCCN QA 76.76 O63 U84 1992. URL

http://obits.mlive.com/
obituaries/grandrapids/
obituary.aspx?pid=180588279;
 http://thenewstack.io/
remembering-roger-faulkner/
;   https://www.usenix.
org/memoriam-roger-faulkner;
 https://www.usenix.org/
sites/default/files/usenix_
winter91_faulkner.pdf.

Frincu:2014:ESV

[FGG14] Marc E. Frincu, Stéphane Genaud, and Julien Gossa. On the efficiency of several VM provisioning strategies for workflows with multithreaded tasks on clouds. Computing, 96(11):1059–1086, November 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s00607-014-0410-0.

Foster:1997:MMC

[FGKT97] Ian Foster, Jonathan Geisler, Carl Kesselman, and Steven Tuecke. Managing multiple communication methods in high-performance networked computing systems. Journal of Parallel and Distributed Computing, 40(1):35–48, January 10, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.1266/production;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.1266/production/pdf;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.1266/production/ref.

Foster:1996:MIW

[FGT96] I. Foster, J. Geisler, and S. Tuecke. MPI on the I-WAY: a wide-area, multimethod implementation of the Message Passing Interface. In IEEE [IEE96], pages 10–17. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Fahringer:1995:UTDb

[FHM95a] T. Fahringer, M. Haines, and P. Mehrotra. On the utility of threads for data parallel programming. In ACM [ACM95a], pages 51–59. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995.

Fahringer:1995:UTDa

[FHM95b] Thomas Fahringer, Matthew Haines, and Piyush Mehrotra. On the utility of threads for data parallel programming. Washington, DC, USA, 1995. ?? pp. Shipping list number 96-0037-M.

Finger:1995:LTC

[Fin95] Jonathan Finger. Lightweight tasks in C. Dr. Dobb’s Journal of Software Tools, 20(5): 48, 50, 102, May 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Fisher:1997:SPS

[Fis97] Michael T. Fisher. A study of the performance of simultaneous multithreading on a superscalar processor. Thesis (M.S.E.E.), State University of New York at Binghamton, Watson School of Engineering and Applied Science, Binghamton, NY, USA, 1997. vi + 98 pp.

Fide:2008:PUS

[FJ08] S. Fide and S. Jenks. Proactive use of shared L3 caches to enhance cache communications in multi-core processors. IEEE Computer Architecture Letters, 7(2):57–60, July 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Farzan:2012:VPC

[FK12] Azadeh Farzan and Zachary Kincaid. Verification of parameterized concurrent programs by modular reasoning about data and control. ACM SIGPLAN Notices, 47(1):297–308, January 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fillo:1997:MMM

[FKD^{+}97] Marco Fillo, Stephen W. Keckler, William J. Dally, Nicholas P. Carter, Andrew Chang, Yevgeny Gurevich, and Whay S. Lee. The Mmachine multicomputer. International Journal of Parallel Programming, 25(3):183–212, June 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Farzan:2015:PSU

[FKP15] Azadeh Farzan, Zachary Kincaid, and Andreas Podelski. Proof spaces for unbounded parallelism. ACM SIGPLAN Notices, 50(1):407–420, January 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Foltzer:2012:MSP

[FKS^{+}12] Adam Foltzer, Abhishek Kulkarni, Rebecca Swords, Sajith Sasidharan, Eric Jiang, and Ryan Newton. A meta-scheduler for the parmonad: composable scheduling for the heterogeneous cloud. ACM SIGPLAN Notices, 47(9):235–246, September 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Foster:1996:NAI

[FKT96] Ian Foster, Carl Kesselman, and Steven Tuecke. The Nexus approach to integrating multithreading and communication. Journal of Parallel and Distributed Computing, 37(1):70–82, August 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0108/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0108/production/pdf.

Faust:1990:POO

[FL90] John E. Faust and Henry M.

Levy. The performance of an object-oriented threads package. ACM SIGPLAN Notices, 25(10):278–288, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frigo:1998:ICM

[FLR98] Matteo Frigo, Charles E.

Leiserson, and Keith H. Randall. The implementation of the Cilk-5 multithreaded language. ACM SIGPLAN Notices, 33(5):212– 223, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/pldi/277650/
                                        p212-frigo/.

Felten:1992:IPM

[FM92] Edward W. Felten and Dylan James McNamee. Improving the performance of message-passing applications by multithreading. Technical report 92-09-07, University of Washington, Dept. of Computer Science and Engineering, Seattle, WA, USA, 1992. 6 pp.

Fang:2015:MMD

[FMY^{+}15] Zhenman Fang, Sanyam Mehta, Pen-Chung Yew, Antonia Zhai, James Greensky, Gautham Beeraka, and Binyu Zang. Measuring microarchitectural details of multi- and many-core memory systems through microbenchmarking. ACM Transactions on Architecture and Code Optimization, 11(4):55:1–55:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Farzan:2017:SDC

[FN17] Azadeh Farzan and Victor Nicolet. Synthesis of divide and conquer parallelism for loops. ACM SIGPLAN Notices, 52(6):540–555, June 2017. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fix:2018:HMT

[FNA^{+}18] Jordan Fix, Nayana P. Nagendra, Sotiris Apostolakis, Hansen Zhang, Sophie Qiu, and David I. August. Hardware multithreaded transactions. ACM SIGPLAN Notices, 53(2):15–29, February 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feliu:2022:VVM

[FNS^{+}22] Josué Feliu, Ajeya Naithani, Julio Sahuquillo, Salvador Petit, Moinuddin Qureshi, and Lieven Eeckhout. VMT: Virtualized multi-threading for accelerating graph workloads on commodity processors. IEEE Transactions on Computers, 71(6):1386–1398, June 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fong:1997:BPM

[Fon97]  Waipang Fong.  Building

a preprocessor for a multithreading compiler. Thesis (M.E.E.), Department of Electrical Engineering, University of Alabama, Tuscaloosa, AL, USA, 1997. ix + 80 pp.

Ford:1995:EDT

[For95a]  Dan Ford.  Event-driven

threads in C++. Dr. Dobb’s Journal of Software Tools, 20 (6):48–50, 52, 54, 98, 100, 102, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Ford:1995:ETC

[For95b]  Dan Ford.  Event-driven

threads in C++. Dr. Dobb’s Journal of Software Tools, 20 (6):48–50, 52, 54, 98, 100, 102, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Forsell:1997:MMV

[For97] M. Forsell. MTAC — a multithreaded VLIW architecture for PRAM simulation. J.UCS: Journal of Universal Computer Science, 3 (9):1037–1055, September 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/medoc.springer.de:8000/
jucs/jucs_3_9/mtac_a_multithreaded_
vliw.

Flanagan:2002:MCM

[FQS02] Cormac Flanagan, Shaz Qadeer, and Sanjit A. Seshia. A modular checker for multithreaded programs. Lecture Notes in Computer Science, 2404:180–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2404/24040180.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2404/24040180.pdf.

Ferreira:1995:PAI

[FR95] Afonso Ferreira and Jose Rolim, editors. Parallel algorithms for irregularly structured problems: second international workshop, IRREGULAR 95, Lyon, France, September, 4–6, 1995: proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-60321-2. LCCN QA76.642.I59 1995.

Forsell:2018:RMM

[FRL18] Martti Forsell, Jussi Roivainen, and Ville Leppänen. REPLICA MBTAC: multithreaded dualmode processor. The Journal of Supercomputing, 74(5): 1911–1933, May 2018. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Field:1995:PPS

[FRT95] John Field, G. Ramalingam, and Frank Tip. Parametric program slicing. In ACM [ACM95b], pages 379–

392. ISBN 0-89791-692-1. LCCN QA 76.7 A11 1995.

URL http://www.acm.org:
80/pubs/citations/proceedings/
plan/199448/p379-field/.

ACM order number: 549950.

Fatouron:1996:SAS

[FS96] P. Fatouron and P. Spirakis. Scheduling algorithms for strict multithreaded computations. Lecture Notes in Computer Science, 1178:407– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feliu:2016:BAL

[FSPD16] J. Feliu, J. Sahuquillo, S. Petit, and J. Duato. Bandwidthaware on-line scheduling in SMT multicores. IEEE Transactions on Computers, 65(2): 422–434, ???? 2016. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feliu:2017:PFP

[FSPD17] J. Feliu, J. Sahuquillo, S. Petit, and J. Duato. Perf Fair: A progress-aware scheduler to enhance performance and fairness in SMT multicores. IEEE Transactions on Computers, 66(5):905–911, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feliu:2020:TII

[FSPE20] J. Feliu, J. Sahuquillo, S. Petit, and L. Eeckhout. Thread isolation to improve symbiotic scheduling on SMT multicore processors. IEEE Transactions on Parallel and Distributed Systems, 31(2):359– 373, February 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Factor:2006:PID

[FSS06] Michael Factor, Assaf Schuster, and Konstantin Shagin. A platform-independent distributed runtime for standard multithreaded Java. International Journal of Parallel Programming, 34(2):113–142, April 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=34&
issue=2&spage=113.

Fung:2009:DWF

[FSYA09] Wilson W. L. Fung, Ivan Sham, George Yuan, and Tor M. Aamodt. Dynamic warp formation: Efficient MIMD control flow on SIMD graphics hardware. ACM Transactions on Architecture and Code Optimization, 6(2): 7:1–7:??, June 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Farcy:1996:ISP

[FT96] A. Farcy and O. Temam.

Improving single-process performance with multithreaded processors. In ACM [ACM96], pages 350–357. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Fabregat-Traver:2014:SSG

[FTAB14] Diego Fabregat-Traver, Yurii S.

Aulchenko, and Paolo Bientinesi. Solving sequences of generalized least-squares problems on multi-threaded architectures. Applied Mathematics and Computation, 234 (??):606–617, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0096300314002951.

Feinbube:2011:JFM

[FTP11] Frank Feinbube, Peter Troger, and Andreas Polze. Joint forces: From multithreaded programming to GPU computing. IEEE Software, 28 (1):51–57, January/February 2011. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Fujita:1997:MPA

[Fuj97] Tetsuya Theodore Fujita. A multithreaded processor architecture for parallel symbolic computation. Technical Report MIT/LCS/TM-338, Laboratory for Computer Science, Massachusetts Institute of Technology, Cambridge, MA, USA, September 1997. 71 pp.

Flautner:2000:TLPa

[FURM00a] Kristián Flautner, Rich Uhlig, Steve Reinhardt, and Trevor Mudge. Thread-level parallelism and interactive performance of desktop applications. ACM SIGARCH Computer Architecture News, 28 (5):129–138, December 2000.

CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Flautner:2000:TLPc

[FURM00b] Kristián Flautner, Rich Uhlig, Steve Reinhardt, and Trevor Mudge. Thread-level parallelism and interactive performance of desktop applications. Operating Systems Review, 34(5):129–138, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Flautner:2000:TLPb

[FURM00c] Krisztián Flautner, Rich Uhlig, Steve Reinhardt, and Trevor Mudge. Thread level parallelism and interactive performance of desktop applications. ACM SIG-PLAN Notices, 35(11):129– 138, November 2000. CO-DEN SINODQ. ISBN 1-58113-317-0. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

http://delivery.acm.org/
10.1145/360000/357001/p129-
flautner.pdf.

Fang:2003:DGO

[FWL03] Weijian Fang, Cho-Li Wang, and Francis C. M. Lau. On the design of global object space for efficient multithreading Java computing on clusters. Parallel Computing, 29(11–12):1563–1587, November/December 2003. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Grant:2009:IEE

[GA09] Ryan E. Grant and Ahmad Afsahi. Improving energy efficiency of asymmetric chip multithreaded multiprocessors through reduced OS noise scheduling. Concurrency and Computation: Practice and Experience, 21(18):2355–2376, December 25, 2009. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guzzi:2014:CPP

[GAC14] P. H. Guzzi, G. Agapito, and M. Cannataro. coreSNP: Parallel processing of microarray data. IEEE Transactions on Computers, 63(12):2961– 2974, December 2014. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gallagher:1994:PLM

[Gal94] William Lynn Gallagher. Performance limitations of the MTS multithreaded architecture. Thesis (M.S. in Engineering), University of Texas at Austin, Austin, TX, USA, 1994. xiv + 101 pp.

Gao:1993:EHD

[Gao93] G. R. Gao. An efficient hybrid dataflow architecture model. Journal of Parallel and Distributed Computing, 19(4):293–307, December 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1113/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1113/production/pdf.

Garber:2001:NBT

[Gar01] Lee Garber. News briefs:

Is tech downturn changing education and employment trends; HTMT promises high-performance computing; controversial software law

[UCITA] hist resistance. Computer, 34(10):19–21, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2001/pdf/rx019.
pdf;   http://www.computer.
org/computer/co2001/rx019abs.
htm.

Giceva:2014:DQP

[GARH14] Jana Giceva, Gustavo Alonso, Timothy Roscoe, and Tim Harris. Deployment of query plans on multicores. Proceedings of the VLDB Endowment, 8(3):233–244, November 2014. CODEN ???? ISSN 2150-8097.

Greiner:1999:PTE

[GB99] John Greiner and Guy E. Blelloch. A provably time-efficient parallel implementation of full speculation. ACM Transactions on Programming Languages and Systems, 21(2): 240–285, March 1999. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/toplas/1999-21-
2/p240-greiner/.

Ghorbani:2020:RDT

[GB20] Mehrdad Ghorbani and Seyed Morteza Babamir. Runtime deadlock tracking and prevention of concurrent multithreaded programs: a learning-based approach. Concurrency and Computation: Practice and Experience, 32(10):e5324:1– e5324:??, May 25, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Giampapa:2005:BGA

[GBB^{+}05] M. E. Giampapa, R. Bellofatto, M. A. Blumrich, D. Chen, M. B. Dombrowa, A. Gara, R. A. Haring, P. Heidelberger, D. Hoenicke, G. V. Kopcsay, B. J. Nathanson, B. D. Steinmacher-Burow, M. Ohmacht, V. Salapura, and P. Vranas. Blue Gene/L advanced diagnostics environment. IBM Journal of Research and Development, 49 (2/):319–331, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/giampapa.
pdf.

Gotsman:2007:TMS

[GBCS07] Alexey Gotsman, Josh Berdine, Byron Cook, and Mooly Sagiv. Thread-modular shape analysis. ACM SIGPLAN Notices, 42(6):266–277, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gao:1995:ATD

[GBG95] Guang R. Gao, Lubomir Bic, and Jean-Luc Gaudiot. Advanced topics in dataflow computing and multithreading. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-6541-6 (hardcover), 0-8186-6540-8 (paperback), 0-8186-6542-4. x + 450 pp. LCCN QA76.9.A73 A356 1995.

Guz:2009:MCV

[GBK^{+}09] Zvika Guz, Evgeny Bolotin, Idit Keidar, Avinoam Kolodny, Avi Mendelson, and Uri C. Weiser. Many-core vs. manythread machines: Stay away from the valley. IEEE Computer Architecture Letters, 8(1):25–28, January/ June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Ghoting:2007:CCF

[GBP^{+}07] Amol Ghoting, Gregory Buehrer, Srinivasan Parthasarathy, Daehyun Kim, Anthony Nguyen, Yen-Kuang Chen, and Pradeep Dubey. Cache-conscious frequent pattern mining on modern and emerging processors. VLDB Journal: Very Large Data Bases, 16(1):77– 96, January 2007. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gokhale:1992:ICI

[GC92] Maya B. Gokhale and William W. Carlson. An introduction to compilation issues for parallel machines. Technical report SRC-TR-92-062, Supercomputing Research Center: IDA, Lanham, MD, USA, September 8, 1992. 38 pp.

Garcia:1999:MMI

[GCC99] F. Garcia, A. Calderon, and J. Carretero. MiMPI: a multithread-safe implementation of MPI. In Dongarra et al. [DLM99], pages 207–214. CO-DEN LNCSD9. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t1697.htm;   http:
//www.springerlink.com/
openurl.asp?genre=issue&
issn=0302-9743&volume=1697.

Ghosh:2015:NCC

[GCC15] Ammlan Ghosh, Rituparna Chaki, and Nabendu Chaki. A new concurrency control mechanism for multi-threaded environment using transactional memory. The Journal of Supercomputing, 71(11):4095– 4115, November 2015. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11227-015-1507-8;
 http://link.springer.com/
content/pdf/10.1007/s11227-
015-1507-8.pdf.

Georges:2004:JPR

[GCRD04] A. Georges, M. Christiaens, M. Ronsse, and K. De Bosschere. JaRec: a portable record/replay environment for multi-threaded Java applications. Software—Practice and Experience, 34(6):523–547, May 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gasiunas:2017:FBA

[GDSA^{+}17] Vaidas Gasiunas, David Dominguez-Sal, Ralph Acker, Aharon Avitzur, Ilan Bronshtein, Rushan Chen, Eli Ginot, Norbert Martinez-Bazan, Michael Müller, Alexander Nozdrin, Weijie Ou, Nir Pachter, Dima Sivov, and Eliezer Levy. Fiber-based architecture for NFV cloud databases. Proceedings of the VLDB Endowment, 10(12):1682–1693, August 2017. CODEN ???? ISSN 2150-8097.

Gravvanis:2008:JMB

[GE08] George A. Gravvanis and Victor N. Epitropou. Java multithreading-based parallel approximate arrowtype inverses. Concurrency and Computation: Practice and Experience, 20(10): 1151–1172, July 2008. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Geary:1998:SM

[Gea98] David Geary. Swing and multithreading. Java Report: The Source for Java Development, 3(??):??, November 1998. CODEN JREPFI. ISSN 1086-4660. URL http:/

                                        /archive.javareport.com/
                                        9811/html/from_pages/ftp_
                                        col1.shtml.

Gravvanis:2007:PPA

[GEG07] George A. Gravvanis, Victor N. Epitropou, and Konstantinos M. Giannoutakis. On the performance of parallel approximate inverse preconditioning using Java multithreading techniques. Applied Mathematics and Computation, 190(1):255–270, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Geiselbrecht:2001:NOS

[Gei01] Travis K. Geiselbrecht. The NewOS operating system. Dr. Dobb’s Journal of Software Tools, 26(12):33, 35, 38, 40, 42, 44, December 2001. CODEN DDJOEB.

ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2001/2001_12/newos.txt;
http://www.ddj.com/ftp/
2001/2001_12/newos.zip. See

correction [TLA^{+}02].

Geppert:2000:MBG

[Gep00] L. Geppert. Microprocessors: the off-beat generation. IEEE Spectrum, 37(7):44–49, July 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gerber:1995:IOX

[Ger95] Bob Gerber. Informix online XPS. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (2):463, May 1995. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gerbessiotis:2018:SIS

[Ger18] Alexandros V. Gerbessiotis.

A study of integer sorting on multicores. Parallel Processing Letters, 28(04):??, December 2018. ISSN 0129-6264 (print), 1793-642X (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S0129626418500147.

Garcia:2000:PTL

[GF00] Felix Garcia and Javier Fernandez. POSIX thread libraries. Linux Journal, 70:??, February 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue/3184.
html.

Gueunet:2019:TBA

[GFJT19] C. Gueunet, P. Fortin, J. Jomier, and J. Tierny. Task-based augmented contour trees with Fibonacci heaps. IEEE Transactions on Parallel and Distributed Systems, 30(8):1889–1905, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gao:1993:DMA

[GGB93a] Guang Gao, Jean-Luc Gaudiot, and Lubomir Bic. Dataflow and multithreaded architectures: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 18(3):271–??, July 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gao:1993:SID

[GGB93b] Guang R. Gao, Jean-Luc Gaudiot, and Lubomir Bic. Special issue on dataflow and multithreaded architectures. Journal of parallel and distributed computing; v. 18, no. 3. Academic Press, New York, USA, 1993. 271–389 pp.

Gruen:1998:NIS

[GH98] T. Gruen and M. A. Hillebrand. NAS integer sort on multi-threaded shared memory machines. Lecture Notes in Computer Science, 1470:999–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gagnon:2003:EIT

[GH03] E. Gagnon and L. Hendren.

Effective inline-threaded interpretation of Java bytecode using preparation sequences. Lecture Notes in Computer Science, 2622:170–184, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Girkar:1998:IIM

[GHG^{+}98] Milind Girkar, Mohammad R.

Haghighat, Paul Grey, Hideki Saito, Nicholas Stavrakos, and Constantine D. Polychronopoulos. Illinois-Intel multithreading library: Multithreading support for Intel architecture based multiprocessor systems. Intel Technology Journal, (Q1):15, 1998. ISSN 1535-766X. URL

http://developer.intel.
com/technology/itj/q11998/
articles/art_5.htm;   http:
//developer.intel.com/technology/
itj/q11998/pdf/iml.pdf.

Gibson:1994:CMC

[Gib94] Ken Gibson. A C++ multitasking class library. Dr. Dobb’s Journal of Software Tools, 19(5):28, 30, 32, 34, 96–98, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Gilbert:1988:DVN

[Gil88] P. D. Gilbert. Development of the VAX NOTES system. Digital Technical Journal, 1 (6):117–124, February 1988. CODEN DTJOEL. ISSN 0898-901X.

Gildea:1993:MTX

[Gil93]  Stephen Gildea.  Multi-

threaded Xlib. The X Resource, 5(1):159–166, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Giloi:1994:PSA

[Gil94] Wolfgang K. Giloi. Parallel supercomputer architectures and their programming models. Parallel Computing, 20 (10–11):1443–1470, November 3, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        parco/cas_sub/browse/browse.
                                        cgi?year=1994&volume=20&
                                        issue=10-11&aid=907.

Gorton:1997:GEI

[GJ97] Ian Gorton and Innes E. Jelly. Guest Editors introduction: Software engineering for parallel and distributed systems: Challenges and opportunities. IEEE Concurrency, 5(3):12–15, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

                                        //dlib.computer.org/pd/

books/pd1997/pdf/p3012.
pdf.

Ganesan:2011:MMP

[GJ11] Karthik Ganesan and Lizy K.

John. MAximum Multicore POwer (MAMPO): an automatic multithreaded synthetic power virus generation framework for multicore systems. In Lathrop et al. [LCK11], pages 53:1–53:12. ISBN 1-4503-0771-X. LCCN ????

Gebhart:2012:HTS

[GJT^{+}12] Mark Gebhart, Daniel R.

Johnson, David Tarjan, Stephen W. Keckler, William J. Dally, Erik Lindholm, and Kevin Skadron. A hierarchical thread scheduler and register file for energy-efficient throughput processors. ACM Transactions on Computer Systems, 30(2):8:1–8:??, April 2012. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Gerlhof:1994:MTA

[GK94] C. A. Gerlhof and A. Kemper. A multi-threaded architecture for prefetching in object bases. Lecture Notes in Computer Science, 779:351– 364, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia:2005:HJA

[GK05] P. Garcia and H. F. Korth. Hash-join algorithms on modern multithreaded computer architectures. Report LUCSE-05-001, Lehigh University, Bethlehem, PA, USA, ???? 2005.

Georgiou:2017:ETD

[GKCE17] Kyriakos Georgiou, Steve Kerrison, Zbigniew Chamski, and Kerstin Eder. Energy transparency for deeply embedded programs. ACM Transactions on Architecture and Code Optimization, 14(1):8:1– 8:??, April 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Granat:2009:NPQ

[GKK09] Robert Granat, Bo Kågström, and Daniel Kressner. A novel parallel QR algorithm for hybrid distributed memory HPC systems. LAPACK Working Note 216, Department of Computing Science and HPC2N, Umeå University, S-901 Umeå, Sweden, April 2009. URL http:

//www.netlib.org/lapack/
lawnspdf/lawn216.pdf.

Garland:2012:DUP

[GKZ12] Michael Garland, Manjunath Kudlur, and Yili Zheng. Designing a unified programming model for heterogeneous machines. In Hollingsworth [Hol12], pages 67:1–67:11. ISBN 1-4673-0804-8. URL http://

conferences.computer.org/
sc/2012/papers/1000a064.
pdf.

Gallmeister:1991:EEP

[GL91] Bill O. Gallmeister and Chris Lanier. Early experience with POSIX 1003.4 and POSIX 1003.4 A. Proceedings — Real-Time Systems Symposium, pages 190–198 (of ix + 307), 1991. CODEN PRSYEA. ISBN 0-8186-2450-7. IEEE catalog number 91CH3090-8.

Golla:1998:CMR

[GL98a] Prasad N. Golla and Eric C.

Lin. Cache memory requirements for multithreaded uniprocessor architecture. Technical paper 98-CSE-03, Dept. of Computer Science and Engineering, Southern Methodist University, Dallas, TX, USA, 1998. 32 pp.

Golla:1998:CEB

[GL98b] Prasad N. Golla and Eric C.

Lin. A comparison of the effect of branch prediction on multithreaded and scalar architectures. ACM SIGARCH Computer Architecture News, 26(4):3–11, September 1998. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Goldwasser:2007:INP

[GL07] Michael H. Goldwasser and David Letscher. Introducing network programming into a CS1 course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):19–22, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Gu:1999:EJT

[GLC99] Yan Gu, B. S. Lee, and Wentong Cai. Evaluation of Java thread performance on two different multithreaded kernels. Operating Systems Review, 33(1):34–46, January 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Glenn:1991:CMH

[Gle91] Ray R. Glenn. Characterizing memory hot spots in a shared memory MIMD machine. Technical report SRC-TR-91-039, Supercomputing Research Center: IDA, Lanham, MD, USA, October 15, 1991. 24 pp.

Grebenshchikov:2012:SSV

[GLPR12] Sergey Grebenshchikov, Nuno P. Lopes, Corneliu Popeea, and Andrey Rybalchenko. Synthesizing software verifiers from proof rules. ACM SIGPLAN Notices, 47(6):405–416, June 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PLDI ’12 proceedings.

Giering:1993:IAF

[GMB93] E. W. Giering, F. Mueller, and T. P. Baker. Implementing Ada 9X features using POSIX threads: Design issues. In ACM [ACM93c], pages 214–228. ISBN 0-89791-621-2. LCCN ???? ACM Order No. 825930.

Gonzalez-Mesa:2014:ETM

[GMGZP14] M. A. Gonzalez-Mesa, Eladio Gutierrez, Emilio L. Zapata, and Oscar Plata. Effective transactional memory execution management for improved concurrency. ACM Transactions on Architecture and Code Optimization, 11 (3):24:1–24:??, October 2014. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gomez:1998:CAM

[GMR98] J. C. Gomez, E. Mascarenhas, and V. Rego. The CLAM approach to multithreaded communication on shared memory multiprocessors: Design and experiments. IEEE Transactions on Parallel and Distributed Systems, 9(1):36–49, January 1998. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//dlib.computer.org/td/
books/td1998/pdf/l0036.
pdf;   http://www.computer.
org/tpds/td1998/l0036abs.
htm.

Ganty:2009:VLA

[GMR09] Pierre Ganty, Rupak Majumdar, and Andrey Rybalchenko. Verifying liveness for asynchronous programs. ACM SIGPLAN Notices, 44(1):102–113, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gabor:2009:SLA

[GMW09] Ron Gabor, Avi Mendelson, and Shlomo Weiss. Service level agreement for multithreaded processors. ACM Transactions on Architecture and Code Optimization, 6(2): 6:1–6:??, June 2009. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Govindarajan:1992:LCM

[GN92] R. Govindarajan and S. S. Nemawarkar. A large context multithreaded architecture. Lecture Notes in Computer Science, 634:423–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grunwald:1996:WPO

[GN96] Dirk Grunwald and Richard Neves. Whole-program optimization for time and space efficient threads. ACM SIG-PLAN Notices, 31(9):50–59, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://

www.acm.org:80/pubs/citations/
proceedings/asplos/237090/
p50-grunwald/. Co-published

as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Gopinath:2000:PSB

[GN00] K. Gopinath and M. K. Krishna Narasimhan. Performance of switch blocking on multithreaded architectures. J.UCS: Journal of Universal Computer Science, 6(10):928–947, October 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_10/
performance_of_switch_blocking.

Goeschl:2001:JTT

[Goe01] Siegfried Goeschl. The JUnit++ testing tool. Dr. Dobb’s Journal of Software Tools, 26(2):34, 36–38, February 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_02/junitpp.txt;   http:
//www.ddj.com/ftp/2001/
2001_02/junitpp.zip.

Goldwasser:1994:PAS

[Gol94] Shafi Goldwasser, editor. Proceedings: 35th Annual Symposium on Foundations of Computer Science, November 20– 22, 1994, Santa Fe, New Mexico. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Gollapudi:1996:MCA

[Gol96] Sreenivas Gollapudi. A multithreaded client-server architecture for distributed multimedia systems. Thesis (M.S.), Dept. of Computer Science, State University of New York at Buffalo, Buffalo, NY, USA, 1996. viii + 72 pp. Also available as technical report 96-13.

Goldstein:1997:LTC

[Gol97] Seth Copen Goldstein. Lazy threads: compiler and runtime structures for fine-grained parallel programming. Thesis (Ph.D.), Computer Science Division, University of California, Berkeley, Berkeley, CA, USA, 1997. xi + 174 pp.

Gonzalez:1990:MSC

[Gon90] Dean W. Gonzalez. Multitasking software components. ACM SIGADA Ada Letters, 10(1):92–96, January/February 1990. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Goossens:1997:MVC

[Goo97] B. Goossens. A multithreaded vector co-processor. Lecture Notes in Computer Science, 1277:311–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gould:2003:GLT

[GOT03] Nicholas I. M. Gould, Dominique Orban, and Philippe L. Toint. GALAHAD, a library of thread-safe Fortran 90 packages for large-scale nonlinear optimization. ACM Transactions on Mathematical Software, 29(4):353–372, December 2003. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Girkar:1995:ETL

[GP95] Milind Girkar and Constantine D. Polychronopoulos. Extracting task-level parallelism. ACM Transactions on Programming Languages and Systems, 17(4):600–634, July 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0164-0925/
210189.html.

Gil:2005:TCS

[GP05] Marisa Gil and Ruben Pinilla.

Thread coloring: a scheduler proposal from user to hardware threads. Operating Systems Review, 39(2):

54–70, April 2005. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gidenstam:2008:LLF

[GP08] Anders Gidenstam and Marina Papatriantafilou. LFTHREADS: a lock-free thread library. ACM SIGARCH Computer Architecture News, 36(5):88– 92, December 2008. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Gupta:2017:DDP

[GPB^{+}17] Ujjwal Gupta, Chetan Arvind Patil, Ganapati Bhat, Prabhat Mishra, and Umit Y. Ogras. DyPO: Dynamic Pareto-optimal configuration selection for heterogeneous MpSoCs. ACM Transactions on Embedded Computing Systems, 16(5s):123:1–123:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gupta:2011:PAR

[GPR11] Ashutosh Gupta, Corneliu Popeea, and Andrey Rybalchenko. Predicate abstraction and refinement for verifying multi-threaded programs. ACM SIGPLAN Notices, 46(1):331–344, January 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gerakios:2014:SSG

[GPS14] Prodromos Gerakios, Nikolaos Papaspyrou, and Konstantinos Sagonas. Static safety guarantees for a lowlevel multithreaded language with regions. Science of Computer Programming, 80(??): 223–263, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167642313001433.

Grossman:2003:TSM

[Gro03]  Dan Grossman.  Type-

safe multithreading in cyclone. ACM SIGPLAN Notices, 38(3):13–25, March 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gomez:2006:STC

[GRR06] Juan Carlos Gomez, Jorge R.

Ramos, and Vernon Rego. Signals, timers, and continuations for multithreaded userlevel protocols. Software— Practice and Experience, 36 (5):449–471, April 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gomez:1997:EMU

[GRS97] Juan Carlos Gomez, Vernon Rego, and V. S. Sunderam. Efficient multithreaded user-space transport for network computing: Design and test of the TRAP protocol. Journal of Parallel and Distributed Computing, 40(1):103–117, January 10, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1269/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1269/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1269/production/ref.

Gomez:2006:SCM

[GRS06] Juan Carlos Gomez, Vernon Rego, and V. S. Sunderam. Scheduling communication in multithreaded programs: experimental results. Concurrency and Computation: Practice and Experience, 18 (1):1–28, January 2006. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gontmakher:2000:JCN

[GS00] Alex Gontmakher and Assaf Schuster. Java consistency: nonoperational characterizations for Java memory behavior. ACM Transactions on Computer Systems, 18(4):333–386, 2000. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic). URL

http://www.acm.org/pubs/

articles/journals/tocs/
2000-18-4/p333-gontmakher/
p333-gontmakher.pdf;   http:
//www.acm.org/pubs/citations/
journals/tocs/2000-18-4/
p333-gontmakher/.

Garg:2002:TOA

[GS02] Rajat P. Garg and Ilya Sharapov. Techniques for optimizing applications: high performance computing. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2002. ISBN 0-13-093476-3. xliii + 616 pp. LCCN QA76.88 .G37 2002.

URL http://www.sun.com/
books/catalog/garg.html/
index.html.

Grelck:2006:SFA

[GS06] Clemens Grelck and Sven-Bodo Scholz. SAC — a functional array language for efficient multi-threaded execution. International Journal of Parallel Programming, 34(4):383–427, August 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=34&
issue=4&spage=383.

Goldstein:1996:LTI

[GSC96] Seth Copen Goldstein, Klaus Erik Schauser, and David E. Culler. Lazy threads: Implementing a fast parallel call. Journal of Parallel and Distributed Computing, 37(1):5–20, August 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0103/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0103/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0104/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0104/production/pdf.

Gu:2018:CCA

[GSK^{+}18] Ronghui Gu, Zhong Shao, Jieung Kim, Xiongnan (Newman) Wu, Jérémie Koenig, Vilhelm Sjöberg, Hao Chen, David Costanzo, and Tahina Ramananandro. Certified concurrent abstraction layers. ACM SIGPLAN Notices, 53(4):646–661, April 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gupta:2010:CSM

[GSL10] M. Gupta, F. Sanchez, and J. Llosa. CSMT: Simultaneous multithreading for clustered VLIW processors. IEEE Transactions on Computers, 59(3):385–399, March 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=5161255.

Gulati:1995:MSM

[Gul95] Manu Gulati. Multithreading on a superscalar microprocessor. Thesis (M.S., Engineering), University of California, Irvine, Irvine, CA, USA, 1995. x + 102 pp.

Gunther:1997:MDF

[Gun97]  B. K. Gunther.  Multi-

threading with distributed functional units. IEEE Transactions on Computers, 46(4):399–411, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=588034.

Gustafsson:2005:TP

[Gus05] Andreas Gustafsson. Threads without the pain. ACM Queue: Tomorrow’s Computing Today, 3(9):42–47, November 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Goossens:1995:FPM

[GV95] B. Goossens and D. T. Vu.

Further pipelining and multithreading to improve RISC processor speed. A proposed architecture and simulation results. Lecture Notes in Computer Science, 964:326– ??, 1995. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Georgakoudis:2017:SSA

[GVT^{+}17] Giorgis Georgakoudis, Hans Vandierendonck, Peter Thoman, Bronis R. De Supinski, Thomas Fahringer, and Dimitrios S. Nikolopoulos. SCALO: Scalability-aware parallelism orchestration for multi-threaded workloads. ACM Transactions on Architecture and Code Optimization, 14(4):54:1–54:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Gibson:2010:FSC

[GW10] Dan Gibson and David A. Wood. Forwardflow: a scalable core for powerconstrained CMPs. ACM SIGARCH Computer Architecture News, 38(3):14– 25, June 2010. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Gabor:2007:FES

[GWM07] Ron Gabor, Shlomo Weiss, and Avi Mendelson. Fairness enforcement in switch on event multithreading. ACM Transactions on Architecture and Code Optimization, 4(3): 15:1–15:??, September 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Haggar:2002:JQD

[Hag02] Peter Haggar. Java Q&A:

Does Java guarantee thread safety? Dr. Dobb’s Journal of Software Tools, 27 (6):91–83, June 2002. CO-DEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2002/
2002_06/jqa0602.txt. Com-

ments on lack of atomicupdate guarantee in Java for objects larger than 32 bits, such as long and double, with sample code to exhibit the failure.

Haines:1997:DLT

[Hai97a] Matthew Haines. On designing lightweight threads for substrate software. Washington, DC, USA, 1997. ?? pp. Shipping list number 98-0847-M.

Haines:1997:OIA

[Hai97b] Matthew Haines. An open implementation analysis and design for lightweight threads. ACM SIGPLAN Notices, 32 (10):229–242, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hamilton:1996:JSN

[Ham96] Marc A. Hamilton. Java and the shift to net-centric computing. Computer, 29(8): 31–39, August 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Mentions Java’s use of Unicode characters.

Hanson:1997:CII

[Han97] David R. Hanson. C Interfaces and Implementations: Techniques for Creating Reusable Software. Addison-Wesley Professional Computing Series. Addison-Wes-ley, Reading, MA, USA, 1997. ISBN 0-201-49841-3. xvii + 519 pp. LCCN QA76.73.C15H37 1997. US$37.95.

URL http://www.cs.princeton.
edu/software/cii/.

Harrington:1999:WMM

[Har99]  John Harrington.  Win32

multithreading made easy. C/C++ Users Journal, 17 (8):48, 50–52, 54–56, August 1999. CODEN CCUJEX. ISSN 1075-2838.

Hayden:1993:BIC

[Hay93] Charles Hayden. A brief introduction to Concurrent Pascal. ACM SIGPLAN Notices, 28(3):353–354, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/plan/154766/
p353-hayden/.

Haines:1992:SMC

[HB92] Matt Haines and Anton Pedro Willem Bohm. Software multithreading in a conventional distributed memory multiprocessor. Technical report CS-92-126, Colorado State University, Dept. of Computer Science, Fort Collins, CO, USA, September 25, 1992. 25 pp.

Hottelier:2015:SLE

[HB15] Thibaud Hottelier and Rastislav Bodik. Synthesis of layout engines from relational constraints. ACM SIGPLAN Notices, 50(10):74–88, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hunt:2013:DTN

[HBCG13] Nicholas Hunt, Tom Bergan, Luis Ceze, and Steven D. Gribble. DDOS: taming nondeterminism in distributed systems. ACM SIGPLAN Notices, 48(4):499–508, April 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hanson:2001:UFI

[HBG01] Richard J. Hanson, Clay P.

Breshears, and Henry A. Gabb. Using a Fortran interface to POSIX threads. In Boisvert and Tang [BT01], pages 257–272. ISBN 0-7923-7339-1. LCCN QA76.758 .I345 2000.

Hanson:2002:AFI

[HBG02] Richard J. Hanson, Clay P.

Breshears, and Henry A. Gabb. Algorithm 821: a Fortran interface to POSIX threads. ACM Transactions on Mathematical Software, 28 (3):354–371, September 2002. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Heber:1998:UMA

[HBTG98] G. Heber, R. Biswas, P. Thulasiraman, and G. R. Gao. Using multithreading for the automatic load balancing of adaptive finite element meshes. Lecture Notes in Computer Science, 1457:132– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hankendi:2017:SCS

[HC17] Can Hankendi and Ayse Kivilcim Coskun. Scale & cap: Scaling-aware resource management for consolidated multi-threaded applications. ACM Transactions on Design Automation of Electronic Systems., 22(2):30:1– 30:??, March 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Halstead:1994:PCR

[HCD^{+}94] Burt Halstead, David Callahan, Jack Dennis, R. S. Nikhil, and Vivek Sarkar. Programming, compilation, and resource management issues for multithreading (panel session II). ACM SIGARCH Computer Architecture News, 22 (1):19–33, March 1994. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Haines:1994:DCT

[HCM94] Matthew Haines, David Cronk, and Piyush Mehrotra. On the design of chant: a talking threads of package: final report. Washington, DC, USA, 1994. ?? pp. Shipping list number 94-0861-M.

Ding:2002:MOP

[HD02] Yun He and Chris H. Q. Ding.

MPI and OpenMP paradigms on cluster of SMP architectures. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL

http://www.sc-2002.org/
paperpdfs/pap.pap325.pdf.

Honarmand:2013:CUA

[HDT^{+}13] Nima Honarmand, Nathan Dautenhahn, Josep Torrellas, Samuel T. King, Gilles Pokam, and Cristiano Pereira. Cyrus: unintrusive application-level record-replay for replay parallelism. ACM SIGPLAN Notices, 48(4):193– 206, April 2013. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heinlein:2003:ATS

[Hei03] C. Heinlein. Advanced thread synchronization in Java using interaction expressions. Lecture Notes in Computer Science, 2591:345–365, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffman:2009:SAT

[HEJ09] Kevin J. Hoffman, Patrick Eugster, and Suresh Jagannathan. Semantics-aware trace analysis. ACM SIG-PLAN Notices, 44(6):453–464, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hroub:2017:EGC

[HEMK17] Ayman Hroub, M. E. S. Elrabaa, M. F. Mudawar, and A. Khayyat. Efficient generation of compact execution traces for multicore architectural simulations. ACM Transactions on Architecture and Code Optimization, 14(3): 27:1–27:??, September 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Halstead:1988:MMP

[HF88] R. H. Halstead, Jr. and T. Fujita. MASA: a multithreaded processor architecture for parallel symbolic computing. ACM SIGARCH Computer Architecture News, 16(2):443–451, May 1988. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Hertzum:1996:BQO

[HF96] Morten Hertzum and Erik Frøkjær. Browsing and querying in online documentation: a study of user interfaces and the interaction process. ACM Transactions on Computer-Human Interaction, 3(2):136–161, June 1996. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tochi/1996-3-2/
p136-hertzum/.

Halappanavar:2012:AWM

[HFV^{+}12] Mahantesh Halappanavar, John Feo, Oreste Villa, Antonino Tumeo, and Alex Pothen. Approximate weighted matching on emerging manycore and multithreaded architectures. The International Journal of High Performance Computing Applications, 26 (4):413–430, November 2012. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://

hpc.sagepub.com/content/
26/4/413.full.pdf+html.

Hum:1991:NHS

[HG91] H. H. J. Hum and G. R.

Gao. A novel high-speed memory organization for finegrain multi-thread computing. Lecture Notes in Computer Science, 505:34–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hum:1992:HSM

[HG92] Herbert H. J. Hum and Guang R. Gao. A high-speed memory organization for hybrid dataflow/von Neumann computing. Future Generation Computer Systems, 8(4):287– 301, September 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hughes:1997:OOM

[HH97] Cameron Hughes and Tracey Hughes. Object-oriented multithreading using C++. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1997. ISBN 0-471-18012-2 (paperback). xvi + 495 pp. LCCN QA76.73.C153H84 1997.

Hong:2011:AMA

[HH11] Bo Hong and Zhengyu He. An asynchronous multithreaded algorithm for the maximum network flow problem with nonblocking global relabeling heuristic. IEEE Transactions on Parallel and Distributed Systems, 22(6):1025– 1033, June 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2016:MCR

[HH16] Shiyou Huang and Jeff Huang. Maximal causality reduction for TSO and PSO. ACM SIG-PLAN Notices, 51(10):447– 461, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hironaka:1991:SVP

[HHOM91] T. Hironaka, T. Hashimoto, K. Okazaki, and K. Murakami. A single-chip vectorprocessor prototype based on multithreaded streaming/ FIFO (MSFV) architecture. In Anonymous [Ano91], pages 77–86. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Hironaka:1992:BVP

[HHOM92] T. Hironaka, T. Hashimoto, K. Okazaki, and K. Murakami. Benchmarking a vector-processor prototype based on multithreaded streaming/FIFO vector (MSFV) architecture. In ACM [ACM92], pages 272–281. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Hussein:2015:DRM

[HHPV15] Ahmed Hussein, Antony L.

Hosking, Mathias Payer, and Christopher A. Vick. Don’t race the memory bus: taming the GC leadfoot. ACM SIG-PLAN Notices, 50(11):15–27, November 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hickey:2020:HC

[Hic20] Rich Hickey. A history of Clojure. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL): 71:1–71:46, June 2020. URL

https://dl.acm.org/doi/
abs/10.1145/3386321.

Hightower:1997:PDD

[Hig97] Lauren Hightower. Publishing dynamic data on the Internet — Allaire’s Cold Fusion is a development tool that provides access (via the Web) to any database the Web server can access using ODBC. Cold Fusion runs as a multithreaded Windows NT system service and works with any ODBC-compliant database. Dr. Dobb’s Journal of Software Tools, 22(1): 70–??, January 1997. CODEN DDJOEB. ISSN 1044-789X.

Hauser:1993:UTI

[HJT^{+}93] Carl Hauser, Christian Jacobi, Marvin Theimer, Brent Welch, and Mark Weiser. Using threads in interactive systems: a case study. Operating Systems Review, 27(5): 94–105, December 1993. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hirata:1992:EPA

[HKN^{+}92] Hiroaki Hirata, Kozo Kimura, Satoshi Nagamine, Yoshiyuki Mochizuki, Akio Nishimura, Yoshimori Nakase, and Teiji Nishizawa. An elementary processor architecture with simultaneous instruction issuing from multiple threads. ACM SIGARCH Computer Architecture News, 20(2): 136–145, May 1992. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Hurson:1996:CMD

[HKSL96] A. R. Hurson, Krishna M.

Kavi, Behrooz Shirazi, and Ben Lee. Cache memories for dataflow systems. IEEE parallel and distributed technology: systems and applications, 4(4):50–64, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http:

//dlib.computer.org/pd/
books/pd1996/pdf/p4050.
pdf;   http://www.computer.
org/concurrency/pd1996/
p4050abs.htm.

Hidaka:1993:MTC

[HKT93] Yasuo Hidaka, Hanpei Koike, and Hidehiko Tanaka. Multiple threads in cyclic register windows. ACM SIGARCH Computer Architecture News, 21(2):131–142, May 1993. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Huelsbergen:1993:CCG

[HL93] Lorenz Huelsbergen and James R.

Larus. A concurrent copying garbage collector for languages that distinguish (im)mutable data. ACM SIG-PLAN Notices, 28(7):73–82, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hur:2007:MSM

[HL07] Ibrahim Hur and Calvin Lin. Memory scheduling for modern microprocessors. ACM Transactions on Computer Systems, 25(4):10:1– 10:??, December 2007. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

He:2008:COD

[HL08] Bingsheng He and Qiong Luo. Cache-oblivious databases: Limitations and opportunities. ACM Transactions on Database Systems, 33(2): 8:1–8:??, June 2008. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Hansen:1990:EPA

[HLB90] G. J. Hansen, C. A. Linthicum, and G. Brooks. Experience with a performance analyzer for multithreaded applications. In IEEE [IEE90], pages 124–131. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Holm:1994:CSP

[HLB94] J. Holm, A. Lain, and P. Banerjee. Compilation of scientific programs into multithreaded and message driven computation. In IEEE [IEE94b], pages 518– 525. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Herdt:2019:CSB

[HLGD19] Vladimir Herdt, Hoang M.

Le, Daniel Große, and Rolf Drechsler. Combining sequentialization-based verification of multi-threaded C programs with symbolic Partial Order Reduction. International Journal on Software Tools for Technology Transfer (STTT), 21(5):545– 565, October 2019. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10009-019-00507-5.

Hu:2016:TDM

[HLH16] Qi Hu, Peng Liu, and Michael C. Huang. Threads and data mapping: Affinity analysis for traffic reduction. IEEE Computer Architecture Letters, 15(2):133–136, July/ December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Helmbold:1996:TRC

[HM96] D. P. Helmbold and C. E. McDowell. A taxonomy of race conditions. Journal of Parallel and Distributed Computing, 33(2):159–164, March 15, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0034/production;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0034/production/pdf.

Haines:1995:RSC

[HMC95] Matthew Haines, Piyush Mehrotra, and David Cronk. Ropes, support for collective operations among distributed threads. Washington, DC, USA, 1995. ?? pp. Shipping list number 96-0037-M.

Haines:1997:DPP

[HMC97] Matthew Haines, Piyush Mehrotra, and David Cronk. Data-parallel programming in a multithreaded environment. Scientific Programming, 6(2): 187–200, Summer 1997. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hashemi:2016:EEB

[HMCP16] Milad Hashemi, Debbie Marr, Doug Carmean, and Yale N. Patt. Efficient execution of bursty applications. IEEE Computer Architecture Letters, 15(2):85–88, July/ December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Harish:2016:PIK

[HMLB16] Pawan Harish, Mentar Mahmudi, Benoît Le Callennec, and Ronan Boulic. Parallel inverse kinematics for multithreaded architectures. ACM Transactions on Graphics, 35(2):19:1–19:??, May 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hirata:1992:MPA

[HMN^{+}92] H. Hirata, Y. Mochizuki, A. Nishmura, Y. Nakase, and T. Nishizawa. A multithreaded processor architecture with simultaneous instruction issuing. Supercomputer, 9(3):23–39, May 1992. CODEN SPCOEL. ISSN 0168-7875.

Hirata:1991:MPA

[HMNN91] H. Hirata, Y. Mochizuki, A. Nishimura, and Y. Nakase. A multithreaded processor architecture with simultaneous instruction issuing. In Anonymous [Ano91], pages 87–96. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Hum:1996:SEM

[HMT^{+}96] Herbert H. J. Hum, Olivier Maquelin, Kevin B. Theobald, Xinmin Tian, Guang R. Gao, and Laurie J. Hendren. A study of the EARTH-MANNA multithreaded system. International Journal of Parallel Programming, 24(4): 319–348, August 1996. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Horiguchi:1991:PEP

[HN91] Susumu Horiguchi and Takeo Nakada. Performance evaluation of parallel fast Fourier transform on a multiprocessor workstation. Journal of Parallel and Distributed Computing, 12(2):158–163, June 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Holub:1998:PJTb

[Hol98a] Allen Holub. Programming Java threads in the real world, Part 2: Common multithreading pitfalls (deadlock, etc.). JavaWorld: IDG’s magazine for the Java community, 3 (10):??, 1998. CODEN ???? ISSN 1091-8906. URL http:

//www.holub.com/goodies/
javaworld/jw_index.html;
 http://www.javaworld.com/
javaworld/jw-10-1998/jw-
10-toolbox.htm.

Holub:1998:PJTc

[Hol98b] Allen Holub. Programming Java threads in the real world, Part 3: Semaphore, lock manager, and mutex. JavaWorld: IDG’s magazine for the Java community, 3 (11):??, 1998. CODEN ????

ISSN 1091-8906. URL http:

//www.holub.com/goodies/
javaworld/jw_index.html;
 http://www.javaworld.com/
javaworld/jw-11-1998/jw-
11-toolbox.htm.

Holub:1998:PJTd

[Hol98c] Allen Holub. Programming Java threads in the real world, Part 4: Condition variables and counting semaphores. JavaWorld: IDG’s magazine for the Java community, 3 (12):??, 1998. CODEN ???? ISSN 1091-8906. URL http:

//www.holub.com/goodies/
javaworld/jw_index.html;
 http://www.javaworld.com/
javaworld/jw-12-1998/jw-
12-toolbox.htm.

Holub:1998:PJTa

[Hol98d] Allen Holub. Programming Java threads in the real world: Threading architectures. JavaWorld: IDG’s magazine for the Java community, 3(9):??, September 1998. CODEN ???? ISSN 1091-8906. URL http:

//www.holub.com/goodies/
javaworld/jw_index.html;
 http://www.javaworld.com/
javaworld/jw-09-1998/jw-
09-threads.htm.

Holub:1999:PJTa

[Hol99a] Allen Holub. Programming Java threads in the real world, Part 5: Timers. JavaWorld: IDG’s magazine for the Java community, 4(2):??, February 1999. CODEN ????

ISSN 1091-8906. URL http:

                                        //www.holub.com/goodies/
                                        javaworld/jw_index.html.

Holub:1999:PJTb

[Hol99b] Allen Holub. Programming Java threads in the real world, Part 6: Mach ’99: Observer and the mysteries of the AWTEventMulticaster. Java-World: IDG’s magazine for the Java community, 4(3):??, March 1999. CODEN ???? ISSN 1091-8906. URL http:

                                        //www.holub.com/goodies/
                                        javaworld/jw_index.html.

Holub:2000:TJT

[Hol00]  Allen I. Holub.  Taming

Java Threads. Apress, Berkeley, CA, USA, 2000. ISBN 1-893115-10-0. x + 300 pp. LCCN QA76.73.J38 H635 2000. US$34.95.

Hollingsworth:2012:SPI

[Hol12] Jeffrey Hollingsworth, editor. SC ’12: Proceedings of the International Conference on High Performance Computing, Networking, Storage and Analysis, Salt Lake Convention Center, Salt Lake City, UT, USA, November 10–16, 2012. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-0804-8.

Hong:1994:FIS

[Hon94] Hoon Hong, editor. First International Symposium on Parallel Symbolic Computation, PASCO ’94, Hagenberg/ Linz, Austria, September 26– 28, 1994, volume 5 of Lecture notes series in computing. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1994. ISBN 981-02-2040-5. LCCN QA76.642.I58 1994.

Hopper:1998:CFM

[Hop98]  Michael A. Hopper.  A

compiler framework for multithreaded parallel systems. Thesis (Ph.D.), School of Electrical and Computer Engineering, Georgia Institute of Technology, Atlanta, GA, USA, 1998. xii + 110 pp.

Howes:1998:TPC

[How98] Brad Howes. Template processing classes for Python. Dr. Dobb’s Journal of Software Tools, 23(2):38, 40, 42, 44–46, 48, 100, February 1998. CO-DEN DDJOEB. ISSN 1044-789X.

Howard:2000:UPW

[How00] David M. Howard. Using predicate waits with Win32 threads. C/C++ Users Journal, 18(5):18–??, May 2000. CODEN CCUJEX. ISSN 1075-2838.

Halappanavar:2015:CLL

[HPA^{+}15] Mahantesh Halappanavar, Alex Pothen, Ariful Azad, Fredrik Manne, Johannes Langguth, and Arif Khan.

Codesign lessons learned from implementing graph matching on multithreaded architectures. Computer, 48(8): 46–55, August 2015. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
co/2015/08/mco2015080046-
abs.html.

Hsu:2011:MSS

[HPB11] Chia-Jui Hsu, José Luis Pino, and Shuvra S. Bhattacharyya. Multithreaded simulation for synchronous dataflow graphs. ACM Transactions on Design Automation of Electronic Systems., 16(3):25:1–25:??, June 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hilton:2010:SDE

[HR10] Andrew Hilton and Amir Roth. SMT-Directory: Efficient load-load ordering for SMT. IEEE Computer Architecture Letters, 9(1):25–28, January/June 2010. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Huang:2016:PMR

[HR16] Jeff Huang and Arun K. Rajagopalan. Precise and maximal race detection from incomplete traces. ACM SIG-PLAN Notices, 51(10):462– 476, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hassanein:2008:AEH

[HRH08] Wessam M. Hassanein, Layali K. Rashid, and Moustafa A. Hammad. Analyzing the effects of hyperthreading on the performance of data management systems. International Journal of Parallel Programming, 36(2):206–225, April 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=36&
issue=2&spage=206.

Hayden:2012:KEG

[HSD^{+}12] Christopher M. Hayden, Edward K. Smith, Michail Denchev, Michael Hicks, and Jeffrey S. Foster. Kitsune: efficient, general-purpose dynamic software updating for C. ACM SIGPLAN Notices, 47(10):249–264, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hayden:2014:KEG

[HSS^{+}14] Christopher M. Hayden, Karla Saur, Edward K. Smith, Michael Hicks, and Jeffrey S. Foster. Kitsune: Efficient, general-purpose dynamic software updating for C. ACM Transactions on Programming Languages and Systems, 36 (4):13:1–13:??, October 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Honarmand:2014:RRR

[HT14] Nima Honarmand and Josep Torrellas. RelaxReplay: record and replay for relaxedconsistency multiprocessors. ACM SIGARCH Computer Architecture News, 42(1):223– 238, March 2014. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Hukerikar:2018:RIA

[HTDL18] Saurabh Hukerikar, Keita Teranishi, Pedro C. Diniz, and Robert F. Lucas. RedThreads: An interface for application-level fault detection/correction through adaptive redundant multithreading. International Journal of Parallel Programming, 46(2): 225–251, April 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Hendren:1997:CCE

[HTZ^{+}97] Laurie J. Hendren, Xinan Tang, Yingchun Zhu, Shereen Ghobrial, Guang R. Gao, Xun Xue, Haiying Cai, and Pierre Ouellet. Compiling C for the EARTH multithreaded architecture. International Journal of Parallel Programming, 25(4):305–338, August 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Huber:2001:EFC

[Hub01] Andreas Huber. Elegant function call wrappers. C/C++ Users Journal, 19(5):8–??, May 2001. CODEN CCUJEX. ISSN 1075-2838.

Hudson:1996:MDA

[Hud96] Greg Hudson. Multithreaded design in the Athena environment. Thesis (M. Eng.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1996. 240 pp.

Halladay:1992:PUM

[HW92] Steve Halladay and Michael Wiebel. A practical use for multiple threads. C Users Journal, 10(1):73–??, January 1992. ISSN 0898-9788.

Hsieh:1993:CME

[HWW93] Wilson C. Hsieh, Paul Wang, and William E. Weihl. Computation migration: enhancing locality for distributedmemory parallel systems. ACM SIGPLAN Notices, 28(7):239–248, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Horwood:2000:DMA

[HWZ00] Peter Horwood, Shlomo Wygodny, and Martin Zardecki. Debugging multithreaded applications. Dr. Dobb’s Journal of Software Tools, 25 (3):32, 34–37, March 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_03/dbgmulti.txt.

Hyde:2000:JTP

[Hyd00] Paul Hyde. Java thread programming. Howard W. Sams, Indianapolis, IN 46268, USA, 2000. ISBN 0-672-31585-8. iv + 510 pp. LCCN QA76.73.J38 H93 1999.

Huang:2015:COM

[HYY^{+}15] Kai Huang, Min Yu, Rongjie Yan, Xiaomeng Zhang, Xiaolang Yan, Lisane Brisolara, Ahmed Amine Jerraya, and Jiong Feng. Communication optimizations for multithreaded code generation from Simulink models. ACM Transactions on Embedded Computing Systems, 14(3):59:1–59:??, May 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2012:EPS

[HZ12] Jeff Huang and Charles Zhang. Execution privatization for scheduler-oblivious concurrent programs. ACM SIGPLAN Notices, 47(10): 737–752, October 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huang:2013:CRL

[HZD13] Jeff Huang, Charles Zhang, and Julian Dolby. CLAP: recording local executions to reproduce concurrency failures. ACM SIGPLAN Notices, 48(6):141–152, June 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Iannucci:1994:MCA

[I^{+}94] Robert A. Iannucci et al., editors. Multithreaded computer architecture: a summary of the state of the art, volume SECS 0281 of The Kluwer international series in engineering and computer science. Kluwer Academic Publishers, Dordrecht, The Netherlands; Boston, MA, USA, 1994. ISBN 0-7923-9477-1. xvi + 400 pp. LCCN QA76.9.A73 M85 1994.

Iannucci:1994:AII

[IAD^{+}94] Robert Iannucci, Anant Agarwal, Bill Dally, Anoop Gupta, Greg Papadopoulos, and Burton Smith. Architectural and implementation issues for multithreading (panel session I). ACM SIGARCH Computer Architecture News, 22 (1):3–18, March 1994. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Iwama:2001:ICB

[IBST01] Chitaka Iwama, Niko Demus Barli, Shuichi Sakai, and Hidehiko Tanaka. Improving conditional branch prediction on speculative multithreading architectures. Lecture Notes in Computer Science, 2150:413–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2150/21500413.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2150/21500413.
pdf.

Illikkal:2010:PQP

[ICH^{+}10] Ramesh Illikkal, Vineet Chadha, Andrew Herdrich, Ravi Iyer, and Donald Newell. PIRATE: QoS and performance management in CMP architectures. ACM SIGMETRICS Performance Evaluation Review, 37(4):3–10, March 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

IEEE:1989:WOS

[IEE89] IEEE, editor. Workstation Operating Systems: Proceedings of the Second Workshop on Workstation Operating Systems (WWOS-II), Pacific Grove, CA, USA, September 27–29, 1989. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. IEEE catalog number 89TH0281-6.

IEEE:1990:PSN

[IEE90] IEEE, editor. Proceedings, Supercomputing ’90: November 12–16, 1990, New York Hilton at Rockefeller Center, New York, New York. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-2056-0 (paperback: IEEE Computer Society), 0-89791-412-0 (paperback: ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

IEEE:1992:PSM

[IEE92] IEEE Computer Society.

Technical Committee on Computer Architecture, editor. Proceedings, Supercomputing ’92: Minneapolis, Minnesota, November 16-20, 1992. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

IEEE:1993:PSP

[IEE93] IEEE, editor. Proceedings, Supercomputing ’93: Portland, Oregon, November 15– 19, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

IEEE:1994:PIW

[IEE94a]  IEEE, editor.  Proceed-

ings 11th IEEE Workshop on Real-Time Operating Systems and Software. RTOSS ’94, Seattle, WA, USA, 18– 19 May 1994. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5710-3. LCCN QA76.54.I173 1994. IEEE catalog number 94TH0639-5.

IEEE:1994:PSH

[IEE94b]  IEEE, editor.  Proceed-

ings of the Scalable High-Performance Computing Conference, May 23–25, 1994, Knoxville, Tennessee. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

IEEE:1994:PSW

[IEE94c]  IEEE, editor.  Proceed-

ings, Supercomputing ’94: Washington, DC, November 14–18, 1994, Supercomputing. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

IEEE:1994:ROS

[IEE94d] IEEE, editor. Real-time operating systems and software: RTOSS ’94: 11th Workshop — May 1994, Seattle, WA, IEEE Workshop on Real Time Operating Systems and Software 1994; 11th. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5710-3. LCCN QA76.54.I173 1994.

IEEE:1995:PCL

[IEE95] IEEE Computer Society.

Technical Committee on Computer Communications, editor. Proceedings: 20th Conference on Local Computer Networks, October 16– 19, 1995, Minneapolis, Minnesota. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

IEEE:1996:PSM

[IEE96] IEEE, editor. Proceedings. Second MPI Developer’s Conference: Notre Dame, IN, USA, 1–2 July 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

IEEE:1997:APD

[IEE97] IEEE, editor. Advances in parallel and distributed computing: March 19–21, 1997, Shanghai, China: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7876-3 (paperback and case), 0-8186-7878-X (microfiche). LCCN QA76.58 .A4 1997.

IEEE:1999:HCS

[IEE99] IEEE, editor. Hot Chips 11: Stanford University, Stanford, California, August 15–17, 1999. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN ???? LCCN ???? URL http://

                                        www.hotchips.org/hotc11_
                                        index.html.

IEEE:2002:STI

[IEE02] IEEE, editor. SC2002: From Terabytes to Insight. Proceedings of the IEEE ACM SC 2002 Conference, November 16–22, 2002, Baltimore, MD, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1524-X. LCCN ????

Iwata:2001:PMT

[III01] Kazunori Iwata, Shingo Itabashi, and Naohiro Ishii. A protocol for multi-threaded processes with choice in πcalculus. Lecture Notes in Computer Science, 2074:138– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2074/20740138.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2074/20740138.
pdf.

Ishihara:2001:CCP

[ILFO01] Takashi Ishihara, Tiejun Li, Eugene F. Fodor, and Ronald A. Olsson. A comparison of concurrent programming and cooperative multithreading. Lecture Notes in Computer Science, 1900:729– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1900/19000729.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1900/19000729.
pdf.

Im:2020:DWF

[IMMP20] Sungjin Im, Benjamin Moseley, Kamesh Munagala, and Kirk Pruhs. Dynamic weighted fairness with minimal disruptions. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1): 19:1–19:18, May 2020. CO-DEN ???? ISSN 2476-1249.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3379485.

Itzkovitz:1998:TMA

[ISS98] Ayal Itzkovitz, Assaf Schuster, and Lea Shalev. Thread migration and its applications in distributed shared memory systems. The Journal of Systems and Software, 42 (1):71–87, July 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Inverso:2022:BVM

[ITF^{+}22] Omar Inverso, Ermenegildo Tomasco, Bernd Fischer, Salvatore La Torre, and Gennaro Parlato. Bounded verification of multi-threaded programs via lazy sequentialization. ACM Transactions on Programming Languages and Systems, 44(1): 1:1–1:50, March 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478536.

Iliakis:2018:DMS

[IXS18] Konstantinos Iliakis, Sotirios Xydis, and Dimitrios Soudris. Decoupled MapReduce for shared-memory multi-core architectures. IEEE Computer Architecture Letters, 17 (2):143–146, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Iliakis:2019:LIG

[IXS19] Konstantinos Iliakis, Sotirios Xydis, and Dimitrios Soudris. LOOG: Improving GPU efficiency with light-weight outof-order execution. IEEE Computer Architecture Letters, 18(2):166–169, July 2019. ISSN 1556-6064.

Jacobs:2018:MTV

[JBK18] Bart Jacobs, Dragan Bosnacki, and Ruurd Kuiper. Modular termination verification of single-threaded and multithreaded programs. ACM Transactions on Programming Languages and Systems, 40 (3):12:1–12:??, August 2018. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jung:2017:LSD

[JCP17] Sungbo Jung, Dar-Jen Chang, and Juw Won Park. Large scale document inversion using a multi-threaded computing system. ACM SIGAPP Applied Computing Review, 17 (2):27–35, August 2017. CO-DEN ???? ISSN 1559-6915 (print), 1931-0161 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3131080.3131083.

Jaisson:2008:IPM

[JD08] Pascal Jaisson and Florian De Vuyst. An innovating PDE model based on fluid flow paradigm for multithread systems. Computer Networks (Amsterdam, Netherlands: 1999), 52(18):3318– 3324, December 22, 2008. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jeffay:1994:LMT

[Jef94] K. Jeffay. On latency management in time-shared operating systems. In IEEE [IEE94a], pages 86–90. ISBN 0-8186-5710-3. LCCN QA76.54.I173 1994. IEEE catalog number 94TH0639-5.

Jensen:1995:DRT

[Jen95] E. Douglas Jensen. Distributed real-time operating systems. Dr. Dobb’s Journal of Software Tools, 20(2): 32–34, 36, 38, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Johnson:2004:MCP

[JEV04] Troy A. Johnson, Rudolf Eigenmann, and T. N. Vijaykumar. Min-cut program decomposition for thread-level speculation. ACM SIGPLAN Notices, 39(6):59–70, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ji:1998:PMM

[JFL98] Minwen Ji, Edward W. Felten, and Kai Li. Performance measurements for multithreaded programs. ACM SIGMET-RICS Performance Evaluation Review, 26(1):161–170, June 1998. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jia:2019:UPD

[JGS^{+}19] Z. Jia, W. Gao, Y. Shi, S. A. McKee, Z. Ji, J. Zhan, L. Wang, and L. Zhang. Understanding processors design decisions for data analytics in homogeneous data centers. IEEE Transactions on Big Data, 5(1):81–94, March 2019. ISSN 2332-7790.

Johnston:2004:ADP

[JHM04] Wesley M. Johnston, J. R. Paul Hanna, and Richard J. Millar. Advances in dataflow programming languages. ACM Computing Surveys, 36(1): 1–34, March 2004. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Jolitz:1991:PUB

[JJ91] W. F. Jolitz and L. G. Jolitz.

Porting UNIX to the 386.

The basic kernel multiprogramming and multitasking. II. Dr. Dobb’s Journal of Software Tools, 16(10):62, 64, 66, 68, 70, 72, 118–120, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Jin:2003:AMP

[JJY^{+}03] Haoqiang Jin, Gabriele Jost, Jerry Yan, et al. Automatic multilevel parallelization using OpenMP. Scientific Programming, 11(2):177– 190, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jung:2016:LPS

[JLA16] Daejin Jung, Sheng Li, and Jung Ho Ahn. Large pages on steroids: Small ideas to accelerate big memory applications. IEEE Computer Architecture Letters, 15(2):101–104, July/December 2016. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Jonsson:1999:NPS

[JLS99] J. Jonsson, H. Loenn, and K. G. Shin. Non-preemptive scheduling of real-time threads on multi-level-context architectures. Lecture Notes in Computer Science, 1586:363– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jang:2010:DTE

[JMS^{+}10] Byunghyun Jang, Perhaad Mistry, Dana Schaa, Rodrigo Dominguez, and David Kaeli. Data transformations enabling loop vectorization on multithreaded data parallel architectures. ACM SIG-PLAN Notices, 45(5):353–354, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joerg:1996:CSP

[Joe96] Christopher F. (Christopher Frank) Joerg. The Cilk system for parallel multithreaded computing. Thesis (Ph.D.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1996. 199 pp.

Jonak:1986:EFL

[Jon86] J. E. Jonak. Experience with a FORTH-like language. ACM SIGPLAN Notices, 21(2):27– 36, February 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:1991:BCL

[Jon91] Michael B. Jones. Bringing the C libraries with us into a multi-threaded future. In USENIX [USE91b], pages 81– 92. LCCN QA 76.76 O63 U84 1992.

Jagannathan:1992:CSC

[JP92] Suresh Jagannathan and Jim Philbin. A customizable substrate for concurrent languages. ACM SIG-PLAN Notices, 27(7):55–67, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/pldi/143095/
p55-jagannathan/.

Jacobs:2008:PMC

[JPS^{+}08] Bart Jacobs, Frank Piessens, Jan Smans, K. Rustan M. Leino, and Wolfram Schulte. A programming model for concurrent object-oriented programs. ACM Transactions on Programming Languages and Systems, 31(1): 1:1–1:48, December 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Joshi:2009:RDP

[JPSN09] Pallavi Joshi, Chang-Seo Park, Koushik Sen, and Mayur Naik. A randomized dynamic program analysis technique for detecting real deadlocks. ACM SIG-PLAN Notices, 44(6):110–120, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joisha:2011:TEA

[JSB^{+}11] Pramod G. Joisha, Robert S. Schreiber, Prithviraj Banerjee, Hans J. Boehm, and Dhruva R. Chakrabarti. A technique for the effective and automatic reuse of classical compiler optimizations on multithreaded code. ACM SIGPLAN Notices, 46(1):623– 636, January 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joisha:2012:TTE

[JSB^{+}12] Pramod G. Joisha, Robert S.

Schreiber, Prithviraj Banerjee, Hans-J. Boehm, and Dhruva R. Chakrabarti. On a technique for transparently empowering classical compiler optimizations on multithreaded code. ACM Transactions on Programming Languages and Systems, 34(2): 9:1–9:??, June 2012. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Joao:2012:BIS

[JSMP12] José A. Joao, M. Aater Suleman, Onur Mutlu, and Yale N. Patt. Bottleneck identification and scheduling in multithreaded applications. ACM SIGARCH Computer Architecture News, 40 (1):223–234, March 2012. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ASPLOS ’12 conference proceedings.

Joao:2013:UBA

[JSMP13] José A. Joao, M. Aater Suleman, Onur Mutlu, and Yale N. Patt. Utility-based acceleration of multithreaded applications on asymmetric CMPs. ACM SIGARCH Computer Architecture News, 41(3): 154–165, June 2013. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ICSA ’13 conference proceedings.

Jeffrey:2011:IBM

[JWTG11] Dennis Jeffrey, Yan Wang, Chen Tian, and Rajiv Gupta. Isolating bugs in multithreaded programs using execution suppression. Software—Practice and Experience, 41(11):1259–1288, October 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jeon:2015:MTH

[JY15] Yongkweon Jeon and Sungroh Yoon. Multi-threaded hierarchical clustering by parallel nearest-neighbor chaining. IEEE Transactions on Parallel and Distributed Systems, 26(9):2534–2548, September 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/trans/
                                        td/2015/09/06893001.pdf.

Jiang:2016:TLH

[JYE^{+}16] Chuntao Jiang, Zhibin Yu, Lieven Eeckhout, Hai Jin, Xiaofei Liao, and Chengzhong Xu. Two-level hybrid sampled simulation of multithreaded applications. ACM Transactions on Architecture and Code Optimization, 12 (4):39:1–39:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kacsuk:1997:MIC

[KA97] P. Kacsuk and M. Amamiya.

A multithreaded implementation concept of Prolog on Datarol-II machine. Lecture Notes in Computer Science, 1336:91–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanalakis:1994:ET

[Kan94] John M. Kanalakis, Jr. Examining OS/2 2.1 threads. Dr. Dobb’s Journal of Software Tools, 19(1):74, 76, 78– 79, 96, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Kongetira:2005:NWM

[KAO05] Poonacha Kongetira, Kathirgamar Aingaran, and Kunle Olukotun. Niagara: a 32-way multithreaded Sparc processor. IEEE Micro, 25(2):21– 29, March/April 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2005/02/m2021abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2005/02/m2021.
pdf.

Kumar:2007:ESI

[KASD07] Nagendra J. Kumar, Vasanth Asokan, Siddhartha Shivshankar, and Alexander G. Dean. Efficient software implementation of embedded communication protocol controllers using asynchronous software thread integration with time- and space-efficient procedure calls. ACM Transactions on Embedded Computing Systems, 6(1):2:1–2:??, February 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Krashinsky:2008:ISV

[KBA08] Ronny Krashinsky, Christopher Batten, and Krste Asanović. Implementing the Scale vector-thread processor. ACM Transactions on Design Automation of Electronic Systems., 13(3):41:1–41:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kyle:2012:EPI

[KBF^{+}12] Stephen Kyle, Igor Böhm, Björn Franke, Hugh Leather, and Nigel Topham. Efficiently parallelizing instruction set simulation of embedded multicore processors using regionbased just-in-time dynamic binary translation. ACM SIG-PLAN Notices, 47(5):21–30, May 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCTES ’12 proceedings.

Koster:2003:TTI

[KBH^{+}03] Rainer Koster, Andrew P.

Black, Jie Huang, Jonathan Walpole, and Calton Pu. Thread transparency in information flow middleware. Software—Practice and Experience, 33(4):321–349, April 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Krashinsky:2004:VTAa

[KBH^{+}04a] Ronny Krashinsky, Christopher Batten, Mark Hampton, Steve Gerding, Brian Pharris, Jared Casper, and Krste Asanovic. The vector-thread architecture. ACM SIGARCH Computer Architecture News, 32(2):52, March 2004. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Krashinsky:2004:VTAb

[KBH^{+}04b] Ronny Krashinsky, Christopher Batten, Mark Hampton, Steve Gerding, Brian Pharris, Jared Casper, and Krste Asanovic. The vectorthread architecture. IEEE Micro, 24(6):84–90, November/December 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/06/m6084.htm;   http:

//csdl.computer.org/dl/
mags/mi/2004/06/m6084.pdf.

Kreuzinger:2003:RTE

[KBP^{+}03] J. Kreuzinger, U. Brinkschulte, M. Pfeffer, S. Uhrig, and T. Ungerer. Real-time eventhandling and scheduling on a multithreaded Java microcontroller. Microprocessors and Microsystems, 27(1):19– 31, 2003. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

Karamcheti:1998:HLB

[KC98] Vijay Karamcheti and Andrew A. Chien. A hierarchical load-balancing framework for dynamic multithreaded computations. In ACM [ACM98d], page ?? ISBN ???? LCCN ???? URL http:

//www.supercomp.org/sc98/
TechPapers/sc98_FullAbstracts/
Karamcheti553/index.htm.

Karamcheti:1999:ASM

[KC99] Vijay Karamcheti and Andrew A. Chien. Architectural support and mechanisms for object caching in dynamic multithreaded computations. Journal of Parallel and Distributed Computing, 58(2):260–300, August 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1999.1555/production;
http://www.idealibrary.

com/links/doi/10.1006/jpdc.
1999.1555/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1999.1555/production/ref.

Kejariwal:2009:PSA

[KC09] Arun Kejariwal and Calin Casçaval. Parallelization spectroscopy: analysis of threadlevel parallelism in HPC programs. ACM SIGPLAN Notices, 44(4):293–294, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kekckler:1999:CEH

[KCCD99] S. W. Kekckler, A. Chang, W. S. L. S. Chatterjee, and W. J. Dally. Concurrent event handling through multithreading. IEEE Transactions on Computers, 48 (9):903–916, September 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=795220.

Kasperink:1997:CDC

[KD97] Harold R. Kasperink and John C. Dekker. Concurrent database commands and C++. Dr. Dobb’s Journal of Software Tools, 22(8):84, 86, 88, 89, 98, August 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Kelefouras:2022:WSM

[KD22] Vasilios Kelefouras and Karim Djemame. Workflow simulation and multi-threading aware task scheduling for heterogeneous computing. Journal of Parallel and Distributed Computing, 168(??): 17–32, October 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0743731522001265.

Keckler:1998:EFG

[KDM^{+}98] Stephen W. Keckler, William J. Dally, Daniel Maskit, Nicholas P. Carter, Andrew Chang, and Whay S. Lee. Exploiting finegrain thread level parallelism on the MIT multi-ALU processor. ACM SIGARCH Computer Architecture News, 26 (3):306–317, June 1998. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Kleiman:1995:IT

[KE95] Steve Kleiman and Joe Eykholt. Interrupts as threads. Operating Systems Review, 29(2):21–26, April 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kerrison:2015:EMS

[KE15] Steve Kerrison and Kerstin Eder. Energy modeling of software for a hardware multithreaded embedded microprocessor. ACM Transactions on Embedded Computing Systems, 14(3):56:1–56:??, May 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kelly:1994:MBC

[Kel94a] Michael Kelly. Multithreading with OS/2 and Borland C++. C/C++ Users Journal, 12(8): 67–??, August 1994. CODEN CCUJEX. ISSN 1075-2838.

Kelly:1994:MOB

[Kel94b] Michael Kelly. Multithreading with OS/2 and Borland C++. C/C++ Users Journal, 12(8): 67–??, August 1994. CODEN CCUJEX. ISSN 1075-2838.

Klasky:2003:GBP

[KEL^{+}03] Scott Alan Klasky, Stephane Ethier, Zhihong Lin, Kevin Martins, Doug McCune, and Ravi Samtaney. Grid-based parallel data streaming implemented for the gyrokinetic toroidal code. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN ???? URL http:/

/www.sc-conference.org/
sc2003/inter_cal/inter_
cal_detail.php?eventid=
10722#2;   http://www.sc-
conference.org/sc2003/paperpdfs/
pap207.pdf.

Kempf:2002:BTL

[Kem02] Bill Kempf. The Boost.Threads library. C/C++ Users Journal, 20(5):6–??, May 2002.

CODEN CCUJEX. ISSN 1075-2838.

Korndorfer:2022:LDL

[KEMC22] Jonas H. Müller Korndörfer, Ahmed Eleliemy, Ali Mohammed, and Florina M. Ciorba. LB4OMP: a dynamic load balancing library for multithreaded applications. IEEE Transactions on Parallel and Distributed Systems, 33(4): 830–841, April 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kepner:2003:MTF

[Kep03] Jeremy Kepner. A multithreaded fast convolver for dynamically parallel image filtering. Journal of Parallel and Distributed Computing, 63(3): 360–372, March 2003. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kyriacou:2006:CCO

[KET06a] Costas Kyriacou, Paraskevas Evripidou, and Pedro Trancoso. CacheFlow: Cache optimizations for data driven multithreading. Parallel Processing Letters, 16(2):229–244, June 2006. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Kyriacou:2006:DDM

[KET06b] Costas Kyriacou, Paraskevas Evripidou, and Pedro Trancoso. Data-driven multithreading using conventional microprocessors. IEEE Transactions on Parallel and Distributed Systems, 17(10): 1176–1188, October 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kougiouris:1997:PMF

[KF97] Panos Kougiouris and Marco Framba. A portable multithreading framework. C/C++ Users Journal, 15(8):??, August 1997. CODEN CCUJEX. ISSN 1075-2838.

Kocberber:2015:AMA

[KFG15] Onur Kocberber, Babak Falsafi, and Boris Grot. Asynchronous memory access chaining. Proceedings of the VLDB Endowment, 9(4):252– 263, December 2015. CODEN ???? ISSN 2150-8097.

Kim:1994:HAM

[KG94] C. Kim and J.-L. Gaudiot. A hierarchical activation management technique for finegrain multithreaded execution. Lecture Notes in Computer Science, 817:577–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keller:2005:TBV

[KG05] Jörg Keller and Andreas Grävinghoff. Thread-based virtual duplex systems in embedded environments. IEEE Micro, 25(2):60–69, March/ April 2005. CODEN IEMIDZ.

ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//csdl.computer.org/comp/
mags/mi/2005/02/m2060abs.
htm;   http://csdl.computer.
org/dl/mags/mi/2005/02/
m2060.pdf.

Kollias:2007:APC

[KG07] Giorgos Kollias and Efstratios Gallopoulos. Asynchronous PageRank computation in an interactive multithreading environment, volume 07071 of Dagstuhl seminar proceedings, page ???? International Begegnungs- und Forschungszentrum für Informatik, Wadern, Germany, 2007. ISBN ???? URL http:/

/drops.dagstuhl.de/opus/
volltexte/2007/1065/pdf/
07071.KolliasGiorgios.Paper.
1065.

Kunal:2009:HDS

[KGGK09] K. Kunal, K. George, M. Gautam, and V. Kamakoti. HTM design spaces: complete decoupling from caches and achieving highly concurrent transactions. Operating Systems Review, 43(2): 98–99, April 2009. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Khan:2012:MAN

[KGPH12] Arif M. Khan, David F. Gleich, Alex Pothen, and Mahantesh Halappanavar. A multithreaded algorithm for network alignment via approximate matching. In Hollingsworth [Hol12], pages 64:1–64:11. ISBN 1-4673-0804-8. URL http://

conferences.computer.org/
sc/2012/papers/1000a054.
pdf.

Kahkonen:2018:TPC

[KH18a] Kari Kähkönen and Keijo Heljanko. Testing programs with contextual unfoldings. ACM Transactions on Embedded Computing Systems, 17 (1):23:1–23:??, January 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kondguli:2018:BUS

[KH18b] Sushant Kondguli and Michael Huang. Bootstrapping: Using SMT hardware to improve single-thread performance. IEEE Computer Architecture Letters, 17(2):205– 208, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Khosla:1997:MAT

[Kho97] Samir Khosla. Multithreading the asynchronous trigger processor. Thesis (M.S.), University of Florida, Gainesville, FL, USA, 1997. ix + 57 pp.

Kavi:1995:DCM

[KHP^{+}95] Krishna M. Kavi, A. R. Hurson, Phenil Patadia, Elizabeth Abraham, and Ponnarasu Shanmugam. Design of cache memories for multithreaded dataflow architecture. ACM SIGARCH Computer Architecture News, 23 (2):253–264, May 1995. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Kawamoto:1995:MTP

[KI95] S.-I. Kawamoto and T. Ito. Multi-threaded PaiLisp with granularity adaptive parallel execution. Lecture Notes in Computer Science, 907:94–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kutsuna:2016:ARM

[KI16] Takuro Kutsuna and Yoshinao Ishii. Abstraction and refinement of mathematical functions toward SMT-based testcase generation. International Journal on Software Tools for Technology Transfer (STTT), 18(1):109–120, February 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10009-015-0389-7.

Kojima:2017:HLG

[KI17] Kensuke Kojima and Atsushi Igarashi. A Hoare logic for GPU kernels. ACM Transactions on Computational Logic, 18(1):3:1–3:??, April 2017. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic).

Kusakabe:1999:INS

[KIAT99] S. Kusakabe, K. Inenaga, M. Amamiya, and X. Tang. Implementing a non-strict functional programming language on a threaded architecture. Lecture Notes in Computer Science, 1586:138– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1994:FPF

[Kim94]  Chinhyun Kim.  Func-

tional programming and finegrain multithreading for highperformance parallel computing. Thesis (Ph.D.), University of Southern California, Los Angeles, CA, USA, 1994. xv + 150 pp.

Keen:2003:CCP

[KIM^{+}03] Aaron W. Keen, Takashi Ishihara, Justin T. Maris, Tiejun Li, Eugene F. Fodor, and Ronald A. Olsson. A comparison of concurrent programming and cooperative multithreading. Concurrency and Computation: Practice and Experience, 15(1):27– 53, January 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2014:SMC

[Kim14] S. Kim. Synthesizing multithreaded code from real-time object-oriented models via schedulability-aware thread derivation. IEEE Transactions on Software Engineering, 40(4):413–426, April 2014. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6617637.

Kranzlmuller:2003:RAP

[KKDV03] Dieter Kranzlmüller, Peter Kacsuk, Jack Dongarra, and Jens Volkert. Recent advances in parallel virtual machine and message passing interface (select papers from the EuroPVMMPI 2002 Conference). The International Journal of High Performance Computing Applications, 17 (1):3–5, Spring 2003. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Kee:2003:POP

[KKH03] Yang-Suk Kee, Jin-Soo Kim, and Soonhoi Ha. ParADE: An OpenMP programming environment for SMP cluster systems. In ACM [ACM03], page ?? ISBN 1-58113-695-

1. LCCN ???? URL http:

//www.sc-conference.org/
sc2003/inter_cal/inter_
cal_detail.php?eventid=
10708#0;   http://www.sc-
conference.org/sc2003/paperpdfs/
pap130.pdf.

Kee:2004:MMM

[KKH04] Yang-Suk Kee, Jin-Soo Kim, and Soonhoi Ha. Memory management for multithreaded software DSM systems. Parallel Computing, 30(1):121–138, January 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kim:2013:DBC

[KKJ^{+}13] Hwanju Kim, Sangwook Kim, Jinkyu Jeong, Joonwon Lee, and Seungryoul Maeng. Demand-based coordinated scheduling for SMP VMs. ACM SIGPLAN Notices, 48(4):369–380, April 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kumar:2008:AVO

[KKS^{+}08] Sanjeev Kumar, Daehyun Kim, Mikhail Smelyanskiy, Yen-Kuang Chen, Jatin Chhugani, Christopher J. Hughes, Changkyu Kim, Victor W. Lee, and Anthony D. Nguyen. Atomic vector operations on chip multiprocessors. ACM SIGARCH Computer Architecture News, 36(3):441–452, June 2008. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Kislal:2018:ECC

[KKT^{+}18] Orhan Kislal, Jagadish Kotra, Xulong Tang, Mahmut Taylan Kandemir, and Myoungsoo Jung. Enhancing computation-to-core assignment with physical location information. ACM SIG-PLAN Notices, 53(4):312–327, April 2018. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaiser:2014:WAM

[KKW14] Alexander Kaiser, Daniel Kroening, and Thomas Wahl. A widening approach to multithreaded program verification. ACM Transactions on Programming Languages and Systems, 36(4):14:1–14:??, October 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kurzak:2009:SLA

[KLDB09] Jakub Kurzak, Hatem Ltaief, Jack Dongarra, and Rosa M. Badia. Scheduling linear algebra operations on multicore processors. LAPACK Working Note 213, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, February 2009. URL http:

//www.netlib.org/lapack/
lawnspdf/lawn213.pdf.

Kleber:2000:TSA

[Kle00] Jeff Kleber. Thread-safe access to collections. C/C++ Users Journal, 18(5):36–??, May 2000. CODEN CCUJEX. ISSN 1075-2838.

Kang:2008:ISE

[KLG08] Dongsoo Kang, Chen Liu, and Jean-Luc Gaudiot. The impact of speculative execution on SMT processors. International Journal of Parallel Programming, 36(4):361–385, August 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=36&
issue=4&spage=361.

Kwak:1997:VMN

[KLH97] H. Kwak, B. Lee, and A. R.

Hurson. Viability of multithreading on networks of workstations. Lecture Notes in Computer Science, 1277:216– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwak:1999:EMC

[KLH^{+}99] H. Kwak, B. Lee, A. R.

Hurson, Suk-Han Yoon, and Woo-Jong Hahn. Effects of multithreading on cache performance. IEEE Transactions on Computers, 48 (2):176–184, February 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=752659.

Koopman:1992:CBC

[KLS92] Philip J. Koopman, Jr., Peter Lee, and Daniel P. Siewiorek.

Cache behavior of combinator graph reduction. ACM Transactions on Programming Languages and Systems, 14(2):265–297, April 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0164-0925/

128867.html. Also see [KLS92].

Koufaty:2003:HTN

[KM03] David Koufaty and Deborah T. Marr. Hyperthreading technology in the netburst microarchitecture. IEEE Micro, 23(2):56–65, March/April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2003/pdf/m2056.
pdf;   http://www.computer.
org/micro/mi2003/m2056abs.
htm.

Kakulavarapu:2001:DLB

[KMAG01] P. Kakulavarapu, O. C. Maquelin, J. N. Amaral, and G. R. Gao. Dynamic load balancers for a multithreaded multiprocessor system. Parallel Processing Letters, 11(1): 169–??, March 2001. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Kavi:2002:MMA

[KMjC02] Krishna M. Kavi, Alireza Moshtaghi, and Deng jyi Chen. Modeling multithreaded applications using Petri nets. International Journal of Parallel Programming, 30(5):353–371, October 2002. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

ipsapp009.lwwonline.com/
content/getfile/4773/29/
1/abstract.htm;   http:/
/ipsapp009.lwwonline.com/
content/getfile/4773/29/
1/fulltext.pdf;   http:
//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=30&
issue=5&spage=353.

Kapil:2004:CMP

[KML04] Sanjiv Kapil, Harlan McGhan, and Jesse Lawrendra. A chip multithreaded processor for network-facing workloads. IEEE Micro, 24(2):20– 30, March/April 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/02/m2020abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2004/02/m2020.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/02/
m2020.pdf.

Klabnik:2017:RPL

[KN17] Steve Klabnik and Carol Nichols. The Rust Programming Language. No Starch Press, San Francisco, CA, USA, 2017. ISBN 1-59327-828-4 (paperback), 1-59327-851-9 (e-pub). xxvii + 519 pp. LCCN QA76.73.R87 K53 2018.

Klabnik:2019:RPL

[KN19] Steve Klabnik and Carol Nichols. The Rust programming language. No Starch Press, San Francisco, CA, USA, second edition, 2019. ISBN 1-09-812253-4, 1-71850-044-0 (paperback). xxix + 526 pp. LCCN QA76.73.R87. URL http://

                                        proquest.safaribooksonline.
                                        com/?fpi=9781098122539;
                                         https://nostarch.com/download/
                                        samples/RustProgrammingLanguage2018_
                                        Sample_ToC.pdf;   https:
                                        //nostarch.com/Rust2018.

Kvatinsky:2014:MBM

[KNE^{+}14] Shahar Kvatinsky, Yuval H. Nacson, Yoav Etsion, Eby G. Friedman, Avinoam Kolodny, and Uri C. Weiser. Memristorbased multithreading. IEEE Computer Architecture Letters, 13(1):41–44, January/ June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Knopp:2014:EMT

[Kno14]  T. Knopp.  Experimental

multi-threading support for the Julia programming language. In 2014 First Workshop for High Performance Technical Computing in Dynamic Languages, pages 1–

5. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2014.

Kim:2016:SEA

[KNPS16] Youngho Kim, Joong Chae Na, Heejin Park, and Jeong Seop Sim. A space-efficient alphabet-independent Four-Russians’ lookup table and a multithreaded Four-Russians’ edit distance algorithm. Theoretical Computer Science, 656 (Part B)(??):173–179, December 20, 2016. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0304397516300676.

Kim:2006:ERI

[KOE^{+}06] Seon Wook Kim, Chong-Liang Ooi, Rudolf Eigenmann, Babak Falsafi, and T. N. Vijaykumar. Exploiting reference idempotency to reduce speculative storage overflow. ACM Transactions on Programming Languages and Systems, 28(5): 942–965, September 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Koniges:2000:ISP

[Kon00] Alice E. Koniges, editor.

Industrial Strength Parallel Computing. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 1-55860-540-1. xxv + 597 pp. LCCN QA76.58 .I483 2000.

Koontz:1993:PBM

[Koo93] K. W. Koontz. Port buffers: a Mach IPC optimization for handling large volumes of small messages. In USENIX [USE93a], pages 89–

102. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86

                                        1993.  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/mobile93/

.

Korty:1989:SLL

[Kor89]  Joseph A. Korty.  Sema:

a Lint-like tool for analyzing semaphore usage in a multithreaded UNIX kernel. In USENIX Association [USE89], pages 113–123.

Karamcheti:1996:RME

[KPC96] Vijay Karamcheti, John Plevyak, and Andrew A. Chien. Runtime mechanisms for efficient dynamic multithreading. Journal of Parallel and Distributed Computing, 37(1):21–40, August 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0105/production;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0105/production/pdf.

Khyzha:2012:AP

[KPP12] Artem Khyzha, Pavel Parízek, and Corina S. Păsăreanu. Abstract pathfinder. ACM SIG-SOFT Software Engineering Notes, 37(6):1–5, November 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kaiser:2006:CJC

[KPPÉR06] Claude Kaiser, Jean-François Pradat-Peyre, Sami Évangelista, and Pierre Rousseau. Comparing Java, C# and Ada monitors queuing policies: a case study and its Ada refinement. ACM SIGADA Ada Letters, 26(2):23–37, August 2006. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kienzle:2001:CTT

[KR01a] Jörg Kienzle and Alexander Romanovsky. Combining tasking and transactions, part II: open multithreaded transactions. ACM SIGADA Ada Letters, 21(1):67–74, March 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kienzle:2001:IEO

[KR01b] Jörg Kienzle and Alexander Romanovsky. Implementing exceptions in open multithreaded transactions based on Ada 95 exceptions. ACM SIGADA Ada Letters, 21(3): 57–63, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Keckler:2012:MMC

[KR12] Stephen W. Keckler and Steven K. Reinhardt. Massively multithreaded computing systems. Computer, 45 (8):24–25, August 2012. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kawaguchi:2012:DPL

[KRBJ12] Ming Kawaguchi, Patrick Rondon, Alexander Bakst, and Ranjit Jhala. Deterministic parallelism via liquid effects. ACM SIGPLAN Notices, 47(6):45–54, June 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PLDI ’12 proceedings.

Krone:1998:LBN

[KRH98] O. Krone, M. Raab, and B. Hirsbrunner. Load balancing for network based multithreaded applications. Lecture Notes in Computer Science, 1497:206–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krinke:1998:SST

[Kri98] Jens Krinke. Static slicing of threaded programs. ACM SIGPLAN Notices, 33(7):35– 42, July 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Klarlund:1993:GT

[KS93] Nils Klarlund and Michael I.

Schwartzbach. Graph types. In ACM [ACM93a], pages 196–205. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15 1993.

URL http://www.acm.org:
80/pubs/citations/proceedings/
plan/158511/p196-klarlund/

. ACM order number 549930.

Krieger:1997:HPO

[KS97] Orran Krieger and Michael Stumm. HFS: a performanceoriented flexible file system based on building-block compositions. ACM Transactions on Computer Systems, 15(3): 286–321, August 1997. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tocs/1997-15-3/
p286-krieger/.

Kalayappan:2016:FRT

[KS16] Rajshekar Kalayappan and Smruti R. Sarangi. Fluid-Check: a redundant threadingbased approach for reliable execution in manycore processors. ACM Transactions on Architecture and Code Optimization, 12(4):55:1–55:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kozicky:2021:JDT

[KS21] Claudio Kozický and Ivan Simecek. Joint direct and transposed sparse matrixvector multiplication for multithreaded CPUs. Concurrency and Computation: Practice and Experience, 33 (13):e6236:1–e6236:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kgil:2008:PUS

[KSB^{+}08] Taeho Kgil, Ali Saidi, Nathan Binkert, Steve Reinhardt, Krisztian Flautner, and Trevor Mudge. PicoServer: Using 3D stacking technology to build energy efficient servers. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(4):16:1–16:??, October 2008. CODEN ???? ISSN 1550-4832.

Kumar:2004:AST

[KSD04] Nagendra J. Kumar, Siddhartha Shivshankar, and Alexander G. Dean. Asynchronous software thread integration for efficient software. ACM SIGPLAN Notices, 39 (7):37–46, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kleiman:1995:PT

[KSS95] Steve Kleiman, Devang Shah, and Bart Smaalders. Programming With Threads. Sun-Soft Press, Mountainview, CA, USA, 1995. ISBN 0-13-172389-8. xxviii and 534 pp. LCCN QA76.58.K59 1996. US$48.00. URL http://www.

amazon.com/exec/obidos/
ISBN=0131723898/sunworldonlineA/
002-4892305-5599452.

Kleiman:1996:PT

[KSS96] Steve Kleiman, Devang Shah, and Bart Smaalders. Programming with threads. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-172389-8. xxviii + 534 pp. LCCN QA76.58 .K53 1996.

Kalla:2004:IPC

[KST04] Ron Kalla, Balaram Sinharoy, and Joel M. Tendler. IBM Power5 chip: a dualcore multithreaded processor. IEEE Micro, 24(2):40– 47, March/April 2004. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2004/02/m2040abs.htm;
 http://csdl.computer.org/
dl/mags/mi/2004/02/m2040.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/02/
m2040.pdf.

Krieger:1994:ASF

[KSU94] Orran Krieger, Michael Stumm, and Ron Unrau. The Alloc Stream Facility: a redesign of application-level stream I/O. Computer, 27 (3):75–82, March 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yu:2011:SDH

[kSYHX^{+}11] Wing kei S. Yu, Ruirui Huang, Sarah Q. Xu, Sung-En Wang, Edwin Kan, and G. Edward Suh. SRAM–DRAM hybrid memory with applications to efficient register files in fine-grained multithreading. ACM SIGARCH Computer Architecture News, 39(3):247–258, June 2011. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Krishnan:1999:CMA

[KT99] V. Krishnan and J. Torrellas. A chip-multiprocessor architecture with speculative multithreading. IEEE Transactions on Computers, 48(9):866–880, September 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&
                                        arnumber=795218.

Kopczynski:2017:LSS

[KT17] Eryk Kopczyński and Szymon Toruńczyk. LOIS: syntax and semantics. ACM SIGPLAN Notices, 52(1):586–598, January 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kambadur:2012:HCA

[KTK12] Melanie Kambadur, Kui Tang, and Martha A. Kim.

Harmony: collection and analysis of parallel block vectors. ACM SIGARCH Computer Architecture News, 40 (3):452–463, June 2012. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ISCA ’12 conference proceedings.

Kambadur:2013:PSP

[KTLK13] Melanie Kambadur, Kui Tang, Joshua Lopez, and Martha A. Kim. Parallel scaling properties from a basic block view. ACM SIGMET-RICS Performance Evaluation Review, 41(1):365–366, June 2013. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Kumar:2004:SIH

[KTR^{+}04] Rakesh Kumar, Dean M.

Tullsen, Parthasarathy Ranganathan, Norman P. Jouppi, and Keith I. Farkas. Single-ISA heterogeneous multicore architectures for multithreaded workload performance. ACM SIGARCH Computer Architecture News, 32(2):64, March 2004. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Keller:2000:JUS

[KU00] J. Keller and T. Ungerer.

J.UCS special issue on multithreaded processors and chip-multiprocessors. J.UCS: Journal of Universal Computer Science, 6(10):906–907, October 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_6_10/
                                        j_ucs_special_issue.

Komosinski:2017:MCE

[KU17] Maciej Komosinski and Szymon Ulatowski. Multithreaded computing in evolutionary design and in artificial life simulations. The Journal of Supercomputing, 73(5): 2214–2228, May 2017. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.
                                        springer.com/content/pdf/
                                        10.1007/s11227-016-1923-
                                        4.pdf.

Kubica:2015:PHT

[Kub15] Bartlomiej Jacek Kubica. Presentation of a highly tuned multithreaded interval solver for underdetermined and welldetermined nonlinear systems. Numerical Algorithms, 70(4): 929–963, December 2015. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s11075-015-9980-y;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s11075-
                                        015-9980-y.pdf.

Kuchlin:1991:MCI

[Küc91] Wolfgang Küchlin. On the multi-threaded computation of integral polynomial greatest common divisors. In Watt [Wat91], pages 333–

342. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

URL http://www.acm.org:
80/pubs/citations/proceedings/
issac/120694/p333-kuchlin/

.

Kuchlin:1992:MTC

[Kuc92] W. Kuchlin. On the multithreaded computation of modular polynomial greatest common divisors. Lecture Notes in Computer Science, 591:369– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kestor:2015:TPD

[KUCT15] Gokcen Kestor, Osman S. Unsal, Adrian Cristal, and Serdar Tasiran. TRADE: Precise dynamic race detection for scalable transactional memory systems. ACM Transactions on Parallel Computing (TOPC), 2(2):11:1–11:??, July 2015. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Kuszmaul:2015:SSF

[Kus15] Bradley C. Kuszmaul. SuperMalloc: a super fast multithreaded malloc for 64bit machines. ACM SIG-PLAN Notices, 50(11):41–55, November 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kejariwal:2009:ELL

[KVN^{+}09] Arun Kejariwal, Alexander V. Veidenbaum, Alexandru Nicolau, Milind Girkar, Xinmin Tian, and Hideki Saito. On the exploitation of loop-level parallelism in embedded applications. ACM Transactions on Embedded Computing Systems, 8(2):10:1–10:??, January 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kleinmann:2017:ACS

[KW17] Amit Kleinmann and Avishai Wool. Automatic construction of statechart-based anomaly detection models for multithreaded industrial control systems. ACM Transactions on Intelligent Systems and Technology (TIST), 8(4):55:1– 55:??, July 2017. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Kwok:2003:EHC

[Kwo03] Yu-Kwong Kwok. On exploiting heterogeneity for cluster based parallel multithreading using task duplication. The Journal of Supercomputing, 25(1):63–72, May 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp009.kluweronline.
com/content/getfile/5189/
43/4/abstract.htm;   http:
//ipsapp009.kluweronline.
com/content/getfile/5189/
43/4/fulltext.pdf.

Kasikci:2015:ACD

[KZC15] Baris Kasikci, Cristian Zamfir, and George Candea. Automated classification of data races under both strong and weak memory models. ACM Transactions on Programming Languages and Systems, 37 (3):8:1–8:??, June 2015. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kandemir:2015:MRR

[KZTK15] Mahmut Kandemir, Hui Zhao, Xulong Tang, and Mustafa Karakoy. Memory row reuse distance and its role in optimizing application performance. ACM SIGMET-RICS Performance Evaluation Review, 43(1):137–149, June 2015. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lim:1993:WAS

[LA93] Beng-Hong Lim and Anant Agarwal. Waiting algorithms for synchronization in large-scale multiprocessors. ACM Transactions on Computer Systems, 11(3): 253–294, August 1993. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tocs/1993-11-3/
p253-lim/.

Lafreniere:2000:SMD

[Laf00] David Lafreniere. State machine design in C++. C/C++ Users Journal, 18(5):58–??, May 2000. CODEN CCUJEX. ISSN 1075-2838.

Liu:2012:FPA

[LAH^{+}12] Gu Liu, Hong An, Wenting Han, Xiaoqiang Li, Tao Sun, Wei Zhou, Xuechao Wei, and Xulong Tang. FlexBFS: a parallelism-aware implementation of breadth-first search on GPU. ACM SIGPLAN Notices, 47(8):279–280, August 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

LakshmanYN:1996:IPI

[Lak96] Lakshman Y. N., editor. IS-SAC ’96: Proceedings of the 1996 International Symposium on Symbolic and Algebraic Computation, July 24– 26, 1996, Zurich, Switzerland. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996.

Lenharth:2009:RDO

[LAK09] Andrew Lenharth, Vikram S. Adve, and Samuel T. King. Recovery domains: an organizing principle for recoverable operating systems. ACM SIGPLAN Notices, 44(3):49– 60, March 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lam:1995:CPC

[Lam95] Richard B. Lam. Crossplatform communication classes. Dr. Dobb’s Journal of Software Tools, 20(3):20, 22, 24, 26, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Lang:1997:MTE

[Lan97] Duncan Walter Temple Lang.

A multi-threaded extension to a high level interactive statistical computing environment. Thesis (Ph.D. in Statistics), University of California, Berkeley, Berkeley, CA, USA, December 1997. vii + 161 pp.

Laneve:2002:TSJ

[Lan02] Cosimo Laneve. A type system for JVM threads. Theoretical Computer Science, 290 (1):741–778, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Larcheveque:1995:OIP

[Lar95]  J.-M. Larchevêque.  Op-

timal incremental parsing. ACM Transactions on Programming Languages and Systems, 17(1):1–15, January 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0164-0925/
200996.html.

Larbi:1997:BRM

[Lar97] Michael Larbi. Book review: Multithreading Applications in Win32. C/C++ Users Journal, 15(7):65–??, July 1997. CODEN CCUJEX. ISSN 1075-2838.

LeSergent:1992:IMT

[LB92] T. Le Sergent and B. Berthomieu. Incremental multi-threaded garbage collection on virtually shared memory architectures. Lecture Notes in Computer Science, 637:179– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lim:1995:LPB

[LB95] Beng-Hong Lim and Ricardo Bianchini. Limits on the performance benefits of multithreading and prefetching. Research report RC 20238 (89547), IBM T. J. Watson Research Center, Yorktown Heights, NY, USA, October 20, 1995. 23 pp.

Lewis:1996:TPG

[LB96a] Bil Lewis and Daniel J. Berg. Threads Primer: a Guide to Multithreaded Programming. Sun BluePrints Program. Sun-Soft Press, Mountainview, CA, USA, 1996. ISBN 0-13-443698-9. xxvi + 319 pp. LCCN QA76.642 .L478 1996.

Lim:1996:LPB

[LB96b] Beng-Hong Lim and Ricardo Bianchini. Limits on the performance benefits of multithreading and prefetching. ACM SIGMETRICS Performance Evaluation Review, 24 (1):37–46, May 1996. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lewis:1998:MPP

[LB98] Bil Lewis and Daniel J. Berg.

Multithreaded programming with pthreads. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, 1998. ISBN 0-13-680729-1 (paperback). xxx + 382 pp. LCCN QA76.76.T55

L49 1998. URL http://www.
amazon.com/exec/obidos/
ASIN/0136807291/ref=sim_
books/002-4892305-5599452;
http://www.sun.com/books/
catalog/lewis2/index.html.

Lewis:2000:MPJ

[LB00] Bil Lewis and Daniel J. Berg.

Multithreaded Programming with Java Technology. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2000. ISBN 0-13-017007-0. xxv + 461 pp. LCCN QA76.73.J38 L488 2000. US$39.99.

URL http://www.sun.com/
books/catalog/lewis3/index.
html.

Lee:2017:MVN

[LB17] Doowon Lee and Valeria Bertacco. MTraceCheck: Validating non-deterministic behavior of memory consistency models in post-silicon validation. ACM SIGARCH Computer Architecture News, 45 (2):201–213, May 2017. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Lo:1998:ADW

[LBE^{+}98] Jack L. Lo, Luiz André Barroso, Susan J. Eggers, Kourosh Gharachorloo, Henry M. Levy, and Sujay S. Parekh. An analysis of database workload performance on simultaneous multithreaded processors. ACM SIGARCH Computer Architecture News, 26 (3):39–50, June 1998. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Ling:2012:HPP

[LBH12] Cheng Ling, Khaled Benkrid, and Tsuyoshi Hamada. High performance phylogenetic analysis on CUDA-compatible GPUs. ACM SIGARCH Computer Architecture News, 40 (5):52–57, December 2012. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). HEART ’12 conference proceedings.

Li:2006:MEMa

[LBvH06a] Xin Li, Marian Boldt, and Reinhard von Hanxleden. Mapping Esterel onto a multithreaded embedded processor. ACM SIGARCH Computer Architecture News, 34(5):303– 314, December 2006. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Li:2006:MEMb

[LBvH06b] Xin Li, Marian Boldt, and Reinhard von Hanxleden. Mapping Esterel onto a multithreaded embedded processor. Operating Systems Review, 40(5):303–314, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:2006:MEMc

[LBvH06c] Xin Li, Marian Boldt, and Reinhard von Hanxleden. Mapping Esterel onto a multithreaded embedded processor. ACM SIGPLAN Notices, 41(11):303–314, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lucia:2013:CEF

[LC13] Brandon Lucia and Luis Ceze.

Cooperative empirical failure avoidance for multithreaded programs. ACM SIGPLAN Notices, 48(4):39–50, April 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:2008:HPP

[LCH^{+}08] Duo Liu, Zheng Chen, Bei Hua, Nenghai Yu, and Xinan Tang. High-performance packet classification algorithm for multithreaded IXP network processor. ACM Transactions on Embedded Computing Systems, 7(2):16:1–16:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lathrop:2011:SPI

[LCK11] Scott Lathrop, Jim Costa, and William Kramer, editors. SC’11: Proceedings of 2011 International Conference for High Performance Computing, Networking, Storage and Analysis, Seattle, WA, November 12–18 2011. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4503-0771-X. LCCN ????

Li:2004:FRT

[LCS04] S. Q. Li, H. Y. Chen, and Y. X. Su. A framework of reachability testing for Java multithread programs. IEEE International Conference on Systems Man and Cybernetics, 3:2730–2734, 2004. CO-DEN ???? ISSN 1062-922X.

Lozi:2016:FPL

[LDT^{+}16] Jean-Pierre Lozi, Florian David, Gaël Thomas, Julia Lawall, and Gilles Muller. Fast and portable locking for multicore architectures. ACM Transactions on Computer Systems, 33(4):13:1– 13:??, January 2016. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Leary:1996:CEH

[Lea96] S. Leary. C++ exception handling in multithreaded programs. C++ Report, 8(2):20– 31, February 1996. CODEN CRPTE7. ISSN 1040-6042.

Lee:1993:TW

[Lee93] David Lee. Threads for Windows 3. Dr. Dobb’s Journal of Software Tools, 18(10):84–??, Fall 1993. CODEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Lee:2006:PT

[Lee06] Edward A. Lee. The problem with threads. Technical Report UCB/EECS-2006-1, Electrical Engineering and Computer Sciences. University of California at Berkeley, Berkeley, CA, USA, January 10, 2006. URL http:

//www.eecs.berkeley.edu/
Pubs/TechRpts/2006/EECS-
2006-1.html.

Legrand:2001:MTD

[Leg01] Iosif Legrand, on behalf of the MONARC Collaboration. Multi-threaded, discrete event simulation of distributed computing systems. Computer Physics Communications, 140 (1–2):274–285, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465501002818.

Leiserson:1997:AAM

[Lei97] C. E. Leiserson. Algorithmic analysis of multithreaded algorithms. Lecture Notes in Computer Science, 1350:132– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lo:1997:CTL

[LEL^{+}97a] Jack L. Lo, Joel S. Emer, Henry M. Levy, Rebecca L. Stamm, and Dean M. Tullsen. Converting thread-level parallelism to instruction-level parallelism via simultaneous multithreading. ACM Transactions on Computer Systems, 15(3):322–354, August 1997. CODEN AC-SYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tocs/1997-15-3/
p322-lo/.

Lo:1997:CTP

[LEL^{+}97b] Jack L. Lo, Joel S. Emer, Henry M. Levy, Rebecca L. Stamm, and Dean M. Tullsen. Converting thread-level parallelism to instruction-level parallelism via simultaneous multithreading. ACM Transactions on Computer Systems, 15(3):322–354, August 1997. CODEN AC-SYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tocs/1997-15-3/
p322-lo/.

Lo:1999:TCO

[LEL^{+}99] Jack L. Lo, Susan J. Eggers, Henry M. Levy, Sujay S. Parekh, and Dean M. Tullsen. Tuning compiler optimizations for simultaneous multithreading. International Journal of Parallel Programming, 27(6):477–503, December 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=27&
issue=6&spage=477.

Leman:2002:EFT

[Lem02] Dmitri Leman. An efficient and flexible tracing technique. C/C++ Users Journal, 20(4): 24–??, April 2002. CODEN CCUJEX. ISSN 1075-2838.

Lenatti:1995:RPM

[Len95] C. Lenatti. Rethinking in Parallel: Multiprocessing is on the rise, despite a dearth of tools to help create multithreaded applications. Unix-World’s Open Computing, 12 (8):57–??, 1995. CODEN OP-COEB. ISSN 1072-4044.

Leppanen:1995:PWO

[Lep95] Ville Leppänen. Performance of work-optimal PRAM simulation algorithms on coated meshes. The Computer Journal, 38(10):801–810, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/Volume_38/Issue_
10/Vol38_10.body.html#AbstractLeppanen.

Leven:1997:MIR

[Lev97] Peter J. Leven. A multithreaded implementation of a Robot Control C Library. Thesis (M.S.), University of Illinois at Urbana-Champaign, Urbana-Champaign, IL, USA, 1997. x + 72 pp.

Lowenthal:1996:UFG

[LFA96] David K. Lowenthal, Vincent W. Freeh, and Gregory R. Andrews. Using finegrain threads and run-time decision making in parallel computing. Journal of Parallel and Distributed Computing, 37(1):41–54, August 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0106/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0106/production/pdf.

Lemon:2004:MCR

[LG04] Oliver Lemon and Alexander Gruenstein. Multithreaded context for robust conversational interfaces: Contextsensitive speech recognition and interpretation of corrective fragments. ACM Transactions on Computer-Human Interaction, 11(3):241–267, September 2004. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Lee:2006:TBR

[LG06] S.-W. Lee and J.-L. Gaudiot. Throttling-based resource management in high performance multithreaded architectures. IEEE Transactions on Computers, 55 (9):1142–1152, September 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=1668042.

Laudon:1994:IMT

[LGH94] James Laudon, Anoop Gupta, and Mark Horowitz. Interleaving: a multithreading technique targeting multiprocessors and workstations. ACM SIGPLAN Notices, 29(11):308–318, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/asplos/195473/

p308-laudon/. Co-published in Operating Systems Review, 28(5).

Lenir:1992:EIL

[LGN92] Philip Lenir, R. Govindarajan, and S. S. Nemawarkar. Exploiting instruction-level parallelism: the multithreaded approach. ACM SIGMI-CRO Newsletter, 23(1–2):189– 192, December 1992. URL

                                        https://dl.acm.org/doi/
                                        10.1145/144965.145798.

Lee:1994:DAM

[LH94] Ben Lee and A. R. Hurson. Dataflow architectures and multithreading. Computer, 27(8):27–39, August 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:2009:MHF

[LH09] Taehee Lee and Tobias Höllerer. Multithreaded hybrid feature tracking for markerless augmented reality. IEEE Transactions on Visualization and Computer Graphics, 15(3):355–368, May/June 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ling:2016:MTH

[LHG^{+}16] Cheng Ling, Tsuyoshi Hamada, Jingyang Gao, Guoguang Zhao, Donghong Sun, and Weifeng Shi. MrBayes tgMC 3++: a high performance and resource-efficient GPUoriented phylogenetic analysis method. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 13 (5):845–854, September 2016. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Liu:2016:PSE

[LHS16] Yongchao Liu, Thomas Hankeln, and Bertil Schmidt. Parallel and space-efficient construction of Burrows–Wheeler transform and suffix array for big genome data. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 13(3):592–598, May 2016. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Li:2005:OSA

[Li05] Xiaoye S. Li. An overview of SuperLU: Algorithms, implementation, and user interface. ACM Transactions on Mathematical Software, 31(3):302– 325, September 2005. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Liedtke:1994:SNIb

[Lie94] Jochen Liedtke. A short note on implementing thread exclusiveness and address space locking. Operating Systems Review, 28(3):38–42, July 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

LaFratta:2013:EEM

[LK13] Patrick A. La Fratta and Peter M. Kogge. Energyefficient multithreading for a hierarchical heterogeneous multicore through localitycognizant thread generation. Journal of Parallel and Distributed Computing, 73(12): 1551–1562, December 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0743731513001494.

LaSalle:2015:MTM

[LK15] Dominique LaSalle and George Karypis. Multi-threaded modularity based graph clustering using the multilevel paradigm. Journal of Parallel and Distributed Computing, 76(??):66–80, February 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0743731514001750.

Langr:2020:RII

[LK20] Daniel Langr and Marin Kočička. Reducing the impact of intensive dynamic memory allocations in parallel multithreaded programs. IEEE Transactions on Parallel and Distributed Systems, 31(5): 1152–1164, May 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2011:LCM

[LKBK11] Sheng Li, Shannon Kuntz, Jay B. Brockman, and Peter M. Kogge. Lightweight Chip Multi-Threading (LCMT):

Maximizing fine-grained parallelism on-chip. IEEE Transactions on Parallel and Distributed Systems, 22(7):1178– 1191, July 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2017:TDS

[LLD17] Hao Luo, Pengcheng Li, and Chen Ding. Thread data sharing in cache: Theory and measurement. ACM SIGPLAN Notices, 52(8):103–115, August 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lakshminarayana:2012:DSP

[LLKS12] Nagesh B. Lakshminarayana, Jaekyu Lee, Hyesoon Kim, and Jinwoo Shin. DRAM scheduling policy for GPGPU architectures based on a potential function. IEEE Computer Architecture Letters, 11(2):33–36, July/December 2012. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lin:2010:TAC

[LLL10] Yi-Neng Lin, Ying-Dar Lin, and Yuan-Cheng Lai. Thread allocation in CMP-based multithreaded network processors. Parallel Computing, 36(2–3):104–116, February/March 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Lai:2015:SAM

[LLLC15] Bo-Cheng Charles Lai, Kun-Chun Li, Guan-Ru Li, and Chin-Hsuan Chiang. Self adaptable multithreaded object detection on embedded multicore systems. Journal of Parallel and Distributed Computing, 78(??):25–38, April 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0743731515000192.

Li:2006:SDH

[LLS06] Tong Li, Alvin R. Lebeck, and Daniel J. Sorin. Spin detection hardware for improved management of multithreaded systems. IEEE Transactions on Parallel and Distributed Systems, 17(6):508–521, June 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2016:SEA

[LMA^{+}16] Qixiao Liu, Miquel Moreto, Jaume Abella, Francisco J. Cazorla, Daniel A. Jimenez, and Mateo Valero. Sensible energy accounting with abstract metering for multicore systems. ACM Transactions on Architecture and Code Optimization, 12(4):60:1–60:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Liu:2014:TAP

[LMC14] Xu Liu and John Mellor-Crummey. A tool to analyze the performance of multithreaded programs on NUMA architectures. ACM SIG-PLAN Notices, 49(8):259–272, August 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2014:PDC

[LMJ14] Yong Li, R. Melhem, and A. K. Jones. A practical data classification framework for scalable and high performance chip-multiprocessors. IEEE Transactions on Computers, 63(12):2905–2918, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ling:2000:AOT

[LML00] Yibei Ling, Tracy Mullen, and Xiaola Lin. Analysis of optimal thread pool size. Operating Systems Review, 34 (2):42–55, April 2000. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:2019:TBH

[LML^{+}19] Bing Li, Mengjie Mao, Xiaoxiao Liu, Tao Liu, Zihao Liu, Wujie Wen, Yiran Chen, and Hai (Helen) Li. Thread batching for high-performance energy-efficient GPU memory design. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(4):39:1–39:??, December 2019. CODEN ???? ISSN 1550-4832. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3330152.

Li:2019:SRM

[LNI^{+}19] Y. Li, K. Nomura, J. A. Insley, V. Morozov, K. Kumaran, N. A. Romero, W. A. Goddard, R. K. Kalia, A. Nakano, and P. Vashishta. Scalable reactive molecular dynamics simulations for computational synthesis. Computing in Science and Engineering, 21(5): 64–75, September 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Lochbihler:2018:MTS

[Loc18] Andreas Lochbihler. Mechanising a type-safe model of multithreaded Java with a verified compiler. Journal of Automated Reasoning, 61(1– 4):243–332, June 2018. CO-DEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10817-018-9452-x.

Loeffler:1997:MJF

[Loe97] G. Loeffler. A multithreaded Java framework for solving linear elliptic partial differential equations in 3D. Lecture Notes in Computer Science, 1343:121–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loepere:2005:STM

[Loe05] Keith Loepere. Stackable thread mechanisms. Operating Systems Review, 39(4): 4–17, October 2005. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Loikkanen:1995:FMS

[Loi95] Matias Loikkanen. A finegrain multithreading superscalar architecture. Thesis (M.S., Engineering), University of California, Irvine, Irvine, CA, USA, 1995. xi + 103 pp.

Lowy:2000:MPO

[Low00] Juval Lowy. Making primitive objects thread safe. C/C++ Users Journal, 18(3):85–??, March 2000. CODEN CCU-JEX. ISSN 1075-2838.

Launchbury:1994:LFS

[LP94] John Launchbury and Simon L. Peyton Jones. Lazy functional state threads. ACM SIGPLAN Notices, 29(6):24– 35, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/pldi/178243/
p24-launchbury/.

Lubbers:2009:RMP

[LP09] Enno Lübbers and Marco Platzner. ReconOS: Multithreaded programming for reconfigurable computers. ACM Transactions on Embedded Computing Systems, 9(1):8:1– 8:??, October 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Li:2011:FSM

[LPD^{+}11] Guodong Li, Robert Palmer, Michael DeLisi, Ganesh Gopalakrishnan, and Robert M. Kirby. Formal specification of MPI 2.0: Case study in specifying a practical concurrent programming API. Science of Computer Programming, 76(2):65– 81, February 1, 2011. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lo:1999:SDR

[LPE^{+}99] J. L. Lo, S. S. Parekh, S. J. Eggers, H. M. Levy, and D. M. Tullsen. Software-directed register deallocation for simultaneous multithreaded processors. IEEE Transactions on Parallel and Distributed Systems, 10(9):922–??, September 1999. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//dlib.computer.org/td/
books/td1999/pdf/l0922.
pdf;   http://www.computer.
org/tpds/td1999/l0922abs.
htm.

Lai:2016:QMD

[LPK16] Bo-Cheng Charles Lai, Luis Garrido Platero, and Hsien-Kai Kuo. A quantitative method to data reuse patterns of SIMT applications. IEEE Computer Architecture Letters, 15(2):73–76, July/ December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Li:2017:EML

[LPM17] Cha V. Li, Vinicius Petrucci, and Daniel Mossé. Exploring machine learning for thread characterization on heterogeneous multiprocessors. Operating Systems Review, 51(1): 113–123, August 2017. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Leadbitter:2007:NM

[LPS07] P. Leadbitter, D. Page, and N. P. Smart. Nondeterministic multithreading. IEEE Transactions on Computers, 56(7):992–998, July 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=4216296.

Lal:2015:DID

[LQ15] Akash Lal and Shaz Qadeer.

DAG inlining: a decision procedure for reachabilitymodulo-theories in hierarchical programs. ACM SIG-PLAN Notices, 50(6):280–290, June 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lu:2016:VCV

[LRZ16] Yaojie Lu, Seyedamin Rooholamin, and Sotirios G. Ziavras. Vector coprocessor virtualization for simultaneous multithreading. ACM Transactions on Embedded Computing Systems, 15(3):57:1–57:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Laudon:2007:CWM

[LS07] James Laudon and Lawrence Spracklen. The coming wave of multithreaded chip multiprocessors. International Journal of Parallel Programming, 35(3):299–330, June 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=35&
issue=3&spage=299.

Liao:2011:AUB

[LS11] Xiongfei Liao and Thambipillai Srikanthan. Accelerating UNISIM-based cyclelevel microarchitectural simulations on multicore platforms. ACM Transactions on Design Automation of Electronic Systems., 16(3):26:1– 26:??, June 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2018:ERD

[LS18] I-Ting Angelina Lee and Tao B. Schardl. Efficient race detection for reducer hyperobjects. ACM Transactions on Parallel Computing (TOPC), 4(4):20:1–20:??, September 2018. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Lashgar:2015:CSR

[LSB15] Ahmad Lashgar, Ebad Salehi, and Amirali Baniasadi. A case study in reverse engineering GPGPUs: Outstanding memory handling resources. ACM SIGARCH Computer Architecture News, 43(4):15– 21, September 2015. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Le:2007:IPM

[LSF^{+}07] H. Q. Le, W. J. Starke, J. S.

Fields, F. P. O’Connell, D. Q. Nguyen, B. J. Ronchetti, W. M. Sauer, E. M. Schwarz, and M. T. Vaden. IBM POWER6 microarchitecture. IBM Journal of Research and Development, 51(6):639– ??, November 2007. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/516/le.html.

Leiserson:2012:DPR

[LSS12] Charles E. Leiserson, Tao B. Schardl, and Jim Sukha. Deterministic parallel randomnumber generation for dynamicmultithreading platforms. ACM SIGPLAN Notices, 47(8):193– 204, August 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Liu:2018:ISI

[LSW^{+}18] Hongyu Liu, Sam Silvestro, Wei Wang, Chen Tian, and Tongping Liu. iReplayer: insitu and identical record-and-replay for multithreaded applications. ACM SIGPLAN Notices, 53(4):344–358, April 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LoCocero:1997:MML

[LT97] Joseph LoCocero and D. E. (Donald E.) Thomas. A multithreaded, multiple language hardware/software cosimulator. Research report CMUCAD-97-13, Center for Electronic Design Automation, Carnegie Mellon University, Pittsburgh, PA, USA, April 1997. 7 pp.

Liu:2014:PPF

[LTHB14] Tongping Liu, Chen Tian, Ziang Hu, and Emery D.

Berger. PREDATOR: predictive false sharing detection. ACM SIGPLAN Notices, 49(8):3–14, August 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2016:HBG

[LTL^{+}16] Jing Li, Hung-Wei Tseng, Chunbin Lin, Yannis Papakonstantinou, and Steven Swanson. HippogriffDB: balancing I/O and GPU bandwidth in big data analytics. Proceedings of the VLDB Endowment, 9(14):1647–1658, October 2016. CODEN ???? ISSN 2150-8097.

Lin:2017:MSP

[LTM^{+}17] Zhongwei Lin, Carl Tropper, Robert A. McDougal, Mohammand Nazrul Ishlam Patoary, William W. Lytton, Yiping Yao, and Michael L. Hines. Multithreaded stochastic PDES for reactions and diffusions in neurons. ACM Transactions on Modeling and Computer Simulation, 27(2): 7:1–7:??, July 2017. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Li:2021:MEC

[LTMK21] Botao Li, Synge Todo, A. C. Maggs, and Werner Krauth. Multithreaded eventchain Monte Carlo with local times. Computer Physics Communications, 261(??): Article 107702, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465520303453.

Lu:1994:MPM

[Lu94] David Ta-Chang Lu. A multithreaded processor for massively parallel architectures. Thesis (M.S.), University of California, Riverside, Riverside, CA, USA, 1994. vii + 42 pp.

Lu:1995:HMC

[Lu95] Howard J. (Howard Jason) Lu. Heterogeneous multithreaded computing. Thesis (M. Eng.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1995. 21 pp.

Lu:1998:ONW

[Lu98]  Honghui Lu.  OpenMP

on networks of workstations. In ACM [ACM98d], page ?? ISBN ???? LCCN ???? URL http:/

                                        /www.supercomp.org/sc98/
                                        papers/.

Luk:2001:TML

[Luk01]  Chi-Keung Luk.  Tolerat-

ing memory latency through software-controlled pre-execution in simultaneous multithreading processors. ACM SIGARCH Computer Architecture News, 29(2):40–51, May 2001. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Lundberg:1997:BMC

[Lun97] L. Lundberg. Bounding the minimal completion time of static mappings of multithreaded Solaris programs. Lecture Notes in Computer Science, 1300:1034–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lundberg:1999:PBS

[Lun99] Lars Lundberg. Predicting and bounding the speedup of multithreaded Solaris programs. Journal of Parallel and Distributed Computing, 57(3):322–333, June 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1999.1536/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1999.1536/production/pdf;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1999.1536/production/ref.

Lobeiras:2013:PSW

[LVA^{+}13] Jacobo Lobeiras, Moisés Viñas, Margarita Amor, Basilio B. Fraguela, Manuel Arenaz, J. A. García, and M. J. Castro. Parallelization of shallow water simulations on current multi-threaded systems. The International Journal of High Performance Computing Applications, 27 (4):493–512, November 2013. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://

hpc.sagepub.com/content/
27/4/493.full.pdf+html.

Li:2012:MRP

[LvH12] Xin Li and Reinhard von Hanxleden. Multithreaded reactive programming — the Kiel Esterel processor. IEEE Transactions on Computers, 61(3):337–349, March 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Laadan:2010:TLA

[LVN10] Oren Laadan, Nicolas Viennot, and Jason Nieh. Transparent, lightweight application execution replay on commodity multiprocessor operating systems. ACM SIG-METRICS Performance Evaluation Review, 38(1):155–166, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lopes:2001:FGM

[LVS01] L. Lopes, V. T. Vasconcelos, and F. Silva. Finegrained multithreading with process calculi. IEEE Transactions on Computers, 50 (8):852–862, August 2001.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=947014.

Laguna:2019:GPD

[LWSB19] Ignacio Laguna, Paul C.

Wood, Ranvijay Singh, and Saurabh Bagchi. GPUMixer: Performance-driven floatingpoint tuning for GPU scientific applications. Report, Lawrence Livermore National Laboratory, Livermore CA 94550, USA, 2019. URL http:

//lagunaresearch.org/docs/
isc-2019.pdf;   https://
www.hpcwire.com/2019/08/
05/llnl-purdue-researchers-
harness-gpu-mixed-precision-
for-accuracy-performance-
tradeoff/.

Lee:2010:REO

[LWV^{+}10] Dongyoon Lee, Benjamin Wester, Kaushik Veeraraghavan, Satish Narayanasamy, Peter M. Chen, and Jason Flinn. Respec: efficient online multiprocessor replay via speculation and external determinism. ACM SIGPLAN Notices, 45(3):77–90, March 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:2016:TAA

[LYH16] Peng Liu, Jiyang Yu, and Michael C. Huang. Threadaware adaptive prefetcher on multicore systems: Improving the performance for multithreaded workloads. ACM Transactions on Architecture and Code Optimization, 13(1): 13:1–13:??, April 2016. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Li:2007:CET

[LZ07] Peng Li and Steve Zdancewic. Combining events and threads for scalable network services implementation and evaluation of monadic, applicationlevel concurrency primitives. ACM SIGPLAN Notices, 42(6):189–199, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lu:2014:EDM

[LZBW14] Kai Lu, Xu Zhou, Tom Bergan, and Xiaoping Wang. Efficient deterministic multithreading without global barriers. ACM SIGPLAN Notices, 49(8):287–300, August 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:2014:TPA

[LZL^{+}14] Bin Liu, Yinliang Zhao, Yuxiang Li, Yanjun Sun, and Boqin Feng. A thread partitioning approach for speculative multithreading. The Journal of Supercomputing, 67(3):

778–805, March 2014. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11227-013-1000-1.

Li:2020:MMT

[LZL^{+}20] Tao Li, Xiankai Zhang, Feng Luo, Fang-Xiang Wu, and Jianxin Wang. MultiMotifMaker: a multi-thread tool for identifying DNA methylation motifs from Pacbio reads. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 17(1): 220–225, January 2020. CO-DEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1109/
TCBB.2018.2861399.

Li:2008:TAN

[LZS^{+}08] Z. Li, C. Zhu, L. Shang, R. Dick, and Y. Sun. Transaction-aware networkon-chip resource reservation. IEEE Computer Architecture Letters, 7(2):53–56, July 2008. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Li:2019:HSG

[LZSS19] Yuxiang Li, Yinliang Zhao, Liyu Sun, and Mengjuan Shen. A hybrid sample generation approach in speculative multithreading. The Journal of Supercomputing, 75(8): 4193–4225, August 2019. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Liu:2015:LRT

[LZTZ15] Peng Liu, Xiangyu Zhang, Omer Tripp, and Yunhui Zheng. Light: replay via tightly bounded recording. ACM SIGPLAN Notices, 50(6):55–64, June 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lu:2013:REM

[LZW^{+}13] Kai Lu, Xu Zhou, Xiaoping Wang, Wenzhe Zhang, and Gen Li. RaceFree: an efficient multi-threading model for determinism. ACM SIG-PLAN Notices, 48(8):297–298, August 2013. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’13 Conference proceedings.

Li:2017:GGB

[LZW17] Yuxiang Li, Yinliang Zhao, and Qiangsheng Wu. GbA: a graph-based thread partition approach in speculative multithreading. Concurrency and Computation: Practice and Experience, 29 (21):??, November 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mushtaq:2014:EHP

[MAAB14] Hamid Mushtaq, Zaid Al-Ars, and Koen Bertels. Efficient and highly portable deterministic multithreading (DetLock). Computing, 96(12):1131–1147, December 2014. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00607-013-0370-9.

Monchiero:2009:HSC

[MAF^{+}09] Matteo Monchiero, Jung Ho Ahn, Ayose Falcón, Daniel Ortega, and Paolo Faraboschi. How to simulate 1000 cores. ACM SIGARCH Computer Architecture News, 37 (2):10–19, May 2009. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Mironov:2019:MPE

[MAF19] Vladimir Mironov, Yuri Alexeev, and Dmitri G. Fedorov. Multithreaded parallelization of the energy and analytic gradient in the fragment molecular orbital method. International Journal of Quantum Chemistry, 119(12):e25937:1– e25937:??, June 15, 2019. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mattson:2021:PPM

[MAG21] Timothy G. Mattson, Todd A.

Anderson, and Giorgis Georgakoudis. PyOMP: Multithreaded parallel programming in Python. Computing in Science and Engineering, 23(6):77–80, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mahafzah:2011:PMI

[Mah11] Basel A. Mahafzah. Parallel multithreaded IDA* heuristic search: algorithm design and performance evaluation. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 26(1):61– 82, 2011. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Mahafzah:2013:PAM

[Mah13] Basel A. Mahafzah. Performance assessment of multithreaded quicksort algorithm on simultaneous multithreaded architecture. The Journal of Supercomputing, 66(1):339–363, October 2013. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11227-013-0910-2.

Muller:2018:CPG

[MAH18] Stefan K. Muller, Umut A. Acar, and Robert Harper. Competitive parallelism: getting your priorities right. Proceedings of the ACM on Programming Languages (PACMPL), 2(ICFP):95:1– 95:30, July 2018. URL https:

//dl.acm.org/doi/abs/10.
1145/3236790.

Man:1991:MLC

[Man91] Richard F. Man. A multithreading library in C for subsumption architecture. C Users Journal, 9(11):42–??, November 1991. ISSN 0898-9788.

Mane:1996:SJP

[Man96] I. Mane. Survey of the Java programming language. Elektronik, 45(17):84–87, ???? 20, 1996. CODEN EKRKAR. ISSN 0013-5658.

Manley:1998:GPT

[Man98] Kevin T. Manley. Generalpurpose threads with I/O completion ports. C/C++ Users Journal, 16(4):??, April 1998. CODEN CCUJEX. ISSN 1075-2838.

Manley:1999:IPT

[Man99] Kevin Manley. Improving performance with thread-private heaps. C/C++ Users Journal, 17(9):50–??, September 1999. CODEN CCUJEX. ISSN 1075-2838.

Mao:1996:PMS

[Mao96] Weihua Mao. Performance modeling of data prefetching and multithreading in scalable multiprocessors. Thesis (Ph.D.), University of Southern California, Los Angeles, CA, USA, 1996. xi + 130 pp.

Marowka:2003:EOT

[Mar03] Ami Marowka. Extending OpenMP for task parallelism. Parallel Processing Letters, 13 (3):341–??, September 2003. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Marowka:2007:PCD

[Mar07] Ami Marowka. Parallel computing on any desktop. Communications of the ACM, 50 (9):74–78, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Masney:1999:IMT

[Mas99] Brian Masney. Introduction to multi-threaded programming. Linux Journal, 61:??, May 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Mateosian:1997:MNT

[Mat97] R. M. Mateosian. Micro news: DARPA aids Tera MTA. IEEE Micro, 17(5):5– 6, September/October 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1997/pdf/m5005.
                                        pdf.

Mattson:2003:HGO

[Mat03] Timothy G. Mattson. How good is OpenMP. Scientific Programming, 11(2):81– 93, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Mendelson:1999:DAM

[MB99] Avi Mendelson and Michael Bekerman. Design alternatives of multithreaded architecture. International Journal of Parallel Programming, 27(3):161–193, June 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=27&
issue=3&spage=161.

McNairy:2005:MDC

[MB05] Cameron McNairy and Rohit Bhatia. Montecito: a dualcore, dual-thread Itanium processor. IEEE Micro, 25(2):10– 20, March/April 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
mi/2005/02/m2010abs.htm;
http://csdl.computer.org/
dl/mags/mi/2005/02/m2010.
pdf.

Madan:2007:PEA

[MB07] Niti Madan and Rajeev Balasubramonian. Power efficient approaches to redundant multithreading. IEEE Transactions on Parallel and Distributed Systems, 18(8):1066– 1079, August 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Moon:2006:TMS

[MC06] Sewon Moon and Byeong-Mo Chang. A thread monitoring system for multithreaded Java programs. ACM SIG-PLAN Notices, 41(5):21–29, May 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCarthy:1997:MTI

[McC97a] Martin McCarthy. Multithreading: Intermediate concepts. Linux Journal, 36: ??, April 1997. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL ftp://ftp.

ssc.com/pub/lj/listings/
issue36/2121.tgz.

McCarthy:1997:WMT

[McC97b] Martin McCarthy. What is multi-threading? Linux Journal, 34:??, February 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Mitchell:1999:ILP

[MCFT99] Nicholas Mitchell, Larry Carter, Jeanne Ferrante, and Dean Tullsen. Instructionlevel parallelism vs. threadlevel parallelism on simultaneous multi-threading processors. In ACM [ACM99b], page ??

McManis:1996:JDSa

[McM96a] Chuck McManis. Java in depth: Synchronizing threads in Java. JavaWorld: IDG’s magazine for the Java community, 1(2):??, April 1996. CODEN ???? ISSN 1091-

8906.  URL http://www.
javaworld.com/javaworld/
jw-04-1996/jw-04-synch.
htm.

McManis:1996:JDSb

[McM96b] Chuck McManis. Java in depth: Synchronizing threads in Java, Part II. JavaWorld: IDG’s magazine for the Java community, 1(3):??, May 1996. CODEN ???? ISSN 1091-8906. URL http://www.

javaworld.com/javaworld/
jw-05-1996/jw-05-mcmanis.
htm.

McManis:1996:JDT

[McM96c] Chuck McManis. Java in depth: Threads and applets and visual controls. JavaWorld: IDG’s magazine for the Java community, 1(5):??, July 1996. CO-DEN ???? ISSN 1091-

8906.  URL http://www.
javaworld.com/javaworld/
jw-07-1996/jw-07-mcmanis.
htm.

McMillan:1997:NSB

[McM97]  Robert McMillan.  News:

Sun boosts Java performance, adding JIT compiler and JVM with multithreading to Solaris 2.6. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CODEN ???? ISSN 1091-

                                        8906.  URL http://www.
                                        javaworld.com/javaworld/
                                        jw-07-1997/jw-07-speedway.
                                        htm.

McManis:1998:DUT

[McM98a] Chuck McManis. In depth: Using threads with collections, Part 1. JavaWorld: IDG’s magazine for the Java community, 3(3):??, March 1998. CODEN ???? ISSN 1091-8906. URL http://www.

                                        javaworld.com/javaworld/
                                        jw-03-1998/jw-03-indepth.
                                        html.

McManis:1998:JDU

[McM98b] Chuck McManis. Java in depth: Using threads with collections, part 2. JavaWorld: IDG’s magazine for the Java community, 3(6):??, June 1998. CODEN ???? ISSN 1091-8906. URL http://www.

                                        javaworld.com/javaworld/
                                        jw-06-1998/jw-06-indepth.
                                        html.

Mannarswamy:2010:CAS

[MCRS10] Sandya Mannarswamy, Dhruva R. Chakrabarti, Kaushik Rajan, and Sujoy Saraswati. Compiler aided selective lock assignment for improving the performance of software transactional memory. ACM SIG-PLAN Notices, 45(5):37–46, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mitchell:2015:GIA

[MCS15] Nathan Mitchell, Court Cutting, and Eftychios Sifakis. GRIDiron: an interactive authoring and cognitive training foundation for reconstructive plastic surgery procedures. ACM Transactions on Graphics, 34(4):43:1–43:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Minutoli:2022:PSH

[MCS^{+}22] Marco Minutoli, Vito Giovanni Castellana, Nicola Saporetti, Stefano Devecchi, Marco Lattuada, Pietro Fezzardi, Antonino Tumeo, and Fabrizio Ferrandi. Svelto: High-level synthesis of multithreaded accelerators for graph analytics. IEEE Transactions on Computers, 71(3): 520–533, March 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Montesinos:2008:DRD

[MCT08] Pablo Montesinos, Luis Ceze, and Josep Torrellas. De-Lorean: Recording and deterministically replaying sharedmemory multiprocessor execution efficiently. ACM SIGARCH Computer Architecture News, 36(3):289– 300, June 2008. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Mikschl:1996:MMS

[MD96] A. Mikschl and W. Datum. MSparc: a multithreaded Sparc. Lecture Notes in Computer Science, 1124:461–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matheou:2015:ASD

[ME15] George Matheou and Paraskevas Evripidou. Architectural support for data-driven execution. ACM Transactions on Architecture and Code Optimization, 11(4):52:1–52:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Matheou:2017:DDC

[ME17] George Matheou and Paraskevas Evripidou. Data-driven concurrency for high performance computing. ACM Transactions on Architecture and Code Optimization, 14(4): 53:1–53:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Mukherjee:1994:MII

[MEG94] Bodhisattwa Mukherjee, Greg Eisenhauer, and Kaushik Ghosh. A machine independent interface for lightweight threads. Operating Systems Review, 28(1):33–47, January 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McDowell:2003:ISS

[MEG03] Luke K. McDowell, Susan J.

Eggers, and Steven D. Gribble. Improving server software support for simultaneous multithreaded processors. ACM SIGPLAN Notices, pages 37– 48, 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mennemeier:1991:HMS

[Men91] Lawrence Mennemeier. Hardware mechanisms to support concurrent threads on RISC and superscalar multiprocessors. Thesis (M.S.), University of California, Santa Cru, 1991. vii + 39 pp.

Metz:1995:IDS

[Met95] David Metz. Interface design and system impact analysis of a message-handling processor for fine-grain multithreading. Thesis (M.S.), Oregon State University, Corvallis, OR, USA, 1995. 63 pp.

Marcuello:1999:EST

[MG99] P. Marcuello and A. Gonzalez. Exploiting speculative threadlevel parallelism on a SMT processor. Lecture Notes in Computer Science, 1593:754– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehta:2015:MTP

[MG15] Kshitij Mehta and Edgar Gabriel. Multi-threaded parallel I/O for OpenMP applications. International Journal of Parallel Programming, 43 (2):286–309, April 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10766-014-0306-9.

Martinsen:2014:HTL

[MGI14] Jan Kasper Martinsen, Hakan Grahn, and Anders Isberg. Heuristics for thread-level speculation in Web applications. IEEE Computer Architecture Letters, 13(2):77– 80, July/December 2014. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mohamed:2000:DDM

[MGK^{+}00] A. S. Mohamed, A. Galal, I. Khalil, K. Sobh, and M. Selim. Dispo: Distributed multithreaded execution of Prolog programs. International Journal of Computer Applications, 22(2):100–108, 2000. ISSN 1206-212X (print), 1925-7074 (electronic). URL https://

www.tandfonline.com/doi/
full/10.1080/1206212X.2000.
11441606.

Marsland:1995:SSM

[MGL95] T. A. Marsland, Yaoqing Gao, and Francis Chi-Moon Lau. A study of software multithreading in distributed systems. Technical report TR 95-23, Dept. of Computing Science, University of Alberta, Edmonton, AB, Canada, 1995. 25 pp.

Madriles:2008:MSM

[MGQS^{+}08] Carlos Madriles, Carlos García-Quiñones, Jesús Sánchez, Pedro Marcuello, Antonio González, Dean M. Tullsen, Hong Wang, and John P. Shen. Mitosis: a speculative multithreaded processor based on precomputation slices. IEEE Transactions on Parallel and Distributed Systems, 19(7):914– 925, July 2008. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Maquelin:1995:CBM

[MHG95] O. C. Maquelin, H. H. J.

Hum, and G. R. Gao. Costs and benefits of multithreading with off-the-shelf RISC processors. Lecture Notes in Computer Science, 966:117– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mauer:2002:FST

[MHW02] Carl J. Mauer, Mark D. Hill, and David A. Wood. Full-system timing-first simulation. ACM SIGMET-RICS Performance Evaluation Review, 30(1):108–116, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Miastkowski:1990:PGG

[Mia90] Stan Miastkowski. PC GUIs go head to head. Byte Magazine, 15(11):82–87, Fall 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Michael:2004:SLF

[Mic04] Maged M. Michael. Scalable lock-free dynamic memory allocation. ACM SIG-PLAN Notices, 39(6):35–46, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maabreh:2018:MHT

[MIGA18] Majdi Maabreh, Hafez Irshid, Ajay Gupta, and Izzat Alasmadi. A multithreading and hashing technique for indexing Target– Decoy peptides databases. Concurrency and Computation: Practice and Experience, 30(9):??, May 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/
doi/abs/10.1002/cpe.4371.

Miller:1995:TPC

[Mil95] Robert C. (Robert Chisolm)

Miller. A type-checking preprocessor for Cilk 2, a multithreaded C language. Thesis (M. Eng.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1995. 38 pp.

Mishra:1996:TIS

[Mis96] Amitabh Mishra. Task and instruction scheduling in parallel multithreaded processors. Thesis (M.S.), Department of Computer Science, Texas A&M University, College Station, TX, USA, 1996. ix + 60 pp.

Mitchell:1996:JTM

[Mit96] John D. Mitchell. Java tips:

More about threads and the resize problem. JavaWorld: IDG’s magazine for the Java community, 1(4):??, June 1996. CODEN ???? ISSN 1091-8906. URL http://www.

javaworld.com/javaworld/
javatips/jw-javatip9.htm.

MixSoftware:1994:UMC

[Mix94]  Mix Software, Inc.  Us-

ing Multi-C: a portable multithreaded C programming library. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-606195-8. vi + 257 pp. LCCN QA76.73.C15 U85 1994. System requirements for computer disk: IBM-compatible PC; DOS; Mix, Borland, or Microsoft-compatible C/C++ compilers.

Meng:2010:AOS

[MJF^{+}10] Lingchuan Meng, Jeremy Johnson, Franz Franchetti, Yevgen Voronenko, Marc Moreno Maza, and Yuzhen Xie. Abstract only: SPIRALgenerated modular FFTs. ACM Communications in Computer Algebra, 44(2):25– 26, June 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Mars:2012:BDS

[MK12] Jason Mars and Naveen Kumar. BlockChop: dynamic squash elimination for hybrid processor architecture. ACM SIGARCH Computer Architecture News, 40(3): 536–547, June 2012. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ISCA ’12 conference proceedings.

Moreno:1997:PMP

[MKC97] E. D. Moreno, S. T. Kofuji, and M. H. Cintra. Prefetching and multithreading performance in bus-based multiprocessors with Petri nets. Lecture Notes in Computer Science, 1300:1017–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maris:2004:CCP

[MKIO04] Justin T. Maris, Aaron W.

Keen, Takashi Ishihara, and Ronald A. Olsson. A comparison of concurrent programming and cooperative multithreading under load balancing applications. Concurrency and Computation: Practice and Experience, 16(4):345– 369, April 10, 2004. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moody:1999:STT

[MKK99] Scott Arthur Moody, Samuel Kwok, and Dale Karr. SimpleGraphics: Tcl/Tk visualization of real-time multithreaded and distributed applications. ACM SIGADA Ada Letters, 19(2):60–66, June 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Maiya:2014:RDA

[MKM14] Pallavi Maiya, Aditya Kanade, and Rupak Majumdar. Race detection for Android applications. ACM SIGPLAN Notices, 49(6):316–325, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marquez:2017:MCH

[MKM17] David Gonzalez Marquez, Adrian Cristal Kestelman, and Esteban Mocskos. Mth: Codesigned hardware/software support for fine grain threads. IEEE Computer Architecture Letters, 16(1):64–67, January/June 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mukherjee:2002:DDE

[MKR02] Shubhendu S. Mukherjee, Michael Kontz, and Steven K. Reinhardt. Detailed design and evaluation of redundant multithreading alternatives. ACM SIGARCH Computer Architecture News, 30 (2):99–110, May 2002. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Muralidhara:2010:IAS

[MKR10] Sai Prashanth Muralidhara, Mahmut Kandemir, and Padma Raghavan. Intraapplication shared cache partitioning for multithreaded applications. ACM SIGPLAN Notices, 45(5):329–330, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marowka:2004:OOA

[MLC04] Ami Marowka, Zhenying Liu, and Barbara Chapman. OpenMP-oriented applications for distributed shared memory architectures. Concurrency and Computation: Practice and Experience, 16 (4):371–384, April 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Madriles:2009:BST

[MLC^{+}09] Carlos Madriles, Pedro López, Josep M. Codina, Enric Gibert, Fernando Latorre, Alejandro Martinez, Raúl Martinez, and Antonio Gonzalez. Boosting single-thread performance in multi-core systems through fine-grain multithreading. ACM SIGARCH Computer Architecture News, 37(3):474–483, June 2009. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Ma:2011:SPC

[MLCW11] Kai Ma, Xue Li, Ming Chen, and Xiaorui Wang. Scalable power control for many-core architectures running multi-threaded applications. ACM SIGARCH Computer Architecture News, 39 (3):449–460, June 2011. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Malakhov:2018:CMT

[MLGW18] Anton Malakhov, David Liu, Anton Gorshkov, and Terry Wilmarth. Composable multi-threading and multiprocessing for numeric libraries. In Fatih Akici, David Lippa, Dillon Niederhut, and M Pacer, editors, Proceedings of the 17th Python in Science Conference, Austin, TX, 9– 15 July 2018, pages 15–21. ????, ????, 2018. URL http:

                                        //conference.scipy.org/
                                        proceedings/scipy2018/anton_
                                        malakhov.html.

Machado:2015:CDD

[MLR15] Nuno Machado, Brandon Lucia, and Luís Rodrigues. Concurrency debugging with differential schedule projections. ACM SIGPLAN Notices, 50(6):586–595, June 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Makreshanski:2015:LSE

[MLS15] Darko Makreshanski, Justin Levandoski, and Ryan Stutsman. To lock, swap, or elide: on the interplay of hardware transactional memory and lockfree indexing. Proceedings of the VLDB Endowment, 8(11): 1298–1309, July 2015. CO-DEN ???? ISSN 2150-8097.

Mauro:2001:SIC

[MM01] Jim Mauro and Richard Mc-Dougall. Solaris Internals: Core Kernel Architecture. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-022496-0. xli + 657 pp. LCCN QA76.76.O63 M37195

2001. URL http://www.sun.
com/books/catalog/mauro/
index.html.

Morandini:2007:UDS

[MM07] Marco Morandini and Paolo Mantegazza. Using dense storage to solve small sparse linear systems. ACM Transactions on Mathematical Software, 33 (1):5:1–5:12, March 2007. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Morishima:2014:PEG

[MM14] Shin Morishima and Hiroki Matsutani. Performance evaluations of graph database using CUDA and OpenMP compatible libraries. ACM SIGARCH Computer Architecture News, 42(4):75–80, 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Mathis:2005:CSM

[MMM^{+}05] H. M. Mathis, A. E. Mericas, J. D. McCalpin, R. J. Eickemeyer, and S. R. Kunkel. Characterization of simultaneous multithreading (SMT) efficiency in POWER5. IBM Journal of Research and Development, 49(4/5):555–564, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/494/mathis.html.

Marino:2009:LES

[MMN09] Daniel Marino, Madanlal Musuvathi, and Satish Narayanasamy. LiteRace: effective sampling for lightweight data-race detection. ACM SIGPLAN Notices, 44(6):134–143, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McKenney:2010:WGM

[MMTW10] Paul E. McKenney, Maged M. Michael, Josh Triplett, and Jonathan Walpole. Why the grass may not be greener on the other side: a comparison of locking vs. transactional memory. Operating Systems Review, 44(3):93–101, July 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Metzner:2000:MMR

[MN00] A. Metzner and J. Niehaus. MSparc: Multithreading in real-time architectures. J.UCS: Journal of Universal Computer Science, 6 (10):1034–1051, October 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_6_10/
                                        msparc_multithreading_in_
                                        real.

McAuley:2003:CVC

[MN03] Derek McAuley and Rolf Neugebauer. A case for virtual channel processors. In Romanow and Mogul [RM03], pages 237–242. ISBN ???? LCCN TK5105.5. ACM order number 534032.

Marinov:2016:PAF

[MNG16] Martin Marinov, Nicholas Nash, and David Gregg. Practical algorithms for finding extremal sets. ACM Journal of Experimental Algorithmics, 21(1):1.9:1–1.9:??, November 2016. CODEN ???? ISSN 1084-6654.

Markovic:2015:TLS

[MNU^{+}15] Nikola Markovic, Daniel Nemirovsky, Osman Unsal, Mateo Valero, and Adrian Cristal. Thread lock sectionaware scheduling on asymmetric single-ISA multi-core. IEEE Computer Architecture Letters, 14(2):160–163, July/ December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Moore:1995:MPD

[Moo95]  Simon W. Moore.  Mul-

tithreaded processor design. Thesis (Ph.D.), University of Cambridge, Computer Laboratory, Cambridge, Cambridgeshire, UK, February 1995. xvi + 125 pp. Available as Technical Report 358.

Moore:1996:MPD

[Moo96] Simon W. (Simon William)

Moore. Multithreaded processor design, volume SECS 358 of The Kluwer international series in engineering and computer science. Kluwer Academic Publishers, Dordrecht, The Netherlands; Boston, MA, USA, 1996. ISBN 0-7923-9718-5. xvi + 142 pp. LCCN QA76.5 .M574 1996.

Mount:2000:ADP

[Mou00]  John Mount.  Automatic

detection of potential deadlock. Dr. Dobb’s Journal of Software Tools, 25(12): 64, 66–70, 72, December 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_12/deadlock.txt;
http://www.ddj.com/ftp/
2000/2000_12/deadlock.zip.

Massalin:1989:TIO

[MP89] H. Massalin and C. Pu. Threads and input/output in the synthesis kernel. Operating Systems Review, 23(5): 191–201, December 1989. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Manson:2001:CSM

[MP01] Jeremy Manson and William Pugh. Core semantics of multithreaded Java. In ACM [ACM01], pages 29–

38. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

2001.  URL http://www.
philippsen.com/JGI2001/
camerareadyabstracts/42.
html;   http://www.philippsen.

com/JGI2001/finalpapers/
18500029.pdf.

McCreesh:2013:MTS

[MP13] Ciaran McCreesh and Patrick Prosser. Multi-threading a state-of-the-art maximum clique algorithm. Algorithms (Basel), 6(4):618–635, December 2013. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https:/

/www.mdpi.com/1999-4893/
6/4/618.

Martin:2004:HPA

[MPD04] María J. Martín, Marta Parada, and Ramón Doallo. High performance air pollution simulation using OpenMP. The Journal of Supercomputing, 28(3):311–321, June 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.kluweronline.
com/IPS/content/ext/x/J/
5189/I/54/A/5/abstract.
htm.

Musuvathi:2007:ICB

[MQ07] Madanlal Musuvathi and Shaz Qadeer. Iterative context bounding for systematic testing of multithreaded programs. ACM SIGPLAN Notices, 42(6):446–455, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Musuvathi:2008:FSM

[MQ08] Madanlal Musuvathi and Shaz Qadeer. Fair stateless model checking. ACM SIGPLAN Notices, 43(6):362–371, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Machado:2016:CDD

[MQLR16] Nuno Machado, Daniel Quinta, Brandon Lucia, and Luís Rodrigues. Concurrency debugging with differential schedule projections. ACM Transactions on Software Engineering and Methodology, 25(2): 14:1–14:??, May 2016. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Mayes:1995:ULT

[MQW95] K. R. Mayes, S. Quick, and B. C. Warboys. User-level threads on a general hardware interface. Operating Systems Review, 29(4):57–62, October 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Marinescu:1994:HLC

[MR94] Dan C. Marinescu and John R. Rice. On high level characterization of parallelism. Journal of Parallel and Distributed Computing, 20(1):107–113, January 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1994.1011/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1994.1011/production/pdf.

Mascarenhas:1998:MTP

[MR98] Edward Mascarenhas and Vernon Rego. Migrant threads on process farms: parallel programming with Ariadne. Concurrency: Practice and Experience, 10(9): 673–698, August 10, 1998. CODEN CPEXEI. ISSN 1040-3108. URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?ID=
10008703;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=10008703&
PLACEBO=IE.pdf.

Mukherjee:2009:PAS

[MR09] Jayanta Mukherjee and Soumyendu Raha. Power-aware speedup for multithreaded numerical linear algebraic solvers on chip multicore processors. Scalable Computing: Practice and Experience, 10(2): 217–228, June 2009. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol10/no2/SCPE_10_
2_07.pdf;   http://www.
scpe.org/vols/vol10/no2/
SCPE_10_2_07.zip.

Meier:2017:PVM

[MRG17] Remigius Meier, Armin Rigo, and Thomas R. Gross. Parallel virtual machines with RPython. ACM SIGPLAN Notices, 52(2):48–59, February 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malan:1991:MA

[MRGB91] G. Malan, R. Rashid, D. Golub, and R. Baron. DOS as a Mach 3.0 application. In USENIX [USE91a], pages 27–40. LCCN QAX 27.

McJones:1987:EUS

[MS87] Paul R. McJones and Garret Frederick Swart. Evolving the UNIX system interface to support multithreaded programs: The Topaz Operating System programmer’s manual, volume 21 of Systems Research Center. Digital Systems Research Center, Palo Alto, CA, USA, September 28, 1987. 100 pp. LCCN QA76.76.O63M42 1987.

McJones:1989:EUS

[MS89] Paul R. McJones and Garret F. Swart. Evolving the UNIX system interface to support multithreaded programs. In USENIX Association [USE89], pages 393–404.

Mahinthakumar:2002:HMO

[MS02] G. Mahinthakumar and F. Saied. A hybrid MPI-OpenMP implementation of an implicit finite-element code on parallel architectures. The International Journal of High Performance Computing Applications, 16(4):371–393, Winter 2002. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Mantel:2003:UAS

[MS03] Heiko Mantel and Andrei Sabelfeld. A unifying approach to the security of distributed and multi-threaded programs. Journal of Computer Security, 11(4):615–676, ???? 2003. CODEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

McCartney:2015:SMT

[MS15] W. P. McCartney and N. Sridhar. Stackless multi-threading for embedded systems. IEEE Transactions on Computers, 64(10):2940–2952, ???? 2015. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Metzger:2021:DHT

[MSFC21] Paul Metzger, Volker Seeker, Christian Fensch, and Murray Cole. Device hopping: Transparent mid-kernel runtime switching for heterogeneous systems. ACM Transactions on Architecture and Code Optimization, 18(4):57:1–57:25, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3471909.

Marsh:1991:FCU

[MSLM91] Brian D. Marsh, Michael L. Scott, Thomas J. LeBlanc, and Evangelos P. Markatos. First-class user-level threads. Operating Systems Review, 25 (5):110–121, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Marino:2010:DSE

[MSM^{+}10] Daniel Marino, Abhayendra Singh, Todd Millstein, Madanlal Musuvathi, and Satish Narayanasamy. DRFX: a simple and efficient memory model for concurrent programming languages. ACM SIGPLAN Notices, 45(6):351– 362, June 2010. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marino:2011:CSP

[MSM^{+}11] Daniel Marino, Abhayendra Singh, Todd Millstein, Madanlal Musuvathi, and Satish Narayanasamy. A case for an SC-preserving compiler. ACM SIGPLAN Notices, 46(6):199–210, June 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marino:2016:DXU

[MSM^{+}16] Daniel Marino, Abhayendra Singh, Todd Millstein, Madanlal Musuvathi, and Satish Narayanasamy. drf x: an understandable, high performance, and flexible memory model for concurrent languages. ACM Transactions on Programming Languages and Systems, 38(4): 16:1–16:??, October 2016. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Moreland:2016:VMA

[MSU^{+}16] Kenneth Moreland, Christopher Sewell, William Usher, Li ta Lo, Jeremy Meredith, David Pugmire, James Kress, Hendrik Schroots, Kwan-Liu Ma, Hank Childs, Matthew Larsen, Chun-Ming Chen, Robert Maynard, and Berk Geveci. VTK-m: Accelerating the visualization toolkit for massively threaded architectures. IEEE Computer Graphics and Applications, 36(3): 48–58, May/June 2016. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cg/2016/03/mcg2016030048-
abs.html.

Morrisett:1993:PLP

[MT93] J. Gregory Morrisett and Andrew P. Tolmach. Procs and locks: a portable multiprocessing platform for Standard ML of New Jersey. ACM SIG-PLAN Notices, 28(7):198–207, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Martinez:2002:SSAa

[MT02a] José F. Martínez and Josep Torrellas. Speculative synchronization: applying threadlevel speculation to explicitly parallel applications. ACM SIGARCH Computer Architecture News, 30(5):18– 29, December 2002. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Martinez:2002:SSAb

[MT02b] José F. Martínez and Josep Torrellas. Speculative synchronization: applying threadlevel speculation to explicitly parallel applications. ACM SIGPLAN Notices, 37(10):18– 29, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Martinez:2002:SSAc

[MT02c] José F. Martínez and Josep Torrellas. Speculative synchronization: applying threadlevel speculation to explicitly parallel applications. Operating Systems Review, 36(5): 18–29, December 2002. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Minh:2007:EHT

[MTC^{+}07] Chi Cao Minh, Martin Trautmann, JaeWoong Chung, Austen McDonald, Nathan Bronson, Jared Casper, Christos Kozyrakis, and Kunle Olukotun. An effective hybrid transactional memory system with strong isolation guarantees. ACM SIGARCH Computer Architecture News, 35 (2):69–80, May 2007. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Matsushita:2000:MSC

[MTN^{+}00] Satoshi Matsushita, Sunao Torii, Masahiko Nomura, Toshiaki Inoue, Atsufumi Shibayama, Sachiko Shimada, Taku Osawa, Hiroaki Inoue, Kouichiro Minami, Junji Sakai, Yoshiyuki Ito, Yuichi Nakamura, Masato Edahiro, Naoki Nishi, and Masakazu Yamashina. Merlot: a singlechip tightly coupled four-way multi-thread processor. In Anonymous [Ano00a], page ??

Miller:2012:VCE

[MTPT12] Timothy N. Miller, Renji Thomas, Xiang Pan, and Radu Teodorescu. VRSync: characterizing and eliminating synchronization-induced voltage emergencies in many-core processors. ACM SIGARCH Computer Architecture News, 40(3):249–260, June 2012. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ISCA ’12 conference proceedings.

Meng:2010:DWS

[MTS10] Jiayuan Meng, David Tarjan, and Kevin Skadron. Dynamic warp subdivision for integrated branch and memory divergence tolerance. ACM SIGARCH Computer Architecture News, 38(3): 235–246, June 2010. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Muller:2003:OCB

[Mül03]  Matthias S. Müller.  An

OpenMP compiler benchmark. Scientific Programming, 11(2):125–131, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Musoll:2009:LSO

[Mus09] Enric Musoll. Leakage-saving opportunities in mesh-based massive multi-core architectures. ACM SIGARCH Computer Architecture News, 37 (5):1–7, December 2009. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Mudigonda:2005:MMA

[MVY05] Jayaram Mudigonda, Harrick M. Vin, and Raj Yavatkar. Managing memory access latency in packet processing. ACM SIGMETRICS Performance Evaluation Review, 33(1):396–397, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

McCann:1993:DPA

[MVZ93] Cathy McCann, Raj Vaswani, and John Zahorjan. A dynamic processor allocation policy for multiprogrammed shared-memory multiprocessors. ACM Transactions on Computer Systems, 11(2): 146–178, May 1993. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tocs/1993-11-2/
p146-mccann/.

Morad:2006:PPE

[MWK^{+}06] T. Y. Morad, U. C. Weiser, A. Kolodnyt, M. Valero, and E. Ayguade. Performance, power efficiency and scalability of asymmetric cluster chip multiprocessors. IEEE Computer Architecture Letters, 5 (1):14–17, January 2006. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Mahesri:2007:HSS

[MWP07] Aqeel Mahesri, Nicholas J.

Wang, and Sanjay J. Patel. Hardware support for software controlled multithreading. ACM SIGARCH Computer Architecture News, 35 (1):3–12, March 2007. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Ma:2021:RTB

[MWP^{+}21] Xiaoxue Ma, Shangru Wu, Ernest Pobee, Xiupei Mei, Hao Zhang, Bo Jiang, and Wing-Kwong Chan. RegionTrack: a trace-based sound and complete checker to debug transactional atomicity violations and nonserializable traces. ACM Transactions on Software Engineering and Methodology, 30 (1):7:1–7:49, January 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3412377.

Myllykoski:2022:ATB

[Myl22]  Mirko Myllykoski.  Algo-

rithm 1019: a task-based multi-shift QR /QZ algorithm with aggressive early deflation. ACM Transactions on Mathematical Software, 48 (1):11:1–11:36, March 2022. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3495005.

Naik:2007:CMA

[NA07] Mayur Naik and Alex Aiken. Conditional must not aliasing for static race detection. ACM SIGPLAN Notices, 42(1):327–338, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nikolopoulos:2001:EMA

[NAAL01] D. S. Nikolopoulos, E. Artiaga, E. Ayguadé, and J. Labarta. Exploiting memory affinity in OpenMP through schedule reuse. ACM SIGARCH Computer Architecture News, 29(5):49– 55, December 2001. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Nagle:2001:MFV

[Nag01] Dan Nagle. Multithreading, Fthreads, and Visual Fortran. Dr. Dobb’s Journal of Software Tools, 26(7):36, 38, 40, July 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_07/fthreads.zip.

Nagler:2021:CSR

[Nag21]  Thomas Nagler.  Code

snippet: R-friendly multithreading in C++. Journal of Statistical Software, 97 (??):??, ???? 2021. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v097c01;
https://www.jstatsoft.
org/index.php/jss/article/
view/v097c01/v97c01.pdf.

Nakhimovsky:2001:ISM

[Nak01]  Greg Nakhimovsky.  Im-

proving scalability of multithreaded dynamic memory allocation. Dr. Dobb’s Journal of Software Tools, 26(7):

44, 46, 48–50, 52, 54, July 2001. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2001/2001_07/mthot.txt;
 http://www.ddj.com/ftp/
2001/2001_07/mthot.zip.

Nakajima:2003:PIS

[Nak03]  Kengo Nakajima.  Paral-

lel iterative solvers of GeoFEM with selective blocking preconditioning for nonlinear contact problems on the Earth Simulator. In ACM [ACM03], page ?? ISBN 1-58113-695-

1. LCCN ???? URL http:

//www.sc-conference.org/
sc2003/inter_cal/inter_
cal_detail.php?eventid=
10703#1;   http://www.sc-
conference.org/sc2003/paperpdfs/
pap155.pdf.

Naik:2006:ESR

[NAW06] Mayur Naik, Alex Aiken, and John Whaley. Effective static race detection for Java. ACM SIGPLAN Notices, 41(6):308– 319, June 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Narlikar:1999:SES

[NB99] Girija J. Narlikar and Guy E. Blelloch. Space-efficient scheduling of nested parallelism. ACM Transactions on Programming Languages and Systems, 21(1): 138–173, January 1999. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/toplas/1999-21-
1/p138-narlikar/.

Nagpal:2012:CGE

[NB12] Rahul Nagpal and Anasua Bhowmik. Criticality guided energy aware speculation for speculative multithreaded processors. Parallel Computing, 38(6–7):329–341, June/ July 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167819112000191.

Nichols:1996:PP

[NBF96] Bradford Nichols, Bick Buttlar, and Jackie Proulx Farrell. Pthreads Programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1996. ISBN 1-56592-115-1. xvi + 267 pp. LCCN QA76.642.N53 1996. US$29.95. URL http://www.

amazon.com/exec/obidos/
ASIN/1565921151/ref=sim_
books/002-4892305-5599452;
http://www.oreilly.com/
catalog/pthread.

Nichols:1998:PP

[NBF98] Bradford Nichols, Dick Buttlar, and Jacqueline Proulx Farrell. Pthreads programming. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 1-56592-115-1. xvi + 267 pp. LCCN QA76.642 .N53 1998.

Najjar:1993:QAD

[NBM93] Walid A. Najjar, A. P. Wim Bohm, and W. Marcus Miller. A quantitative analysis of dataflow program execution — preliminaries to a hybrid design. Journal of Parallel and Distributed Computing, 18(3):314–326, July 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1067/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1067/production/pdf.

Nagarakatte:2012:MAP

[NBMM12] Santosh Nagarakatte, Sebastian Burckhardt, Milo M. K. Martin, and Madanlal Musuvathi. Multicore acceleration of priority-based schedulers for concurrency bug detection. ACM SIGPLAN Notices, 47(6):543–554, June 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PLDI ’12 proceedings.

Nelson:2015:RGH

[NBS^{+}15] Thomas Nelson, Geoffrey Belter, Jeremy G. Siek, Elizabeth Jessup, and Boyana Norris. Reliable generation of high-performance matrix algebra. ACM Transactions on Mathematical Software, 41(3): 18:1–18:27, June 2015. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Natarajan:1993:PVM

[NCA93] Venkat Natarajan, Derek Chiou, and Boon Seong Ang. Performance visualization on Monsoon. Journal of Parallel and Distributed Computing, 18(2):169–180, June 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1054/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1054/production/pdf.

Norton:1996:TTM

[ND96] Scott J. Norton and Mark D.

DiPasquale. Thread time: the multithreaded programming guide. Hewlett–Packard professional books. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-190067-6 (paperback). xx + 538 pp. LCCN QA76.642.N67

1996.  URL http://www.
amazon.com/exec/obidos/
ASIN/0131900676/ref=sim_
books/002-4892305-5599452.

Norris:2013:CCC

[ND13] Brian Norris and Brian Demsky. CDSChecker: checking concurrent data structures written with C/C++ atomics. ACM SIGPLAN Notices, 48(10):131–150, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Norris:2016:PAM

[ND16] Brian Norris and Brian Demsky. A practical approach for model checking C/C++11 code. ACM Transactions on Programming Languages and Systems, 38(3):10:1–10:??, May 2016. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Nemeth:2000:AMD

[Ném00] Zsolt Németh. Abstract machine design on a multithreaded architecture. Future Generation Computer Systems, 16(6):705–716, April 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.elsevier.com/gej-ng/
                                        10/19/19/41/29/36/abstract.
                                        html.

Nevison:1999:SSC

[Nev99] Christopher H. Nevison. Seminar: safe concurrent programming in Java with CSP. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (1):367, March 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nazarpour:2017:CPS

[NFBB17] Hosein Nazarpour, Yliès Falcone, Saddek Bensalem, and Marius Bozga. Concurrencypreserving and sound monitoring of multi-threaded component-based systems: theory, algorithms, implementation, and evaluation. Formal Aspects of Computing, 29(6): 951–986, November 2017. CO-DEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00165-017-0422-6.

Nemawarkar:1994:PIN

[NGGA94] S. S. Nemawarkar, R. Govindarajan, G. R. Gao, and V. K. Agarwal. Performance of interconnection network in multithreaded architectures. Lecture Notes in Computer Science, 817:823–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neamtiu:2009:STU

[NH09] Iulian Neamtiu and Michael Hicks. Safe and timely updates to multi-threaded programs. ACM SIGPLAN Notices, 44(6):13–24, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Neamtiu:2008:CEV

[NHFP08] Iulian Neamtiu, Michael Hicks, Jeffrey S. Foster, and Polyvios Pratikakis. Contextual effects for versionconsistent dynamic software updating all and safe concurrent programming. ACM SIG-PLAN Notices, 43(1):37–49, January 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nikhil:1994:MII

[Nik94] Rishiyur S. Nikhil. A multithreaded implementation of Id using P-RISC graphs. Lecture Notes in Computer Science, 768:390–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielsen:2000:MTN

[NJ00] Ida M. B. Nielsen and Curtis L. Janssen. Multithreading: a new dimension to massively parallel scientific computation. Computer Physics Communications, 128 (1–2):238–244, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S001046550000062X.

Narayanaswamy:2016:VCA

[NJK16] Ganesh Narayanaswamy, Saurabh Joshi, and Daniel Kroening. The virtues of conflict: analysing modern concurrency. ACM SIGPLAN Notices, 51(8):25:1–25:??, August 2016. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nicolau:2009:TEP

[NLK09] Alexandru Nicolau, Guangqiang Li, and Arun Kejariwal. Techniques for efficient placement of synchronization primitives. ACM SIGPLAN Notices, 44(4):199–208, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nakaike:2010:LER

[NM10] Takuya Nakaike and Maged M.

Michael. Lock elision for read-only critical sections in Java. ACM SIGPLAN Notices, 45(6):269–278, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nordstrom:1990:TL

[Nor90] D. J. Nordstrom. Threading Lisp. ACM SIGPLAN Notices, 25(2):17–24, February 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Northrup:1996:PUT

[Nor96] Charles J. Northrup. Programming with UNIX Threads. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1996.

ISBN 0-471-13751-0 (paperback). xv + 399 pp. LCCN QA76.76.O63 N674 1996.

Nikhil:1992:MMP

[NPA92] R. S. Nikhil, G. M. Papadopoulos, and Arvind. T: a multithreaded massively parallel architecture. ACM SIGARCH Computer Architecture News, 20(2):156–167, May 1992. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Narayanasamy:2006:RSM

[NPC06] Satish Narayanasamy, Cristiano Pereira, and Brad Calder. Recording shared memory dependencies using strata. ACM SIGPLAN Notices, 41(11):229–240, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nebro:1998:EMR

[NPT98] A. J. Nebro, E. Pimentel, and J. M. Troya. Evaluating a multithreaded runtime system for concurrent object-oriented languages. Lecture Notes in Computer Science, 1505:167– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nanda:2006:ISM

[NR06] Mangala Gowri Nanda and S. Ramesh. Interprocedural slicing of multithreaded programs with applications to Java. ACM Transactions on Programming Languages and Systems, 28(6): 1088–1144, November 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Neves:1997:TRS

[NS97] Richard Neves and Robert B.

Schnabel. Threaded runtime support for execution of fine grain parallel code on coarse grain multiprocessors. Journal of Parallel and Distributed Computing, 42(2):128–142, May 1, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1997.1322/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1997.1322/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1997.1322/production/ref.

Ngo:2014:EVC

[NSH14] Tri Minh Ngo, Mariëlle Stoelinga, and Marieke Huisman. Effective verification of confidentiality for multithreaded programs. Journal of Computer Security, 22(2): 269–300, ???? 2014. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Niewiadomski:2014:SVG

[NSP^{+}14] Artur Niewiadomski, Jaroslaw Skaruz, Wojciech Penczek, Maciej Szreter, and Mariusz Jarocki. SMT versus genetic and OpenOpt algorithms: Concrete planning in the PlanICS framework. Fundamenta Informaticae, 135(4): 451–466, October 2014. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Niu:2014:MCF

[NT14] Ben Niu and Gang Tan. Modular control-flow integrity. ACM SIGPLAN Notices, 49(6):577–587, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nemeth:1999:MLK

[NTKA99] Z. Nemeth, H. Tomiyasu, P. Kacsuk, and M. Amamiya. Multithreaded LOGFLOW on KUMP/ D. Lecture Notes in Computer Science, 1615:320– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nogueira:2016:BBW

[NTR16] David Nogueira, Pedro Tomas, and Nuno Roma. BowMapCL: Burrows–Wheeler mapping on multiple heterogeneous accelerators. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 13 (5):926–938, September 2016. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Norwood:1994:SMP

[NV94] John Norwood and Shankar Vaidyanathan. Symmetric multiprocessing for PCs. Dr. Dobb’s Journal of Software Tools, 19(1):80, 82–85, 98– 99, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Nguyen:2015:RCC

[NV15] Phúc C. Nguyěn and David Van Horn. Relatively complete counterexamples for higher-order programs. ACM SIGPLAN Notices, 50(6):446– 456, June 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Narayanasamy:2007:ACB

[NWT^{+}07] Satish Narayanasamy, Zhenghao Wang, Jordan Tigani, Andrew Edwards, and Brad Calder. Automatically classifying benign and harmful data races all using replay analysis. ACM SIGPLAN Notices, 42(6):22–31, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nutaro:2017:HAA

[NZ17] James Nutaro and Bernard Zeigler. How to apply Amdahl’s law to multithreaded multicore processors. Journal of Parallel and Distributed Computing, 107(??): 1–2, September 2017. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0743731517300941.

Ottoni:2008:COGa

[OA08a] Guilherme Ottoni and David I. August. Communication optimizations for global multithreaded instruction scheduling. ACM SIGARCH Computer Architecture News, 36 (1):222–232, March 2008. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Ottoni:2008:COGb

[OA08b] Guilherme Ottoni and David I. August. Communication optimizations for global multithreaded instruction scheduling. Operating Systems Review, 42(2):222–232, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ottoni:2008:COGc

[OA08c] Guilherme Ottoni and David I. August. Communication optimizations for global multithreaded instruction scheduling. ACM SIGPLAN Notices, 43(3):222–232, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oz:2019:SMA

[OA19] Isil Oz and Sanem Arslan.

A survey on multithreading alternatives for soft error fault tolerance. ACM Computing Surveys, 52(2): 27:1–27:??, May 2019. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3302255.

Olszewski:2009:KED

[OAA09] Marek Olszewski, Jason Ansel, and Saman Amarasinghe. Kendo: efficient deterministic multithreading in software. ACM SIGPLAN Notices, 44(3):97–108, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ossner:2013:GMB

[OB13] Christopher Oßner and Klemens Böhm. Graphs for mining-based defect localization in multithreaded programs. International Journal of Parallel Programming, 41(4):570–593, August 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10766-012-0237-2.

Ostler:2007:IHT

[OCRS07] Chris Ostler, Karam S.

Chatha, Vijay Ramamurthi, and Krishnan Srinivasan. ILP and heuristic techniques for system-level design on network processor architectures. ACM Transactions on Design Automation of Electronic Systems., 12(4):48:1– 48:??, September 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ozer:2001:WMT

[ÖCS01] Emre Özer, Thomas M. Conte, and Saurabh Sharma. Weld: a multithreading technique towards latencytolerant VLIW processors. Lecture Notes in Computer Science, 2228:192–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2228/22280192.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2228/22280192.
pdf.

Odaira:2014:EGI

[OCT14] Rei Odaira, Jose G. Castanos, and Hisanobu Tomari. Eliminating global interpreter locks in Ruby through hardware transactional memory. ACM SIGPLAN Notices, 49(8):131– 142, August 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Olivier:2012:CMW

[OdSSP12] Stephen L. Olivier, Bronis R. de Supinski, Martin Schulz, and Jan F. Prins. Characterizing and mitigating work time inflation in task parallel programs. In Hollingsworth [Hol12], pages 65:1–65:12. ISBN 1-4673-0804-8. URL http:

//conferences.computer.
org/sc/2012/papers/1000a066.
pdf.

Ogata:1992:DIH

[OKID92] Kazuhiro Ogata, Satoshi Kurihara, Mikio Inari, and Norihisa Doi. The design and implementation of HoME. ACM SIG-PLAN Notices, 27(7):44–54, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/pldi/143095/
p44-ogata/.

Oplinger:2002:ESRa

[OL02a] Jeffrey Oplinger and Monica S. Lam. Enhancing software reliability with speculative threads. ACM SIGARCH Computer Architecture News, 30(5):184– 196, December 2002. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Oplinger:2002:ESRb

[OL02b] Jeffrey Oplinger and Monica S. Lam. Enhancing software reliability with speculative threads. ACM SIGPLAN Notices, 37(10):184–196, October 2002. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oplinger:2002:ESRc

[OL02c] Jeffrey Oplinger and Monica S. Lam. Enhancing software reliability with speculative threads. Operating Systems Review, 36(5):184– 196, December 2002. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Omma:2004:BMA

[Omm04] M. Omma. On building multithreaded applications. IEEE Distributed Systems Online, 5 (4):1–3, April 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/iel5/8968/28913/01301256.
                                        pdf?isnumber=28913&prod=
                                        JNL&arnumber=1301256&arSt=
                                        +1&ared=+3&arAuthor=Omma%
                                        2C+M.;   http://ieeexplore.
                                        ieee.org/xpls/abs_all.jsp?
                                        isnumber=28913&arnumber=
                                        1301256&count=5&index=3.

Ongwattanakul:1997:RDM

[Ong97] Songpol Ongwattanakul. A runtime distributed multithreading library for the PARC language. Thesis (M.E.E.), Department of Electrical Engineering, University of Alabama, Tuscaloosa, AL, USA, 1997. viii + 71 pp.

Onion:1997:MM

[Oni97] F. Onion. Multithreading in MFC. C++ Report, 9(3):50– 53, 56, March 1997. CODEN CRPTE7. ISSN 1040-6042.

Oh:2012:MTS

[OR12] Doohwan Oh and Won W.

Ro. Multi-threading and suffix grouping on massive multiple pattern matching algorithm. The Computer Journal, 55 (11):1331–1346, November 2012. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/content/55/11/1331.
full.pdf+html.

Odersky:1993:CNA

[ORH93] Martin Odersky, Dan Rabin, and Paul Hudak. Call by name, assignment, and the lambda calculus. In ACM [ACM93a], pages 43–

56. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15 1993.

URL http://www.acm.org:
80/pubs/citations/proceedings/
plan/158511/p43-odersky/.

ACM order number 549930.

Ottoni:2006:SPC

[ORS^{+}06] G. Ottoni, R. Rangan, A. Stoler, M. J. Bridges, and D. I. August. From sequential programs to concurrent threads. IEEE Computer Architecture Letters, 5(1):6–9, January 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Oikawa:1995:RDU

[OT95] Shuichi Oikawa and Hideyuki Tokuda. Reflection of developing user-level real-time thread packages. Operating Systems Review, 29(4):63–76, October 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Oyama:2000:OCC

[OTY00] Yoshihiro Oyama, Kenjiro Taura, and Akinori Yonezawa. Online computation of critical paths for multithreaded languages. Lecture Notes in Computer Science, 1800:301– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1800/18000301.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1800/18000301.
pdf.

Oaks:1997:JT

[OW97] Scott Oaks and Henry Wong. Java threads. Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-216-6. xiii + 252 pp. LCCN QA76.73.J38 O25 1997.

Oaks:1999:JT

[OW99] Scott Oaks and Henry Wong.

Java threads. Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1999. ISBN 1-56592-418-

5. xiii + 319 pp. LCCN QA76.73.J38 O25 1999.

Peternier:2014:IEU

[PAB^{+}14] Achille Peternier, Danilo Ansaloni, Daniele Bonetta, Cesare Pautasso, and Walter Binder. Improving execution unit occupancy on SMT-based processors through hardwareaware thread scheduling. Future Generation Computer Systems, 30(??):229–241, January 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X13001295.

Pereira:2017:SBC

[PAdS^{+}17] Phillipe Pereira, Higo Albuquerque, Isabela da Silva, Hendrio Marques, Felipe Monteiro, Ricardo Ferreira, and Lucas Cordeiro. SMTbased context-bounded model checking for CUDA programs. Concurrency and Computation: Practice and Experience, 29(22):??, November 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pant:1999:TCP

[Pan99] Lalit Pant. Thread communication in parallel algorithms: Enabling efficient interaction between threads. Dr. Dobb’s Journal of Software Tools, 24 (4):32, 34, 36, 38–39, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/1999/
                                        1999_04/parallel.txt.

Park:1991:PTM

[Par91] Won Woo Park. Performancearea trade-offs in multithreaded processing unit. Thesis (Ph.D.), University of Texas at Austin, Austin, TX, USA, 1991. xvii + 165 pp.

Papadopoulos:1992:MCS

[PBDO92] G. M. Papadopoulos, A. P. W. Bohm, A. T. Dahbura, and R. R. Oldehoeft. Multithreaded computer systems. In IEEE Computer Society. Technical Committee on Computer Architecture [IEE92], pages 772–775. ISBN 0-8186-2632-1 (case), 0-8186-2630-5 (paper), 0-8186-2631-3 (microfiche), 0-89791-537-2 (ACM Library series). LCCN QA76.5 .S894 1992. Cover title: Supercomputing ’91. ACM order number 415922. IEEE Computer Society Press order number 2630 IEEE catalog number 92CH3216-9.

Park:2017:HHC

[PBL^{+}17] Jaehyun Park, Seungcheol Baek, Hyung Gyu Lee, Chrysostomos Nicopoulos, Vinson Young, Junghee Lee, and Jongman Kim. HoPE: Hot-cacheline prediction for dynamic early decompression in compressed LLCs. ACM Transactions on Design Automation of Electronic Systems., 22(3):40:1–40:??, May 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Porter:2015:PFG

[PBR^{+}15] Donald E. Porter, Michael D.

Bond, Indrajit Roy, Kathryn S. Mckinley, and Emmett Witchel. Practical fine-grained information flow control using Laminar. ACM Transactions on Programming Languages and Systems, 37(1): 4:1–4:??, January 2015. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Park:2016:CJP

[PCM16] Hyukwoo Park, Myungsu Cha, and Soo-Mook Moon. Concurrent JavaScript parsing for faster loading of Web apps. ACM Transactions on Architecture and Code Optimization, 13(4):41:1–41:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Perez:2015:ECR

[PCPS15] J. F. Pérez, G. Casale, and S. Pacheco-Sanchez. Estimating computational requirements in multi-threaded applications. IEEE Transactions on Software Engineering, 41(3):264–278, March 2015. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6926798.

Papadopoulos:2016:TAD

[PDMM16] Stavros Papadopoulos, Kushal Datta, Samuel Madden, and Timothy Mattson. The TileDB array data storage manager. Proceedings of the VLDB Endowment, 10(4): 349–360, November 2016. CO-DEN ???? ISSN 2150-8097.

Pokam:2013:QPI

[PDP^{+}13] Gilles Pokam, Klaus Danne, Cristiano Pereira, Rolf Kassa, Tim Kranich, Shiliang Hu, Justin Gottschlich, Nima Honarmand, Nathan Dautenhahn, Samuel T. King, and Josep Torrellas. QuickRec: prototyping an Intel architecture extension for record and replay of multithreaded programs. ACM SIGARCH Computer Architecture News, 41 (3):643–654, June 2013. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ICSA ’13 conference proceedings.

Peacock:1992:FSM

[Pea92] J. Kent Peacock. File system multithreading in System V Release 4 MP. In USENIX [USE92a], pages 19–30. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Philbin:1996:TSC

[PEA^{+}96] James Philbin, Jan Edler, Otto J. Anshus, Craig C. Douglas, and Kai Li. Thread scheduling for cache locality. ACM SIGPLAN Notices, 31(9):60–71, September 1996. CODEN SIN-ODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/asplos/237090/

p60-philbin/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Peterson:2000:CCT

[Pet00] Mark Peterson. C/C++ tips:

Tip #4: Self destructing threads. C/C++ Users Journal, 18(12):44–??, December 2000. CODEN CCUJEX. ISSN 1075-2838.

Petitpierre:2003:JTC

[Pet03] C. Petitpierre. Java threads can be very useful building blocks. Lecture Notes in Computer Science, 2604:204, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plakal:2001:CGC

[PF01] Manoj Plakal and Charles N. Fischer. Concurrent garbage collection using program slices on multithreaded processors. ACM SIGPLAN Notices, 36(1):94–100, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pratikakis:2006:LCS

[PFH06] Polyvios Pratikakis, Jeffrey S. Foster, and Michael Hicks. LOCKSMITH: context-sensitive correlation analysis for race detection. ACM SIGPLAN Notices, 41(6):320–331, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pons:2022:EHT

[PFP^{+}22] Lucía Pons, Josué Feliu, José Puche, Chaoyi Huang, Salvador Petit, Julio Pons, María E. Gómez, and Julio Sahuquillo. Effect of hyperthreading in latency-critical multithreaded cloud applications and utilization analysis of the major system resources. Future Generation Computer Systems, 131(??):194–208, June 2022. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0167739X22000334.

Park:2003:IMP

[PFV03] Il Park, Babak Falsafi, and T. N. Vijaykumar. Implicitly-multithreaded processors. ACM SIGARCH Computer Architecture News, 31(2):39–51, May 2003. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Pham:1992:MDA

[PG92] Thuan Q. Pham and Pankaj K.

Garg. On migrating a distributed application to a multithreaded environment. In USENIX [USE92a], pages 45– 54. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Pham:1996:MPW

[PG96] Thuan Q. Pham and Pankaj K.

Garg. Multithreaded programming with Windows NT. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-120643-5. xviii + 227 pp. LCCN QA76.642 .P52 1996.

Pham:1999:MPW

[PG99] Thuan Q. Pham and Pankaj K.

Garg. Multithreaded Programming with Win32. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-010912-6. xix + 219 pp. LCCN QA76.642.P518

1998.  URL http://www.
phptr.com/ptrbooks/ptr_

                                        0130109126.html. Includes

CD-ROM.

Parcerisa:2001:ILT

[PG01] J.-M. Parcerisa and A. Gonzalez. Improving latency tolerance of multithreading through decoupling. IEEE Transactions on Computers, 50(10):1084–1094, October 2001. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&
                                        arnumber=956093.

Pinilla:2003:UJT

[PG03] Ruben Pinilla and Marisa Gil. ULT: a Java threads model for platform independent execution. Operating Systems Review, 37(4):48–62, October 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pusukuri:2012:TTD

[PGB12] Kishore Kumar Pusukuri, Rajiv Gupta, and Laxmi N. Bhuyan. Thread tranquilizer: Dynamically reducing performance variation. ACM Transactions on Architecture and Code Optimization, 8(4):46:1– 46:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Pusukuri:2014:LCA

[PGB14] Kishore Kumar Pusukuri, Rajiv Gupta, and Laxmi Narayan Bhuyan. Lock contention aware thread migrations. ACM SIGPLAN Notices, 49 (8):369–370, August 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pusukuri:2016:TEL

[PGB16] Kishore Kumar Pusukuri, Rajiv Gupta, and Laxmi N. Bhuyan. Tumbler: an effective load-balancing technique for multi-CPU multicore systems. ACM Transactions on Architecture and Code Optimization, 12(4):36:1–36:??, January 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Park:1997:HPM

[PH97] Sung-Yong Park and Salim Hariri. A high performance message passing system for Network of Workstations. The Journal of Supercomputing, 11(2):159–180, October 1997. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=11&
issue=2&spage=159;   http:
//www.wkap.nl/oasis.htm/
149826.

Pham:1991:EMD

[Pha91] Thuan Quang Pham. The experimental migration of a distributed application to a multithreaded environment. Thesis (M.S.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1991. 51 pp.

Pham:2018:TSM

[PHBC18] Binh Pham, Derek Hower, Abhishek Bhattacharjee, and Trey Cain. TLB shootdown mitigation for low-power many-core servers with L1 virtual caches. IEEE Computer Architecture Letters, 17(1): 17–20, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Pichel:2009:IDR

[PHCR09] J. C. Pichel, D. B. Heras, J. C. Cabaleiro, and F. F. Rivera. Increasing data reuse of sparse algebra codes on simultaneous multithreading architectures. Concurrency and Computation: Practice and Experience, 21(15):1838–1856, October 2009. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ponamgi:1991:DMP

[PHK91] M. Krish Ponamgi, Wenwey Hseush, and Gail E. Kaiser. Debugging multithreaded programs with MPD. IEEE Software, 8(3):37–43, May 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Pager:2015:SSM

[PJS15] Jared Pager, Reiley Jeyapaul, and Aviral Shrivastava. A software scheme for multithreading on CGRAs. ACM Transactions on Embedded Computing Systems, 14 (1):19:1–19:??, January 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Park:2007:MEP

[PJZA07] Soyeon Park, Weihang Jiang, Yuanyuan Zhou, and Sarita Adve. Managing energyperformance tradeoffs for multithreaded applications on multiprocessor architectures. ACM SIGMETRICS Performance Evaluation Review, 35 (1):169–180, June 2007. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Powell:1991:SMT

[PKB^{+}91] M. L. Powell, S. R. Kleiman, S. Barton, D. Shah, D. Stein, and M. Weeks. SunOS multi-thread architecture. In USENIX [USE91b], pages 65– 80. LCCN QA 76.76 O63 U84 1992.

Price:2003:CAF

[PL03] Gregory W. Price and David K.

Lowenthal. A comparative analysis of fine-grain threads packages. Journal of Parallel and Distributed Computing, 63(11):1050–1063, November 2003. CODEN JPDCER.

ISSN 0743-7315 (print), 1096-0848 (electronic).

Plauger:1993:MCS

[Pla93] Dave Plauger. Making C++ save for threads. C Users Journal, 11(2):58–??, February 1993. ISSN 0898-9788.

Plauger:1998:SCCl

[Pla98]  P. J. Plauger.  Standard

C/C++: Thread safety. C/C++ Users Journal, 16 (12):??, December 1998. CO-DEN CCUJEX. ISSN 1075-2838.

Plauger:1999:SCCg

[Pla99]  P. J. Plauger.  Standard

C/C++: a better red-black tree. C/C++ Users Journal, 17(7):10–??, July 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Plachetka:2002:QTS

[Pla02] Tomas Plachetka. (quasi-) thread-safe PVM and (quasi-) thread-safe MPI without active polling. Lecture Notes in Computer Science, 2474:296– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2474/24740296.htm;   http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2474/24740296.pdf.

Porter:2015:MMS

[PLT^{+}15] Leo Porter, Michael A. Laurenzano, Ananta Tiwari, Adam Jundt, William A. Ward, Jr., Roy Campbell, and Laura Carrington. Making the most of SMT in HPC: Systemand application-level perspectives. ACM Transactions on Architecture and Code Optimization, 11(4):59:1–59:??, January 2015. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Plyler:1989:AMC

[Ply89] Kevin Brian Plyler. Adding multithreaded capabilities to the process manager of the BIGSAM distributed operating system. Thesis (M.S.), Arizona State University, Tempe, AZ, USA, 1989. x + 105 + 2 pp.

Pricopi:2014:TSA

[PM14] M. Pricopi and T. Mitra. Task scheduling on adaptive multicore. IEEE Transactions on Computers, 63(10):2590– 2603, October 2014. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Perrin:2022:EWF

[PMCP22] Matthieu Perrin, Achour Mostéfaoui, and Ludmila Courtillat-Piazza. Extending the wait-free hierarchy to multi-threaded systems. Distributed Computing, 35(4): 375–398, August 2022. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s00446-022-00425-x.

Prabhu:2003:UTL

[PO03] Manohar K. Prabhu and Kunle Olukotun. Using thread-level speculation to simplify manual parallelization. ACM SIGPLAN Notices, pages 1–12, 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Polychronopoulos:1990:ASC

[Pol90] C. D. (Constantine D.) Polychronopoulos. Auto scheduling: control flow and data flow come together. Technical Report CSRD 1058, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1990. 28 pp.

Pomerantz:1998:CNS

[Pom98] Dave Pomerantz. C++ notifiers: Simplifying system development. Dr. Dobb’s Journal of Software Tools, 23(8): 26, 28, 30–31, 89–90, August 1998. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/1998/
                                        1998_08/notifier.txt;
                                        http://www.ddj.com/ftp/
                                        1998/1998_08/notifier.zip.

Parashar:2013:TIC

[PPA^{+}13] Angshuman Parashar, Michael Pellauer, Michael Adler, Bushra Ahsan, Neal Crago, Daniel Lustig, Vladimir Pavlov, Antonia Zhai, Mohit Gambhir, Aamer Jaleel, Randy Allmon, Rachid Rayess, Stephen Maresh, and Joel Emer. Triggered instructions: a control paradigm for spatiallyprogrammed architectures. ACM SIGARCH Computer Architecture News, 41(3): 142–153, June 2013. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ICSA ’13 conference proceedings.

Prieto:2011:MCM

[PPG11] Pablo Prieto, Valentin Puente, and Jose-Angel Gregorio. Multilevel cache modeling for chip-multiprocessor systems. IEEE Computer Architecture Letters, 10(2):49–52, July/ December 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Puche:2020:ECF

[PPGS20] José Puche, Salvador Petit, María E. Gómez, and Julio Sahuquillo. An efficient cache flat storage organization for multithreaded workloads for low power processors. Future Generation Computer Systems, 110(??): 1037–1054, September 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X1930384X.

Piumarta:1998:ODT

[PR98] Ian Piumarta and Fabio Riccardi. Optimizing directthreaded code by selective inlining. ACM SIG-PLAN Notices, 33(5):291–300, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/pldi/277650/
                                        p291-piumarta/.

Petric:2005:EEP

[PR05] Vlad Petric and Amir Roth. Energy-effectiveness of preexecution and energy-aware P-thread selection. ACM SIGARCH Computer Architecture News, 33(2):322–333, May 2005. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Prabhakar:1995:IDO

[Pra95a] Ernest N. Prabhakar. Implementing distributed objects. Dr. Dobb’s Journal of Software Tools, 20(8):80, 82, 84– 85, 105–106, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Prasad:1995:WTS

[Pra95b] Shashi Prasad. Weaving a thread — Solaris and Windows NT bring the power, speed, and efficiency of multithreading and symmetric multiprocessing to the desktop.

Byte Magazine, 20(10):173– ??, October 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Prasad:1995:WNT

[Pra95c] Shashi Prasad. Windows NT threads — a multithreaded application may actually run slower on an SMP machine than on its single-threaded equivalent. Here’s how to avoid that. Byte Magazine, 20 (11):253–??, November 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Prasad:1997:MPT

[Pra97] Shashi Prasad. Multithreading programming techniques. The J. Ranade workstation series. McGraw-Hill, New York, NY, USA, 1997. ISBN 0-07-912250-7, 0-07-050710-4 (Computer disk). xix + 410 pp. LCCN QA76.76.D47 P72 1997.

Permandla:2007:TSP

[PRB07] Pratibha Permandla, Michael Roberson, and Chandrasekhar Boyapati. A type system for preventing data races and deadlocks in the Java Virtual Machine language: 1. ACM SIGPLAN Notices, 42(7):10, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Presotto:1990:MSP

[Pre90] D. L. Presotto. Multiprocessor streams for Plan 9. In UKUUG. UNIX - The Legend Evolves. Proceedings of the Summer 1990 UKUUG Conference, pages 11–19 (of xi + 260). UK Unix Users Group, Buntingford, Herts, UK, ???? 1990. ISBN 0-9513181-7-9. LCCN ????

Petrovic:2014:LHM

[PRS14] Darko Petrović, Thomas Ropars, and André Schiper. Leveraging hardware message passing for efficient thread synchronization. ACM SIG-PLAN Notices, 49(8):143–154, August 2014. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Protopopov:2001:MMP

[PS01] Boris V. Protopopov and Anthony Skjellum. A multithreaded Message Passing Interface (MPI) architecture: Performance and program issues. Journal of Parallel and Distributed Computing, 61(4):449–466, April 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1674;   http://www.
idealibrary.com/links/doi/
10.1006/jpdc.2000.1674/
pdf;   http://www.idealibrary.

com/links/doi/10.1006/jpdc.
2000.1674/ref.

Pozniansky:2003:EFD

[PS03] Eli Pozniansky and Assaf Schuster. Efficient on-the-fly data race detection in multithreaded C++ programs. ACM SIGPLAN Notices, pages 179–190, 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pozniansky:2007:MEF

[PS07] Eli Pozniansky and Assaf Schuster. MultiRace: efficient on-the-fly data race detection in multithreaded C++ programs. Concurrency and Computation: Practice and Experience, 19(3):327– 340, March 10, 2007. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pyarali:2001:EOT

[PSCS01] Irfan Pyarali, Marina Spivak, Ron Cytron, and Douglas C. Schmidt. Evaluating and optimizing thread pool strategies for real-time CORBA. ACM SIGPLAN Notices, 36(8):214–222, August 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parashar:2006:SSBa

[PSG06a] Angshuman Parashar, Anand Sivasubramaniam, and Sudhanva Gurumurthi. SlicK: slice-based locality exploitation for efficient redundant multithreading. ACM SIGARCH Computer Architecture News, 34(5):95– 105, December 2006. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Parashar:2006:SSBb

[PSG06b] Angshuman Parashar, Anand Sivasubramaniam, and Sudhanva Gurumurthi. SlicK: slice-based locality exploitation for efficient redundant multithreading. Operating Systems Review, 40(5):95– 105, December 2006. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Parashar:2006:SSBc

[PSG06c] Angshuman Parashar, Anand Sivasubramaniam, and Sudhanva Gurumurthi. SlicK: slice-based locality exploitation for efficient redundant multithreading. ACM SIG-PLAN Notices, 41(11):95–105, November 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pang:2001:PSR

[PSM01] James Pang, Gholamali Shoja, and Eric Manning. Providing soft real-time QoS guarantees for Java threads. In ACM [ACM01], pages 39–46. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

2001.  URL http://www.
philippsen.com/JGI2001/
camerareadyabstracts/21.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500039.pdf.

Pang:2003:PSR

[PSM03] James C. Pang, Gholamali C.

Shoja, and Eric G. Manning. Providing soft real-time quality of service guarantees for Java threads. Concurrency and Computation: Practice and Experience, 15(3–5):521– 538, March/April 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peacock:1992:EMS

[PST^{+}92] J. Kent Peacock, Sunil Saxena, Dean Thomas, Fred Yang, and Wilfred Yu. Experiences from multithreading system V release 4. In USENIX [USE92b], pages 77–

92. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Papadopoulos:1991:MRV

[PT91] Gregory M. Papadopoulos and Kenneth R. Traub. Multithreading: a revisionist view of dataflow architectures. ACM SIGARCH Computer Architecture News, 19 (3):342–351, May 1991. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Prvulovic:2003:RUT

[PT03] Milos Prvulovic and Josep Torrellas. ReEnact: using thread-level speculation mechanisms to debug data races in multithreaded codes. ACM SIGARCH Computer Architecture News, 31(2):110–121, May 2003. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Piringer:2009:MTA

[PTMB09] Harald Piringer, Christian Tominski, Philipp Muigg, and Wolfgang Berger. A multithreading architecture to support interactive visual exploration. IEEE Transactions on Visualization and Computer Graphics, 15(6):1113– 1120, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pfeffer:2004:RTG

[PUF^{+}04] M. Pfeffer, T. Ungerer, S. Fuhrmann, J. Kreuzinger, and U. Brinkschulte. Realtime garbage collection for a multithreaded Java microcontroller. Real-Time Systems, 26(1):89–106, 2004. CODEN RESYE9. ISSN 0922-6443.

Pulleyn:2000:EPM

[Pul00] Ivan Pulleyn. Embedding Python in multi-threaded C/ C++ applications. Linux Journal, 73:??, May 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Pickett:2006:SSF

[PV06] Christopher J. F. Pickett and Clark Verbrugge. Sable-SpMT: a software framework for analysing speculative multithreading in Java. ACM SIGSOFT Software Engineering Notes, 31(1):59–66, January 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pathania:2017:DTM

[PVS^{+}17] Anuj Pathania, Vanchinathan Venkataramani, Muhammad Shafique, Tulika Mitra, and Jörg Henkel. Defragmentation of tasks in many-core architecture. ACM Transactions on Architecture and Code Optimization, 14(1):2:1– 2:??, April 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Preissl:2012:CSS

[PWD^{+}12] Robert Preissl, Theodore M.

Wong, Pallab Datta, Myron Flickner, Raghavendra Singh, Steven K. Esser, William P. Risk, Horst D. Simon, and Dharmendra S. Modha. Compass: a scalable simulator for an architecture for cognitive computing. In Hollingsworth [Hol12], pages 54:1–54:11. ISBN 1-4673-0804-8. URL http:

//conferences.computer.

org/sc/2012/papers/1000a085.
pdf.

Preissl:2011:MGA

[PWL^{+}11] Robert Preissl, Nathan Wichmann, Bill Long, John Shalf, Stephane Ethier, and Alice Koniges. Multithreaded global address space communication techniques for gyrokinetic fusion applications on ultrascale platforms. In Lathrop et al. [LCK11], pages 12:1– 12:11. ISBN 1-4503-0771-X. LCCN ????

Polap:2018:MTL

[PWWD18] Dawid Polap, Marcin Woźniak, Wei Wei, and Robertas Damasevicius. Multi-threaded learning control mechanism for neural networks. Future Generation Computer Systems, 87(??):16–34, October 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S0167739X18300931.

Park:2010:ISP

[PYP^{+}10] Jung-Wook Park, Hoon-Mo Yang, Gi-Ho Park, Shin-Dug Kim, and Charles C. Weems. An instruction-systolic programmable shader architecture for multi-threaded 3D graphics processing. Journal of Parallel and Distributed Computing, 70(11): 1110–1118, November 2010. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Quintana-Orti:2012:RSP

[QOIM^{+}12] Gregorio Quintana-Ortí, Francisco D. Igual, Mercedes Marqués, Enrique S. Quintana-Ortí, and Robert A. van de Geijn. A runtime system for programming out-of-core matrix algorithms-by-tiles on multithreaded architectures. ACM Transactions on Mathematical Software, 38(4):25:1– 25:25, August 2012. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Quintana-Orti:2009:PMA

[QOQOV^{+}09] Gregorio Quintana-Ortí, Enrique S. Quintana-Ortí, Robert A. Van De Geijn, Field G. Van Zee, and Ernie Chan. Programming matrix algorithmsby-blocks for thread-level parallelism. ACM Transactions on Mathematical Software, 36(3):14:1–14:26, July 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Qian:2016:EFS

[QSaS^{+}16] Junjie Qian, Witawas Srisaan, Sharad Seth, Hong Jiang, Du Li, and Pan Yi. Exploiting FIFO scheduler to improve parallel garbage collection performance. ACM SIG-PLAN Notices, 51(7):109–121, July 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Qian:2016:ODG

[QSHI16] Xuehai Qian, Koushik Sen, Paul Hargrove, and Costin Iancu. OPR: deterministic group replay for one-sided communication. ACM SIG-PLAN Notices, 51(8):47:1– 47:??, August 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Qian:2014:PRR

[QSQ14] Xuehai Qian, Benjamin Sahelices, and Depei Qian. Pacifier: record and replay for relaxed-consistency multiprocessors with distributed directory protocol. ACM SIGARCH Computer Architecture News, 42(3):433– 444, June 2014. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Rajagopal:1993:DMI

[Raj93] Arjun Rajagopal. Design of a multithreaded instruction cache for a hyperscalar processor. Thesis (M.S.), Department of Electrical Engineering, Texas A&M University, College Station, TX, USA, 1993. ix + 84 pp.

Ramsey:1994:CTB

[Ram94] Norman Ramsey. Correctness of trap-based breakpoint implementations. In ACM [ACM94b], pages 15– 24. ISBN 0-89791-636-0.

LCCN QA76.7 .A15 1994.

URL http://www.acm.org:
80/pubs/citations/proceedings/
plan/174675/p15-ramsey/.

Roberts:2018:MID

[RB18] Malcolm Roberts and John C.

Bowman. Multithreaded implicitly dealiased convolutions. Journal of Computational Physics, 356(??): 98–114, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0021999117308641.

Ritchie:2022:DPF

[RB22] Robert Ritchie and Khodakhast Bibak. DOTMIX-Pro: faster and more efficient variants of DOTMIX for dynamic-multithreading platforms. The Journal of Supercomputing, 78(1):945– 961, January 2022. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s11227-021-03904-3.

Rufai:2005:MPO

[RBAA05] Raimi Rufai, Muslim Bozyigit, Jaralla Alghamdi, and Moataz Ahmed. Multithreaded parallelism with OpenMP. Parallel Processing Letters, 15(4): 367–378, December 2005. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Rashid:1989:MFO

[RBF^{+}89] R. Rashid, R. Baron, A. Forin, D. Golub, M. Jones, D. Orr, and R. Sanzi. Mach: a foundation for open systems (operating systems). In IEEE [IEE89], pages 109–113. IEEE catalog number 89TH0281-6.

Ratanaworabhan:2009:DTA

[RBK^{+}09] Paruj Ratanaworabhan, Martin Burtscher, Darko Kirovski, Benjamin Zorn, Rahul Nagpal, and Karthik Pattabiraman. Detecting and tolerating asymmetric races. ACM SIGPLAN Notices, 44(4):173– 184, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ranganathan:2000:AMT

[RBPM00] M. Ranganathan, Mark Bednarek, Fernand Pors, and Doug Montgomery. AGNI: a multi-threaded middleware for distributed scripting. In USENIX [USE00b], page ?? ISBN 1-880446-24-3. LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/tcl2k/
                                        ranganathan.html.

Reda:2012:APC

[RCC12] Sherief Reda, Ryan Cochran, and Ayse K. Coskun. Adaptive power capping for servers with multithreaded workloads. IEEE Micro, 32(5):64– 75, September/October 2012.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rahman:2014:CCO

[RCC14] Musfiq Rahman, Bruce R.

Childers, and Sangyeun Cho. COMeT+: Continuous online memory testing with multithreading extension. IEEE Transactions on Computers, 63(7):1668–1681, July 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ro:2006:DEH

[RCDG06] Won W. Ro, Stephen P.

Crago, Alvin M. Despain, and Jean-Luc Gaudiot. Design and evaluation of a hierarchical decoupled architecture. The Journal of Supercomputing, 38(3):237–259, December 2006. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=38&
issue=3&spage=237.

Rakvic:2010:TMT

[RCG^{+}10] R. Rakvic, Q. Cai, J. González, G. Magklis, P. Chaparro, and A. González. Threadmanagement techniques to maximize efficiency in multicore and simultaneous multithreaded microprocessors. ACM Transactions on Architecture and Code Optimization, 7(2):9:1–9:??, September 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Radojkovic:2012:OTA

[RCM^{+}12] Petar Radojković, Vladimir Cakarević, Miquel Moretó, Javier Verdú, Alex Pajuelo, Francisco J. Cazorla, Mario Nemirovsky, and Mateo Valero. Optimal task assignment in multithreaded processors: a statistical approach. ACM SIGARCH Computer Architecture News, 40(1):235–248, March 2012. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ASPLOS ’12 conference proceedings.

Radojkovic:2016:TAM

[RCM^{+}16] P. Radojkovic, P. M. Carpenter, M. Moreto, V. Cakarevic, J. Verdu, A. Pajuelo, F. J. Cazorla, M. Nemirovsky, and M. Valero. Thread assignment in multicore/multithreaded processors: A statistical approach. IEEE Transactions on Computers, 65(1):256–269, ???? 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rogers:1995:SDD

[RCRH95] Anne Rogers, Martin C. Carlisle, John H. Reppy, and L. J. Hendren. Supporting dynamic data structures on distributed-memory machines. ACM Transactions on Programming Languages and Systems, 17(2): 233–263, March 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0164-0925/
201065.html.

Radojkovic:2010:TSB

[RČV^{+}10] Petar Radojković, Vladimir Čakarević, Javier Verdú, Alex Pajuelo, Francisco J. Cazorla, Mario Nemirovsky, and Mateo Valero. Thread to strand binding of parallel network applications in massive multithreaded systems. ACM SIG-PLAN Notices, 45(5):191–202, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ruddock:1996:MPG

[RD96] David E. Ruddock and Balakrishnan Dasarathy. Multithreading programs: Guidelines for DCE applications. IEEE Software, 13(1):80– 90, January 1996. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Ronsse:1999:RFI

[RD99] Michiel Ronsse and Koen De Bosschere. RecPlay: a fully integrated practical record/ replay system. ACM Transactions on Computer Systems, 17(2):133–152, May 1999. CODEN ACSYEC.

ISSN 0734-2071 (print), 1557-7333 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/tocs/
                                        1999-17-2/p133-ronsse/.

Russell:2006:ESRa

[RD06] Kenneth Russell and David Detlefs. Eliminating synchronizationrelated atomic operations with biased locking and bulk rebiasing. ACM SIGPLAN Notices, 41(10):263–272, October 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reck:1998:TSR

[Rec98]  Bill Reck.  Thread syn-

chronization with referencecounting handles. C/C++ Users Journal, 16(2):??, February 1998. CODEN CCUJEX. ISSN 1075-2838.

Reich:1995:DHP

[Rei95] David E. Reich. Designing high-powered OS/2 Warp applications: the anatomy of multithreaded programs. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-11586-X (paperback). xxxi + 336 pp. LCCN QA76.76.O63R437 1995.

Reilly:2001:TNF

[Rei01]  Douglas Reilly.  Thread-

ing and the .Net framework. Dr. Dobb’s Journal of Software Tools, 26(8):

30, 32–33, 36, 38, August 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_08/thrednet.txt.

Redstone:2000:AOSa

[REL00a] Joshua A. Redstone, Susan J.

Eggers, and Henry M. Levy. An analysis of operating system behavior on a simultaneous multithreaded architecture. ACM SIGARCH Computer Architecture News, 28 (5):245–256, December 2000. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Redstone:2000:AOSb

[REL00b] Joshua A. Redstone, Susan J.

Eggers, and Henry M. Levy. An analysis of operating system behavior on a simultaneous multithreaded architecture. ACM SIGPLAN Notices, 35(11):245–256, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Redstone:2000:AOSc

[REL00c] Joshua A. Redstone, Susan J.

Eggers, and Henry M. Levy. An analysis of operating system behavior on a simultaneous multithreaded architecture. Operating Systems Review, 34(5):245–256, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rajwar:2003:TET

[RG03] Ravi Rajwar and James Goodman. Transactional execution: Toward reliable, highperformance multithreading. IEEE Micro, 23(6):117–125, November/December 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        mi/2003/06/m6117abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/mi/2003/06/m6117.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mi/2003/06/
                                        m6117.pdf.

Radojkovic:2012:EIS

[RGG^{+}12] Petar Radojković, Sylvain Girbal, Arnaud Grasset, Eduardo Quiñones, Sami Yehia, and Francisco J. Cazorla. On the evaluation of the impact of shared resources in multithreaded COTS processors in time-critical environments. ACM Transactions on Architecture and Code Optimization, 8(4):34:1–34:??, January 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Rodgers:1999:TSN

[RGK99] Jeremy B. Rodgers, Rhonda Kay Gaede, and Jeffrey H. Kulick. IN-Tune: an In-Situ noninvasive performance tuning tool for multi-threaded Linux on symmetric multiprocessing Pentium workstations. Software—Practice and Experience, 29(9): 775–792, July 25, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?ID=
62501865;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=62501865&
PLACEBO=IE.pdf.

Radulovic:2017:LLI

[RGT17] Milan B. Radulović, Sylvain Girbal, and Milo V. Tomasević. Low-level implementation of the SISC protocol for thread-level speculation on a multi-core architecture. Parallel Computing, 67(??):1–19, September 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167819117300972.

Rashid:2010:AEP

[RHH10] Layali Rashid, Wessam M.

Hassanein, and Moustafa A. Hammad. Analyzing and enhancing the parallel sort operation on multithreaded architectures. The Journal of Supercomputing, 53 (2):293–312, August 2010. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=53&
issue=2&spage=293.

Richman:1991:EHC

[Ric91]  Scott Richman.  Examin-

ing the Hamilton C shell (Unix power for OS/2). Dr. Dobb’s Journal of Software Tools, 16(1):98, 100, 102, 104– 106, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Richards:1999:ALT

[Ric99] Etienne Richards. Adding level-2 thread safety to existing objects. C/C++ Users Journal, 17(2):??, February 1999. CODEN CCUJEX. ISSN 1075-2838.

Ringle:1999:SCT

[Rin99] Jonathan Ringle. Singleton creation the thread-safe way. C/C++ Users Journal, 17 (10):??, October 1999. CO-DEN CCUJEX. ISSN 1075-2838.

Rinard:2001:AMP

[Rin01]  Martin Rinard.  Analy-

sis of multithreaded programs. Lecture Notes in Computer Science, 2126:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2126/21260001.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2126/21260001.
                                        pdf.

Reddy:2011:BFH

[RKBH11] Dheeraj Reddy, David Koufaty, Paul Brett, and Scott Hahn. Bridging functional heterogeneity in multicore architectures. Operating Systems Review, 45(1):21– 33, January 2011. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Reus:1998:VCO

[RKCW98] B. Reus, A. Knapp, P. Cenciarelli, and M. Wirsing. Verifying a compiler optimization for multi-threaded Java. Lecture Notes in Computer Science, 1376:402–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reiche:2017:AVI

[RKHT17] Oliver Reiche, Christof Kobylko, Frank Hannig, and Jürgen Teich. Auto-vectorization for image processing DSLs. ACM SIGPLAN Notices, 52(4):21– 30, May 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rodrigues:2015:DSE

[RKK15] Rance Rodrigues, Israel Koren, and Sandip Kundu. Does the sharing of execution units improve performance/power of multicores? ACM Transactions on Embedded Computing Systems, 14(1):17:1–17:??, January 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Raman:2010:SPUa

[RKM^{+}10a] Arun Raman, Hanjun Kim, Thomas R. Mason, Thomas B. Jablin, and David I. August. Speculative parallelization using software multi-threaded transactions. ACM SIGARCH Computer Architecture News, 38(1):65–76, March 2010. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Raman:2010:SPUb

[RKM^{+}10b] Arun Raman, Hanjun Kim, Thomas R. Mason, Thomas B. Jablin, and David I. August. Speculative parallelization using software multi-threaded transactions. ACM SIGPLAN Notices, 45(3):65–76, March 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ribic:2014:EEW

[RL14] Haris Ribic and Yu David Liu. Energy-efficient workstealing language runtimes. ACM SIGARCH Computer Architecture News, 42(1):513– 528, March 2014. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Raghavan:2009:DLC

[RLJ^{+}09] P. Raghavan, A. Lambrechts, M. Jayapala, F. Catthoor, and D. Verkest. Distributed loop controller for multithreading in unithreaded ILP architectures. IEEE Transactions on Computers, 58(3):311–321, March 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=4624249.

Roe:1999:PMI

[RM99] Kevin Roe and Piyush Mehrotra. Parallelization of a multigrid incompressible viscous cavity flow solver using openMP. NASA contractor report NASA/CR-1999-209551, NASA Langley Research Center, Hampton, VA, USA, 1999. ???? pp. Also ICASE report 99-36.

Reinhardt:2000:TFD

[RM00] Steven K. Reinhardt and Shubhendu S. Mukherjee. Transient fault detection via simultaneous multithreading. ACM SIGARCH Computer Architecture News, 28(2): 25–36, May 2000. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

ACM:2003:ATA

[RM03] Allyn Romanow and Jeff Mogul, editors. Proceedings of the ACM SIGCOMM Workshop on Network-I/O Convergence: experience, Lessons, Implications 2003, Karlsruhe, Germany, August 25–27, 2003. ACM Press, New York, NY 10036, USA, 2003. ISBN ???? LCCN TK5105.5. ACM order number 534032.

Raad:2022:EIX

[RMV22] Azalea Raad, Luc Maranget, and Viktor Vafeiadis. Extending Intel-x86 consistency and persistency: formalising the semantics of Intel-x86 memory types and non-temporal stores. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 22:1–22:31, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3498683.

Roh:1996:GOE

[RNSB96] Lucas Roh, Walid A. Najjar, Bhanu Shankar, and A. P. Wim Böhm. Generation, optimization, and evaluation of multithreaded code. Journal of Parallel and Distributed Computing, 32(2):188–204, February 1, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0013/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0013/production/pdf.

Rogers:2014:LYL

[ROA14] Timothy G. Rogers, Mike O’Connor, and Tor M. Aamodt. Learning your limit: managing massively multithreaded caches through scheduling. Communications of the ACM, 57(12): 91–98, December 2014. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://cacm.
acm.org/magazines/2014/
12/180789/fulltext.

Robison:2003:MCN

[Rob03] Arch D. Robison. Memory consistency and .NET. Dr. Dobb’s Journal of Software Tools, 28(4):46, 48–50, April 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/documents/
s=7827/ddj0304e/.

Rodley:1994:UIC

[Rod94]  John Rodley.  OS/2 and

UnixWare interprocess communication. Dr. Dobb’s Journal of Software Tools, 19(5): 78–82, 84, 107–109, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Rodens:1995:ESC

[Rod95a]  Ira Rodens.  Examining

Symantec C++ 7.0. Dr. Dobb’s Journal of Software Tools, 20(8):86–89, 106–107, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Rodley:1995:TPU

[Rod95b] John Rodley. Thread programming in UnixWare 2.0. Dr. Dobb’s Journal of Software Tools, 20(6):56, 58–61, 102, 104, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Roh:1995:CGE

[Roh95] Lucas J. Roh. Code generations, evaluations, and optimizations in multithreaded executions. Thesis (Ph.D.), Colorado State University, Fort Collins, CO, USA, 1995. ix + 154 pp.

Rodriguez:2022:EHB

[ROPdlT22] Alfonso Rodríguez, Andrés Otero, Marco Platzner, and Eduardo de la Torre. Exploiting hardware-based dataparallel and multithreading models for smart edge computing in reconfigurable FP-GAs. IEEE Transactions on Computers, 71(11):2903– 2914, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Roth:2019:AOC

[Rót19] Ágoston Róth. Algorithm 992: An OpenGL- and C++-based function library for curve and surface modeling in a large class of extended Chebyshev spaces. ACM Transactions on Mathematical Software, 45(1): 13:1–13:32, March 2019. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3284979.

Roy:2009:LPF

[RPB^{+}09] Indrajit Roy, Donald E.

Porter, Michael D. Bond, Kathryn S. McKinley, and Emmett Witchel. Laminar: practical fine-grained decentralized information flow control. ACM SIGPLAN Notices, 44(6):63–74, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ruan:2005:EIS

[RPNT05] Yaoping Ruan, Vivek S. Pai, Erich Nahum, and John M. Tracey. Evaluating the impact of simultaneous multithreading on network servers using real hardware. ACM SIG-METRICS Performance Evaluation Review, 33(1):315–326, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ruan:2008:DCS

[RPNT08] Yaoping Ruan, Vivek S. Pai, Erich Nahum, and John M. Tracey. Do commodity SMT processors need more OS research? Operating Systems Review, 42(1):21–25, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Raghunath:1993:DIN

[RR93] M. T. Raghunath and Abhiram Ranade. Designing interconnection networks for multi-level packaging. In IEEE [IEE93], pages 772– 781. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Robbins:1996:PUP

[RR96] Kay A. Robbins and Steven Robbins. Practical UNIX programming: a guide to concurrency, communication, and multithreading. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-443706-3. xiv + 658 pp. LCCN QA76.76.O63 R615 1996.

Rugina:1999:PAM

[RR99] Radu Rugina and Martin Rinard. Pointer analysis for multithreaded programs. ACM SIGPLAN Notices, 34(5):77–90, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/pldi/301122/
                                        p77-rugina/. See PLDI’99

proceedings [ACM99a].

Robbins:2003:USP

[RR03] Kay A. Robbins and Steven Robbins. UNIX Systems programming: communication, concurrency, and threads. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2003. ISBN 0-13-042411-0. xvii + 893 pp. LCCN QA76.76.O63 R6215 2003.

Roy:2011:SRP

[RRK11] Soumyaroop Roy, Nagarajan Ranganathan, and Srinivas Katkoori. State-retentive power gating of register files in multicore processors featuring multithreaded inorder cores. IEEE Transactions on Computers, 60 (11):1547–1560, November 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=5669257.

Rivara:2012:MPL

[RRMJ12] Maria-Cecilia Rivara, Pedro Rodriguez, Rafael Montenegro, and Gaston Jorquera. Multithread parallelization of Lepp-bisection algorithms. Applied Numerical Mathematics: Transactions of IMACS, 62(4):473–488, April 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0168927411001292.

Reddy:2006:UPB

[RRP06] Vimal K. Reddy, Eric Rotenberg, and Sailashri Parthasarathy.

Understanding predictionbased partial redundant threading for low-overhead, highcoverage fault tolerance. ACM SIGPLAN Notices, 41(11):83– 94, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosu:2007:ITO

[RS07] Grigore Roşu and Koushik Sen. An instrumentation technique for online analysis of multithreaded programs. Concurrency and Computation: Practice and Experience, 19(3):311–325, March 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rounce:2008:DIS

[RS08] Peter A. Rounce and Alberto F. De Souza. Dynamic instruction scheduling in a trace-based multithreaded architecture. International Journal of Parallel Programming, 36(2):184–205, April 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=36&
                                        issue=2&spage=184.

Riccobene:2009:SCB

[RSB^{+}09] Elvinia Riccobene, Patrizia Scandurra, Sara Bocchio, Alberto Rosti, Luigi Lavazza, and Luigi Mantellini. SystemC/C-based model-driven design for embedded systems. ACM Transactions on Embedded Computing Systems, 8(4): 30:1–30:??, July 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Roh:2001:RMD

[RSBN01] Lucas Roh, Bhanu Shankar, Wim Böhm, and Walid Najjar. Resource management in dataflow-based multithreaded execution. Journal of Parallel and Distributed Computing, 61(5):581–608, May 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2001.1708;   http://www.
idealibrary.com/links/doi/
10.1006/jpdc.2001.1708/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2001.1708/ref.

Rangan:2008:PSD

[RVOA08] Ram Rangan, Neil Vachharajani, Guilherme Ottoni, and David I. August. Performance scalability of decoupled software pipelining. ACM Transactions on Architecture and Code Optimization, 5(2):8:1– 8:??, August 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Roth:2004:MTC

[RVR04] Marcus Roth, Gerrit Voss, and Dirk Reiners. Multithreading and clustering for scene graph systems. Computers and Graphics, 28(1): 63–66, February 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Raychev:2013:ERD

[RVS13] Veselin Raychev, Martin Vechev, and Manu Sridharan. Effective race detection for event-driven programs. ACM SIGPLAN Notices, 48(10):151–166, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Ravoor:1997:MTP

[RW97] Suresh B. Ravoor and Johnny S. K. Wong. Multithreaded transaction processing in distributed systems. The Journal of Systems and Software, 38(2):107–117, August 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Romanous:2021:ELL

[RWT21] Bashar Romanous, Skyler Windh, and Vassilis Tsotras. Efficient local locking for massively multithreaded in-memory hash-based operators. VLDB Journal: Very Large Data Bases, 30(3): 333–359, May 2021. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s00778-020-00642-5.

Robatmili:2004:TSI

[RYSN04] B. Robatmili, N. Yazdani, S. Sardashti, and M. Nourani. Thread-sensitive instruction issue for SMT processors. IEEE Computer Architecture Letters, 3(1):5, January 2004. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Shaw:1998:CIP

[SAC^{+}98] Andrew Shaw, Arvind, Kyoo-Chan Cho, Christopher Hill, R. Paul Johnson, and John Marshall. A comparison of implicitly parallel multithreaded and data-parallel implementations of an ocean model. Journal of Parallel and Distributed Computing, 48(1):1–51, January 10, 1998. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1997.1390/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1997.1390/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1997.1390/production/ref.

Samorodin:1999:SFS

[Sam99] Steven Howard Samorodin.

Supporting flexible safety and sharing in multi-threaded environments. Thesis (M.S.), Computer Science Department, University of California, Davis, Davis, CA, USA, 1999. 39 pp.

Sanden:2004:CJT

[San04] B. Sanden. Coping with Java threads: Java works for many kinds of concurrent software, but it was not designed for safety-critical realtime applications and does not protect the programmer from the pitfalls associated with multithreading. Computer, 37(4):20–27, 2004. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sato:2002:SJL

[Sat02] Y. Sato. A study of Java language for effective thread migration. Record of Electrical and Communication Engineering Conversazione Tohoku University, 71(1):597– 598, 2002. CODEN ???? ISSN 0385-7719.

Smith:1980:ASD

[SB80] Connie Smith and J. C. Browne. Aspects of software design analysis: Concurrency and blocking. ACM SIG-METRICS Performance Evaluation Review, 9(2):245–253, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sah:1996:PIS

[SBB96] A. Sah, K. Brown, and E. Brewer. Programming the Internet from the serverside with Tcl and Audience1. In Association [Ass96], pages 235–??, 183– 188. ISBN 1-880446-78-2. LCCN QA76.73.T44 T44 1996.

Saavedra-Barrera:1991:ASM

[SBC91] Rafael H. Saavedra-Barrera and David E. Culler. An analytical solution for a Markov chain modeling multithreaded execution. Report UCB/CSD 91/623, University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, April 1991. 24 pp.

Saavedra-Barrera:1990:AMA

[SBCV90] Rafael H. Saavedra-Barrera, David E. Culler, and Thorsten Von Eiken. Analysis of multithreaded architectures for parallel computing. Report UCB/CSD 90/569, University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, April 1990. 10 pp. To appear in the 2nd Annual ACM Symposium on Parallel Algorithms and Architectures, Crete, Greece, July 1990.

Spoto:2019:SII

[SBE^{+}19] Fausto Spoto, Elisa Burato, Michael D. Ernst, Pietro Ferrara, Alberto Lovato, Damiano Macedonio, and Ciprian Spiridon. Static identification of injection attacks in Java. ACM Transactions on Programming Languages and Systems, 41(3): 18:1–18:??, July 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3332371.

Storino:1999:MTB

[SBKK99] Salvatore Storino, John M. Borkenhagen, Ronald N. Kalla, and Steven R. Kunkel. A multi-threaded 64-bit PowerPC commercial RISC processor design. In IEEE [IEE99], page ?? ISBN ???? LCCN ???? URL http://

www.hotchips.org/hotc11_
index.html.

Savage:1997:EDD

[SBN^{+}97] Stefan Savage, Michael Burrows, Greg Nelson, Patrick Sobalvarro, and Thomas Anderson. Eraser: a dynamic data race detector for multithreaded programs. ACM Transactions on Computer Systems, 15(4):391– 411, November 1997. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tocs/1997-15-4/

p391-savage/. Co-published in Operating Systems Review, 31(5).

Sanderson:2017:PGP

[SC17] Conrad Sanderson and Ryan Curtin. gmm diag and gmm full: C++ classes for multi-threaded Gaussian mixture models and expectationmaximisation. Journal of Open Source Software, 2(18): 365:1–365:2, October 2017. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00365.

Saillard:2015:SDV

[SCB15] Emmanuelle Saillard, Patrick Carribault, and Denis Barthou. Static/dynamic validation of MPI collective communications in multi-threaded context. ACM SIGPLAN Notices, 50(8):279–280, August 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Saez:2013:DFP

[SCCP13] Juan Carlos Sáez, Fernando Castro, Daniel Chaver, and Manuel Prieto. Delivering fairness and priority enforcement on asymmetric multicore systems via OS scheduling. ACM SIGMETRICS Performance Evaluation Review, 41 (1):343–344, June 2013. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schweitzer:2015:PEM

[SCD^{+}15] P. Schweitzer, S. Cipière, A. Dufaure, H. Payno, Y. Perrot, D. R. C. Hill, and L. Maigne. Performance evaluation of multithreaded Geant4 simulations using an Intel Xeon Phi cluster. Scientific Programming, 2015 (??):980752:1–980752:10, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/980752/.

Schauser:1995:SCP

[SCG95] Klaus E. Schauser, David E. Culler, and Seth C. Goldstein. Separation constraint partitioning: a new algorithm for partitioning non-strict programs into sequential threads. In ACM [ACM95b], pages 259–271. ISBN 0-89791-692-

1. LCCN QA 76.7 A11 1995.

URL http://www.acm.org:
80/pubs/citations/proceedings/
plan/199448/p259-schauser/

. ACM order number: 549950.

Schonberg:1989:FDA

[Sch89] Edith Schonberg. On-the-fly detection of access anomalies. ACM SIGPLAN Notices, 24(7):285–297, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/pldi/73141/
p285-schonberg/.

Schmitt:1990:CEM

[Sch90] David A. Schmitt. C extensions for multi-threading. C Users Journal, 8(8):33–??, August 1990. ISSN 0898-9788.

Schauser:1991:CDT

[Sch91] Klaus Erik Schauser. Compiling dataflow into threads: efficient compiler-controlled multithreading for lenient parallel languages. Thesis (M.S.), University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, July 2, 1991. 71 pp. Also available as Report UCB/CSD 91/644.

Schmidt:1998:EAM

[Sch98] Douglas C. Schmidt. Evaluating architectures for multithreaded object request brokers. Communications of the ACM, 41(10):54–60, October 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1998-41-10/
p54-schmidt/.

Schildt:2014:JCR

[Sch14] Herbert Schildt, editor. Java:

The Complete Reference. Mc-Graw-Hill, New York, NY, USA, ninth edition, 2014. ISBN 0-07-180855-8 (paperback), 0-07-180925-2, 0-07-180856-6. xxxiv + 1274 pp. LCCN QA76.73.J38 S332 2014eb.

Schafer:2017:PHL

[Sch17] Benjamin Carrion Schafer. Parallel high-level synthesis design space exploration for behavioral IPs of exact latencies. ACM Transactions on Design Automation of Electronic Systems., 22(4): 65:1–65:??, July 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sendag:2005:IIS

[SCL05] Resit Sendag, Ying Chen, and David J. Lilja. The impact of incorrectly speculated memory operations in a multithreaded architecture. IEEE Transactions on Parallel and Distributed Systems, 16(3): 271–285, March 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Steinke:2005:NPF

[SCM05] Robert Steinke, Micah Clark, and Elihu McMahon. A new pattern for flexible worker threads with in-place consumption message queues. Operating Systems Review, 39 (2):71–73, April 2005. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schauser:1991:CCM

[SCv91a] Klaus Erik Schauser, David E. Culler, and Thorsten von Eicken. Compiler-controlled multithreading for lenient parallel languages. Lecture Notes in Computer Science, 523:50– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schauser:1991:CML

[SCv91b] Klaus Erik Schauser, David E.

Culler, and Thorsten von Eicken. Compiler-controlled multithreading for lenient parallel languages. Report UCB/CSD 91/640, University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, July 30, 1991. 21 pp. A version of this report is to appear in the Proceedings of FPCA ’91 Conference on Functional Programming Languages and Computer Architecture, Aug. 1991, Springer-Verlag.

Steffan:2000:SAT

[SCZM00] J. Greggory Steffan, Christopher B. Colohan, Antonia Zhai, and Todd C. Mowry. A scalable approach to threadlevel speculation. ACM SIGARCH Computer Architecture News, 28(2):1–12, May 2000. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Spertus:1995:ELB

[SD95] Ellen Spertus and William J.

Dally. Evaluating the locality benefits of active messages. ACM SIGPLAN Notices, 30(8):189–198, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

So:2013:STI

[SD13] Won So and Alexander G. Dean. Software thread integration for instruction-level parallelism. ACM Transactions on Embedded Computing Systems, 13(1):8:1–8:??, August 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sartor:2012:EMT

[SE12] Jennfer B. Sartor and Lieven Eeckhout. Exploring multithreaded Java application performance on multicore hardware. ACM SIGPLAN Notices, 47(10):281–296, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Seiden:1998:ROM

[Sei98] S. S. Seiden. Randomized online multi-threaded paging. Lecture Notes in Computer Science, 1432:264–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seiden:1999:ROM

[Sei99] Steven S. Seiden. Randomized online multi-threaded paging. Nordic Journal of Computing, 6(2):148–??, Summer 1999. CODEN NJCOFR.

ISSN 1236-6064. URL http:/

/www.cs.helsinki.fi/njc/
References/seiden1999:148.
html.

Sen:2008:RDR

[Sen08] Koushik Sen. Race directed random testing of concurrent programs. ACM SIG-PLAN Notices, 43(6):11–21, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sengupta:2019:JHP

[Sen19] Avik Sengupta. Julia high performance optimizations, distributed computing, multithreading, and GPU programming with Julia 1.0 and beyond. Packt Publishing, Birmingham, UK, second edition, 2019. ISBN 1-78829-230-8, 1-78829-811-X. 218 pp. LCCN

???? URL http://portal.
igpublish.com/iglibrary/
search/PACKT0005341.html.

Severance:1996:MOB

[SEP96] Charles Severance, Richard Enbody, and Paul Petersen. Managing the overall balance of operating system threads on a multiprocessor using automatic self-allocating threads (ASAT). Journal of Parallel and Distributed Computing, 37(1):106–112, August 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.

                                        1996.0111/production;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0111/production/pdf.

Sundaresan:1996:COO

[SG96] Neelakantan Sundaresan and Dennis Gannon. Coir: An object-oriented system for control and dynamic data parallelism. Journal of Parallel and Distributed Computing, 37(1):98–105, August 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0110/production;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0110/production/pdf.

Sahin:2018:CSC

[SG18] Semih Sahin and Bugra Gedik. C-Stream: a coroutine-based elastic stream processing engine. ACM Transactions on Parallel Computing (TOPC), 4(3): 15:1–15:??, April 2018. CO-DEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Sung:2014:PTR

[SGLGL^{+}14] I-Jui Sung, Juan Gómez-Luna, José María González-Linares, Nicolás Guil, and Wen-Mei W. Hwu. Inplace transposition of rectangular matrices on accelerators. ACM SIGPLAN Notices, 49(8):207–218, August 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sodan:1997:ENN

[SGM^{+}97] Angela Sodan, Guang R. Gao, Olivier Maquelin, Jens-Uwe Schultz, and Xin-Min Tian. Experiences with non-numeric applications on multithreaded architectures. ACM SIG-PLAN Notices, 32(7):124–135, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sridharan:2014:AEP

[SGS14] Srinath Sridharan, Gagan Gupta, and Gurindar S. Sohi. Adaptive, efficient, parallel execution of parallel programs. ACM SIGPLAN Notices, 49(6):169–180, June 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shahnaz:1995:DMD

[Sha95a] Munira Shahnaz. Design of a multithreaded data cache for a hyperscalar processor. Thesis (M.S.), Department of Electrical Engineering, Texas A&M University, College Station, TX, USA, 1995. xi + 80 pp.

Shankar:1995:STI

[Sha95b] Bhanu Shankar. The spectrum of thread implementations on hybrid multithreaded architectures. Thesis (Ph.D.), Colorado State University, Fort Collins, CO, USA, 1995. xi + 176 pp.

Shaw:1998:CPM

[Sha98] Andrew Shaw. Compiling for parallel multithreaded computation on symmetric multiprocessors. Thesis (Ph.D.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1998. 149 pp.

Shene:1998:MPI

[She98] Chin-Kuang Shene. Multithreaded programming in an introduction to operating systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):242–246, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shene:2002:TST

[She02] Ching-Kuang Shene. Thread-Mentor: a system for teaching multithreaded programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (3):229, September 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shinjo:2000:DCEb

[Shi00]  Yasushi Shinjo.  Develop-

ing correct and efficient multithreaded programs with thread-specific data and a partial evaluator. Operating Systems Review, 34(2):40, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shi:2015:CLM

[SHK15] Qingchuan Shi, Henry Hoffmann, and Omer Khan. A cross-layer multicore architecture to tradeoff program accuracy and resilience overheads. IEEE Computer Architecture Letters, 14(2):85– 89, July/December 2015. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Shoffner:1997:JSSa

[Sho97a] Michael Shoffner. Java step by step: Write your own threaded discussion forum. JavaWorld: IDG’s magazine for the Java community, 2 (2):??, February 1997. CO-DEN ???? ISSN 1091-

8906.  URL http://www.
javaworld.com/javaworld/
jw-02-1997/jw-02-step.htm.

Shoffner:1997:JSSb

[Sho97b]  Michael Shoffner.  Java

step by step: Write your own threaded discussion forum: The communications and server components. Java-World: IDG’s magazine for the Java community, 2(3): ??, March 1997. CO-DEN ???? ISSN 1091-

                                        8906.  URL http://www.
                                        javaworld.com/javaworld/
                                        jw-03-1997/jw-03-step.htm.

Shomron:2019:SSS

[SHW19] G. Shomron, T. Horowitz, and U. Weiser. SMT-SA: Simultaneous multithreading in systolic arrays. IEEE Computer Architecture Letters, 18 (2):99–102, July 2019. ISSN 1556-6056 (print), 1556-6064 (electronic).

Sime:1997:GPM

[Sim97] J. Sime. Guarded pointers: moving smart pointers into multithreaded systems. C++ Report, 9(4): 32–41, April 1997. CODEN CRPTE7. ISSN 1040-6042.

Sinharoy:1997:OTC

[Sin97]  Balaram Sinharoy.  Opti-

mized thread creation for processor multithreading. The Computer Journal, 40(6): 388–??, ???? 1997. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /www.oup.co.uk/computer_
                                        journal/Volume_40/Issue_
                                        06/Vol40_06.body.html#AbstractSinharoy;
                                         http://www3.oup.co.uk/
                                        computer_journal/Volume_
                                        40/Issue_06/Vol40_06.body.
                                        html#AbstractSinharoy.

Sinharoy:1999:COI

[Sin99] Balaram Sinharoy. Compiler optimization to improve data locality for processor multithreading. Scientific Programming, 7(1):21–37, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.asp%
3Fwasp=64cr5a4mg33tuhcbdr02%
26referrer=parent%26backto=
issue%2C2%2C7%3Bjournal%
2C8%2C9%3Blinkingpublicationresults%
2C1%2C1.

Steensgaard:1995:ONC

[SJ95] B. Steensgaard and E. Jul.

Object and native code thread mobility among heterogeneous computers (includes sources). Operating Systems Review, 29 (5):68–77, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sharafeddine:2012:DOE

[SJA12] Mageda Sharafeddine, Komal Jothi, and Haitham Akkary. Disjoint out-of-order execution processor. ACM Transactions on Architecture and Code Optimization, 9(3):19:1– 19:??, September 2012. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Singh:1992:DRS

[SJB92a] Gurjot Singh, Moses Joseph, and Dave Barnett. Debugging real-time systems. Dr. Dobb’s Journal of Software Tools, 17 (9):70, 72, 74, 76–77, 116– 117, September 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Singh:1992:DRT

[SJB92b] Gurjot Singh, Moses Joseph, and Dave Barnett. Debugging real-time systems. Dr. Dobb’s Journal of Software Tools, 17 (9):70, 72, 74, 76–77, 116– 117, September 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Stewart:1997:MDH

[SK97] David B. Stewart and Pradeep K. Khosla. Mechanisms for detecting and handling timing errors. Communications of the ACM, 40(1):87–93, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/cacm/
                                        1997-40-1/p87-stewart/.

Shirole:2012:TCU

[SK12] Mahesh Shirole and Rajeev Kumar. Testing for concurrency in UML diagrams. ACM SIGSOFT Software Engineering Notes, 37(5):1–8, September 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sung:2001:MDA

[SKA01] Michael Sung, Ronny Krashinsky, and Krste Asanović.

Multithreading decoupled architectures for complexityeffective general purpose computing. ACM SIGARCH Computer Architecture News, 29(5):56–61, December 2001. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Smaragdakis:2007:TIC

[SKBY07] Yannis Smaragdakis, Anthony Kay, Reimer Behrends, and Michal Young. Transactions with isolation and cooperation. ACM SIGPLAN Notices, 42(10):191–210, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schonherr:2011:MTI

[SKG^{+}11] M. Schönherr, K. Kucher, M. Geier, M. Stiebler, S. Freudiger, and M. Krafczyk. Multithread implementations of the lattice Boltzmann method on non-uniform grids for CPUs and GPUs. Computers and Mathematics with Applications, 61(12):3730–3743, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122111002999.

Sohn:2001:CTC

[SKK^{+}01] Andrew Sohn, Yuetsu Kodama, Jui-Yuan Ku, Mitsuhisa Sato, and Yoshinori Yamaguchi. Chapter 15. Tolerating communication latency through dynamic thread invocation in a multithreaded architecture. Lecture Notes in Computer Science, 1808:525– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1808/18080525.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1808/18080525.
pdf.

Son:2009:CDD

[SKKC09] Seung Woo Son, Mahmut Kandemir, Mustafa Karakoy, and Dhruva Chakrabarti. A compiler-directed data prefetching scheme for chip multiprocessors. ACM SIG-PLAN Notices, 44(4):209–218, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sung:2002:CPE

[SKP^{+}02] Minyoung Sung, Soyoung Kim, Sangsoo Park, Naehyuck Chang, and Heonshik Shin. Comparative performance evaluation of Java threads for embedded applications: Linux Thread vs. Green Thread. Information Processing Letters, 84(4):221– 225, November 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sato:1992:TBP

[SKS^{+}92] Mitsuhisa Sato, Yuetsu Kodama, Shuichi Sakai, Yoshinori Yamaguchi, and Yasuhito Koumura. Thread-based programming for the EM-4 hybrid dataflow machine. ACM SIGARCH Computer Architecture News, 20(2):146–155, May 1992. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Sonenberg:2021:PAW

[SKV21] Nikki Sonenberg, Grzegorz Kielanski, and Benny Van Houdt. Performance analysis of work stealing in largescale multithreaded computing. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 6(2):6:1– 6:28, June 2021. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

https://dl.acm.org/doi/
10.1145/3470887.

Steele:2014:FSP

[SLF14] Guy L. Steele, Jr., Doug Lea, and Christine H. Flood. Fast splittable pseudorandom number generators. ACM SIGPLAN Notices, 49(10): 453–472, October 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shin:2004:NAD

[SLG04] Chulho Shin, Seong-Won Lee, and Jean-Luc Gaudiot. The need for adaptive dynamic thread scheduling in simultaneous multithreading. Parallel Processing Letters, 14(3/4): 327–??, September/December 2004. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Shin:2006:ADT

[SLG06] Chulho Shin, Seong-Won Lee, and Jean-Luc Gaudiot. Adaptive dynamic thread scheduling for simultaneous multithreaded architectures with a detector thread. Journal of Parallel and Distributed Computing, 66(10):1304–1321, October 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Scherer:1999:TAP

[SLGZ99] Alex Scherer, Honghui Lu, Thomas Gross, and Willy Zwaenepoel. Transparent adaptive parallelism on NOWs using OpenMP. ACM SIGPLAN Notices, 34(8):96– 106, August 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        proceedings/ppopp/301104/
                                        p96-scherer/.

Sangaiah:2018:SSA

[SLJ^{+}18] Karthik Sangaiah, Michael Lui, Radhika Jagtap, Stephan Diestelhorst, Siddharth Nilakantan, Ankit More, Baris Taskin, and Mark Hempstead. SynchroTrace: Synchronizationaware architecture-agnostic traces for lightweight multicore simulation of CMP and HPC workloads. ACM Transactions on Architecture and Code Optimization, 15(1):2:1– 2:??, April 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Su:2019:SSC

[SLJ^{+}19] Xing Su, Xiangke Liao, Hao Jiang, Canqun Yang, and Jingling Xue. SCP: Shared cache partitioning for highperformance GEMM. ACM Transactions on Architecture and Code Optimization, 15 (4):43:1–43:??, January 2019. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3274654.

Sharkey:2008:RRP

[SLP08] Joseph J. Sharkey, Jason Loew, and Dmitry V. Ponomarev. Reducing register pressure in SMT processors through L2-miss-driven early register release. ACM Transactions on Architecture and Code Optimization, 5(3):13:1– 13:??, November 2008. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Sidiroglou:2009:AAS

[SLP^{+}09] Stelios Sidiroglou, Oren Laadan, Carlos Perez, Nicolas Viennot, Jason Nieh, and Angelos D. Keromytis. ASSURE: automatic software self-healing using rescue points. ACM SIG-PLAN Notices, 44(3):37–48, March 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Solihin:2002:UUL

[SLT02] Yan Solihin, Jaejin Lee, and Josep Torrellas. Using a userlevel memory thread for correlation prefetching. ACM SIGARCH Computer Architecture News, 30(2):171–182, May 2002. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Solihin:2003:CPU

[SLT03] Yan Solihin, Jaejin Lee, and Josep Torrellas. Correlation prefetching with a userlevel memory thread. IEEE Transactions on Parallel and Distributed Systems, 14(6): 563–580, June 2003. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/trans/
                                        td/2003/06/l0563abs.htm;
                                         http://csdl.computer.org/
                                        dl/trans/td/2003/06/l0563.
                                        pdf.

Shea:2019:HSD

[SM19] Colin Shea and Tinoosh Mohsenin. Heterogeneous scheduling of deep neural networks for low-power realtime designs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(4):36:1–36:??, December 2019. CODEN ???? ISSN 1550-4832. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3358699.

Sodan:2010:PMM

[SMD^{+}10] Angela C. Sodan, Jacob Machina, Arash Deshmeh, Kevin Macnaughton, and Bryan Esbaugh. Parallelism via multithreaded and multicore CPUs. Computer, 43 (3):24–32, March 2010. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:1992:MTX

[Smi92] John Allen Smith. The multithreaded X server. The X Resource, 1(1):73–89, January 1992. CODEN XRESEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Smith:2001:CMM

[Smi01] Burton Smith. Cray MTA:

Multithreading for latency response. Computer, 34 (4):69, April 2001. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/

books/co2001/pdf/r4059.
pdf;   http://www.computer.
org/computer/co2001/r4059abs.
htm.

Smith:2006:ITP

[Smi06] Geoffrey Smith. Improved typings for probabilistic noninterference in a multithreaded language. Journal of Computer Security, 14(6): 591–623, ???? 2006. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Sanchez:2010:ACI

[SMK10] Daniel Sanchez, George Michelogiannakis, and Christos Kozyrakis. An analysis of onchip interconnection networks for large-scale chip multiprocessors. ACM Transactions on Architecture and Code Optimization, 7(1):4:1–4:??, April 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Storey:2019:SDP

[SMP19] Kyle Storey, Eric Mercer, and Pavel Parizek. A sound dynamic partial order reduction engine for Java Pathfinder. ACM SIGSOFT Software Engineering Notes, 44(4):15, December 2019. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3364452.3364457.

Suleman:2009:ACS

[SMQP09] M. Aater Suleman, Onur Mutlu, Moinuddin K. Qureshi, and Yale N. Patt. Accelerating critical section execution with asymmetric multicore architectures. ACM SIG-PLAN Notices, 44(3):253–264, March 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Swanson:2003:ESI

[SMS^{+}03] Steven Swanson, Luke K. Mc-Dowell, Michael M. Swift, Susan J. Eggers, and Henry M. Levy. An evaluation of speculative instruction execution on simultaneous multithreaded processors. ACM Transactions on Computer Systems, 21(3):314–340, August 2003. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Scionti:2018:EMM

[SMZ18] Alberto Scionti, Somnath Mazumdar, and Stephane Zuckerman. Enabling massive multi-threading with fast hashing. IEEE Computer Architecture Letters, 17(1):1–4, January/June 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Singh:2012:EES

[SNM^{+}12] Abhayendra Singh, Satish Narayanasamy, Daniel Marino, Todd Millstein, and Madanlal Musuvathi. End-to-end sequential consistency. ACM SIGARCH Computer Architecture News, 40(3): 524–535, June 2012. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ISCA ’12 conference proceedings.

Sodan:2002:AMA

[Sod02] Angela C. Sodan. Applications on a multithreaded architecture: a case study with EARTH-MANNA. Parallel Computing, 28(1):3– 33, January 2002. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
35/21/60/27/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/35/21/60/27/
28/00001684.pdf.

Samorodin:2000:SFS

[SP00a] Steven H. Samorodin and Raju Pandey. Supporting flexible safety and sharing in multi-threaded environments. Lecture Notes in Computer Science, 1800:1184–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1800/18001184.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1800/18001184.
pdf.

Shinjo:2000:DCEa

[SP00b] Yasushi Shinjo and Calton Pu.

Developing correct and efficient multithreaded programs with thread-specific data and a partial evaluator. Operating Systems Review, 34 (2):33, April 2000. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shinjo:2005:AEP

[SP05] Y. Shinjo and C. Pu. Achieving efficiency and portability in systems software: a case study on POSIXcompliant multithreaded programs. IEEE Transactions on Software Engineering, 31(9):785–800, September 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1514446.

Sharkey:2007:EOA

[SP07] Joseph J. Sharkey and Dmitry V.

Ponomarev. Exploiting operand availability for efficient simultaneous multithreading. IEEE Transactions on Computers, 56 (2):208–223, February 2007. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=4042681.

Saarikivi:2017:MTS

[SPDLK^{+}17] Olli Saarikivi, Hernán Ponce-De-León, Kari Kähkönen, Keijo Heljanko, and Javier Esparza. Minimizing test suites with unfoldings of multithreaded programs. ACM Transactions on Embedded Computing Systems, 16(2): 45:1–45:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Spero:1994:MMD

[Spe94] Simon E. Spero. MDMA — multithreaded daemon for multimedia access. In Anonymous [Ano94d], page ?? CO-DEN ONCDEW. ISSN 0309-

                                        314X. URL http://www.
                                        ncsa.uiuc.edu/SDG/IT94/
                                        Proceedings/WWW2_Proceedings.
                                        html.

Skjellum:1996:TTM

[SPH96] A. Skjellum, B. Protopopov, and S. Hebert. A thread taxonomy for MPI. In IEEE [IEE96], pages 50–57. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Saxena:1993:PMS

[SPY^{+}93] Sunil Saxena, J. Kent Peacock, Fred Yang, Vijaya Verma, and Mohan Krishnan. Pitfalls in multithreading SVR4 STREAMS and other weightless processes. In USENIX [USE93b], pages 85–96. ISBN 1-880446-48-

0. LCCN QA 76.76 O63

U84 1993. URL http://www.
usenix.org/publications/
library/proceedings/sd93/

.

Suleman:2008:FDTa

[SQP08a] M. Aater Suleman, Moinuddin K. Qureshi, and Yale N. Patt. Feedback-driven threading: power-efficient and highperformance execution of multi-threaded workloads on CMPs. ACM SIGARCH Computer Architecture News, 36(1):277–286, March 2008. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Suleman:2008:FDTb

[SQP08b] M. Aater Suleman, Moinuddin K. Qureshi, and Yale N. Patt. Feedback-driven threading: power-efficient and highperformance execution of multi-threaded workloads on CMPs. Operating Systems Review, 42(2):277–286, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Suleman:2008:FDTc

[SQP08c] M. Aater Suleman, Moinuddin K. Qureshi, and Yale N. Patt. Feedback-driven threading: power-efficient and highperformance execution of multi-threaded workloads on CMPs. ACM SIGPLAN Notices, 43(3):277–286, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Squillante:1994:AMP

[Squ94] Mark S. Squillante. Analytic modeling of processor utilization in multithreaded processor architectures. Research report RC 19543 (84999), IBM T. J. Watson Research Center, Yorktown Heights, NY, USA, April 1994. 9 pp.

Salcianu:2001:PEA

[SR01a] Alexandru Salcianu and Martin Rinard. Pointer and escape analysis for multithreaded programs. ACM SIGPLAN Notices, 36(7):12– 23, July 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/proceedings/ppopp/
                                        379539/p12-salcianu/p12-
                                        salcianu.pdf;   http://
                                        www.acm.org/pubs/citations/
                                        proceedings/ppopp/379539/
                                        p12-salcianu/.

Sohi:2001:SMP

[SR01b] Gurindar S. Sohi and Amir Roth. Speculative multithreaded processors. Computer, 34(4):66–73, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //dlib.computer.org/co/
                                        books/co2001/pdf/r4066.
                                        pdf;   http://www.computer.

org/computer/co2001/r4066abs.

htm. See errata [Ano01].

Samak:2014:MTS

[SR14] Malavika Samak and Murali Krishna Ramanathan. Multithreaded test synthesis for deadlock detection. ACM SIGPLAN Notices, 49 (10):473–489, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sen:2006:OEP

[SRA06] Koushik Sen, Grigore Rosu, and Gul Agha. Online efficient predictive safety analysis of multithreaded programs. International Journal on Software Tools for Technology Transfer (STTT), 8(3):248– 260, June 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=1433-2779&volume=8&
issue=3&spage=248.

Srinivasan:1993:SDS

[Sri93] Sumathi Srinivasan. System design and simulation for the Demus-2 multithreaded processor. Thesis (M. Eng.), Department of Electrical and Computer Engineering, Mc-Master University, Hamilton, ON, Canada, 1993. x + 109 pp.

Srinivasan:1995:MMX

[Sri95] Murali V. Srinivasan. A methodology for multithreaded X client development. The X Resource, 13(1):181, January 1995. CODEN XRE-SEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Samak:2015:SRT

[SRJ15] Malavika Samak, Murali Krishna Ramanathan, and Suresh Jagannathan. Synthesizing racy tests. ACM SIG-PLAN Notices, 50(6):175–185, June 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Saghi:1998:MSH

[SRS98] Gene Saghi, Kirk Reinholtz, and Paul A. Savory. A multithreaded scheduler for a high-speed spacecraft simulator. Software—Practice and Experience, 28(6): 641–656, May 1998. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        1802;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=1802&PLACEBO=IE.pdf.

Silc:1998:APC

[SRU98] J. Silc, B. Robic, and T. Ungerer. Asynchrony in parallel computing: From dataflow to multithreading. Parallel and Distributed Computing Practices, 1(1):??, ???? 1998. CODEN ???? ISSN 1097-

2803. URL http://www.cs.
okstate.edu/~pdcp/vols/
vol01/vol01no1abs.html#
silc.

Speer:1991:DTP

[SS91] Thomas G. Speer and Mark W.

Storm. Digital’s transaction processing monitors. Digital Technical Journal, 3(1): 18–32, Winter 1991. CO-DEN DTJOEL. ISSN 0898-

901X.  URL ftp://ftp.
digital.com/pub/Digital/
info/DTJ/v3n1/Digitals_
Transaction_Processi_01oct1991DTJ102P8.
ps;   http://www.digital.
com:80/info/DTJ102/DTJ102SC.
TXT.

Small:1995:SAB

[SS95] Christopher Small and Margo Seltzer. Scheduler activations on BSD: sharing thread management between kernel and application. Technical Report 31-95, Center for Research in Computing Technology, Harvard University, Cambridge, MA, USA, 1995. 12 pp.

Szymanski:1996:LCR

[SS96] Boleslaw K. Szymanski and Balaram Sinharoy, editors. Languages, Compilers and Run-Time Systems for Scalable Computers, Troy, NY, USA, May 22–24, 1995. Kluwer Academic Publishers, Dordrecht, The Netherlands; Boston, MA, USA, 1996. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Sutherland:2010:CTC

[SS10] Dean F. Sutherland and William L. Scherlis. Composable thread coloring. ACM SIGPLAN Notices, 45(5):233– 244, May 2010. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schwab:2022:SSV

[SSB^{+}22] Michail Schwab, David Saffo, Nicholas Bond, Shash Sinha, Cody Dunne, Jeff Huang, James Tompkin, and Michelle A. Borkin. Scalable scalable vector graphics: Automatic translation of interactive SVGs to a multithread VDOM for fast rendering. IEEE Transactions on Visualization and Computer Graphics, 28(9):3219–3234, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Schmitt:2018:RHG

[SSK^{+}18] Christian Schmitt, Moritz Schmid, Sebastian Kuckuk, Harald Köstler, Jürgen Teich, and Frank Hannig. Reconfigurable hardware generation of multigrid solvers with conjugate gradient coarse-grid solution. Parallel Processing Letters, 28(04):??, December 2018. ISSN 0129-6264 (print), 1793-642X (elec-

                                        tronic). URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0129626418500160.

Shi:2007:CCP

[SSkP^{+}07] Xudong Shi, Feiqi Su, Jih kwon Peir, Ye Xia, and Zhen Yang. CMP cache performance projection: accessibility vs. capacity. ACM SIGARCH Computer Architecture News, 35(1):13– 20, March 2007. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Soundararajan:2010:CSE

[SSN10] Niranjan Soundararajan, Anand Sivasubramaniam, and Vijay Narayanan. Characterizing the soft error vulnerability of multicores running multithreaded applications. ACM SIGMETRICS Performance Evaluation Review, 38(1):379– 380, June 2010. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Saito:1999:MRS

[SSP99] H. Saito, N. Stavrakos, and C. Polychronopoulos. Multithreading runtime support for loop and functional parallelism. Lecture Notes in Computer Science, 1615:133– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sohn:1997:DWD

[SSYG97] Andrew Sohn, Mitsuhisa Sato, Namhoon Yoo, and Jean-Luc Gaudiot. Data and workload distribution in a multithreaded architecture. Journal of Parallel and Distributed Computing, 40(2):256–264, February 1, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1262/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1262/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1262/production/ref.

Skillicorn:1998:MLP

[ST98] David B. Skillicorn and Domenico Talia. Models and languages for parallel computation. ACM Computing Surveys, 30(2): 123–169, June 1998. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/surveys/1998-30-
2/p123-skillicorn/.

Snavely:2000:SJSa

[ST00a] Allan Snavely and Dean M. Tullsen. Symbiotic job scheduling for a simultaneous multithreaded processor. ACM SIGARCH Computer Architecture News, 28(5):234– 244, December 2000. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Snavely:2000:SJSc

[ST00b] Allan Snavely and Dean M.

Tullsen. Symbiotic jobscheduling for a simultaneous multithreaded processor. Operating Systems Review, 34 (5):234–244, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Snavely:2000:SJSb

[ST00c] Allan Snavely and Dean M.

Tullsen. Symbiotic jobscheduling for a simultaneous multithreading processor. ACM SIGPLAN Notices, 35(11):234–244, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sundell:2005:FLF

[ST05] Håkan Sundell and Philippas Tsigas. Fast and lock-free concurrent priority queues for multi-thread systems. Journal of Parallel and Distributed Computing, 65(5):609–627, May 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Stapleton:1990:DSS

[Sta90] Joseph Francis Stapleton. Dynamic server selection in a multithreaded network computing environment. Thesis (M.S.), Iowa State University, Ames, IA, USA, 1990. 66 pp.

Stark:2005:FSV

[Stä05] Robert F. Stärk. Formal specification and verification of the C# thread model. Theoretical Computer Science, 343(3): 482–508, October 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Steensgaard:2001:TSH

[Ste01] Bjarne Steensgaard. Threadspecific heaps for multithreaded programs. ACM SIGPLAN Notices, 36(1):18– 24, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stoller:2002:MCM

[Sto02] Scott D. Stoller. Modelchecking multi-threaded distributed Java programs. International Journal on Software Tools for Technology Transfer (STTT), 4(1):71–91, October 2002. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Samak:2016:DSF

[STR16] Malavika Samak, Omer Tripp, and Murali Krishna Ramanathan. Directed synthesis of failing concurrent executions. ACM SIGPLAN Notices, 51(10):430–446, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stuckey:1995:FCI

[Stu95] Richard Stuckey. A fully conformant implementation of ECMA-162. Ada User, 16(2): 83–94, June 1995. CODEN AUJOET. ISSN 0268-652X.

Snavely:2002:SJP

[STV02] Allan Snavely, Dean M.

Tullsen, and Geoff Voelker. Symbiotic jobscheduling with priorities for a simultaneous multithreading processor. ACM SIGMETRICS Performance Evaluation Review, 30 (1):66–76, June 2002. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Schmidtmann:1993:DIM

[STW93] Carl Schmidtmann, Michael Tao, and Steven Watt. Design and implementation of a multi-threaded Xlib. In USENIX [USE93b], pages 193–203. ISBN 1-880446-48-0. LCCN QA 76.76 O63

U84 1993. URL http://www.
usenix.org/publications/
library/proceedings/sd93/

.

Shen:1999:ATL

[STY99] Kai Shen, Hong Tang, and Tao Yang. Adaptive twolevel thread management for fast MPI execution on shared memory machines. In ACM [ACM99b], page ??

Sigmund:1996:IBM

[SU96] U. Sigmund and T. Ungerer. Identifying bottlenecks in a multithreaded superscalar microprocessor. Lecture Notes in Computer Science, 1124:797– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sigmund:2001:SCS

[SU01] U. Sigmund and T. Ungerer. On speculation control in simultaneous multithreaded processors. J.UCS: Journal of Universal Computer Science, 7(9):848–868, September 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_7_9/
on_speculation_control_
in.

Suito:2012:DRM

[SUF^{+}12] Kazutoshi Suito, Rikuhei Ueda, Kei Fujii, Takuma Kogo, Hiroki Matsutani, and Nobuyuki Yamasaki. The dependable responsive multithreaded processor for distributed real-time systems. IEEE Micro, 32(6):52–61, November/December 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

SunSoft:1995:SMP

[Sun95] SunSoft. Solaris multithreaded programming guide. Sun-Soft Press, Mountainview, CA, USA, 1995. ISBN 0-13-160896-7. xviii + 158 pp. LCCN QA76.76.O63 S635 1995.

Sutter:1999:OAM

[Sut99]  Herb Sutter.  Optimiza-

tions that aren’t (in a multithreaded world). C/C++ Users Journal, 17(6):??, June 1999. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:1996:CAPb

[SV96a] D. C. Schmidt and S. Vinoski. Comparing alternative programming techniques for multithreaded CORBA servers. C++ Report, 8(4): 56–66, April 1996. CODEN CRPTE7. ISSN 1040-6042.

Schmidt:1996:CAPc

[SV96b] D. C. Schmidt and S. Vinoski.

Comparing alternative programming techniques for multithreaded CORBA servers. C++ Report, 8(7):47–56, July 1996. CODEN CRPTE7. ISSN 1040-6042.

Schmidt:1996:CAPa

[SV96c] D. C. Schmidt and S. Vinoski. Comparing alternative programming techniques for multithreaded servers. C++ Report, 8(2):50–59, February 1996. CODEN CRPTE7. ISSN 1040-6042.

Smith:1998:SIF

[SV98] Geoffrey Smith and Dennis Volpano. Secure information flow in a multi-threaded imperative language. In ACM [ACM98b], pages 355–

364. ISBN 0-89791-979-3. LCCN QA76.7 .A15 1998.

URL http://www.acm.org:
80/pubs/citations/proceedings/
plan/268946/p355-smith/.

ACM order number: 549981.

Sabarimuthu:2019:ADC

[SV19] J. M. Sabarimuthu and T. G. Venkatesh. Analytical derivation of concurrent reuse distance profile for multithreaded application running on chip multi-processor. IEEE Transactions on Parallel and Distributed Systems, 30(8): 1704–1721, August 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Steele:2021:PLB

[SV21] Guy L. Steele Jr. and Sebastiano Vigna. LXM: better splittable pseudorandom number generators (and almost as fast). Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):148:1–148:31, October 2021. CODEN ???? ISSN 2475-1421 (electronic).

URL https://dl.acm.org/
doi/10.1145/3485525.

Shepherd:1997:UCA

[SW97] George Shepherd and Scot Wingo. Undocumented corner: ATL and the IUknown interface. Dr. Dobb’s Journal of Software Tools, 22(8):119– 123, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Schaffer:2008:UHM

[SW08] Kevin Schaffer and Robert A.

Walker. Using hardware multithreading to overcome broadcast/reduction latency in an associative SIMD processor. Parallel Processing Letters, 18(4):491–509, December 2008. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Sleiman:2016:ESO

[SW16] Faissal M. Sleiman and Thomas F. Wenisch. Efficiently scaling out-of-order cores for simultaneous multithreading. ACM SIGARCH Computer Architecture News, 44(3):431–443, June 2016. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Sweetman:2007:SMR

[Swe07] Dominic Sweetman. See MIPS Run. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 2007. ISBN 0-12-088421-6. xix + 492 pp. LCCN QA76.9.A73 S88 2007.

Swinnen:2009:APA

[Swi09]  Gérard Swinnen.  Ap-

prendre á programmer avec Python: objet, multithreading, événements, bases de données, programmation web, programmation réseau, Unicode. Editions Eyrolles, Paris, France, 2009. xviii + 341 pp. LCCN ????

Shee:1994:DMA

[SWYC94] Jang Chung Shee, Chao Chin Wu, Lin Wen You, and Cheng Chen. Design of a multithread architecture and its parallel simulation and evaluation environment. In Anonymous [Ano94a], pages 69–76 (vol. 1). ISBN ???? LCCN ???? 2 vol.

Shih:2014:COR

[SYHL14] Wen-Li Shih, Yi-Ping You, Chung-Wen Huang, and Jenq Kuen Lee. Compiler optimization for reducing leakage power in multithread BSP programs. ACM Transactions on Design Automation of Electronic Systems., 20(1): 9:1–9:??, November 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Schwan:1992:MRT

[SZ92] Karsten Schwan and Hongyi Zhou. Multiprocessor realtime threads. Operating Systems Review, 26(1):54– 65, January 1992. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sterling:2002:GMP

[SZ02] Thomas L. Sterling and Hans P. Zima. Gilgamesh: a multithreaded processorin-memory architecture for petaflops computing. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN

???? URL http://www.sc-
2002.org/paperpdfs/pap.
pap105.pdf.

Schwan:1991:RTT

[SZG91] Karsten Schwan, Hongyi Zhou, and Ahmed Gheith. Real-time threads. Operating Systems Review, 25(4): 35–46, October 1991. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sinenian:2013:MMS

[SZM^{+}13] Nareg Sinenian, Alex B. Zylstra, Mario J.-E. Manuel, Johan A. Frenje, Atma D. Kanojia, Joshua Stillerman, and Richard D. Petrasso. A multithreaded modular software toolkit for control of complex experiments. Computing in Science and Engineering, 15 (1):66–75, January/February 2013. CODEN CSENFA. ISSN 1521-9615.

Taft:2013:TPS

[Taf13] S. Tucker Taft. Tutorial: proving safety of parallel /multithreaded programs. ACM SIGADA Ada Letters, 33(3): 1–2, December 2013. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Theobald:2000:LCE

[TAK^{+}00] Kevin B. Theobald, Gagan Agrawal, Rishi Kumar, Gerd Heber, Guang R. Gao, Paul Stodghill, and Keshav Pingali. Landing CG on EARTH: a case study of fine-grained multithreading on an evolutionary path. In ACM [ACM00],

page 47. URL http://www.
sc2000.org/proceedings/
techpapr/papers/pap293.
pdf.

Tamasanis:1995:MMW

[Tam95] Doug Tamasanis. Mathematica meets Warp. Byte Magazine, 20(5), May 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Thoziyoor:2008:CMM

[TAM^{+}08] Shyamkumar Thoziyoor, Jung Ho Ahn, Matteo Monchiero, Jay B. Brockman, and Norman P. Jouppi. A comprehensive memory modeling tool and its application to the design and analysis of future memory hierarchies. ACM SIGARCH Computer Architecture News, 36 (3):51–62, June 2008. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Tanner:1987:MTI

[Tan87] P. P. Tanner. Multi-thread input. Computer Graphics, 21(2):142–145, April 1987. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Tolmach:2004:IFL

[TAN04] Andrew Tolmach, Sergio Antoy, and Marius Nita. Implementing functional logic languages using multiple threads and stores. ACM SIG-PLAN Notices, 39(9):90–102, September 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tam:2007:TCS

[TAS07] David Tam, Reza Azimi, and Michael Stumm. Thread clustering: sharing-aware scheduling on SMP–CMP–SMT multiprocessors. Operating Systems Review, 41(3):47–58, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Thompson:1997:THP

[TB97a] P. Thompson and G. Bumgardner. Threads.h++: a portable C++ library for multithreaded programming. C++ Report, 9(3):24– 37, March 1997. CODEN CRPTE7. ISSN 1040-6042.

Thompson:1997:TPC

[TB97b] P. Thompson and G. Bumgardner. Threads.h++: a portable C++ library for multithreaded programming. C++ Report, 9(3):24– 37, March 1997. CODEN CRPTE7. ISSN 1040-6042.

Trotter:2022:MTO

[TCF22] James D. Trotter, Xing Cai, and Simon W. Funke. On memory traffic and optimisations for low-order finite element assembly algorithms on multi-core CPUs. ACM Transactions on Mathematical Software, 48(2):19:1–19:31, June 2022. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/doi/
10.1145/3503925.

Toulouse:1995:CID

[TCG95] Michel Toulouse, Teodor Gabriel Crainic, and Michel Gendreau. Communication issues in designing cooperative multi-thread parallel searches. Report CRT-95-47, Centre de recherche sur les transports, Université de Montréal, Montréal, Québec, Canada, 1995.

Thornley:1998:SSH

[TCI98] John Thornley, K. Mani Chandy, and Hiroshi Ishii. A system for structured highperformance multithreaded programming in Windows NT. In USENIX [USE98a], page ?? ISBN 1-880446-95-

2. LCCN QA76.76.O63 U885

1998.  URL http://www.
usenix.org/publications/
library/proceedings/usenix-
nt98/thornley.html;   http:
//www.usenix.org/publications/
library/proceedings/usenix-
nt98/thornley_slides.

Tino:2020:SXE

[TCS20] Anita Tino, Caroline Collange, and André Seznec. SIMT-X: Extending singleinstruction multi-threading to out-of-order cores. ACM Transactions on Architecture and Code Optimization, 17(2): 15:1–15:23, June 2020. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3392032.

Tseng:2003:DST

[TDW03] Y. Tseng, R. F. DeMara, and P. J. Wilder. Distributed-sum termination detection supporting multithreaded execution. Parallel Computing, 29 (7):953–968, July 2003. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Thekkath:1994:ISB

[TE94a] R. Thekkath and S. J. Eggers. Impact of sharingbased thread placement on multithreaded architectures. ACM SIGARCH Computer Architecture News, 22(2): 176–186, April 1994. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Thekkath:1994:EMH

[TE94b] Radhika Thekkath and Susan J. Eggers. The effectiveness of multiple hardware contexts. ACM SIGPLAN Notices, 29(11):328–337, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/asplos/195473/
p328-thekkath/.

Tullsen:1996:ECI

[TEE^{+}96] Dean M. Tullsen, Susan J. Eggers, Joel S. Emer, Henry M. Levy, Jack L. Lo, and Rebecca L. Stamm. Exploiting choice: instruction fetch and issue on an implementable simultaneous multithreading processor. ACM SIGARCH Computer Architecture News, 24(2):191–202, May 1996. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Tullsen:1995:SMM

[TEL95] Dean M. Tullsen, Susan J. Eggers, and Henry M. Levy. Simultaneous multithreading: maximizing on-chip parallelism. ACM SIGARCH Computer Architecture News, 23 (2):392–403, May 1995. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Tullsen:1998:RSM

[TEL98a] Dean M. Tullsen, Susan J. Eggers, and Henry M. Levy. Retrospective: Simultaneous multithreading: maximizing on-chip parallelism. In ACM [ACM98c], pages 115– 116. ISBN 0-8186-8491-7, 0-8186-8492-5, 0-8186-8493-3. LCCN QA76.9.A73 S97 1998.

URL http://portal.acm.
org/toc.cfm?id=279358;
http://portal.acm.org/toc.
cfm?id=285930. ACM Order

Number 414984. IEEE Computer Society Order Number PR08491; IEEE Order Plan Catalog Number 98CB36235.

Tullsen:1998:SMM

[TEL98b] Dean M. Tullsen, Susan J.

Eggers, and Henry M. Levy. Simultaneous multithreading: maximizing on-chip parallelism. In ACM [ACM98c], pages 533–544. ISBN 0-8186-8491-7, 0-8186-8492-5, 0-8186-8493-3. LCCN QA76.9.A73 S97 1998. URL http:/

/portal.acm.org/toc.cfm?
id=279358;   http://portal.
acm.org/toc.cfm?id=285930.

ACM Order Number 414984. IEEE Computer Society Order Number PR08491; IEEE Order Plan Catalog Number 98CB36235.

TempleLang:1997:MTE

[Tem97] Duncan Walter Temple Lang.

A multi-threaded extension to a high level interactive statistical computing environment. Thesis (Ph.D. in Statistics), Dept. of Statistics, University of California, Berkeley, Berkeley, CA, USA, December 1997. vii + 161 pp.

Tennberg:1998:CAD

[Ten98] Patrick Tennberg. Creating active data types via multithreading. C/C++ Users Journal, 16(1):??, January 1998. CODEN CCUJEX. ISSN 1075-2838.

Tennberg:2002:RGO

[Ten02] Patrick Tennberg. Refactoring global objects in multithreaded applications. C/C++ Users Journal, 20(5): 20–??, May 2002. CODEN CCUJEX. ISSN 1075-2838.

Trancoso:2006:CCM

[TESK06] Pedro Trancoso, Paraskevas Evripidou, Kyriakos Stavrou, and Costas Kyriacou. A case for chip multiprocessors based on the data-driven multithreading model. International Journal of Parallel Programming, 34(3):213–235, June 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=34&
                                        issue=3&spage=213.

Tetewsky:1994:GDR

[Tet94] Avram K. Tetewsky. GUI development for real-time applications. Dr. Dobb’s Journal of Software Tools, 19(6): 28, 30, 32, 36, 38, 40–41, June 1994. CODEN DDJOEB. ISSN 1044-789X.

Tian:2010:SPU

[TFG10] Chen Tian, Min Feng, and Rajiv Gupta. Speculative parallelization using state separation and multiple value prediction. ACM SIGPLAN Notices, 45(8):63–72, August 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tang:1999:APT

[TG99] Xinan Tang and Guang R.

Gao. Automatically partitioning threads for multithreaded architectures. Journal of Parallel and Distributed Computing, 58(2):159–189, August 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1999.1551/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1999.1551/production/pdf;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1999.1551/production/ref.

Thakur:2009:TSE

[TG09] Rajeev Thakur and William Gropp. Test suite for evaluating performance of multithreaded MPI communication. Parallel Computing, 35 (12):608–617, December 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tian:2005:PCT

[TGBS05] Xinmin Tian, Milind Girkar, Aart Bik, and Hideki Saito. Practical compiler techniques on efficient multithreaded code generation for OpenMP programs. The Computer Journal, 48(5):588–601, September 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/5/588;   http://comjnl.
oxfordjournals.org/cgi/
reprint/48/5/588.

Tan:1999:OFN

[TGO99] Kian-Lee Tan, Cheng Hian Goh, and Beng Chin Ooi. Online feedback for nested aggregate queries with multithreading. In Atkinson et al. [AOV^{+}99], pages 18–29. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL http:

//www.vldb.org/dblp/db/
conf/vldb/TanGO99.html. Also

known as VLDB’99.

Tan:2000:PEN

[TGO00] Kian-Lee Tan, Cheng Hian Goh, and Beng Chin Ooi. Progressive evaluation of nested aggregate queries. VLDB Journal: Very Large Data Bases, 9(3):261–278, December 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Terechko:2012:BPS

[THA^{+}12] Andrei Terechko, Jan Hoogerbrugge, Ghiath Alkadi, Surendra Guntur, Anirban Lahiri, Marc Duranton, Clemens Wüst, Phillip Christie, Axel Nackaerts, and Aatish Kumar. Balancing programmability and silicon efficiency of heterogeneous multicore architectures. ACM Transactions on Embedded Computing Systems, 11S(1):14:1–14:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Thekkath:1995:DPM

[The95] Radhika Thekkath. Design and performance of multithreaded architectures. Thesis (Ph.D.), University of Washington, Seattle, WA, USA, 1995. x + 100 pp.

Throop:1999:SOS

[Thr99]  Joe Throop.  Standards:

OpenMP: Shared-memory parallelism from the ashes. Computer, 32(5):108–109, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co1999/pdf/r5108.
pdf.

Timmerman:2003:EWC

[Tim03] Martin Timmerman. Examining Windows CE .NET. Dr. Dobb’s Journal of Software Tools, 28(2):62, 64, February 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/documents/
s=7790/ddj0302h/.

Tsai:1998:POC

[TJY98] J.-Y. Tsai, Z. Jiang, and P.-C. Yew. Program optimization for concurrent multithreaded architectures. Lecture Notes in Computer Science, 1366:146–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tu:2011:MBM

[TJY^{+}11] Xuping Tu, Hai Jin, Zhibin Yu, Jie Chen, Yabin Hu, and Xie Xia. MT-BTRIMER: A master-slave multi-threaded dynamic binary translator. International Journal of Computer Systems Science and Engineering, 26(5):??, September 2011. CODEN CSSEEI. ISSN 0267-6192.

Thitikamol:1998:PNM

[TK98] K. Thitikamol and P. Keleher. Per-node multithreading and remote latency. IEEE Transactions on Computers, 47(4):414–426, April 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&
                                        arnumber=675711.

Theobald:2001:DCI

[TKA^{+}01] Kevin B. Theobald, Rishi Kumar, Gagan Agrawal, Gerd Heber, Ruppa K. Thulasiram, and Guang R. Gao. Developing a communication intensive application on the EARTH multithreaded architecture (distinguished paper). Lecture Notes in Computer Science, 1900:625–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1900/19000625.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1900/19000625.
pdf.

Theobald:2002:IEC

[TKA^{+}02] Kevin B. Theobald, Rishi Kumar, Gagan Agrawal, Gerd Heber, Ruppa K. Thulasiram, and Guang R. Gao. Implementation and evaluation of a communication intensive application on the EARTH multithreaded system. Concurrency and Computation: Practice and Experience, 14 (3):183–201, March 2002. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/93513486/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=93513486{\&}PLACEBO=
IE.pdf.

Thulasiraman:2004:FGL

[TKHG04] Parimala Thulasiraman, Ashfaq A. Khokhar, Gerd Heber, and Guang R. Gao. A finegrain load-adaptive algorithm of the 2D discrete wavelet transform for multithreaded architectures. Journal of Parallel and Distributed Computing, 64(1):68–78, January 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tang:2021:MMR

[TKK21] Xulong Tang, Mahmut Taylan Kandemir, and Mustafa Karakoy. Mix and match: Reorganizing tasks for enhancing data locality. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2): 20:1–20:24, June 2021. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3460087.

Tang:2018:CND

[TKZ^{+}18] Xulong Tang, Mahmut Taylan Kandemir, Hui Zhao, Myoungsoo Jung, and Mustafa Karakoy. Computing with near data. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(3):42:1– 42:30, December 2018. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3287321.

Editors:2002:LUC

[TLA^{+}02] The Editors, Kim Reidar Lantz, Ze’ev Atlas, Pete Nelson, and Gus J. Grubba. Letters: URL correction [“The NewOS Operating System”]; passing context to threads; compiling Perl/Tk scripts; standing by Al’s principles; understanding photomosaics. Dr. Dobb’s Journal of Software Tools, 27(1):10, 12, January 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/.

See [Gei01].

Turakhia:2017:TPE

[TLGM17] Yatish Turakhia, Guangshuo Liu, Siddharth Garg, and Diana Marculescu. Thread progress equalization: Dynamically adaptive powerconstrained performance optimization of multi-threaded applications. IEEE Transactions on Computers, 66(4): 731–744, ???? 2017. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tian:2016:ETR

[TLZ^{+}16] Zhenzhou Tian, Ting Liu, Qinghua Zheng, Ming Fan, Eryue Zhuang, and Zijiang Yang. Exploiting threadrelated system calls for plagiarism detection of multithreaded programs. The Journal of Systems and Software, 119(??):136–148, September 2016. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0164121216300838.

Tian:2017:RSP

[TLZ^{+}17] Z. Tian, T. Liu, Q. Zheng, E. Zhuang, M. Fan, and Z. Yang. Reviving sequential program birthmarking for multithreaded software plagiarism detection. IEEE Transactions on Software Engineering, PP(99):1, ???? 2017. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
7888597.

Tian:2018:RSP

[TLZ^{+}18] Zhenzhou Tian, Ting Liu, Qinghua Zheng, Eryue Zhuang, Ming Fan, and Zijiang Yang. Reviving sequential program birthmarking for multithreaded software plagiarism detection. IEEE Transactions on Software Engineering, 44(5):491–511, May 2018. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

https://ieeexplore.ieee.
org/document/7888597/.

Tremblay:2003:IEP

[TMAG03] G. Tremblay, C. J. Morrone, J. N. Amaral, and G. R. Gao. Implementation of the EARTH programming model on SMP clusters: a multithreaded language and runtime system. Concurrency and Computation: Practice and Experience, 15(9):821– 844, August 10, 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tallent:2009:EPM

[TMC09] Nathan R. Tallent and John M.

Mellor-Crummey. Effective performance measurement and analysis of multithreaded applications. ACM SIGPLAN Notices, 44(4):229– 240, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tallent:2010:ALC

[TMCP10] Nathan R. Tallent, John M.

Mellor-Crummey, and Allan Porterfield. Analyzing lock contention in multithreaded applications. ACM SIGPLAN Notices, 45(5):269–280, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taylor:1995:CSA

[TNB^{+}95] Richard N. Taylor, Kari A.

Nies, Gregory Alan Bolcer, Craig A. MacFarlane, Kenneth M. Anderson, and Gregory F. Johnson. Chiron-1: a software architecture for user interface development, maintenance, and run-time support. ACM Transactions on Computer-Human Interaction, 2(2):105–144, June 1995. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tochi/1995-2-2/
p105-taylor/.

Trott:2010:AVI

[TO10] Oleg Trott and Arthur J. Olson. AutoDock Vina: Improving the speed and accuracy of docking with a new scoring function, efficient optimization, and multithreading. Journal of Computational Chemistry, 31(2):455– 461, January 30, 2010. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Todiwala:1995:DRT

[Tod95] Khushroo Rustom Todiwala. A distributed ray tracing implementation using multithreaded RPC. Thesis (M.S.), University of Texas at El Paso, El Paso, TX, USA, 1995. xi + 140 pp.

Thebault:2018:AMC

[TP18] Loïc Thébault and Eric Petit. Asynchronous and multithreaded communications on irregular applications using vectorized divide and conquer approach. Journal of Parallel and Distributed Computing, 114(??):16–27, April 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0743731517303350.

Tao:2021:CDS

[TPZ21] Xiaohan Tao, Jianmin Pang, and Yu Zhu. Compilerdirected scratchpad memory data transfer optimization for multithreaded applications on a heterogeneous many-core architecture. The Journal of Supercomputing, 77(12):14502– 14524, December 2021. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s11227-021-03853-x.

Tarvo:2014:AAM

[TR14] Alexander Tarvo and Steven P.

Reiss. Automated analysis of multithreaded programs for performance modeling. ACM SIGMETRICS Performance Evaluation Review, 42(1):557– 558, June 2014. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Traub:1991:MTC

[Tra91] Kenneth R. Traub. Multithread code generation for dataflow architectures from non-strict programs. Lecture Notes in Computer Science, 523:73–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Troutwine:2018:HCR

[Tro18] Brian L. Troutwine. Handson Concurrency with Rust: Confidently Build Memorysafe, Parallel, and Efficient Software in Rust. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78839-997-8 (paperback), 1-78847-835-5. v + 449 pp. LCCN QA76.76.A65. URL http://

                                        proquest.safaribooksonline.
                                        com/?fpi=9781788399975.

Tsai:1997:PSC

[Tsa97a] Jenn-Yuan Tsai. Performance study of a concurrent multithreaded processor. Technical report TR 97-034, University of Minnesota, Dept. of Computer Science and Engineering, Minneapolis, MN, USA, 1997. 24 pp.

Tsai:1997:SIC

[Tsa97b] Jenn-Yuan Tsai. Superthreading: integrating compilation technology and processor architecture for costeffective concurrent multithreading. Technical report TR 97-033, University of Minnesota, Dept. of Computer Science and Engineering, Minneapolis, MN, USA, January 29, 1997. 16 pp.

Torrant:1999:SMS

[TSCH99] Marc Torrant, Muhammad Shaaban, Roy Czernikowski, and Ken Hsu. A simultaneous multithreading simulator. ACM SIGARCH Computer Architecture News, 27 (5):1–5, December 1999. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Tumeo:2012:DNG

[TSV12] Antonino Tumeo, Simone Secchi, and Oreste Villa. Designing next-generation massively multithreaded architectures for irregular applications. Computer, 45(8): 53–61, August 2012. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tang:1999:CRT

[TSY99] Hong Tang, Kai Shen, and Tao Yang. Compile/runtime support for threaded MPI execution on multiprogrammed shared memory machines. ACM SIGPLAN Notices, 34(8):107–118, August 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
proceedings/ppopp/301104/
p107-tang/.

Tang:2000:PTR

[TSY00] Hong Tang, Kai Shen, and Tao Yang. Program transformation and runtime support for threaded MPI execution on shared-memory machines. ACM Transactions on Programming Languages and Systems, 22(4): 673–700, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/toplas/2000-22-
4/p673-tang/.

Thulasiram:2003:PEM

[TT03] Ruppa K. Thulasiram and Parimala Thulasiraman. Performance evaluation of a multithreaded Fast Fourier Transform algorithm for derivative pricing. The Journal of Supercomputing, 26(1): 43–58, August 2003. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

                                        /ipsapp009.kluweronline.
                                        com/content/getfile/5189/
                                        46/4/abstract.htm;   http:
                                        //ipsapp009.kluweronline.
                                        com/content/getfile/5189/
                                        46/4/fulltext.pdf.

Thulasiraman:2002:EMA

[TTKG02] Parimala Thulasiraman, Kevin Theobald, Ashfaq A. Khokhar, and Guang R. Gao. Efficent multithreaded algorithms for the Fast Fourier Transform. Parallel and Distributed Computing Practices, 5(2):239– 258, June 2002. CODEN ???? ISSN 1097-2803.

Taura:1999:SMI

[TTY99] Kenjiro Taura, Kunio Tabata, and Akinori Yonezawa. Stack-Threads/MP: integrating futures into calling standards. ACM SIGPLAN Notices, 34(8):60–71, August 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/

proceedings/ppopp/301104/
p60-taura/.

Tullsen:1996:SM

[Tul96] Dean Michael Tullsen. Simultaneous multithreading. Thesis (Ph.D.), University of Washington, Seattle, WA, USA, 1996. vi + 99 pp.

Tentyukov:2010:MVF

[TV10] M. Tentyukov and J. A. M.

Vermaseren. The multithreaded version of FORM. Computer Physics Communications, 181(8):1419–1427, August 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465510001207.

Tembey:2013:SSS

[TVB^{+}13] Priyanka Tembey, Augusto Vega, Alper Buyuktosunoglu, Dilma Da Silva, and Pradip Bose. SMT switch: Software mechanisms for power shifting. IEEE Computer Architecture Letters, 12(2):67–70, July/December 2013. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Torlak:2010:MCA

[TVD10] Emina Torlak, Mandana Vaziri, and Julian Dolby. MemSAT: checking axiomatic specifications of memory models. ACM SIGPLAN Notices, 45(6):341–350, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Turon:2014:GNW

[TVD14] Aaron Turon, Viktor Vafeiadis, and Derek Dreyer. GPS: navigating weak memory with ghosts, protocols, and separation. ACM SIGPLAN Notices, 49(10):691–707, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taura:1997:FGM

[TY97] Kenjiro Taura and Akinori Yonezawa. Fine-grain multithreading with minimal compiler support — a cost effective approach to implementing efficient multithreading languages. ACM SIG-PLAN Notices, 32(5):320–333, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/pldi/258915/
p320-taura/.

Utterback:2017:POR

[UALK17] Robert Utterback, Kunal Agrawal, I-Ting Angelina Lee, and Milind Kulkarni. Processor-oblivious record and replay. ACM SIGPLAN Notices, 52(8):145–161, August 2017. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Utterback:2019:POR

[UALK19] Robert Utterback, Kunal Agrawal, I-Ting Angelina Lee, and Milind Kulkarni. Processor-oblivious record and replay. ACM Transactions on Parallel Computing (TOPC), 6(4):20:1– 20:??, December 2019. CO-DEN ???? ISSN 2329-4949 (print), 2329-4957 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3365659.

Ungerer:2002:MP

[URŠ02a] Theo Ungerer, Borut Robič, and Jurij Šilc. Multithreaded processors. The Computer Journal, 45(3):320– 348, ???? 2002. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_
journal/hdb/Volume_45/Issue_
03/450320.sgm.abs.html;
 http://www3.oup.co.uk/
computer_journal/hdb/Volume_
45/Issue_03/pdf/450320.
pdf.

Ungerer:2002:SPE

[URŠ02b] Theo Ungerer, Borut Robič, and Jurij Šilc. A survey of processors with explicit multithreading. ACM Computing Surveys, 35(1):29–63, March 2002. CODEN CMSVAN.

ISSN 0360-0300 (print), 1557-7341 (electronic).

Ungerer:2003:SPE

[URŠ03] Theo Ungerer, Borut Robič, and Jurij Šilc. A survey of processors with explicit multithreading. ACM Computing Surveys, 35(1):29–63, March 2003. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

USENIX:1989:PWU

[USE89] USENIX Association, editor. Proceedings of the Winter 1989 USENIX Conference: January 30–February 3, 1989, San Diego, California, USA. USENIX Association, Berkeley, CA, USA, 1989.

USENIX:1991:PUM

[USE91a] USENIX, editor. Proceedings of the USENIX Mach Symposium: November 20–22, 1991, Monterey, California, USA. USENIX Association, Berkeley, CA, USA, 1991. LCCN QAX 27.

USENIX:1991:PWU

[USE91b] USENIX, editor. Proceedings of the Winter 1991 USENIX Conference: January 21– January 25, 1991, Dallas, TX, USA. USENIX Association, Berkeley, CA, USA, 1991. LCCN QA 76.76 O63 U84 1992.

USENIX:1992:PSU

[USE92a] USENIX, editor. Proceedings of the Summer 1992 USENIX Conference: June 8–12, 1992, San Antonio, Texas, USA. USENIX Association, Berkeley, CA, USA, Summer 1992. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

USENIX:1992:SED

[USE92b]  USENIX, editor.  Sympo-

sium on Experiences with Distributed and Multiprocessor Systems (SEDMS III), March 26–27, 1992. Newport Beach, CA. USENIX Association, Berkeley, CA, USA, March 26–27, 1992. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

USENIX:1993:PUMb

[USE93a] USENIX, editor. Proceedings of the USENIX Mobile and Location-Independent Computing Symposium: August 2–3, 1993, Cambridge, Massachusetts, USA. USENIX Association, Berkeley, CA, USA, 1993. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86

1993.  URL http://www.
usenix.org/publications/
library/proceedings/mobile93/

.

USENIX:1993:PWU

[USE93b] USENIX, editor. Proceedings of the Winter 1993 USENIX Conference: January 25–29, 1993, San Diego, California, USA. USENIX Association, Berkeley, CA, USA, 1993. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84

1993.  URL http://www.
usenix.org/publications/
library/proceedings/sd93/

.

USENIX:1996:PFA

[USE96] USENIX, editor. Proceedings of the fourth annual Tcl/Tk Workshop, July 10– 13, 1996, Monterey, California. USENIX Association, Berkeley, CA, USA, 1996. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35

1996.  URL http://www.
usenix.org/publications/
library/proceedings/tcl96/

.

USENIX:1998:PUWa

[USE98a] USENIX, editor. Proceedings of the 2nd USENIX Windows NT Symposium: August 3–5, 1998, Seattle, Washington. USENIX Association, Berkeley, CA, USA, 1998. ISBN 1-880446-95-2. LCCN QA76.76.O63 U885

1998.  URL http://db.
usenix.org/publications/
library/proceedings/usenix-
nt98.

USENIX:1998:PSA

[USE98b] USENIX, editor. Proceedings of the sixth annual Tcl/Tk Conference, September 18– 24 [i.e. 14–18], 1998, San Diego, California. USENIX Association, Berkeley, CA, USA, 1998. ISBN 1-880446-98-7. LCCN QA76.73.T44 T34 1998. URL http://db.

usenix.org/publications/
library/proceedings/tcl98/

.

USENIX:2000:UAT

[USE00a]  USENIX, editor.  2000

USENIX Annual Technical Conference: San Diego, CA, USA, June 18–23, 2000. USENIX Association, Berkeley, CA, USA, 2000. ISBN 1-880446-22-7. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/usenix2000.

USENIX:2000:PUT

[USE00b] USENIX, editor. Proceedings of the 7th USENIX Tcl/Tk Conference (Tcl/2k): February 14–18, 2000, Austin, Texas, USA. USENIX Association, Berkeley, CA, USA, 2000. ISBN 1-880446-24-3. LCCN ???? URL http://db.

usenix.org/publications/
library/proceedings/tcl2k/

.

USENIX:2001:PJV

[USE01] USENIX, editor. Proceedings of the Java Virtual Machine Research and Technology Sy[m]posium (JVM ’01): April 23–24, 2001, Monterey, California, USA. Berkeley, CA. USENIX Association, Berkeley, CA, USA, 2001. ISBN 1-880446-11-

1. LCCN QA76.73.J38 J42

2001.  URL http://www.
usenix.org/publications/

                                        library/proceedings/jvm01/

.

USENIX:2002:PBF

[USE02] USENIX, editor. Proceedings of BSDCon 2002: February 11–14, 2002, Cathedral Hill Hotel, San Francisco, CA. USENIX Association, Berkeley, CA, USA, 2002. ISBN 1-880446-02-

2. LCCN QA76.76.O63 B736

                                        2002.  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/bsdcon02/
                                        tech.html.

Unger:2000:CCA

[UZU00] A. Unger, E. Zehendner, and Th. Ungerer. A combined compiler and architecture technique to control multithreaded execution of branches and loop iterations. ACM SIGARCH Computer Architecture News, 28 (1):53–61, March 2000. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Vera:2009:SRL

[VACG09] Xavier Vera, Jaume Abella, Javier Carretero, and Antonio González. Selective replication: a lightweight technique for soft errors. ACM Transactions on Computer Systems, 27(4):8:1–8:30, December 2009. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

vanHoff:1995:JIP

[van95] Arthur van Hoff. Java and Internet programming. Dr. Dobb’s Journal of Software Tools, 20(8):56, 58, 60–61, 101–102, August 1995. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ddj/issues/j508a.htm.

Vanhelsuwe:1997:BRJ

[Van97a] Laurence Vanhelsuwé. Book review: The Java Threads API makes it to print media. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CO-DEN ???? ISSN 1091-

8906.  URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-threads.
htm.

Vanhelsuwe:1997:JPE

[Van97b] Laurence Vanhelsuwé. JavaBeans: properties, events, and thread safety. JavaWorld: IDG’s magazine for the Java community, 2(9):??, September 1997. CODEN ???? ISSN 1091-8906. URL http://www.

javaworld.com/javaworld/
jw-09-1997/jw-09-raceconditions.
htm.

Vckovski:2000:MTS

[VB00] Andrej Vckovski and Jason Brazile. A multi-threaded server for shared hash table access. In USENIX [USE00b], page ?? ISBN 1-880446-24-3. LCCN ???? URL http://db.

usenix.org/publications/

library/proceedings/tcl2k/
vckovski.html.

Venkataramani:2019:SMM

[VCM19] Vanchinathan Venkataramani, Mun Choon Chan, and Tulika Mitra. Scratchpadmemory management for multi-threaded applications on many-core architectures. ACM Transactions on Embedded Computing Systems, 18(1):10:1–10:??, February 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3301308.

Volkov:2008:LQC

[VD08] Vasily Volkov and James W. Demmel. LU, QR and Cholesky factorizations using vector capabilities of GPUs. LAPACK Working Note 202, Department of Electrical Engineering and Computer Science, University of California, Berkeley, Berkeley, CA, USA, May 2008. URL http:

//www.netlib.org/lapack/
lawnspdf/lawn202.pdf.

Vishkin:1998:EMT

[VDBN98] Uzi Vishkin, Shlomit Dascal, Efraim Berkovich, and Joseph Nuzman. Explicit multi-threading (XMT) bridging models for instruction parallelism (extended abstract). In SPAA ’98: 10th Annual ACM Symposium on Parallel Algorithms and Architectures, June 28–July 2, 1998, Puerto Vallarta, Mexico. ACM Press, New York, NY 10036, USA, 1998. ISBN 0-89791-989-0. LCCN QA76.58 .A26 1998. URL

http://delivery.acm.org/
10.1145/280000/277680/p140-
vishkin.pdf. ACM order

number 417980.

Volkman:1993:CDB

[VE93] Victor R. Volkman and John English. Class DOSThread: a base class for multithreaded DOS programs. C Users Journal, 11(12):113–??, December 1993. ISSN 0898-9788.

Venners:1997:UHH

[Ven97] Bill Venners. Under the hood:

How the Java virtual machine performs thread synchronization. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CO-DEN ???? ISSN 1091-

8906.  URL http://www.
javaworld.com/javaworld/
jw-07-1997/jw-07-hood.htm.

Venners:1998:DTS

[Ven98] Bill Venners. Design for thread safety. JavaWorld: IDG’s magazine for the Java community, 3(8):??, August 1998. CODEN ???? ISSN 1091-8906. URL http://www.

javaworld.com/javaworld/
jw-08-1998/jw-08-techniques.
htm.

Verriello:1996:MSM

[Ver96] Anthony Verriello. Memory sharing in multithreaded transaction environments. Thesis (M.S.), Hofstra University, Westport, CT, USA, 1996. 180 pp.

Vermeulen:1997:JDW

[Ver97] Alain Vermeulen. Java deadlock: The woes of multithreaded design. Dr. Dobb’s Journal of Software Tools, 22 (9):52, 54–56, 88, 89, September 1997. CODEN DDJOEB. ISSN 1044-789X.

Vlachos:2010:PEAa

[VGK^{+}10a] Evangelos Vlachos, Michelle L. Goodstein, Michael A. Kozuch, Shimin Chen, Babak Falsafi, Phillip B. Gibbons, and Todd C. Mowry. Para-Log: enabling and accelerating online parallel monitoring of multithreaded applications. ACM SIGARCH Computer Architecture News, 38 (1):271–284, March 2010. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Vlachos:2010:PEAb

[VGK^{+}10b] Evangelos Vlachos, Michelle L. Goodstein, Michael A. Kozuch, Shimin Chen, Babak Falsafi, Phillip B. Gibbons, and Todd C. Mowry. ParaLog: enabling and accelerating online parallel monitoring of multithreaded applications. ACM SIGPLAN Notices, 45(3):271– 284, March 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vasconcelos:2006:TCM

[VGR06] Vasco T. Vasconcelos, Simon J. Gay, and António Ravara. Type checking a multithreaded functional language with session types. Theoretical Computer Science, 368(1–2):64–87, December 5, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vachharajani:2005:CMP

[VIA^{+}05] Neil Vachharajani, Matthew Iyer, Chinmay Ashok, Manish Vachharajani, David I. August, and Daniel Connors. Chip multi-processor scalability for single-threaded applications. ACM SIGARCH Computer Architecture News, 33(4):44–53, November 2005. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Vlassov:1999:QMM

[VK99] V. Vlassov and A. Kraynikov.

A queuing model of a multithreaded architecture: a case study. Lecture Notes in Computer Science, 1662:306– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Volkman:1993:CCP

[Vol93] Victor R. Volkman. Convert C programs into multithreaded applications. C Users Journal, 11(4):87–??, April 1993. ISSN 0898-9788.

Verdu:2016:PSA

[VP16] Javier Verdu and Alex Pajuelo. Performance scalability analysis of JavaScript applications with Web Workers. IEEE Computer Architecture Letters, 15(2):105–108, July/ December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Vijaykumar:2002:TFR

[VPC02] T. N. Vijaykumar, Irith Pomeranz, and Karl Cheng. Transient-fault recovery using simultaneous multithreading. ACM SIGARCH Computer Architecture News, 30 (2):87–98, May 2002. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

vonPraun:2003:SCA

[vPG03] Christoph von Praun and Thomas R. Gross. Static conflict analysis for multithreaded object-oriented programs. ACM SIGPLAN Notices, 38(5):115–128, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vitali:2012:LSO

[VPQ12] Roberto Vitali, Alessandro Pellegrini, and Francesco Quaglia. Load sharing for optimistic parallel simulations on multi core machines. ACM SIGMETRICS Performance Evaluation Review, 40(3):2– 11, December 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Vrenios:2004:PPC

[Vre04] A. Vrenios. Parallel Programming in C with MPI and OpenMP [book review]. IEEE Distributed Systems Online, 5(1):7.1– 7.3, ???? 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://ieeexplore.ieee.
org/iel5/8968/28452/01270716.
pdf?isnumber=28452&prod=
JNL&arnumber=1270716&arSt=
+7.1&ared=+7.3&arAuthor=
Vrenios%2C+A.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=28452&
arnumber=1270716&count=
8&index=5.

Vinoski:1996:DCD

[VS96] S. Vinoski and D. C. Schmidt.

Distributed callbacks and decoupled communication in CORBA. C++ Report, 8(9): 48–56, 77, October 1996. CO-DEN CRPTE7. ISSN 1040-6042.

Vandierendonck:2011:FMM

[VS11a] Hans Vandierendonck and Andre Seznec. Fairness metrics for multi-threaded processors. IEEE Computer Architecture Letters, 10(1):4–7, January/June 2011. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Vandierendonck:2011:MSR

[VS11b] Hans Vandierendonck and André Seznec. Managing SMT resource usage through speculative instruction window weighting. ACM Transactions on Architecture and Code Optimization, 8(3):12:1– 12:??, October 2011. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Vander-Swalmen:2009:CAM

[VSDK09] Pascal Vander-Swalmen, Gilles Dequen, and Michaël Krajecki. A collaborative approach for multi-threaded SAT solving. International Journal of Parallel Programming, 37(3):324–342, June 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=37&
                                        issue=3&spage=324.

Vale:2016:PDT

[VSDL16] Tiago M. Vale, João A. Silva, Ricardo J. Dias, and João M. Lourenço. Pot: Deterministic transactional execution. ACM Transactions on Architecture and Code Optimization, 13(4):52:1–52:??, December 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Vantrease:2008:CSI

[VSM^{+}08] Dana Vantrease, Robert Schreiber, Matteo Monchiero, Moray McLaren, Norman P. Jouppi, Marco Fiorentino, Al Davis, Nathan Binkert, Raymond G. Beausoleil, and Jung Ho Ahn. Corona: System implications of emerging nanophotonic technology. ACM SIGARCH Computer Architecture News, 36 (3):153–164, June 2008. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

VanZee:2016:BFE

[VSM^{+}16] Field G. Van Zee, Tyler M.

Smith, Bryan Marker, Tze Meng Low, Robert A. Van De Geijn, Francisco D. Igual, Mikhail Smelyanskiy, Xianyi Zhang, Michael Kistler, Vernon Austel, John A. Gunnels, and Lee Killough. The BLIS framework: Experiments in portability. ACM Transactions on Mathematical Software, 42(2): 12:1–12:19, June 2016. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Vlassov:1996:AMM

[VT96] V. Vlassov and L.-E. Thorelli.

Analytical models of multithreading with data prefetching. Lecture Notes in Computer Science, 1124:714–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Volos:2012:ATM

[VTSL12] Haris Volos, Andres Jaan Tack, Michael M. Swift, and Shan Lu. Applying transactional memory to concurrency bugs. ACM SIGARCH Computer Architecture News, 40 (1):211–222, March 2012. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). ASPLOS ’12 conference proceedings.

Villa:2012:FAS

[VTSM12] Oreste Villa, Antonino Tumeo, Simone Secchi, and Joseph B. Manzano. Fast and accurate simulation of the Cray XMT multithreaded supercomputer. IEEE Transactions on Parallel and Distributed Systems, 23(12): 2266–2279, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Vishkin:2000:ELR

[VV00] Dascal Vishkin and Uzi Vishkin. Experiments with list ranking for explicit multithreaded (XMT) instruction parallelism. ACM Journal of Experimental Algorithmics, 5: 10:1–10:??, ???? 2000. CO-DEN ???? ISSN 1084-6654.

VanDeGeijn:2011:HPD

[VV11] Robert A. Van De Geijn and Field G. Van Zee. High-performance up-and-downdating via Householderlike transformations. ACM Transactions on Mathematical Software, 38(1):4:1–4:17, November 2011. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Winter:2008:ATN

[WA08] Jonathan A. Winter and David H. Albonesi. Addressing thermal nonuniformity in SMT workloads. ACM Transactions on Architecture and Code Optimization, 5(1):4:1– 4:??, May 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Walter:1995:PMS

[Wal95] Stephen Walter. Put multiprocessing systems to work. II. UNIX review, 13(1):39– ??, January 1995. CODEN UNRED5. ISSN 0742-3136.

Walmsley:2000:MTP

[Wal00]  Mark Walmsley.  Multi-

threaded programming in C++. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 1-85233-146-1. x + 223 pp. LCCN QA76.73.C153 W3148 2000. US$49.95.

Wang:1994:MAD

[Wan94]  Xiaobao Wang.  Multi-

threaded architecture: design and performance analysis. Thesis (M. S.), Department of Electrical Engineering, University of Hawaii at Manoa, Manoa, HI, USA, 1994. 59 pp.

Watt:1991:IPI

[Wat91] Stephen M. Watt, editor. ISSAC ’91: proceedings of the 1991 International Symposium on Symbolic and Algebraic Computation, July 15– 17, 1991, Bonn, Germany. ACM Press, New York, NY 10036, USA, 1991. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Wayner:1995:FAN

[Way95] Peter Wayner. Free agents: a new generation of lightweight, multithreaded operating environments provide security and interoperability for agent developers. Byte Magazine, 20(3):105–??, March 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wu:1999:GMC

[WC99] C.-C. Wu and C. Chen. Grouping memory consistency model for parallelmultithreaded shared-memory multiprocessor systems. International Journal of High Speed Computing (IJHSC), 10 (1):53–82, March 1999. CO-DEN IHSCEZ. ISSN 0129-0533.

Wang:2007:EAP

[WCC^{+}07] Perry H. Wang, Jamison D. Collins, Gautham N. Chinya, Hong Jiang, Xinmin Tian, Milind Girkar, Nick Y. Yang, Guei-Yuan Lueh, and Hong Wang. EXOCHI: architecture and programming environment for a heterogeneous multi-core multithreaded system. ACM SIGPLAN Notices, 42(6):156–166, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wallace:1998:TMP

[WCT98] Steven Wallace, Brad Calder, and Dean M. Tullsen. Threaded multiple path execution. ACM SIGARCH Computer Architecture News, 26(3):238– 249, June 1998. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Wilde:1998:RES

[WCV^{+}98] Norman Wilde, Christopher Casey, Joe Vandeville, Gary Trio, and Dick Hotz. Reverse engineering of software threads: a design recovery technique for large multiprocess systems. The Journal of Systems and Software, 43 (1):11–17, October 1998. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2004:HTVd

[WCW^{+}04a] Perry H. Wang, Jamison D.

Collins, Hong Wang, Dongkeun Kim, Bill Greene, Kai-Ming Chan, Aamir B. Yunus, Terry Sych, Stephen F. Moore, and John P. Shen. Helper threads via virtual multithreading. IEEE Micro, 24(6):74–82, November/December 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2004/06/m6074.htm;   http:
//csdl.computer.org/dl/
mags/mi/2004/06/m6074.pdf.

Wang:2004:HTVa

[WCW^{+}04b] Perry H. Wang, Jamison D. Collins, Hong Wang, Dongkeun Kim, Bill Greene, Kai-Ming Chan, Aamir B. Yunus, Terry Sych, Stephen F. Moore, and John P. Shen. Helper threads via virtual multithreading on an experimental Itanium-2 processor-based platform. ACM SIGARCH Computer Architecture News, 32(5):144– 155, December 2004. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Wang:2004:HTVb

[WCW^{+}04c] Perry H. Wang, Jamison D. Collins, Hong Wang, Dongkeun Kim, Bill Greene, Kai-Ming Chan, Aamir B. Yunus, Terry Sych, Stephen F. Moore, and John P. Shen. Helper threads via virtual multithreading on an experimental Itanium-2 processorbased platform. ACM SIG-PLAN Notices, 39(11):144– 155, November 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2004:HTVc

[WCW^{+}04d] Perry H. Wang, Jamison D.

Collins, Hong Wang, Dongkeun Kim, Bill Greene, Kai-Ming Chan, Aamir B. Yunus, Terry Sych, Stephen F. Moore, and John P. Shen. Helper threads via virtual multithreading on an experimental Itanium-2 processor-based platform. Operating Systems Review, 38 (5):144–155, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2007:OSC

[WCZ^{+}07] Qin Wang, Junpu Chen, Weihua Zhang, Min Yang, and Binyu Zang. Optimizing software cache performance of packet processing applications. ACM SIGPLAN Notices, 42(7):227–236, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wester:2013:PDR

[WDC^{+}13] Benjamin Wester, David Devecsery, Peter M. Chen, Jason Flinn, and Satish Narayanasamy. Parallelizing data race detection. ACM SIGPLAN Notices, 48(4):27– 38, April 2013. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weaver:2008:OIO

[Wea08] David L. Weaver, editor. OpenSPARC Internals: OpenSPARC T1/T2 Chip Multithreaded Throughput Computing. Lulu, Inc., 860 Aviation Parkway, Suite 300, Morrisville, NC 27560, USA, 2008. ISBN 0-557-01974-5. xviii + 369 pp. LCCN ???? US$20.00. URL http://www.

                                        opensparc.net/publications/
                                        books/opensparc-internals.
                                        html.

Weisz:1997:MFA

[Wei97] Russell Weisz. More first aid for the thread impaired: Cool ways to take advantage of multithreading. Microsoft Systems Journal, 12(7):33–??, July 1997. CODEN MSJOED. ISSN 0889-9932.

Weissman:1998:ATT

[Wei98a]  Boris Weissman.  Active

threads: towards efficient finegrained parallelism in objectoriented systems. Thesis (Ph.D. in Computer Science), Department of Computer Science, University of California, Berkeley, Berkeley, CA, USA, 1998.

Weissman:1998:PCS

[Wei98b]  Boris Weissman.  Perfor-

mance counters and state sharing annotations: a unified approach to thread locality. ACM SIGPLAN Notices, 33(11):127–138, November 1998. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/asplos/291069/
p127-weissman/. Co-published

in Operating Systems Review, 32(5).

Wong:1994:SSI

[WG94] W. F. Wong and E. Goto.

A simulation study on the interactions between multithreaded architectures and the cache. International Journal of High Speed Computing (IJHSC), 6(2):343–??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Weissman:1999:HPT

[WG99] B. Weissman and B. Gomes.

High performance thread migration on clusters of SMPs. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999. CO-DEN ???? ISSN 1097-

2803. URL http://www.cs.
okstate.edu/~pdcp/vols/
vol02/vol02no2abs.html#
boris.

Walcott:2007:DPA

[WHG07] Kristen R. Walcott, Greg Humphreys, and Sudhanva Gurumurthi. Dynamic prediction of architectural vulnerability from microarchitectural state. ACM SIGARCH Computer Architecture News, 35 (2):516–527, May 2007. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

White:2003:UTL

[Whi03] Tom White. Using threadlocal variables in Java. Dr. Dobb’s Journal of Software Tools, 28(7):42, 44–46, July 2003. CODEN DDJOEB. ISSN 1044-789X. URL

                                        http://www.ddj.com/ftp/
                                        2003/2003_07/thread.txt;
                                         http://www.ddj.com/ftp/
                                        2003/2003_07/thread.zip.

Wallach:1995:OAM

[WHJ^{+}95] Deborah A. Wallach, Wilson C. Hsieh, Kirk L. Johnson, M. Frans Kaashoek, and William E. Weihl. Optimistic active messages: a mechanism for scheduling communication with computation. ACM SIG-PLAN Notices, 30(8):217–226, August 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Williams:1994:NST

[Wil94a]  Al Williams.  NT-Style

threads for MS-DOS. Dr. Dobb’s Journal of Software Tools, 19(2):74, 76–77, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Williams:1994:NTM

[Wil94b]  Al Williams.  NT-Style

threads for MS-DOS. Dr. Dobb’s Journal of Software Tools, 19(2):74, 76–77, February 1994. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1997:BTP

[Wil97]  Greg Wilson.  Bookshelf:

Threads primer: a guide to multithreaded programming. IEEE Software, 14 (5):116, September/October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:

//dlib.computer.org/so/
books/so1997/pdf/s5115.
pdf.

Wilmot:1998:DTM

[Wil98] Dick Wilmot. Data threaded microarchitecture. ACM SIGARCH Computer Architecture News, 26(5):22– 32, December 1998. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Wilson:2000:PBC

[Wil00] Gregory V. Wilson. Programmer’s bookshelf: Classics old and new. Dr. Dobb’s Journal of Software Tools, 25(11):159– 160, November 2000. CODEN DDJOEB. ISSN 1044-789X.

Wei:2012:OLL

[WJ12] Zheng Wei and Joseph Jaja.

Optimization of linked list prefix computations on multithreaded GPUs using CUDA. Parallel Processing Letters, 22 (4):1250012, December 2012. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Wang:2019:MEM

[WJA^{+}19] L. Wang, M. Jahre, A. Adileh, Z. Wang, and L. Eeckhout. Modeling emerging memorydivergent GPU applications. IEEE Computer Architecture Letters, 18(2):95–98, July 2019. ISSN 1556-6056 (print), 1556-6064 (electronic).

Wegiel:2008:MCVa

[WK08a] Michal Wegiel and Chandra Krintz. The mapping collector: virtual memory support for generational, parallel, and concurrent compaction. ACM SIGARCH Computer Architecture News, 36(1):91– 102, March 2008. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Wegiel:2008:MCVb

[WK08b] Michal Wegiel and Chandra Krintz. The Mapping Collector: virtual memory support for generational, parallel, and concurrent compaction. Operating Systems Review, 42(2): 91–102, March 2008. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wegiel:2008:MCVc

[WK08c] Michal Wegiel and Chandra Krintz. The mapping collector: virtual memory support for generational, parallel, and concurrent compaction. ACM SIGPLAN Notices, 43(3):91–102, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2017:JRJ

[WKG17] Kaiyuan Wang, Sarfraz Khurshid, and Milos Gligoric. JPR: Replaying JPF traces using standard JVM. ACM SIGSOFT Software Engineering Notes, 42(4):1–5, October 2017. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wadden:2014:RWD

[WLG^{+}14] Jack Wadden, Alexander Lyashevsky, Sudhanva Gurumurthi, Vilas Sridharan, and Kevin Skadron. Realworld design and evaluation of compiler-managed GPU redundant multithreading. ACM SIGARCH Computer Architecture News, 42 (3):73–84, June 2014. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Wang:2009:TDA

[WLK^{+}09] Yin Wang, Stéphane Lafortune, Terence Kelly, Manjunath Kudlur, and Scott Mahlke. The theory of deadlock avoidance via discrete control. ACM SIGPLAN Notices, 44(1):252–263, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Won:2015:MMC

[WLM15] Youjip Won, Kyeongyeol Lim, and Jaehong Min. MUCH: Multithreaded content-based file chunking. IEEE Transactions on Computers, 64(5): 1375–1388, ???? 2015. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2019:HUI

[WLT19] Jimmy Ming-Tai Wu, Jerry Chun-Wei Lin, and Ashish Tamrakar. High-utility itemset mining with effective pruning strategies. ACM Transactions on Knowledge Discovery from Data (TKDD), 13(6):58:1–58:??, December 2019. CODEN ???? ISSN 1556-4681 (print), 1556-472X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3363571.

Watcharawitch:2003:MME

[WM03] Panit Watcharawitch and Simon W. Moore. Mul-TEP: MulTithreaded Embedded Processors. In Anonymous [Ano03], page ?? ISBN ???? LCCN ????

Wendykier:2010:PCH

[WN10] Piotr Wendykier and James G. Nagy. Parallel Colt: a high-performance Java library for scientific computing and image processing.

ACM Transactions on Mathematical Software, 37(3):31:1– 31:22, September 2010. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Wismuller:1996:IDP

[WOKH96] Roland Wismüller, Michael Oberhuber, Johann Krammer, and Olav Hansen. Interactive debugging and performance analysis of massively parallel applications. Parallel Computing, 22(3):415–442, April 29, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
parco/cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=3&aid=1049.

Welch:2010:SCF

[WP10] Peter H. Welch and Jan B.

Pedersen. Santa Claus: Formal analysis of a processoriented solution. ACM Transactions on Programming Languages and Systems, 32 (4):14:1–14:37, April 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Wang:2018:TWB

[WQLJ18] Jui-Hsien Wang, Ante Qu, Timothy R. Langlois, and Doug L. James. Toward wave-based sound synthesis for computer animation. ACM Transactions on Graphics, 37(4):109:1–109:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Watt:2019:WW

[WRPP19] Conrad Watt, Andreas Rossberg, and Jean Pichon-Pharabod. Weakening WebAssembly. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):133:1–133:28, October 2019. URL https:

//dl.acm.org/doi/abs/10.
1145/3360559.

Wang:2006:RAA

[WS06] L. Wang and S. D. Stoller. Runtime analysis of atomicity for multithreaded programs. IEEE Transactions on Software Engineering, 32(2):93–110, February 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1599419.

Warg:2008:DTS

[WS08] Fredrik Warg and Per Stenstrom. Dual-thread speculation: a simple approach to uncover thread-level parallelism on a simultaneous multithreaded processor. International Journal of Parallel Programming, 36(2):166–183, April 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/

openurl.asp?genre=article&
issn=0885-7458&volume=36&
issue=2&spage=166.

Whittaker:1997:TML

[WSKS97] Steve Whittaker, Jerry Swanson, Jakov Kucan, and Candy Sidner. TeleNotes: managing lightweight interactions in the desktop. ACM Transactions on Computer-Human Interaction, 4(2):137–168, June 1997. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tochi/1997-4-2/
p137-whittaker/.

Wheeler:2010:VMM

[WT10] Kyle B. Wheeler and Douglas Thain. Visualizing massively multithreaded applications with ThreadScope. Concurrency and Computation: Practice and Experience, 22 (1):45–67, January 2010. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2012:SPA

[WTH^{+}12] Jingyue Wu, Yang Tang, Gang Hu, Heming Cui, and Junfeng Yang. Sound and precise analysis of parallel programs through schedule specialization. ACM SIGPLAN Notices, 47(6):205–216, June 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PLDI ’12 proceedings.

Wong:2008:TAF

[WTKW08] Chee Siang Wong, Ian Tan, Rosalind Deena Kumari, and Fun Wey. Towards achieving fairness in the Linux scheduler. Operating Systems Review, 42(5):34–43, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Waldspurger:1993:RRF

[WW93] Carl A. Waldspurger and William E. Weihl. Register relocation: flexible contexts for multithreading. ACM SIGARCH Computer Architecture News, 21(2):120–130, May 1993. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Wise:1996:SDP

[WW96] David S. Wise and Joshua Walgenbach. Static and dynamic partitioning of pointers as links and threads. ACM SIGPLAN Notices, 31(6):42– 49, June 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2002:SPE

[WWW^{+}02] Hong Wang, Perry H. Wang, Ross Dave Weldon, Scott M. Ettinger, Hideki Saito, Milind Girkar, Steve Shih wei Liao, and John P. Shen. Speculative precomputation: Exploring the use of multithreading for latency tools. Intel Technology Journal, 6(1):22–35, February 2002. ISSN 1535-766X. URL

http://developer.intel.
com/technology/itj/2002/
volume06issue01/vol6iss1_
hyper_threading_technology.
pdf.

Wang:2022:ASM

[WXAL22] Zhe Wang, Chen Xu, Kunal Agrawal, and Jing Li. Adaptive scheduling of multiprogrammed dynamic-multithreading applications. Journal of Parallel and Distributed Computing, 162(??):76–88, April 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0743731522000144.

Wenjie:2020:APW

[WYT^{+}20] Tang Wenjie, Yao Yiping, Li Tianlin, Song Xiao, and Zhu Feng. An adaptive persistence and work-stealing combined algorithm for load balancing on parallel discrete event simulation. ACM Transactions on Modeling and Computer Simulation, 30(2): 12:1–12:26, April 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3364218.

Wang:2019:SSS

[WZSK19] Wenlu Wang, Ji Zhang, Min-Te Sun, and Wei-Shinn Ku. A scalable spatial skyline evaluation system utilizing parallel independent region groups. VLDB Journal: Very Large Data Bases, 28(1):73– 98, February 2019. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wang:2008:PIM

[WZWS08] Kun Wang, Yu Zhang, Huayong Wang, and Xiaowei Shen. Parallelization of IBM Mambo system simulator in functional modes. Operating Systems Review, 42(1):71–76, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xu:2006:RTR

[XHB06] Min Xu, Mark D. Hill, and Rastislav Bodik. A regulated transitive reduction (RTR) for longer memory race recording. ACM SIGPLAN Notices, 41(11):49–60, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xekalakis:2012:MSM

[XIC12] Polychronis Xekalakis, Nikolas Ioannou, and Marcelo Cintra. Mixed speculative multithreaded execution models. ACM Transactions on Architecture and Code Optimization, 9(3):18:1–18:??, September 2012. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Xu:1999:DIT

[XMN99] Zhichen Xu, Barton P. Miller, and Oscar Naim. Dynamic instrumentation of threaded applications. ACM SIGPLAN Notices, 34(8):49–59, August 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
proceedings/ppopp/301104/
p49-xu/.

Xian:2008:CAS

[XSaJ08] Feng Xian, Witawas Srisa-an, and Hong Jiang. Contentionaware scheduler: unlocking execution parallelism in multithreaded Java programs. ACM SIGPLAN Notices, 43(10):163–180, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xue:2012:RJC

[Xue12] Jingling Xue. Rethinking Java call stack design for tiny embedded devices. ACM SIGPLAN Notices, 47(5):1– 10, May 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCTES ’12 proceedings.

Xu:2014:STM

[XWG^{+}14] Yunlong Xu, Rui Wang, Nilanjan Goswami, Tao Li, and Depei Qian. Software transactional memory for GPU architectures. IEEE Computer Architecture Letters, 13(1):49– 52, January/June 2014. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Yam:1995:CFD

[Yam95] Michael Yam. A C++ framework for DCE threads. Dr. Dobb’s Journal of Software Tools, 20(??):27–??, July/ August 1995. CODEN DDJOEB. ISSN 1044-789X.

Yam:1996:DPV

[Yam96] Michael Yam. DCE pthreads versus NT threads. Michael ports PTF, a C++ class library for DCE pthreads, from HP-UX System 9 to Windows NT. In doing so, he examines the differences between pthreads and NT threads, and describes the porting experience. Dr. Dobb’s Journal of Software Tools, 21(12):16– ??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Yang:1997:MUA

[Yan97] Chia Wei Yang. A multicontext uniprocessor: another multithreaded architecture. Thesis (M.S.), California Polytechnic State University, San Luis Obispo, CA, USA, 1997. viii + 129 pp.

Yan:2002:RCC

[Yan02] C. Yan. Race condition and concurrency safety of multithreaded object-oriented programming in Java. IEEE International Conference on Systems Man and Cybernetics, 6:??, 2002. CODEN ???? ISSN 1062-922X.

Yasrebi:1995:EDO

[Yas95]  M. Yasrebi.  Experience

with distributed objects in a portable and multithreaded library for a LAN/WAN gateway application. In IEEE Computer Society. Technical Committee on Computer Communications [IEE95], pages 164–173. ISBN 0-8186-7163-7 (microfiche), 0-8186-7162-9. LCCN TK5105.7 .C66 1995 Bar. IEEE Computer Society Press order number PR07162. IEEE catalog number 95TB100005.

Yiapanis:2016:CDS

[YBL16] Paraskevas Yiapanis, Gavin Brown, and Mikel Luján. Compiler-driven software speculation for thread-level parallelism. ACM Transactions on Programming Languages and Systems, 38(2):5:1–5:??, January 2016. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Yang:2014:MPP

[YCW^{+}14] Junfeng Yang, Heming Cui, Jingyue Wu, Yang Tang, and Gang Hu. Making parallel programs reliable with stable multithreading. Communications of the ACM, 57(3):

58–69, March 2014. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yin:2020:SCA

[YDLW20] L. Yin, W. Dong, W. Liu, and J. Wang. On scheduling constraint abstraction for multithreaded program verification. IEEE Transactions on Software Engineering, 46(5):549– 565, 2020. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Yee:2020:CMT

[Yee20] Alexander J. Yee. y-cruncher: a multi-threaded pi-program. Web site, March 30, 2020.

                                        URL http://www.numberworld.
                                        org/y-cruncher/.

Yamashita:2012:APS

[YFF^{+}12] Makoto Yamashita, Katsuki Fujisawa, Mituhiro Fukuda, Kazuhide Nakata, and Maho Nakata. Algorithm 925: Parallel solver for semidefinite programming problem having sparse Schur complement matrix. ACM Transactions on Mathematical Software, 39(1): 6:1–6:22, November 2012. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Yi:2010:NAS

[YG10] Kyueun Yi and J.-L. Gaudiot. Network applications on simultaneous multithreading processors. IEEE Transactions on Computers, 59(9):1200–1209, September 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=5374374.

Yu:2013:GDS

[YKL13] Hongtao Yu, Hou-Jen Ko, and Zhiyuan Li. General data structure expansion for multithreading. ACM SIGPLAN Notices, 48(6):243–252, June 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yao:2016:OCO

[YL16] Yuan Yao and Zhonghai Lu. Opportunistic competition overhead reduction for expediting critical section in NoC based CMPs. ACM SIGARCH Computer Architecture News, 44(3): 279–290, June 2016. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Yu:2016:DLR

[YLLS16] Hairong Yu, Guohui Li, Jianjun Li, and Lihchyun Shu. DO_{cyclical}: a latency-resistant cyclic multi-threading approach for automatic program parallelization. The Computer Journal, 59(8):1155–1173, August 2016. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/content/59/8/1155.

Young-Myers:1992:DTC

[YM92] Helene Wen-Hsin Young-Myers. Database transitive closure: a performance study of multithreaded algorithms. Thesis (Ph.D.), College of Business and Management, University of Maryland at College Park, College Park, MD, USA, 1992. ix + 198 pp.

Young-Myers:1993:ESTa

[YMR93a] Helene Young-Myers and Louiqa Raschid. An experimental study of three dataflow paradigms in multithreaded database transitive closure algorithms on shared memory multiprocessors. Technical report CS-TR-3060; UMIACS-TR-93-33, University of Maryland, College Park, MD, USA, April 1993. 21 pp. To appear in a special issue of the Journal of Parallel and Distributed Computing on Dataflow and Multithreaded Architectures, July, 1993.

Young-Myers:1993:ESTb

[YMR93b] Helene Young-Myers and Louiqa Raschid. An experimental study of three dataflow paradigms in multithreaded database transitive closure algorithms on shared memory multiprocessors. Journal of Parallel and Distributed Computing, 18(3):371–389, July 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1071/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1071/production/pdf.

Yu:2009:CIC

[YN09] Jie Yu and Satish Narayanasamy.

A case for an interleaving constrained shared-memory multi-processor. ACM SIGARCH Computer Architecture News, 37(3):325–336, June 2009. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Yu:2012:MCD

[YNPP12] Jie Yu, Satish Narayanasamy, Cristiano Pereira, and Gilles Pokam. Maple: a coveragedriven testing tool for multithreaded programs. ACM SIGPLAN Notices, 47(10): 485–502, October 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yoo:1996:CAA

[Yoo96a]  H. Chuck Yoo.  Com-

parative analysis of asynchronous I/O in multithreaded UNIX. Software— Practice and Experience, 26 (9):987–997, September 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        16832.

Yoo:1996:PCM

[Yoo96b] Namhoon Yoo. Parallelism control in multithreaded multiprocessors. Thesis (Ph.D.), University of Southern California, Los Angeles, CA, USA, 1996. x + 86 pp.

Yeh:2017:PFG

[YSS^{+}17] Tsung Tai Yeh, Amit Sabne, Putt Sakdhnagool, Rudolf Eigenmann, and Timothy G. Rogers. Pagoda: Fine-grained GPU resource virtualization for narrow tasks. ACM SIG-PLAN Notices, 52(8):221–234, August 2017. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yeh:2019:PGR

[YSS^{+}19] Tsung Tai Yeh, Amit Sabne, Putt Sakdhnagool, Rudolf Eigenmann, and Timothy G. Rogers. Pagoda: a GPU runtime system for narrow tasks. ACM Transactions on Parallel Computing (TOPC), 6(4): 21:1–21:??, November 2019. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Youseff:2009:PES

[YSY^{+}09] Lamia Youseff, Keith Seymour, Haihang You, Dmitrii Zagorodnov, Jack Dongarra, and Rich Wolski. Paravirtualization effect on singleand multi-threaded memoryintensive linear algebra software. The Journal of Networks, Software Tools, and Cluster Computing, 12(2): 101–122, ???? 2009. ISSN 1386-7857.

Yong:2003:AMC

[YWJ03] Xie Yong and Hsu Wen-Jing.

Aligned multithreaded computations and their scheduling with FAB performance guarantees. Parallel Processing Letters, 13(3):353– ??, September 2003. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Yan:2007:HMC

[YZ07] Jun Yan and Wei Zhang.

Hybrid multi-core architecture for boosting singlethreaded performance. ACM SIGARCH Computer Architecture News, 35(1):141– 148, March 2007. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Yang:2014:CNR

[YZ14] Yi Yang and Huiyang Zhou.

CUDA-NP: realizing nested thread-level parallelism in GPGPU applications. ACM SIGPLAN Notices, 49(8):93– 106, August 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yang:2007:RUL

[YZYL07] Jin-Min Yang, Da-Fang Zhang, Xue-Dong Yang, and Wen-Wei Li. Reliable user-level rollback recovery implementation for multithreaded processes on windows. Software —Practice and Experience, 37 (3):331–346, March 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zoppetti:2001:IDD

[ZAK01] Gary Zoppetti, Gagan Agrawal, and Rishi Kumar. Impact of data distribution on performance of irregular reductions on multithreaded architectures. Lecture Notes in Computer Science, 2110:483– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2110/21100483.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2110/21100483.
                                        pdf.

Zhang:2015:DMB

[ZBS15] Junchao Zhang, Babak Behzad, and Marc Snir. Design of a multithreaded Barnes– Hut algorithm for multicore clusters. IEEE Transactions on Parallel and Distributed Systems, 26(7):1861– 1873, July 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://www.
computer.org/csdl/trans/
td/2015/07/06837521-abs.
html.

Zhang:2010:FTS

[ZCO10] Yao Zhang, Jonathan Cohen, and John D. Owens. Fast tridiagonal solvers on the GPU. ACM SIGPLAN Notices, 45(5):127–136, May 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhai:2002:COSa

[ZCSM02a] Antonia Zhai, Christopher B.

Colohan, J. Gregory Steffan, and Todd C. Mowry. Compiler optimization of scalar value communication between speculative threads. ACM SIGARCH Computer Architecture News, 30(5):171–183, December 2002. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Zhai:2002:COSb

[ZCSM02b] Antonia Zhai, Christopher B.

Colohan, J. Gregory Steffan, and Todd C. Mowry. Compiler optimization of scalar value communication between speculative threads. ACM SIG-PLAN Notices, 37(10):171– 183, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhong:2019:SHS

[ZDTM19] Guanwen Zhong, Akshat Dubey, Cheng Tan, and Tulika Mitra. Synergy: an HW/SW framework for high throughput CNNs on embedded heterogeneous SoC. ACM Transactions on Embedded Computing Systems, 18(2):13:1–13:??, April 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3301278.

Zhou:1998:LST

[ZG98] Honbo Zhou and Al Geist. LPVM: a step towards multithread PVM. Concurrency: Practice and Experience, 10 (5):407–416, April 25, 1998. CODEN CPEXEI. ISSN 1040-3108. URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?ID=
5385;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5385&PLACEBO=IE.pdf.

Zhang:2000:WMH

[Zha00] Peter Zhang. Webrelay: a multithreaded HTTP relay server. Dr. Dobb’s Journal of Software Tools, 25(2): 86, 88, 90–94, 96, February 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_02/webrelay.txt;
http://www.ddj.com/ftp/
2000/2000_02/webrelay.zip.

Zhang:2015:LOS

[ZHCB15] Minjia Zhang, Jipeng Huang, Man Cao, and Michael D. Bond. Low-overhead software transactional memory with progress guarantees and strong semantics. ACM SIG-PLAN Notices, 50(8):97–108, August 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zignin:1996:TDM

[Zig96] Bernard Zignin. Techniques du multithread: du parallèlisme dans les processus (French) [Multithreading techniques: parallelism in processes]. CNAM. Syntheses informatiques. Hermes, Paris, France, 1996. ISBN 2-86601-562-2. 72 pp. LCCN ????

Ziarek:2009:SWB

[ZJFA09] Lukasz Ziarek, Suresh Jagannathan, Matthew Fluet, and Umut A. Acar. Speculative N-way barriers (abstract only). ACM SIG-PLAN Notices, 44(5):8, May 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2010:DCS

[ZJS10] Eddy Z. Zhang, Yunlian Jiang, and Xipeng Shen. Does cache sharing on modern CMP matter to the performance of contemporary multithreaded programs? ACM SIGPLAN Notices, 45(5):203– 212, May 2010. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhu:2011:TPS

[ZJS^{+}11] David (Yu) Zhu, Jaeyeon Jung, Dawn Song, Tadayoshi Kohno, and David Wetherall. TaintEraser: protecting sensitive data leaks using application-level taint tracking. Operating Systems Review, 45(1):142–154, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2012:SCC

[ZJS12] Eddy Zheng Zhang, Yunlian Jiang, and Xipeng Shen. The significance of CMP cache sharing on contemporary multithreaded applications. IEEE Transactions on Parallel and Distributed Systems, 23(2): 367–374, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2011:DCC

[ZKR^{+}11] Qin Zhao, David Koh, Syed Raza, Derek Bruening, Weng-Fai Wong, and Saman Amarasinghe. Dynamic cache contention detection in multithreaded applications. ACM SIGPLAN Notices, 46(7):27– 38, July 2011. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2015:DPO

[ZKW15] Naling Zhang, Markus Kusano, and Chao Wang. Dynamic partial order reduction for relaxed memory models. ACM SIGPLAN Notices, 50(6):250–259, June 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zier:2010:PED

[ZL10] David A. Zier and Ben Lee.

Performance evaluation of dynamic speculative multithreading with the Cascadia architecture. IEEE Transactions on Parallel and Distributed Systems, 21(1):47–59, January 2010. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2016:TED

[ZLJ16] Tong Zhang, Dongyoon Lee, and Changhee Jung. TxRace: Efficient data race detection using commodity hardware transactional memory. Operating Systems Review, 50(2): 159–173, June 2016. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2016:SAN

[ZLW^{+}16] Mingzhe Zhang, Francis C. M.

Lau, Cho-Li Wang, Luwei Cheng, and Haibo Chen. Scalable adaptive NUMA-aware lock: combining local locking and remote locking for efficient concurrency. ACM SIG-PLAN Notices, 51(8):50:1– 50:??, August 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zebchuk:2007:BBC

[ZM07] J. Zebchuk and A. Moshovos. A building block for coarsegrain optimizations in the onchip memory hierarchy. IEEE Computer Architecture Letters, 6(2):33–36, February 2007. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Zhuang:2004:BRA

[ZP04] Xiaotong Zhuang and Santosh Pande. Balancing register allocation across threads for a multithreaded network processor. ACM SIGPLAN Notices, 39(6):289–300, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhuang:2011:CST

[ZP11] Xiaotong Zhuang and Santosh Pande. Compiler-supported thread management for multithreaded network processors. ACM Transactions on Embedded Computing Systems, 10(4):44:1–44:??, November 2011. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zarrabi:2013:LSF

[ZSA13] Amirreza Zarrabi, Khairulmizam Samsudin, and Wan Azizun Wan Adnan. Linux support for fast transparent general purpose checkpoint/ restart of multithreaded processes in loadable kernel module. Journal of Grid Computing, 11(2):187–210, June 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10723-013-9248-5.

Zhuravlev:2012:SST

[ZSB^{+}12] Sergey Zhuravlev, Juan Carlos Saez, Sergey Blagodurov, Alexandra Fedorova, and Manuel Prieto. Survey of scheduling techniques for addressing shared resources in multicore processors. ACM Computing Surveys, 45(1): 4:1–4:??, November 2012. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Ziarek:2006:SMC

[ZSJ06] Lukasz Ziarek, Philip Schatz, and Suresh Jagannathan. Stabilizers: a modular checkpointing abstraction for concurrent functional programs. ACM SIGPLAN Notices, 41 (9):136–147, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zois:2019:EMM

[ZTN19] Vasileios Zois, Vassilis J. Tsotras, and Walid A. Najjar. Efficient main-memory topk selection for multicore architectures. Proceedings of the VLDB Endowment, 13(2): 114–127, October 2019. CO-DEN ???? ISSN 2150-8097.

Zuberek:2002:APB

[Zub02]  W. M. Zuberek.  Anal-

ysis of performance bottlenecks in multithreaded multiprocessor systems. Fundamenta Informaticae, 50(2): 223–241, February 2002. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zou:2022:BSP

[ZWGX22] Changwei Zou, Xudong Wang, Yaoqing Gao, and Jingling Xue. Buddy stacks: Protecting return addresses with efficient thread-local storage and runtime rerandomization. ACM Transactions on Software Engineering and Methodology, 31 (2):35e:1–35e:37, April 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3494516.

Zheng:2015:ACC

[ZWL15] Zhong Zheng, Zhiying Wang, and Mikko Lipasti. Adaptive cache and concurrency allocation on GPGPUs. IEEE Computer Architecture Letters, 14(2):90–93, July/December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).